Olympia: The Age of Gods

Rules of Play

by Scott R. Turner and Rich Skrenta Questions and comments: srt@pbm.com

Distribution

Olympia: The Age of Gods, its rules, documentation and game text are Copyright \begin{matrix}c \\ ◯ \end{matrix} 1998 Scott R. Turner. All rights reserved.

Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

1 Introduction

Welcome to Olympia, the Age of Gods!

Olympia is an closed-ended computer moderated fantasy simulation. Characters move, battle, explore and study in the Olympian world. Each week, players submit orders for their units. After the turn runs, Olympia sends reports to the players detailing what happened.

Olympia turn reports tend to be long, detailed and dense. Since your player character may hire other units which can study, move and fight on their own, large factions of characters may be organized. Supplying orders for many units in a large faction can be a lot of work!

A turn report for a new player controlling three characters is about five (66 line) pages. Average turn reports tend to be 15-25 pages. The size of the report will depend on whether a player chooses to concentrate development in a few characters, or to pursue an empire- building strategy of acquiring many units and controlling as much territory as possible.

1.1 Overview

Each Olympia:The Age of Gods game is based on a scenario, which defines the rules of the game.

Each player joins as a member of a nation. The player begins with control of one noble and a sum of gold, the Olympian currency. Nobles may hire subordinates, study skills, trade items, build things, and explore. Each turn a player submits orders for all of the nobles they control.

The Olympia program will execute orders for noble in parallel. Each order may take no time or some number of game days to complete. As many orders will be processed for each noble as game time allows. Orders not completed by the end of the turn may continue into the next. Each turn covers one month of game time. An Olympian month lasts 30 game days.

1.2 Important terms

Entity Unit Everything in the Olympian world has a unique code. Nobles,

items, locations, and skills all are referenced with an entity number. The code is shown in brackets after the name. Some examples:

Rich Skrenta [xt8] a player

Osswid the Destroyer [5499] a noble Shipcraft [1000]

a skill

City of the Lost [gx14] a place

Gold [1] an item

Scroll [yq12] an item

Noble Character Used interchangeably. These are the individuals under the con-

trol of players. All player orders are given to nobles. Players start with one noble. Others may be hired or persuaded to join the player’s faction. Nobles may possess items, travel through locations, learn skills, engage in combat, cast magical spells, etc.

Faction All of the units controlled by a player are called the player’s faction. A player starts with only one noble, but the faction may grow to have many units.

Player character The player character, or PC, is the character the player starts with. The PC begins with a loyalty of oath-2. The PC may later FORM other characters. Nothing is special about the PC other than being the player’s first noble; if the PC is killed, play continues with the player’s other nobles.

Nation Each faction belongs to a particular nation, such as the "Imperial Empire". Nations are political structures that are usually in conflict with each other.

Scenario Each game of Olympia:The Age of Gods is based upon a particular scenario. The scenario defines what nations are available, what skills are available, the win conditions, and so on. The scenario for a particular game is describe on the game’s web page, and in the newsletter. For more information, see ⟨undefined⟩ [Scenarios], page ⟨undefined⟩.

Item Possession Nobles may hold items, such as gold, scrolls, weapons, magic

potions, jewels, lumber, rugs, etc.

Men Nobles may also have non-descript men in their employ. These men are represented as possessions for simplicity. They include peasants, workers, sailors, and different kinds of soldiers. These men may not learn skills, hold any items, or act independently from the noble they are with. For example, one might see:

    Seen here:
       Law Netexus [2020], with three peasants

Law Netexus is a noble; the three peasants are non-descript men accompanying him. Nobles obtain peasants with the recruit order.

Skills Nobles may learn skills, which are used to perform tasks. For instance, Sailing [1001] must be known in order to sail a ship. Skills are grouped into the following categories:

• Alchemy

• Combat

• Construction

• Forestry

• Mining

• Persuasion

• Shipcraft

• Stealth

• Ranger

• Trading

There are also six schools of magic, and eight churches. For information about learning skills, see ⟨undefined⟩ [study], page ⟨undefined⟩.

Noble Points A player starts with some number of Noble Points (NP). Each player gets an additional NP every eight turns (one each game year). NPs are used to buy nobles with the form command. They are also required to learn some advanced skills, and to swear nobles to oath loyalty.

Stack A group of nobles joined such that they move and fight together.

Province A location on the map. Provinces may have sub-locations within them, such as cities, bogs, caves, etc. Provinces are either forest, swamp, mountain, desert, plains, or ocean.

Month Each turn is a game month, or 30 game days.

Safe Haven Some cities are designated safe havens. Combat, magic and building are not allowed in safe havens. New players may choose to start in a safe haven so that they have some chance to acclimate themselves before becoming a target for existing players.

Olympian Calendar The Olympia calendar has two months for each season, for a total of eight months per Olympian year. Each month is 30 game days long.

Fierce winds Spring

Snowmelt Spring

Blossom bloom Summer

Sunsear Summer

Thunder and rain Fall

Harvest Fall

Waning days Winter

Dark night Winter

1.3 Game Mechanics

This section describes how games are set up, how a player submits orders, receives turn reports, the schedule of games, and other important details about how to play TAG.

1.3.1 Scenarios

A scenario defines the rules for a particular game of Olympia:The Age of Gods. Several different scenarios are available, and are described here. The scenarios for current games is listed on the Olympia:The Age of Gods web page, and in the newsletters for each game.

There are also several "alternate rules" available which do not change the basic scenarios, but provided a slightly different flavor to the game. If a particular game uses any of the alternate rules, that will be noted in both the web pages and the newsletter. (See also ⟨undefined⟩ [Alternate Rules], page ⟨undefined⟩.)

1.3.1.1 The Age of Gods Scenario

The Age of Gods scenario takes place at the time in Olympia history when the Imperial and Mandor empires were re-discovering each other. (See ⟨undefined⟩ [Appendix - History], page ⟨undefined⟩.) Players have the choice of joining the Imperial Empire, the Mandor Empire, or the Pen. The game is won when one nation conquers the rest (see ⟨undefined⟩ [Winning the Game], page ⟨undefined⟩.

1.3.1.2 Novice Scenario

The novice scenario takes place in a small world with only one nation. Players are limited to 12 turns of free play. This scenario is intended to give the new TAG player some experience with the game.

1.3.1.3 The Never-Ending Kingdom Scenario

The Never-Ending Kingdom scenario takes place long after the Age of Gods. Olympia has been unified into a single nation, which after a long fall into barbarism is now setting out to explore and conquer the world. All skills are available to all players. This is an open-ended scenario.

1.3.1.4 The Twelve Kings Scenario

Twelve islands in the Endless Sea surround an extinct volcano. On each island reigns a single king. Prophets have foretold a great disaster if the islands are not unified under a single king by the turn of the millenium in ten years (80 turns). Unfortunately, no king is willing to let another wear his crown...

This scenario is limited to twelve players. Each player begins with a castle, a city, and his own island. Each player has access to all the skills and one religion of his choice. Normal win conditions apply, with the addition that if no one wins by turn 80, the game ends and all are deemed to have lost.

1.3.2 Alternate Rules

Some TAG games will have particular rules turned off to create a different atmosphere or challenge for the player. For example, the conflict between priests and magicians may be disabled, to set a TAG game in an era where priests and magicians are peaceful. This section describes the rule options available and what they affect.

1.3.2.1 Mage/Priest Antipathy

The antipathy between mages and priests can be turned off. This has the following effects:

Stacks containing priests do not receive a combat attack bonus against stacks containing a magician or undead.

Bless Follower can be used on nobles possessing undead.

The Staff of the Sun is not in play.

Magicians in a province do not prevent peasants from making offerings at temples.

Priests do not receive a piety bonus for performing Last Rites on a magician.

Magicians do not receive an aura bonus for eating the body of a priest.

Priests and magicians may stack together.

1.3.2.2 Turn Limit

Some TAG games have a turn limit. After a player reaches this number of turns, he is automatically dropped from the game.

This is used primarily in the free novice game.

1.3.2.3 Autodrop

Some TAG games automatically drop a player if he does not submit orders for his units for 3 turns in a row. This is used primarily in free games, to limit the number of inactive players.

1.3.3 Winning the Game

Some TAG scenarios are "close-ended", which means there is a specific way that a player (or group of players) can win the game and cause it to end. Other scenarios are "open-ended", which means they will continue until no one is interested in playing any longer.

Most closed-ended TAG scenarios use the same win condition. The game ends when one nation conquers the world. In particular, to win the game a nation must achieve the following conditions for two months in a row:

(1) No other nations may control any provinces or cities. Control of a province or city is given by garrisoning the province.

(2) The winning nation must have twice as many NPs as all the other nations combined.

The win condition is checked at the end of each month.

If a scenario uses an alternate win condition, it will be stated in the scenario description (see ⟨undefined⟩ [Scenarios], page ⟨undefined⟩).

1.3.4 Submitting orders

Orders should be sent to orders@game.pbm.com. The game part of this address will normally be something like "tag7" and will be published in the game newsletter.

The ‘Reply-To:’ header on turn reports is set to the correct address, so using the reply feature of your mailer should send orders to the right place.

The orders address is read by an automated daemon which scans the orders for syntax errors and queues them for your units. It will send a reply as soon as it receives mail from you, showing whether or not there were any errors with the orders it received.

Orders must be of the following form:

begin player-number password

      .
      .  email, password commands
      .

unit player-number

      .
      .  commands for player entity:  name, format or quit
      .

unit unit-number

      .
      .  commands for unit
      .

unit unit-number

      .
      .  commands for unit
      .
    end

The ‘Subject:’ line on your message is ignored.

The begin keyword tells the order scanner what your player number is. If you have not set a password, you do not need to supply one.

The unit command replaces a set of orders for a unit. Any pending orders for the unit will be cleared, and the new orders sent in will queue up. Orders that are still executing for the unit will not be interrupted unless the first order queued is the stop order (see ⟨undefined⟩ [stop], page ⟨undefined⟩).

Do not match an end for every unit command! There should only be one end, at the end of all of the unit sections. The Olympia order parser will not read beyond the end.

For example, here is a set of orders for player Fate [812], who has two nobles, Osswid [5499] and Candide [1269]:

    begin 812
      password sneaky
    unit 5499
      explore
      move east
      explore
      study 1000
    unit 1269
      move north
      stack 5499
    end

The parser tries to be as flexible as possible. It is case insensitive and is not strict about spaces on a line, so you may use indentation to make your orders more readable.

Orders may be commented with the ‘#’ character. Everything from a ‘#’ to the end of the line will be ignored by the parser:

    move north    # Head to Drassa to meet up with Osswid
    stack 5499    # stack with him

No one will read the comments but you. Neither the GM nor the Olympia engine will try to interpret them for any reason.

Note that arguments must be enclosed in quotes if they are more than one word:

    name "Osswid the Constructor"

The acknowledgement will show any errors that occurred while the orders were being parsed, and list the current pending commands for all of your units.

1.3.4.1 Interrupting orders

Suppose the turn report shows the following orders queued:

    unit 5499
       # > study 2800 (executing for three more days)
       recruit 10
       explore

Sending in new orders for this unit will not disturb the still-running study command unless the first order is stop. For example:

If this were sent in:

    unit 5499 move south

This would be the result:

    unit 5499
       # > study 2800 (executing for three more days)
       move south

To interrupt the study and get on with the move right away, instead send in:

    unit 5499 stop move south

This will show:

    unit 5499
       # > study 2800 (executing for three more days)
       stop
       move south

Note that the stop queues like any other order; it does not actually interrupt the executing command until the turn runs. This means that the stop itself can be replaced by sending in another set of orders later.

1.3.4.2 Units not controlled by you (yet)

Orders may be sent in for units which are not yet under control, such as nobles that you intend to bribe or terrorize into switching to your faction.

As soon as the unit comes under your control, the orders queued for it will begin to execute.

Orders may also be sent in for new nobles which will be formed during the turn. First choose one of the possible unit numbers from the choices listed near the beginning of the turn report:

    The next five nobles formed will be:  5717 3215 4902 4489 5628

Supply one of these numbers as the first parameter to the form order:

    form 5628 "Feasel the Wicked"

Then queue some orders for Feasel to execute as soon as he appears:

    unit 5628 unstack study 2800 move out recruit ...

1.3.4.3 Use the order template

An order template listing all of the units for a player and showing any pending orders for those units, appears at the bottom of the turn sheet.

Order template
--------------
begin 812  # Master Bogomil's Family
unit 2508  # Tudor
   # > make 74 (still executing)
unit 2947  # Milo
unit 4375  # Beorn
   # > move s (executing for one more day)
   pillage
   recruit
unit 4763  # Sylvia
unit 5977  # Drango
   # > collect 87 0 0 (still executing)
   sail e
   sail s
   fish
   explore
unit 5418  # Comte de le Sang
end

Note that the layout of the order template matches the syntax the order scanner expects. Many players find it convenient to edit this template to add or change commands for their units. Mail everything from the begin to end (inclusive) to the order scanner. It is wise to save a copy of the orders you submit in case there are errors and they need to be resent.

1.3.4.4 Be careful

Against stupidity the very gods Themselves contend in vain.

Maid of Orleans, Schiller

Beware of sending in different sets of orders too quickly. Sometimes messages sent within a short time of each other will arrive out of order. This can wreak havoc on your turn if the wrong orders arrive last. Compose your orders offline and review them before mailing. A simple typographical error in your orders could ruin your whole turn!

1.3.4.5 Failed orders

Commands that fail generally take zero time. For instance, if study is issued for a skill which the location does not offer, it will fail immediately, and take zero time. The failed study order will not take a week, and it will not count toward the diminishing study returns for the month.

Production commands fail immediately if none of their input resources are available. For instance, recruit in a location with no peasants will immediately fail, taking zero time. However, resources may sometimes become depleted while the command is being executed. In such cases, the command may fail even after it has spent some time executing.

1.3.4.6 Player Entity

Each Olympia player faction has a number. This number is represented by an entity in the game. However, unlike a noble, this entity is mostly used as a place holder for the faction. No one can see the faction entity, and it can issue very few orders.

For example, suppose player Fate [xt8] has one noble Osswid [5499]. Fate does not exist in any location, so it does not receive a location report, and no one can see it. However, Osswid, being the player noble for the faction, is sworn to Fate [xt8]. Fate may execute only a limited set of administrative orders.

Nobles, not factions, issue most orders. Do not try to form or recruit with the player entity. For most turns, the player entity will have no orders queued for it.

The orders a player entity may issue are:

• ⟨undefined⟩ [admit], page ⟨undefined⟩

• ⟨undefined⟩ [format], page ⟨undefined⟩

• ⟨undefined⟩ [name], page ⟨undefined⟩

• ⟨undefined⟩ [quit], page ⟨undefined⟩

1.3.4.7 Quitting the Game

To drop out of the game, issue the quit order for your player entity. For instance, player xt8 would quit by sending in the following orders:

    begin xt8 password
    unit xt8               # don't forget unit for the player unit!
       quit
    end

No turn report will be sent for the turn in which a player quits.

1.3.4.8 More examples

You can replace orders for some units, but leave pending orders for other units alone, by only including unit sections for the ones you want to change.

If you want to see what orders are queued, but not change anything send in:

 begin 999               # whatever your player number is ...
 end

To change your email address, send in

 begin 999
 email new@address.com   # give your new address
 end

As a security measure, the confirmation will be sent to both the new and old addresses.

If you want to send your turn report to two or more addresses, use both addresses in the email command, separated with a comma:

 begin 999
 email new@address.com,other@address.com
 end

To change the name of your faction, issue the name order for the faction’s player entity:

 begin 999
 unit 999
    name "Seekers of Fame and Power"
    ...

Important: Don’t forget the unit command for the player entity.

1.3.4.9 Order of nobles in a location

When a noble enters a location, it is added to the end of the list of nobles already there. The unit that has been in a location the longest will appear at the top of the list.

If a noble leaves a location and later returns, it will be put at the end of the list again.

For example:

 Seen here:
    Candide [1269]
    Osswid [5499]
    Feasel the Wicked [1109]

Candide has been here longest, followed by Osswid, then Feasel. If Candide were to leave and return, he would appear at the end of the list.

If a noble unstacks from beneath another unit, the noble will appear just after the unit, rather than at the end of the list.

For example:

    Seen here:
       Candide [1269], accompanied by:
          Osswid [5499]
       Feasel the Wicked [1109]

Osswid is stacked beneath Candide. If Osswid unstacks, he will appear after Candide, not after Feasel:

    Seen here:
       Candide [1269]
       Osswid [5499]
       Feasel the Wicked [1109]

1.3.4.10 Command priority

All orders have a priority of 0-4.

• Permission commands (admit, hostile, etc.) are priority 0.

• Zero time commands^{1}, and wait, are priority 1.

• move and fly are priority 2.

• The sail and study commands are priority 4.

• All other commands are priority 3.

Primarily, the order scheduler processes commands on a day-by-day basis. Priority only orders commands within a day, and only between different nobles.

Each day, the order scheduler will first executes all priority 1 orders scheduled for that day. Only when all priority 1 orders are processed, will priority 2 orders for that day be processed, then priority 3 orders, etc. In other words, the order scheduler will not start an order at a higher priority when an order may be started at a lower priority.

Orders at the same priority are resolved in location order. If two units in a location are both waiting to start a move order, the first unit in the location will go first.

The above description of order priorities may seem complicated, but the intent is to let players ignore same-day synchronization issues in most cases. Rather that needing wait to guarantee that give happens before move, the lower priority of give makes this happen naturally.

For example, consider three units stacked together, top, mid and bot:

1 A zero time command is an order which always takes zero time. This does not include an order which may sometimes take zero time. For instance, unstack is always a zero time order. However, recruit is not, even though recruit may terminate immediately under some conditions.

top:    move ec69
        yew
mid:    unstack
        recruit
bot:    recruit

These should be executed in the following order:

mid: unstack                    # unstack is prio-1
top: move ec69                  # move is prio-2
mid: recruit                    # recruit is prio-3
[top and bot arrive at ec69]
top: yew                        # yew is prio-3
bot: recruit                    # recruit is prio-3

The unstack happened first since it’s a priority 1 command. The move went second. When top and bot finished moving, there were only priority three commands left, so they ran in location order.

1.3.4.11 Command precedence within a location

Seen here:
   Candide [1269]
   Osswid [5499]
   Feasel the Wicked [1109]

Order precedence within a location is an advantage for commands or skill uses which obtain resources from the location. For instance, if Candide and Osswid both attempted to harvest all of the lumber available in their location, Candide would have precedence, since his harvest order would finish before Osswid’s, if they were started on the same day.

1.3.4.12 At the same time …

No two things ever happen at exactly the same instant in Olympia. Someone always goes first.

Suppose two nobles were outside of a building (which nobody is inside), and both wanted to enter, to claim it:

Seen here:
   Osswid [5499]
   Candide [1269]
Inner locations:
   Hooting Own Inn [ep76], inn

Both Osswid and Candid issue ‘move ep76’ as their first order on day 1 of the month. What happens?

Osswid’s command begins before Candide’s, since Osswid appears before Candid in the location list. Therefore, Osswid will enter the inn first.

Order precedence between units in different provinces is undefined.

    Osswid             ?            Candide
    ------- 8 days -------- 8 days --------
    city A          city B          city C

If Osswid and Candide both leave for city B on the same day, we cannot predict who will get their first.

1.3.5 Turn reports

A turn report will be e-mailed to you shortly after the turn is run. Turn reports usually arrive within an hour.

The Age of Gods can produce turn reports in several different formats. The default is text format. This format can be read by any mail program.

You may also receive your turn in HTML format. This format can be read by many mail programs, particularly Outlook and the mail readers in most web browsers. In this format, your turn contains many useful features such as links between parts of the report, links to the rules, etc.

You can also receive your turn in TAGS format. Tags are XML-like markers in the turn intended to be read and processed by a computer program. You could, for instance, write a program to read this tags and build a database of information automatically. The TAGS format is the same as the TEXT format, with the TAGS embedded.

Finally, you can receive your turn in RAW format. This is format is completely unprocessed and is intended only for people who want to build their own post-processing tools.

For details on the various formats, see ⟨undefined⟩ [Appendix - Output Formats], page ⟨undefined⟩.

Your turn report will also be available over the WWW. To use this facility, you must first set a ⟨undefined⟩ [password], page ⟨undefined⟩. Then you may access your turn reports at:

         http://www.pbm.com/oly/tag/reports/<your faction id>

When you access this, you’ll be prompted for a user name and password. You user name is your faction id. If you play multiple TAG games, then your password will correspond to the last game run.

1.3.6 Turn schedule

The turn schedule for each game is shown on the game’s web page, and in each issue of the game newsletter. Players should receive their turn reports by the morning following the run.

Late orders will queue for the next turn. Beware! Even if mailed before the deadline, your message may take some time to get to the order scanner. Send your orders in a day early to avoid grief. This will also give you time to correct mistakes reported by the order scanner.

New player additions are performed just after the turn is run.

1.3.7 Game Newsletter

The game newsletter, (called The Gods Speak) is published with the game turn and mailed to each player in the game. The newsletter has three sections: any comments or notices from the GM (if there are any), signed press, and rumors.

Players should submit items to the newsletter with the press and rumor commands.

The newsletter may pay gold to the player for submissions. Check the newsletter banner for a particular game to see if it is paying for submissions.

1.3.8 Compensation for errors

What if there’s a bug which ruins your whole turn? What if an error in the combat system lets all of your men be killed in one day?

• I will not re-run turns. Turns can’t be rerun for one player, the entire simulation has to be run for everyone. This takes some time and produces a large amount of data which is mailed out. Only for the most serious of game disasters would I consider rerunning a turn.

• If you are stung by a bug, I may grant items for compensation, but I won’t alter the universe. I might give you 1,000 gold if you are the victim of some heinous bug, or perhaps even an NP, but I won’t move a unit that should have been in Pactra to Drassa.

1.3.9 Cheating

1. A player may not have more than one faction in the game. Also, multiple players from the same account are not allowed. Each player must have their own unique email address.

2. Players should not send in orders for another player’s faction, to ruin that person’s turn or otherwise benefit.

3. Players must inform the GM of any game bugs found.

4. Anti-social behavior, including harassing telephone calls, sending obscene/obnoxious unwanted communications, mail bombing, etc. will not be tolerated.

5. It is not cheating to form alliances or cooperate with players of another nation. It’s even acceptable to occasionally act against the interests of your nation, particularly if that serves some specific goal. For example, if you’re trying to become a great merchant, and the enemy offers you a bunch of gold to reveal the location of one of your nation’s large armies, you might accept. But joining one nation with the sole purpose of helping another nation win by being a "mole" in the midst of the other nation is not permitted. If you want the Imperial Empire to win, join the Imperial Empire. If you do choose to betray your nation, do not complain when they turn upon you and rend your faction. You still cannot re-join a different nation.

Punishment for serious violations is generally banishment from Olympia. So don’t cheat, do play fair, and be a good sport.

All decisions of the GM are final.

1.3.10 Game additions and rule changes

Features may be added to Olympia every so often, such as new areas in the map, NPC races, magical artifacts, skills, spells, etc. It is also sometimes necessary to make alterations to existing game mechanics to correct bugs or flaws in game balance.

While it is inevitable that some players will be affected by rule changes, every effort is made to not disrupt existing game positions.

This is mentioned only as a warning that the rules may evolve over time. Changes are announced via email to all players.

1.3.11 How do I join?

It’s easiest to join through the WWW form available at http://www.pbm.com/oly/tag/join.html. If that’s impossible, send mail to srt@pbm.com for instructions on establishing an account via e-mail.

Some notes about starting:

1. Players select a nation when starting the game (see ⟨undefined⟩ [Political Geography], page ⟨undefined⟩). Depending upon the game, there may be several nations available, or only one.

2. Players select a city in which to create their initial noble. The initial noble may be created in any city controlled by the new player’s nation. (A city is controlled by a nation if the city’s province is garrisoned by the nation.) The list of cities controlled by a nation is not published. To get a list of cities, new players should contact an established player in the nation they are joining and negotiate for that information. Optionally, the new player can choose to start randomly in some city controlled by the nation he is joining. If all else fails, the initial noble will be created in the nation’s capital, even if that city has been conquered by another nation.

If a player attempts to create an initial noble in a city that is not controlled by his nation, the noble will be created instead in the nation’s capital.

3. New players start with a variable number of noble points and gold. The amounts are determined by how well each nation is doing. Nations with few noble points start with more noble points; nations with little gold start with more gold. The approximate starting amounts for each nation are published each month in the Game Newsletter, but the actual amounts will vary from the numbers published there.

1.3.12 How do I quit?

To quit the game, use the quit command (see ⟨undefined⟩ [quit], page ⟨undefined⟩).

When you quit the game, all your nobles (and their possessions) will "dissolve" into the ground.

If you choose to restart, you must join the same nation. To close your Shadow Island Games account, contact support@pbm.com.

1.3.13 Email address, etc.

Joining, info info@pbm.com

Order scanner orders@game.pbm.com

Diplomatic forwarding entity@game.pbm.com

The latest rules, back issues of the Game Newsletter, and other game related material may are available on the Web at http://www.pbm.com/, or may be downloaded via anonymous ftp from ftp.pbm.com.

Shadow Island Games
P.O. Box 390983
Mountain View, CA 94039-0983
USA

2 The Olympia World

Much of TAG is about interacting with the world – both the units and items in Olympia’s fantasy setting and the other players and their political entities. This chapter describes the world in both those aspects.

2.1 Olympian geography

Olympia’s map is a large grid of locations called provinces. Groups of provinces form continents, islands and oceans. These collections are called regions, and are usually named.

A province’s description will include a list of the directions in which a noble may travel:

    Plain [ae48], plain, in region Tollus
    Routes leaving Plain [ae48]:
       North, to Plain [ad48], 7 days
       East, to Plain [ae49], 7 days
       South, to Ocean [af48], Tymaerian Sea, 1 day
       West, to Ocean [ae47], Tymaerian Sea, 1 day

This is a non-descript province in the Tollus region (Tollus is a small continent in the northeast corner of the map.) From this province, a noble may travel north or east on foot or by horse, or may sail by ship to the south or west.

    move north  -or-   move n   -or-   move ad48
    move east   -or-   move e   -or-   move ae49
    sail south  -or-   sail e   -or-   sail af48
    sail west   -or-   sail w   -or-   sail ae47

Land movement will automatically use the fastest available mode. For example, if a noble has enough horses for all of the members in the party to ride, then the travelers will go on horseback.

Ocean movement requires that the noble be in a ship.

Route distances are rated for the number of days it normally takes to traverse them. Land distances are rated for a lightly loaded noble walking, and ocean distances are given for an ordinary ship traveling in normal weather.

Actual travel times may differ from times given in the route listing. Land distances depend on the surrounding terrain and the modes of transport available. For example, horses often speed up movement, but over especially rough or treacherous terrain, they may actually slow travel because they must be led and managed. A stiff wind may speed ocean vessels, while lack of wind may slow their progress.

2.1.1 Inner Locations

A province may contain sublocations within its borders. Sublocations may usually only be entered from the surrounding province. They will be listed separately in the location description:

    Inner locations:
       Carim [em28], city, 1 day

The city Carim may be entered with the ‘move em28’ order. Travel into a city requires one day.

(move in may be used to enter a sublocation, although this order may be ambiguous if the location contains more than one sublocation. In such a case, the first sublocation in the Inner locations list will be entered. Using move in is not recommended if the entity number of the sublocation is known.)

Nobles in a sublocation will receive a report for the surrounding province. However, nobles in the outer province will not normally be able to see into an inner location without entering it.

2.1.2 Inside a City

    Carim [em28], city, in province Plain [ae48]
    Routes leaving Carim [em28]:
       Out, to Plain [ae48], 1 day
    Inner locations:
       Hooting Own Inn [ep76], inn

Nobles in the city Carim may move out (or ‘move ae48’) to the surrounding province. They may also attempt to enter the inn, which is a sublocation of the city. Notice that no travel time rating is listed for the inn; entering it takes no time (zero days).

    move out    -or-    move ae48
    move ep76

A noble in Carim will receive a location report both for the city as well as the surrounding province Plain [ae48]. Nobles in the Hooting Own Inn will receive a location report for the inn and one for Carim, but will not get a report for Plain [ae48].

Nobles in a sublocation receive a report for the immediate surrounding location.

Nobles see into inner locations without going into them.

Cities also contain markets (see ⟨undefined⟩ [Markets], page ⟨undefined⟩), where nobles can buy and sell goods.

2.1.3 Who else is here?

Nobles spotted will be listed in the location report:

    Seen here:
       Fighters of Pelenth [2019], "carrying a gold banner"
       Osswid the Constructor [5499]

All nobles listed as being seen in the location may interact without requiring any travel. Thus, the Fighters of Pelenth and Osswid are considered to be in essentially the same place.

This is true whether the nobles are in a province, a city, a ship, an inn, or some other sublocation.

However, a noble in a sublocation may not interact with nobles in the surrounding area. A noble in the city must first enter the inn before he may interact with those inside.

2.1.4 More about geography

Olympian provinces are arranged in a square grid. Travel is possible in the four main compass points. Thus, to move diagonally, two move orders are required. To move northwest, for instance, one would first need to ‘move n’, then ‘move w’.

The map coordinates for a province may be read from the province’s entity number. The row is represented by the leading two letters, the column by the two digits. The northwest corner is [aa00], with rows increasing to the south, and columns increasing to the east.

For example, a world 100 by 100 provinces would be numbered:

        00   01   02  ... 99
      +----------------------
    aa|
    ab|         ab02
    ac|    ac01 ac02     ac99
    ...
    ...
    dv|         dv02

Entity numbers for sublocations do not correspond to any coordinate system.

The map wraps at all edges. Traveling west from ac00 will lead to ac99, for instance.

2.1.5 Holes in the Map

The map has some holes, representing impassable provinces. Routes into some provinces may also be hidden.

    Plain [cd21], plain, in region Tollus
    Routes leaving Plain [cd21]:
       North, to Plain [cc21], 7 days
       East, to Plain [cd22], 7 days
       West, to Plain [cd20], 7 days

Notice the lack of a southern exit. This means that there is no known southern route from Plain [cd21], into what should be Plain [ce21]. Exploration may find a southern route, but it is possible that none may ever be found, and the terrain to the south is completely impassable.

2.1.6 Hidden routes

If exploration finds a hidden route, any noble in the player’s faction will be able to use it.

    > explore
    A hidden route has been found!
       South, to Plain [ce21], 7 days

The location description for this place will now include the hidden route:

    Plain [cd21], plain, in region Tollus
    Routes leaving Plain [cd21]:
       North, to Plain [cc21], 7 days
       South, to Plain [ce21], 7 days, hidden
       East, to Plain [cd22], 7 days
       West, to Plain [cd20], 7 days

However, units from other factions, even if they know that the hidden route’s entity number is [ce21], will not be able to travel across it.

All factions with units in a stack traveling across a hidden route, with the exception of units being held prisoner, will learn of its existence. Nobles from factions wanting to learn how to use the hidden route can stack with a noble about to move across the route.

2.1.7 Ocean ports

A ship in an ocean province may sail into an adjoining land province.

Ocean [cw12], ocean, in South Sea
Routes leaving Ocean [cw12]:
North, to Ocean [cv12], Atnos Sea, 4 days
East, to Mountain [cw13], West Camaris, impassable
South, to Plain [cx12], West Camaris, 1 day
West, to Ocean [cw11], 4 days
Inner locations:
Island [eb97], island, 1 day

A ship sailing in this ocean province may dock by sailing to Plain [cx12] or Island [eb97].

Ships may not dock in mountain provinces, as the rocky cliffs are too dangerous to approach. Routes between ocean and mountain provinces are

marked ‘impassable’.

2.1.8 Port Cities

A city in a province adjoining an ocean will have been founded on the best spot for an ocean port. The ocean will only be accessible through the port city in this case, and not through the surrounding region.

Plain [ae48], plain, in region Tollus
Routes leaving Plain [ae48]:
...
West, to Ocean [ae47], Tymaerian Sea, impassable
Inner locations:
Carim [em28], port city, 1 day

Note that from the province surrounding the port city, access to the ocean is not possible.

Carim [em28], port city, in province Plain [ae48]
Routes leaving Carim [em28]:
West, to Ocean [ae47], Tymaerian Sea, 1 day
Out, to Plain [ae48], 1 day

However, ships may sail into and out of the port city itself. From the Tymaerian Sea, this looks like:

Ocean [af48], ocean, in Tymaerian Sea
Routes leaving Ocean [af48]:
North, city, to Carim [em28], Tollus, 1 day
South, to Ocean [ag48], 3 days
...

2.1.9 An example city description

Drassa [ew66], port city, in province Forest [cu26], safe haven
Routes leaving Drassa [ew66]:
East, to Ocean [cu27], Atnos Sea, 1 day
South, to Ocean [cv26], Atnos Sea, 1 day
Out, to Forest [cu26], 1 day
Skills taught here:
Shipcraft [1000]
Combat [1100]
Construction [1400]
Seen here:
Kosar the Indefectible [2022], with six peasants, one archer,
two soldiers, accompanied by:
   Dr. Pangloss [3682]
   Law Netexus [2020], prisoner
   Alion Krysaka [2785], prisoner
Ships docked at port:
HMS Pinafore [ib18], owner:
   Captain McCook [2019], with five workers
Market report:
No goods offered for trade.

2.1.10 Wilderness and civilization

Every province has a peasant population. Provinces with less than 100 peasants resident are considered wilderness. If a province has a population greater than 100, the population is shown in the province description:

Barbary Coast [ap21], plain, in West Camaris, peasants: 127

2.1.11 Common Dangers

Players should take care when exploring the Olympian world. There are many dangers, both from non-player characters (NPC’s) as well as from other players. While the threat of death to nobles is always present, the following dangers warrant extra caution.

2.1.12 Monsters

Olympia is roamed by many sorts of monsters. Monsters often attack player stacks.

Wilderness provinces (with less than 100 peasants) have more monsters than civilized provinces. Garrisoned provinces are even safer, as the mere presence of an active garrison keeps some monsters away. Very few monsters will enter a city, making cities the safest places of all.

A stack with 25 soldiers will be safe from most attacks, but very dangerous monsters exist that can challenge even a very powerful stack. When possible, "double up" with other players to give yourself as much protection as possible, especially early in the game. Missile weapons are also valuable against monsters, as few monsters have the intelligence to take advantage of missile combat.

Dumb monsters will attack any stack they outnumber. Smarter monsters will evaluate the combat strength of a stack before attacking. Certain monsters (such as balrogs) attack anything that moves, and can be very dangerous. Fortunately such monsters are rare.

All monsters fight to the death when they attack or are attacked.

Many monsters will kill and eat any prisoners they take, although they may release prisoners if already well-fed. A few, like bandits, will merely strip prisoners of all their valuables and turn them loose again.

Some monsters will rush to join in when they see other monsters attack. For instance, wolves run in packs and will usually aid other wolves. Bandits are quick to take advantage of any combat, and may join in a battle if they see an opportunity. Smart monsters may learn combat skills, making them tougher than expected. Players should keep these and similar possibilities in mind when attacking monster stacks.

Monsters spend much of their time breeding. Players are wise to eliminate small, weak stacks of monsters before they became large, strong stacks.

Monsters will also pillage, attack garrisons, raze structures and do various other things to stop the incursion of humans into their territory.

Monsters get no special advantages in Olympia. With a couple of minor exceptions, monsters use the same commands as nobles. The only difference between monsters and nobles is that nobles only queue commands at the beginning of the month, while monsters can queue commands at any time. However, an effort is made to avoid giving monsters a "special advantage" through this.

Cities offer bounties on some monsters. For example, a city might offer 1 gold for each rat tail. When a rat (or other monster) is killed, the rat tail (or appropriate monster part) will automatically be added to the inventory of the stack leader. These can then be sold at city markets to collect the bounty. Note that not all cities offer bounties on all monsters; the bounties offered will appear in the city’s market description.

Bounties are initially offered (for sale or purchase) at their "base price". For example, the base price for elf ears is 2 gold, the base price for nazgul tails is 40 gold, and the base price for centaur hides is 5 gold. Here’s the complete list:

Part                Base Price
----                ----------
elf ear [317]            1
savage ear [318]         1
nazgul tail [319] 40
centaur hide [320]  5
minotaur hide [321]  5
spider eye [322]  2
rat tail [323]  1
lion mane [324]          5
bird feather [325]  10
lizard tail [326]  5
bandit ear [327]  1
chimera eye [328] 10
harpie feather [329]  5
dragon scale [330] 40
orc scalp [331]  4
gorgon liver [332] 10
wolf hide [333]  3
cyclops eye [334]  8
giant's tongue [335] 10
balrog horn [336]      300

Bounty prices rise and fall with demand, as with any other good. An unfulfilled bounty will eventually be dropped and replaced with a different bounty.

2.1.13 Hades

No one goes to Hades with all his immense wealth.

Maxims, Theognis. Hades, being the sphere of death, is inimical to living nobles. Any noble

in Hades loses 1-10 aura or piety each week. A noble who falls to zero aura or piety while in Hades is immediately destroyed.

Hades is also populated with demons, ghouls and spirits thirsty for the blood of the living. Being already dead, they fear no living creature.

    name                    combat rating
    ----                    -------------
    Greater Demon            (500,500,0)
    Lesser Demon             (250,250,0)
    Ghostly presence         (100,100,0)
    Tortured spirits         (50,50,0)

Necromancers with the "Transcend Death" skill are immune to the spiritdraining effects of Hades, and need not fear attack from its denizens.

2.1.14 Faery

Come away, O human child! To the waters and the wild With a faery, hand in hand, For the world’s more full of weeping than you can understand.

Stolen Child, William Butler Yeats

The Faery world lies in a nearby, but separate reality. Occasionally a Faery hill will exist simultaneously in both Faery and the outside world. During this time, mortals may cross between the two worlds.

Faery is protected by Faery Hunts, tough bands of elves armed with magical bows. Each hunt group consists of 48-200 elves, each with a combat rating of (50,50,100).

Rumors speak of a magical talisman, the elfstone, which allows mortals to pass unharmed in Faery, and to summon Faery hills to the mortal world.

Also associated with Faery is the Wild Hunt. The Wild Hunt is a group of High Elves and their hounds, who ride in the dark of the night across the land. Occasionally they abduct a noble, spiriting him away from all his companions and whirling him about the countryside, sometimes depositing him back where he started, and sometimes leaving him far away.

2.1.15 The Cloudlands

Boats sail on the rivers, And ships sail on the seas; But clouds that sail across the sky Are prettier far than these.

There are bridges on the rivers,

As pretty as you please;

But rainbows that bridge heaven, And overtop the trees, And build a road from earth to sky, Are prettier far than these.

Boats Sail on the Rivers, Christina Giorgina Rossetti

Rumor tells of a magical floating country in the clouds, the work of some long-forgotten wizards. Blown hither and thither over the countryside by the prevailing winds, the Cloudlands can be reached only by flight – and then only from the tallest of mountains.

The Cloudlands is home to three cities: Nimbus, Stratos and Aerovia.

2.1.16 Guard monsters

Monsters are sometimes found guarding sublocations, such as a lairs, caves, or graveyards. These locations may be hidden, requiring the noble to explore to find them. Often these monsters will have some rare or useful possession, or in some cases may have a prisoner. Nobles who attack and defeat these monsters can expect to obtain the monster’s treasure.

Such treasures include:

• gold

• a quantity of some rare tradegood

• a magical artifact

• a book which teaches a rare skill

• a prisoner who is grateful for being rescued

2.2 Political Geography

The players in Olympia: The Age of Gods are divided into several nations. Each player selects a nation when joining the game. In most scenarios, nations compete to win the game by conquering the other nations (see ⟨undefined⟩ [Winning the Game], page ⟨undefined⟩).

The nations available in a game depends upon the ⟨undefined⟩ [Scenarios], page ⟨undefined⟩. The following sections describe some of the commonly available nations.

2.2.1 The Imperial Empire

The Imperial Empire is the oldest nation on Olympia, and is rich in tradition and history. The nobles of the Imperial Empire bear impressive titles ("The Baron of Casington Lake") and are identified by a confusingly complex system of heraldry. The noble houses of the Imperial Empire are closely tied to the Eight Churches, and younger sons of the noble houses often serve the Churches. Since the Eight Churches view magic as the evil legacy of Tor, the Imperial Empire has made it a holy mission to eliminate all use of magic from Olympia.

Nobles in the Imperial Empire may not learn any magic skills (e.g., [2800] Magic, [2900] Scrying, [3000] Gatecraft, [3100] Artifact Construction, [3200] Necromancy, or [3300] Advanced Sorcery).

2.2.2 The Mandor Confederacy.

The Mandor Confederacy is a loose collection of small kingdoms and citystates. Founded by refugees from the Imperial Empire, the Mandorians reject many of the principals of that nation. They eschew the formal trappings of titles and manners in favor of a free and independent society. They also reject the teachings of the Eight Churches, and embrace instead the study of magic. In times of conflict, they band together to meet foes with fierce resistance.

Nobles in the Mandor Confederacy may not learn any religion skills (e.g.,

[2000] Church of Eres, [2100] Church of Anteus, [2200] Church of Dol, [2300] Church of Timeid, [2400] Church of Ham, [2500] Church of Kireus, [2600] Church of Halon, or [2700] Church of Domingo).

2.2.3 The Pen

Some time in the gap of history that followed the Time of Flames, solitary ships began appearing in human ports bearing stocky gnomes who called themselves "Pen". Treated cautiously or with fear at first, the Pen soon became accepted and welcomed throughout the known lands. For the most part the Pen were merchants and traders. They were honest to a fault, and strictly neutral in Human affairs. As contact between Pen and Man grew, the Pen began to offer other services – such as labourers and mercenaries – but always remained honest and neutral.

Amongst Men, the Pen are sometimes called "Sea Gnomes".

The Pen live on a small, rocky island they call Penhome. Few in number, they are divided into five families. Like all gnomes, they covet riches, and the five families compete to see who can hoard the most treasure. Once a year the tally is taken and the patriarch of the richest family is crowned High King of Pen. The High King of Pen receives a magical artifact called the Pen Crown.

The Pen are limited to five players. Nobles in the Pen may not learn any religion or magic skills (e.g., [2000] Church of Eres, [2100] Church of Anteus,

[2200] Church of Dol, [2300] Church of Timeid, [2400] Church of Ham, [2500] Church of Kireus, [2600] Church of Halon, [2700] Church of Domingo, [2800] Magic, [2900] Scrying, [3000] Gatecraft, [3100] Artifact Construction, [3200] Necromancy, or [3300] Advanced Sorcery).

A Pen player must always pay the NP cost when converting a noble from another faction. Likewise, when another faction converts a Pen noble, they must always repay the NP cost.

Note that because of their limitations (no use of magic or religion) and difficult starting position (a small, isolated island), it is impossible for the Pen to compete on equal terms with the other nations. Players choosing to play the Pen should accept that any major role they play in world politics must be through diplomacy and cleverness, not by might or power. In reality, the Pen are "apart" from the competition between the other nations, and concentrate on competing amongst each other to be the greatest merchant.

3 Nobles & What They Can Do

3.1 Health

Nobles have a health rating of 1-100, which indicates how wounded they are. A noble with health 100 has no wounds; a noble with a health of zero dies. Nobles also have a flag which indicates whether they are suffering from an illness. A sick noble will lose 3-15 health each week, while a wounded noble who is otherwise free of illness will recover somewhat each week.

Health is shown in the turn report for each noble that the player controls. This is a noble in perfect health:

    Health:         100%

This is a rather sick noble:

    Health:         38% (getting worse)

If the noble were cured of illness, this would instead show:

    Health:         38% (getting better)

Medical technology is rather crude in the age of Olympia. Sanitation and hygiene are not the best. Even a minor wound runs a risk of developing into a serious, possibly life-threatening problem.

When a noble receives a new injury, their health is reduced by the amount of the injury, and a check is made to see whether they get sick. The chance that a noble falls ill is (100 - health). Thus, the more seriously the noble is wounded, the greater the probability that infection will set in.

Health is updated at the end of each game week (on days 7, 14, 21 and 28). Sick nobles lose 3-15 health each week. Healthy but wounded nobles recover by a like amount. Each week sick nobles have a 5% chance of fighting off their illness.

Nobles located in an inn benefit from the rest and relaxation that the inn provides. Inns increase the chance of fighting off infection to 10% each week.

Illness can also be cured by spells, special skills, or magical potions. Sick nobles would be wise to seek out those practiced in healing, and recover in a nearby inn.

Notes:

1. Nobles who recover from illness will not suddenly become sick again. The illness check is only made when a new wound is received.

2. Since the chance that a noble becomes sick is based on the unit’s health after deducting a hit, an already-wounded noble stands a greater chance of becoming ill from a wound than a healthy one.

3. Health is only tracked for nobles. Peasants, workers, sailors, soldiers, etc. are either dead or alive; they have no health rating.

4. A wound received in combat will be randomly chosen between 1 and 100. Wounds received in other activities are scaled to match the level of danger involved.

5. Some NPCs are not rated for health. Health for these nobles appears as n/a in the turn report. For means of directly inflicting damage against a noble such as terrorize, Aura blasts, and Lightning bolts, such a unit requires a hit of 50 or more points in order to be killed.

3.1.1 Death

When a noble is killed, his body falls in the province where he was killed. If he was killed in a sublocation, such as a sacred grove, his body will still fall in the province. If he is killed at sea, his body will wash ashore into the nearest land province.

The body remains in the province where it falls for 12 months. It can be found and recovered during that time with the ⟨undefined⟩ [explore], page ⟨undefined⟩ command. At the end of that time, the body will decay and be lost forever.

Priests may learn a skill to administer last rites, hastening the passing of the dead noble’s body. Some exceptionally skilled priests possess the ability to resurrect dead nobles.

3.2 Loyalty bonds

Immortal gods, I crave no pelf; I pray for no man but myself; Grant I may never prove so fond, To trust man on his oath or bond.

Timon of Athens, William Shakespeare

Nobles are bound to their lords by one of three kinds of loyalty:

contract The noble is being paid for his service

oath The noble has taken an oath to serve his master

fear The noble serves out of fear

The loyalty bond is rated. For example, a noble’s loyalty could be oath-1, contract-500, or fear-50.

A player’s initial noble has oath-2 loyalty.

Newly hired nobles have loyalty contract-500. Nobles are paid for their service with the honor command. A noble who issues ‘honor 50’ will spend 50 gold to raise his own contract loyalty rating by 50 points.

A noble may take an oath of loyalty, pledging one or two Noble Points to secure it. This would yield loyalty oath-1 or oath-2. The oath order secures an oath of loyalty for a noble.

Nobles may be terrorized by their masters. The severity of their treatment accumulates in the loyalty rating: fear-10, for instance. Fear loyalty is maintained with the terrorize order.

Only one kind of loyalty may be active at a time.

Contract and fear loyalty decay over time. Contract loyalty loses the greater of 50 points or 10% of the current rating each month. Fear loyalty loses 1-2 rating points each month. Oath loyalty does not decay.

A contract noble who is about to fall to contract-0 will automatically honor himself to keep from deserting. He will spend 50 gold from his own inventory (if available) or from claim gold. He will do this even if he is being held prisoner.

Units which fall to contract-0 or fear-0 have a 50% chance of deserting each month.

Summary:

contract Amount of gold invested in noble Decays by max(50, 10% of current rating) each month

fear Severity of terrorize used on noble Decays 1-2 points each month

oath 1 or 2 NPs may be invested in an oath bond Does not decay

Commands dealing with loyalty bonds:

  • bribe
  • honor
  • oath
  • terrorize

3.2.1 Converting Nobles to a New Faction

Several commands (such as ⟨undefined⟩ [bribe], page ⟨undefined⟩) and skills are available which can be used to convert a noble from one faction to another faction.

Nobles serving through contract or fear are susceptible to bribes, which may induce them to renounce loyalty to their lord, and pledge their service to the bribing faction. For details on bribing nobles, see ⟨undefined⟩ [bribe], page ⟨undefined⟩.

Oath-1 nobles ignore all bribes. There is a persuasion skill which may cause an oath-1 noble to defect, although its use is difficult and rarely succeeds. Oath-1 nobles may reveal their factional affiliation if tortured.

Oath-2 nobles will not renounce loyalty to their lord under any circumstances, nor can they be forced to reveal any information about themselves.

When a noble is converted from one faction to another through the use of a bribe, persuasion, or any other means, the faction that receives the noble may have to pay the NP cost of the noble. The NP cost of the noble is simply the number of NPs that have been invested in the noble by the other faction (e.g., 1 to form the noble, 1 to change his loyalty to oath-1, 1 to learn a magic skill for a total of 3 NPs). If the receiving faction does not have sufficient NPs to pay the NP cost, then the attempt to convert the noble will fail. There will be no indication that the failure occurred from a lack of NPs.

There are several cases where the converting faction does not have to pay the NP cost.

The first case is when the noble comes from another, non-neutral nation. For example, if a Mandorian player converts an Imperial noble, he does not have to pay the NP cost of that noble. Capturing nobles is one of the greatest prizes of war. You cannot capture NPs from a neutral nation (e.g., the Pen), and players of neutral nations cannot capture NPs from anyone.

The second case is when the noble is a refugee. A refugee is a noble whose original player has quit the game. When a player quits the game, all his nobles die, except those that are being held prisoner by another nation. Those nobles become "refugees" and can be converted at no NP cost by anyone.

The third case is when you are reclaiming a deserted noble from your own faction. If a noble deserts, and some time later you re-capture him and convert him back to your faction, there is no NP cost (even though he comes from your nation).

3.2.2 When a Noble Deserts

When a noble deserts your faction because his loyalty runs out, he joins the "deserted noble" faction. You do not receive back his noble points when this happens.

What happens when you capture a deserted noble and convert him? If he was your noble originally, you pay no NP cost. If he belonged to another player in your nation before deserting, you must

pay the normal NP cost.

If he belonged to a player in another nation before deserting, you pay no NP cost.

3.2.3 When a Noble Dies

When a noble dies and his body decays, his NPs are lost. The moral is to be careful not to let your nobles that have a lot of NPs invested in them be killed!

3.2.4 When a Player Quits

When a player quits, all his nobles die, except those that are being held prisoner by another nation. Those nobles become "refugees" and can be converted at no NP cost by anyone.

Most of the goods being held by the nobles who die are lost. Unique items (such as artifacts) will typically fall into the location where the noble dies, and can be found with ⟨undefined⟩ [explore], page ⟨undefined⟩.

3.3 Skills

No profit grows where is no pleasure taken; In brief, sir, study what you most affect.

William Shakespeare

Skills represent knowledge that Olympian nobles may know. Shipbuilding, thievery, kidnaping, training soldiers, castle building, mixing potions, and forging magical artifacts are just a few of the possible actions which skills allow a noble to perform.

Skills are organized into categories. Before a noble can learn any of the skills in a category he must first learn the category skill. The category skills are:

    num  name                       time to learn
    ---  ----                       -------------
    1000  Shipcraft                  one week
    1100  Combat                     one week
    1200  Stealth                    two weeks
    1300  Persuasion                 two weeks
    1400  Construction               one week
    1500  Alchemy                    two weeks
    1600  Forestry                   one week
    1700  Mining                     one week
    1800  Trading                    one week
    1900  Ranger                     one week

There are also six schools of magic and eight religions. For more details on learning and casting magical spells, see ⟨undefined⟩ [Magical Arts], page ⟨undefined⟩. For more details on religion skills, see ⟨undefined⟩ [Religion], page ⟨undefined⟩. Note that magicians cannot learn religion skills, and priests and their followers cannot learn magic skills.

Most skills are invoked with the ‘use skill’ order. Specific information about arguments to use and and requirements or limitations for using all Olympia skills may be found in ⟨undefined⟩ [Appendix - Skill Dictionary], page ⟨undefined⟩.

To begin study of a skill requires the following:

• A source of instruction (possibly).

• Payment of a fee for necessary materials.

• Payment of Noble Points to learn advanced skills (possibly).

The following sections describe the process of learning new skills in more detail.

3.3.1 Skill Types

There are four types of skills in Olympia:

• Category skills (e.g., Combat [1100]). A category skill represents the basic foundation of skills and knowledge in a particular area. Learning a category requires a teacher — another noble who has already mastered the skill, or a book of instruction.

• Teachable subskills (e.g., Train Swordsmen [1105]). These skills represent extensions of the category skill which may require several weeks of practice to master, but which are otherwise straightforward. Learning these skills does not require a teacher, but the time required can be reduced significantly if a teacher is available.

• Unteachable subskills (e.g., Use Beasts in Battle [1195]). These skills are more complicated and difficult skills which do not follow directly from the category skill. Each student must go through an individual process of discovery to master these skills. These skills cannot be taught.

• Guild skills (e.g., Train Mounted [1190]). Many skill categories have guilds which zealously protect the secrets to one or two important and powerful skills. To learn these skills, a noble must be a member of the appropriate guild (see ⟨undefined⟩ [Guild Halls], page ⟨undefined⟩).

The complete list of Olympia skills can be found in ⟨undefined⟩ [Appendix - Skill Dictionary], page ⟨undefined⟩.

Each skill a noble learns costs him $100. Certain difficult skills (such as all the magic and religious category skills) also cost 1 or more noble points (NPs). In addition, each category skill beyond the third costs a noble point. (However, this charge is waived for Basic Religion [3500], the strength skill when learned in conjunction with a religious skill, and Advanced Sorcery [3300]) This represents the difficulty of becoming an expert in a wide variety of subject areas.

3.3.2 Studying

I will study and get ready, and perhaps my chance will come.

Abraham Lincoln

All skills require a study period before they can be put to use. The length of study required depends upon the complexity of the skill being practiced. Generally, the study time is one to four weeks.

To study a skill, the noble should use the study command (see ⟨undefined⟩ [study], page ⟨undefined⟩). The argument to the study command is the id number of the skill to be studied. For example, to study Combat [1100], a noble would give the command:

1: > study 1100
1: Will study Combat for seven days.
7: Learned Combat [1100].
Skills known:
   Combat [1100]

If a source of instruction (such as a teacher or a book of instruction) is available and useful, the study command will automatically make use of it.

3.3.3 Instruction

Instruction is necessary to learn category skills and speeds the learning of teachable subskills. There are four sources of instruction in Olympia:

• Some cities teach skills. Shipcraft [1000], for example, is sometimes taught in port cities. The turn report lists skills taught by the location:

    Skills taught here:
       Shipcraft [1000]

To make use of this instruction, the noble must issue the study command while in the city. (Note that you cannot be in a tower in the city; you must be directly in the city.)

• Guilds and temples teach the appropriate skill. For example, a Guild of Shipcraft teaches Shipcraft [1000]. A non-guild member can enter a guild by being stacked with a guild member to take advantage of this teaching. Note, however, that he still cannot learn guild skills (even though he is inside a guild hall) unless he becomes a guild member.

• Books of instruction are sometimes available. Items which teach skills are shown in the inventory listing:

    Inventory:
                 qty  name
                 ---  ----
                   1  Old book [6001]
    Old book [6001] permits study of the following skills:
       Alchemy [1500], 20 days instruction

Books of instruction may be created by nobles, discovered in the possession of monsters, or found for sale in cities. Each book provides a specific number of days of instruction. The number of days instruction remaining in a book is shown in the inventory listing for the book. Note that books provide days of instruction; not learning. Each day a noble uses a book consumes one day of instruction from the book, even if the noble doesn’t get a full day’s benefit from that study (because, for example, that it is his third week of study for the month). A book will be used automatically when a study command is issued and no other form of instruction is available. To avoid using a book (as, for instance, when you wish to save the book for another noble), the noble must give the book to another noble or otherwise remove it from his inventory.

• Nobles can be teachers. A noble who knows a skill may use the teach command to teach other nobles the skill. However, the teach command requires a quiet, uninterrupted environment, so teaching must be done in a tower or an appropriate guild. There can only be one class in a tower/guild at a time. The teacher is the top noble issuing a teach command. The students are the top five nobles in the tower/guild who benefit from the taught subject. Any other nobles are too far from the teacher to follow his instruction. (See ⟨undefined⟩ [teach], page ⟨undefined⟩) The ideal teaching situation is for the teacher to be the owner of the tower/guild (the top noble) and the students be five nobles stacked under him. However, it is often difficult to coordinate this when arranging teaching between multiple players. The rules for teaching are a compromise so that players aren’t punished for simple mistakes in timing or order. For example, suppose that on a particular day, the following nobles are in a tower executing the indicated commands:

       Newark Academy [z38], tower, defense 20, owner:
          Fenris [re7]                      (study 1001)  *
          Kosar the Indefectible [aw0]      (study 1193)
          Captain McCook [ea0]              (study 1000)   *
          Oleg [o1eg]                       (teach 1000)   +
          Dr. Pangloss [ac6y]               (study 1005)  *
          Law Netexus [au0]                 (study 1000)   *
          Alion Krysaka [e1y]               (study 1100)
          Dr. Pain [em28]                   (teach 1100)
          S.K. Renta [ep76]                 (study 1000)   *
          Armweak [em78] (study 1000)

The plus in this example indicates the teacher, the asterisks indicate students who benefit from his teaching.

The only teacher in this tower is Oleg, teach [1000]. Dr. Pain is not considered a teacher, because he is not the first teacher in the tower. Fenris benefits from Oleg’s teaching because [1001] is a subskill of [1000] and he is one of the first five students. Kosar does not benefit, because 1193 is not a subskill of [1000]. Likewise Alion does not benefit because he is studying [1100]. Note that Armweak does not benefit because, although he is studying the proper subject, he is not one of the first five students in the tower. A source of instruction is required to learn a category skill. For example,

to learn Combat [1100], a noble must be in a city that teaches that skill, a tower where another noble is teaching that skill, or possess a book of instruction for that skill.

Instruction speeds the learning of teachable subskills. If instruction in the category skill for a teachable subskill is available, then the student will learn the skill in half the normal time. If instruction in the specific subskill is available, then the student will learn the skill in one-quarter the normal time.

For example, Train Swordsmen [1105] is a teachable subskill for the category skill Combat [1100], with a learning time of two weeks. If a noble studies this skill without any instruction, it will take him two weeks of study to learn the skill. If he studies it with a teacher of Combat [1100], he will learn the skill in one week. If he studies it with a teacher of Train Swordsmen [1105], he will learn it in 3.5 days.

Instruction has no effect for unteachable subskills or guild skills.

In some cases, a noble may have more than one source of instruction available. For example, he might be studying Train Swordsmen [1105] in a city that offers Combat [1100] while holding a book of instruction for Train Swordsmen [1105]. Or he might be holding two books, one of which gives instruction in Combat [1100] while the other gives instruction in Train Swordsmen [1105]. In these cases, the source of instruction is selected according to the following rules:

• If the noble is studying a subskill, then instruction in that specific subskill is selected over instruction in the category skill.

• When a book and some other source offer the same type of instruction, then the other source is selected. Accordingly, if a noble studies Train Swordsmen [1105] in a city that

offers Combat [1100] while holding a book of instruction for Train Swordsmen [1105], he will use the book for instruction — because the book offers specific instruction. If he were to study Combat [1100] in a city that offers Combat [1100] while holding a book of instruction for Combat [1100], then he will use the city for instruction.

The benefits of instruction are calculated daily. If a noble begins studying with a book which only contains two days of instruction, then the first two days of the noble’s study will be at 2x or 4x the normal rate (depending on whether the book teaches the specific subskill being studied or the category skill) and the remaining five days of study will be at the normal rate. If instruction ends prematurely for a category skill that requires instruction, then the noble’s study command also ends.

3.3.4 Diminishing returns from study

Of making many books there is no end; and much study is a weariness of the flesh.

Ecclesiastes 12:12

Study is a mentally taxing activity. The more a noble studies during a month, the less advancement each attempt will yield.

The first seven days of study each month are fully effective. The second week yields 4/7th the normal advancement, the third yields 2/7th, and the fourth, 1/7th.

    days spent
    studying        yield
    ----------      -----
       1-7          full
       8-14         4/7th
      15-21         2/7th
       22+          1/7th

For example, a noble could learn a skill requiring two weeks of study by either issuing four study order for one month, or two study orders for two separate months.

Another way to view diminishing study results is that a noble who studies four weeks per month will progress twice as fast one studying one week per month.

This limit exists on the act of studying, and is not related to the study of a particular skill. For example, a noble who studied Shipbuilding the first week, and Combat the second, would have 7/7th effectiveness for learning Shipbuilding, but 4/7th for learning Combat.

Intervening tasks during the month do not affect the tally of study time. One could study the first week, engage in other tasks the second and third, and study again on the fourth week. The study command issued on the fourth week would yield 4/7th.

Study results are kept track of daily. For instance, if a noble studies a skill for three days, and then is interrupted, the three days of study are not lost. If the skill requires a week to learn, four more days of study are needed.

A study command which continues into the next month will revert to the fully effective rate of return. For instance, suppose a noble spent all month studying skills. The last study command doesn’t finish at the end of the turn:

    unit 5499  # Osswid
       # > study 2900 (executing for three more days)

Although this study command was yielding 1/7th per day at the end of this turn, the remaining three days will yield 7/7th. The total advancement from this command would therefore be 3 4/7 days of study credit.

    Partially known skills:
       Scrying [2900], 3.57/21

3.3.5 Learning Bonuses

Some other factors can speed or slow the learning of a skill.

Nobles who are dedicated to a god (see ⟨undefined⟩ [dedicate], page ⟨undefined⟩) have a strength skill and a weakness skill. All skills in their strength skill category require one less day of study than normal, and skills in their weakness skill category two more days study than normal. For example, a follower of Eres has a strength skill of Shipcraft [1000] and a weakness skill of Mining [1100]. Both skills normally take one week of study. For a follower of Eres, Shipcraft requires only six days of study, while Mining requires eight days of study.

Nobles who study alone in a tower (possibly using a book for instruction) benefit from the greater concentration the isolation of the tower provides, and have the study time required for any skill reduced by one day. For example, a noble who studied Train Swordsmen [1105] in a tower without instruction would require 13 days of study instead of the normal 14 days. Studying alone means that (1) there are no other nobles or units in the tower during the study period, and (2) that the noble does not have any men or beasts in his inventory during the study period.

Note that the diminishing returns from studying more than one week a month, the benefits of instruction, and the special learning bonuses are all cumulative.

3.3.6 Costs to begin study

A fee of 100 gold is charged when study is first issued for a skill. The payment is used to acquire various materials needed for the study and eventual practice of the skill.

Some skills also require one or more noble points to learn. These skills are clearly marked in ⟨undefined⟩ [Appendix - Skill Dictionary], page ⟨undefined⟩.

In addition to these fixed costs, there is a charge of one additional NP for the category skills a noble learns in excess of three skills, i.e., for the fourth, fifth, etc., category skills a noble learns there will be an additional 1 NP charge. (Basic Religion [3500] and Advanced Sorcery [3300] are ignored when counting a noble’s skills.)

Note that religion and magic skills count toward the three skill limit, and if you study religion or magic as your fourth skill, it will cost 2 NP (one for the normal cost and one for the fourth skill).

3.3.7 Experience and Practice

Once a skill is known, further study of that skill has no effect. Gains in proficiency come from using or practicing the skill.

Each time a noble uses a skill successfully, his experience level for the skill is increased by one. If a skill is used more than once per turn, only the first success will count towards experience. Projects which take more than one month to complete, such as shipbuilding or castle construction, only count towards experience when the project is finished.

Some skills may also be practiced (see ⟨undefined⟩ [practice], page ⟨undefined⟩). Skills that may be practiced list a practice time and cost. Each completion of practice raises a noble’s skill level by one. There is no limit to the amount of time a noble can spend practicing.

    Level          Description
    -----          -----------
     0-1            apprentice
     2-3            journeyman
     4-5            adept
     6-8            master
      9+            grand master

Experience will speed work with some skills. For example, a master shipbuilder will be able to construct a hull somewhat faster than an apprentice. Some skill uses benefit more from experience than others.

The skill listing for a noble shows his level of experience in each skill.

3.3.8 Skills not rated for experience

A few skills are not rated for experience. These may be skills which are not directly used, or ones for which experience has no meaning. If a skill is not rated for experience, the skill level will not be shown in the skill listing.

For instance:

    Skills known:
       1001  Sailing, apprentice
       9999  Strength bonus

Since experience is not applicable to Strength bonus [9999], its level is not shown.

3.3.9 Guild Skills

Some skill categories have guilds which zealously protect the secrets to one or two important and powerful skills. To learn these skills, a noble must join the guild and study these skills in a guild hall (see ⟨undefined⟩ [Guild Halls], page ⟨undefined⟩).

A noble may be a member of only one guild, and once he joins a guild he can never quit. To join a guild, a noble should either (1) enter an appropriate guild hall, or (2) create a new guild hall using the guild command (see ⟨undefined⟩ [guild], page ⟨undefined⟩).

Note that whether a noble joins a guild by entering an existing guild hall or creating his own, he must first be ready to join the guild, i.e., he must have learned all the non-guild skills.

3.3.10 Study Points

A faction starts with some number of study points. The number of study points available to a faction is shown in the beginning of the turn report. Study points permit a noble to learn a common category skill (Combat, Shipcraft, Construction, Forestry, Mining, Ranger or Trading) without a source of instruction.

When a noble studies one of the above category skills with no other source of instruction available, he uses one study point for each day he studies. For example, a noble in the wilderness with no book who studied Forestry for seven days would use seven study points. Study terminates when the study points run out.

3.3.11 Summary of study

• The category skill must be learned before a sub-skill within the category may be studied.

• Category skills must be learned from a teacher. Teachable subskills are learned faster when a teacher is available. Unteachable subskills cannot be taught.

• Teachers include: some cities, other nobles using the teach command, and books of instruction.

• Guild skills are only taught within guilds.

• Beginning study of any skill costs 100 gold.

• Payment of Noble Points may be required to begin learning some advanced skills.

• An additional Noble Point is required for every category skill in excess of three known by a noble. However, no charge is made when learning Basic Religion [3500] or Advanced Sorcery [3300], and they are ignored when counting the category skills known by a noble.

• Diminishing returns from multiple studies within a month:

− First week: 7/7

− Second week: 4/7

− Third week: 2/7

− Fourth week: 1/7

• Once a skill is known, further study of that skill has no effect.

• Study points substitute for a source of instruction for the skills Combat, Shipcraft, Construction, Forestry, Mining, Ranger and Trading.

3.3.12 Religion

Be Thou my vision, oh Lord of my heart Nought be all else to me, save that thou art Thou my best thought in day and in night Waking or sleeping, Thou presence my light

Mary E. Byrne

There are eight religions in Olympia, corresponding to the eight sons of Sol that remained faithful during the Rebellion (see ⟨undefined⟩ [Appendix - Mythology of Olympia], page ⟨undefined⟩). Priests of these religions gather followers and are rewarded with holy energy (called “piety”) that can be used to perform various religious miracles.

Because magic was taught to mankind by Tor and the other rebellious sons of Sol, priests are fundamentally opposed to the practice of magic. Once a noble learns a religion skill, he can never study magic. Similarly, a priest’s follower cannot learn magic skills. Priests also refuse to stack with magicians, and compete with them in assembling the Staff of the Sun (see ⟨undefined⟩ [Staff of the Sun], page ⟨undefined⟩).

3.3.12.1 The Eight Religions

Each of the Olympia religions reflects in many ways the character of its god. For example, each religion considers one type of terrain holy, and only builds temples in that terrain. The table below summarizes some of the characteristics of the eight religions. The meaning of each column is described below.

God     Strength   Weakness   Plant       Animal Color   Terrain
---     --------   --------   -----       ------ -----   -------
Eres    Shipcraft  Mining     Seagrass    Otter  Blue    Near ocean
Anteus  Mining     Shipcraft  Firewort    Mole   Brown   Mountain
Dol     Alchemy    Forestry   Beastnip    Bull   Red     Plains
Timeid  Forestry   Alchemy    Green Rose  Eagle  Green   Woods
Ham     Persuasion Construct  Elf Poppy   Monkey Gold    City
Kireus  Construct  Persuasion Ironwood    Hare   Granite Swamp
Halon   Combat     Stealth    King's Fern Wardog Silver  City
Domingo Stealth    Combat     Moon Palm   Rat    Shadows Hidden

Strength Each religion is strong in a particular Olympia skill category. When a priest joins a religion, he automatically learns this strength skill if he does not already know it (see ⟨undefined⟩ [Becoming a Priest], page ⟨undefined⟩). (Learning the strength skill is free of the normal gold and NP costs.) Nobles dedicated to the god who are not priests learn skills in this category one day faster than normal. Priests may only join the guild of their strength skill (e.g., a Priest of Ham may only learn guild skills from Persuasion, and a noble who joins the Forestry guild can only become a priest of Timeid.)

Weakness Each religion also has a corresponding skill category in which it is weak. When a priest joins a religion, he automatically forgets this skill if he knows it, and he cannot learn it as long as he remains a priest (see ⟨undefined⟩ [Becoming a Priest], page ⟨undefined⟩). Nobles dedicated to the god who are not priests learn skills in this category two days slower than normal.

Plant Each god has placed on Olympia one plant of special holy significance. Many of the miracles given to priests require these plants.

Animal Each god has also placed on Olympia one holy animal. Currently there is no significance to these animals in game terms.

Color The gods normally appear to mortals in particular guises, often wearing clothes of a particular design or color. Worshippers often emulate their gods by wearing clothes or tokens of the same color.

Terrain Each god considers a particular type of terrain holy, and only permits his temples to be built in that holy terrain (see ⟨undefined⟩ [Temples], page ⟨undefined⟩). Eres, the God of the Sea, requires his temples to be built in a province adjoining the ocean (or on an island). Domingo, the Hidden One, requires that his temples be built in hidden locations.

3.3.12.2 Becoming a Priest

To become a priest, a noble must study the base skill for the appropriate religion:

• [2000] Church of Eres

• [2100] Church of Anteus

• [2200] Church of Dol

• [2300] Church of Timeid

• [2400] Church of Ham

• [2500] Church of Kireus

• [2600] Church of Halon

• [2700] Church of Domingo The base skill for a religion can be studied in a temple of the religion,

from a priest providing instruction via ⟨undefined⟩ [teach], page ⟨undefined⟩ or from a holy book of the religion.

Only priests of the appropriate religion can enter a temple, so to learn the base skill for a religion at a temple, a noble must be escorted inside the temple by a priest of that religion. Specifically, the noble must be stacked underneath the priest when he enters the temple. Once inside, the noble may study the base skill and become a priest himself.

Holy books occasionally appear for sale in city markets, are found in ancient ruins, as monster treasure, etc. A noble may learn the base skill for a religion from a holy book without being in a temple.

Learning the base skill for a religion costs 1 noble point (NP). A noble may only be a priest of one religion at a time. Furthermore, because the magical arts derive from the rebellious children of Sol (see ⟨undefined⟩ [Appendix - Mythology of Olympia], page ⟨undefined⟩), a priest may not learn any of the magical arts. If a noble knows any of the magical arts when he commits himself to a god, he will forget those skills.

When a noble learns the base skill for a religion, he will also automatically learn the strength skill for the religion and forget the weakness skill of the religion. He will also learn Basic Religion [3500]. He will not be charged gold or noble points for learning these skills (even if they are the 4th or later skills). Furthermore, he cannot learn the weakness skill so long as he remains a priest. For a list of the strength and weakness skills for the religions, see ⟨undefined⟩ [The Eight Religions], page ⟨undefined⟩.

3.3.12.3 Miracles

Once a noble has become a priest, he may learn a variety of prayers specific to his religion. These prayers permit the priest to perform various miracles. Each religion has specific prayers, but each religion includes these common prayers:

• Devotional Prayer

• Gather Holy Plant

• Bless Follower

• Proselytise

• Create Holy Symbol

• Heal

• Last Rites

• Resurrect dead noble

For specific descriptions of these prayers, see ⟨undefined⟩ [Appendix - Skill Dictionary], page ⟨undefined⟩.

Prayers require differing amounts of piety to use. Most prayers also require that the priest possess a holy symbol (which can be made through the prayer “Create Holy Symbol”), and some prayers require a holy plant (which can be gathered in the priest’s holy terrain with the prayer “Gather Holy Plant”).

The effects of prayers can never cross oceans.

3.3.12.4 Piety

How strangely high endeavours may be blest, Where piety and valour jointly go.

John Dryden

Piety is holy energy that the gods share with their priests so that the priests may perform miracles. The gods grant piety to the priests who most please them. There are two ways to please the gods: through devotional prayer, and by gathering followers for the god.

In devotional prayer, the priest removes himself to a holy place and spends a week contemplating the nature of holiness, praising his deity, and performing holy rituals. His god rewards this activity by granting him 2 additional piety. To perform devotional prayer, a priest should be in one of his god’s temples, and use the “Devotional Prayer” prayer.

The gods are also pleased by priests who recruit new worshippers. Each month, the gods grant their priests piety based on how many followers each priest has. Followers are non-priest (and non-mage) nobles who have dedicated themselves to a priest (see ⟨undefined⟩ [Followers], page ⟨undefined⟩).

For each religion, each month the priest with the most followers receives the designation of “High Priest” and 12 piety. He also receives 10 angels (see ⟨undefined⟩ [Monster combat ratings], page ⟨undefined⟩) from his god.

These angels will be in the High Priest’s control until the end of the month, when they are recalled to the spiritual realm.

The two priests with the next most numerous followers receive the designation of “Bishop” and 6 piety. They also receive 5 angels each.

For skills and prayers which require a priest to be alone, angels are not counted. That is, a priest with 10 angels is still considered to be "alone".

Of the remaining priests, the half with the most followers receives 3 piety each. The other half receives no piety.

3.3.12.5 Followers

A follower is a non-priest noble who has dedicated himself to a priest (and thus to the priest’s god). To become a follower, a noble should use the dedicate command (see ⟨undefined⟩ [dedicate], page ⟨undefined⟩). The dedication ceremony takes seven days, and requires that the noble be in the presence of the priest of the appropriate religion.

Mages cannot become followers.

Once a noble has dedicated himself to a priest, he may not change that dedication unless the priest dies and is laid to rest.

There are several advantages to becoming a follower of a priest.

First, a noble may only be blessed or resurrected by the priest he follows. A noble who does not follow any priest cannot be resurrected. The exception is priests themselves, who can be resurrected by any other priest of their religion.

Second, a follower receives bonuses when using his religion’s strength skill. In particular, he will learn skills in the strength skill category one day faster than normal, and uses some skills that take at least 5 days one day faster than normal.

There are also some disadvantages.

A follower receives penalties when using his religion’s weakness skill. In particular, he will learn skills in the weakness skill category two days slower than normal, and uses any weakness skill two days slower than normal.

Followers also cannot learn magic skills.

3.3.12.6 Special Religious Troops

Postulants are peasants dedicated to a particular priest. They are recruited by use of the "Proselytise" prayer. Postulants have the same statistics as peasants. However, because of their fervent dedication to their god, postulants are willing to serve the priest who recruited them for nothing (i.e., they have no maintenance costs). By the same token, though, they are unsuitable for training into workers, soldiers, etc.

Postulants also cannot be transfered between nobles.

Priests of Halon may convert soldiers [12] in their employ into fanatics [28]. Fanatics have twice the attack factor of normal soldiers, and no maintenance costs, but cannot be transfered between nobles.

Priests of Domingo may convert soldiers [12] in their employ into ninjas [29]. Ninjas have the same attack and defense ratings as normal soldiers, but are stealthy. Ninjas cannot be transfered between nobles. For skills and prayers which require a noble to be alone, ninjas are not counted. That is, a priest with 10 ninjas is still considered to be "alone".

3.3.12.7 Holy Symbols

The holy symbol is a specially created and blessed artifact that permits a priest to more easily contact his god. It is a necessary component for many prayers.

Holy symbols are specific to each religion, but one priest of a religion may create a holy symbol for use by another priest of the same religion.

Holy symbols are created with the prayer “Create Holy Symbol”. This prayer requires 1 noble point (NP) to learn.

3.3.12.8 Holy Plants

The gods have blessed certain plants with spiritual powers. Each religion has a different holy plant. These plants are necessary components of some prayers.

Holy plants are gathered through the prayer “Gather Holy Plant.” Holy plants are rare and must be gathered using a strict ritual. For these reasons, there’s only a 50% chance that a priest will find and successfully gather a holy plant with each 7 day use of “Gather Holy Plant.”

3.3.12.9 Temples

Temples are sacred places of worship. Priests use temples to study prayers, to perform devotional prayer, and to perform certain difficult miracles, such as resurrection.

To create a new temple, a priest must perform a special dedication ceremony in an undedicated temple. (Temples may be built by any noble with the skill Construction [1400].) This is done through the prayer “Dedicate Temple”. In a week long ceremony, the priest prays himself into a religious ecstasy and then blesses the structure, inscribes it with various holy symbols, and decorates it with icons. The decorations and other required implements cost 500 gold.

To be dedicated, a temple must be located in the priest’s holy terrain. Only one temple (of any particular religion) may be built in any location.

Only priests of the appropriate religion may enter a temple. Non-priests wishing to enter a temple must be stacked underneath a priest when he enters. Temples may be attacked like other buildings, but cannot be occupied unless the attacker is an appropriate priest (see ⟨undefined⟩ [attack], page ⟨undefined⟩). Temples, like any building, may be razed from the outside if they are unoccupied (see ⟨undefined⟩ [raze], page ⟨undefined⟩).

Temples produce monthly income in offerings from devoted followers in the nearby countryside. 1 gold is produced for each 10 peasants in the province in which the temple is located, up to a maximum of 100 gold per month income. (The magical artifact ⟨undefined⟩ [Improved Income], page ⟨undefined⟩ can increase this maximum.) The temple income is received only if a priest of the proper religion is in the temple at the end of the month. Undedicated temples do not produce income. If there is a magician in the province at the end of the month no income will be generated, as the local peasants will be too fearful to come to the temple.

3.3.12.10 Miscellaneous Notes

The average peasant distrusts anything unusual, and so will only work for familiar religions. For this reason, priests may only recruit peasants in provinces which contain a temple of their religion.

Because of the basic enmity between magicians and priests, neither will stack with the other. Of course, if either or both are cloaking their professions, stacking is possible. Also, priests will not stack with any unit possessing undead.

Priests also give troops great encouragement when attacking magicians or undead. This encouragement adds +10 to the attack factor of any unit attacking from a stack containing a priest against a stack containing a magician or undead.

3.3.13 Magic

Do not meddle in the affairs of wizards, for they are subtle and quick to anger.

J.R.R. Tolkien Magic is the mysterious and powerful art by which events are influenced

outside of the normal boundaries of cause-and-effect. Rather than the glamorous ideal of shining wizards casting powerful fireballs at wicked foes, the reality of magic instead tends to be base, tedious work which earns few friends.

Because magic was taught to mankind by Tor and the other rebellious sons of Sol (see ⟨undefined⟩ [Appendix - Mythology of Olympia], page ⟨undefined⟩), magicians are fundamentally opposed to the religious orders. Once a noble studies magic, he can never study any religion skill. Magicians also refuse to stack with priests, and compete with them in assembling the Staff of the Sun (see ⟨undefined⟩ [Staff of the Sun], page ⟨undefined⟩).

The casting of a magical spell is accomplished with three ingredients: Knowledge of the spell, possession of any items necessary to fuel the spell, and a sufficient level of magical aura to perform the ritual or ceremony.

As with religious miracles, the effects of magic spells cannot cross open water (i.e., oceans).

3.3.13.1 Aura

Aura is a mystical force necessary to cast spells. Powerful mages will have a high aura rating, while apprentice sorcerers may only command a few points worth of aura.

Nobles are rated for their current aura level and their maximum aura level. With each magical spell learned, a noble will gain one point of maximum and current aura. Current aura is depleted by casting spells, and is naturally replenished at a rate of two points per turn (four points per turn with an auraculum). Current aura will increase until it reaches the maximum aura level.

Other ways of gaining aura include meditation (⟨undefined⟩ [Meditate [2801]], page ⟨undefined⟩), tapping your health (⟨undefined⟩ [Tap health for aura [2832]], page ⟨undefined⟩), destroying artifacts (⟨undefined⟩ [Destroy artifact [3103]], page ⟨undefined⟩), and power jewels (⟨undefined⟩ [Power Jewel], page ⟨undefined⟩).

Current Aura may be at most 5 times the maximum effective aura (maximum aura plus aura from item bonuses).

Spells are rated on the amount of aura which the casting mage must possess. Minor spells may demand only one point of aura to cast. Powerful spells may require a current aura level of ten or higher.

3.3.13.2 Magician status

Nobles receive a rating in the turn reports based on their magical abiliity:

    maximum aura            label
    ------------            -----
       6-10                 conjurer
      11-15                 mage
      16-20                 wizard
      21-30                 sorcerer
      30+                   ??

For example:

    Osswid the Brave [5639], wizard, with three workers

The Basic magic [2800] spell Appear common [2803] allows magicians to prevent this label from displaying.

3.3.13.3 Required Items

Many spells will require the magician to possess a rare or obscure item in order for the cast to succeed. Many of these items exist, of interest chiefly to sorcerers. Roots from plants found only in dense forests, bat’s wings, and a dark blue powder which produces a brilliant cobalt flame when burned are only a few. Usually the required item is consumed by the attempt to cast the spell.

3.3.13.4 Study of Spells

Magic is divided into six schools of study:

    num  name                       time to learn
    ---  ----                       -------------
    2800  Magic                      two weeks     1 NP req'd
    2900  Scrying                    three weeks   1 NP req'd
    3000  Gatecraft                  three weeks   1 NP req'd
    3100  Artifact construction      four weeks    1 NP req'd
    3200  Necromancy                 four weeks    1 NP req'd
    3300  Advanced sorcery           four weeks

Magical spells are simply sub-skills of one of the magical skill categories. An aspiring mage will issue the study order to learn the basics of a

particular school of magic, known as the category skill. For example, one wishing to pursue knowledge of Magic [2800] would order:

    study 2800

A noble must know the category skill for a school of magic before a spell in that school can be known.

Knowledge of individual spells in a school of magic is not possible without having learned the category skill.

3.3.13.5 Schools of Magic

Magic [2800] The most common and well-known of the magical schools, Magic nonetheless has many useful and powerful spells.

Scrying [2900] Scrying is the study of magical far-seeing, the ability to view images or learn information at a distance.

Gatecraft [3000] The study of ancient portals of teleportation which long ago connected distant regions of the land. Knowledge of this lost art tends to be difficult to acquire and jealously guarded when found.

Artifact construction [3100] The realm of advanced sorcerers, artifact construction concerns the making of physical items to focus, amplify or otherwise enhance a mage’s power. Most spells in this school require the magician to possess high levels of aura.

Necromancy [3200] The darkest of the Dark Arts, necromancy involves trafficking with undead or demon spirits to gain knowledge. Hated by civilized men, feared by other magicians, the necromancer seeks power and domination over the physical world.

Advanced sorcery [3300] The wizard who has mastered the other five schools of Magic may attempt to learn these strange, advanced arts. The only way to learn this art is to forge an auraculum.

Monsters may sometimes be found guarding ancient books which teach the rare magical skill categories (see ⟨undefined⟩ [Guard monsters], page ⟨undefined⟩).

3.3.13.6 Magical Combat

Some spells are classified as "combat spells". These are used automatically in combat, during the "special attacks" phase of each round. Combat spells may strike opposing forces with lightning or fireballs, drain opposing magicians of aura, or have other effects.

For a description of how combat spells are used in battle, see See ⟨undefined⟩ [Combat Spells], page ⟨undefined⟩.

3.4 Training men

Nobles may be accompanied by different kinds of men. Men may be peasants obtained with the recruit order, or peasants who have been trained into other kinds of men such as sailors, soldiers or workers.

The different kinds of men:

• peasant [10]

• worker [11]

• soldier [12]

• archer [13]

• knight [14]

• paladin [15]

• pikeman [16]

• sailor [19]

• lt. foot [20] • lt. xbowman [21]

• elite archer [22]

• pirate [24]

• skirmisher [36]

• hvy. foot [37]

• cavalier [43]

• hvy. xbowman [38]

• horse archer [41]

The peasant is the basis of all these types of men. Starting with a peasant, a noble with the proper skills can slowly train him to be any one of these types of men. The possibilities are shown in this chart:

            +--- sailor ------- pirate
           +---- worker              +--- ninja
          /                         +---- fanatic
         /                         +----- lt. xbowman
        /                         +------ hvy. xbowman
       /                         /
    peasant --- skirmisher --- soldier --- lt. foot --- hvy. foot
                                 \
                                  +--- cavalier
                                   +--- pikeman
                                    +--- archer
                                          \
                                           +--- horse archer
                                            +--- elite archer
    hvy. foot --- knight--- paladin

Training a man may require that the noble have a certain skill, or possess some item. For instance, training soldiers into light foot requires a longsword [74] for each footsman produced. The training noble must also know Train Swordsmen [1105].

Some men may only be trained in certain kinds of locations. For example, pirates may only be trained on board a ship.

Training takes one day per man. Training five archers into five elite archers would take five days, for example.

num  kind            skill  input man       input item        where
---  -----------     -----  --------------  ----------------  ------
11  worker           none  peasant [10]
12  soldier          1108  skirmisher [36]
13  archer           1902  soldier [12]    longbow [72]
14  knight           1190  hvy. foot [37]  warmount [53]
15  paladin          1191  knight [14]                       guild
16  pikeman          1100  soldier [12]    pike [75]
19  sailor           1001  peasant [10]
20  lt. foot         1105  soldier [12]    longsword [74]
21  lt. xbowman      1100  soldier [12]    lt. crossbow [85]
22  elite archer     1902  archer [13]     elvish arrow [39]
24  pirate           1105  sailor [19]     longsword [74]    ship
28  fanatic          2635  soldier [12]
29  ninja            2733  soldier [12]
37  hvy. foot        1192  lt. foot [20]   plate armor [73]
36  skirmisher       1100  peasant [10]
38  hvy. xbowman     1100  soldier [12]    hvy. crossbow [40]
41  horse archer     1190  archer [13]     riding horse [52]
43  cavalier       1190  soldier [12]    mithril [82], warmount[53]

For order for more information and examples, see ⟨undefined⟩ [train], page ⟨undefined⟩.

3.4.1 Maintenance cost

Men such as soldiers, workers, archers, etc. must be paid in gold monthly or they will leave the service of their noble. Peasants do not willingly leave a noble’s service, but will starve if they are not paid. This cost is charged to the noble holding them at the end of each month.

If the noble does not have enough gold to pay his men, he will ask other nobles in his stack (provided they belong to the same player) for gold. Thus, only one member of a stack need carry gold for maintenance costs for the entire stack. Nobles will not share gold with units from other players.

If the noble can only afford to pay some of his men, one-third of those not paid will leave service at the end of the month. The computer chooses which men remain and which leave or starve.

num   kind            cost
---   -----------     ----
28   fanatic          0
27   postulant        0
10   peasant          1
36   skirmisher       1
11   worker           2
19   sailor           2
12   soldier          2
29   ninja            2
21   lt. xbowman      2
38   hvy. xbowman     2
13   archer           3
41   horse archer     3
16   pikeman          3
20   lt. foot         3
24   pirate           3
37   hvy. foot        4
22   elite archer     4
43   cavalier        4
14   knight           5
15   paladin          2

Note that nobles may drop men to release them from service deliberately.

3.4.2 Making weapons and armor

Weapons and armor are required for the training of some kinds of fighters. Archers require longbows, for instance, hvy. foot require plate armor, etc.

Weapons and armor are made with the make command (see ⟨undefined⟩ [make], page ⟨undefined⟩). The Weaponsmithing [1106] subskill of Combat [1100] is required to make weapons and armor.

num   item           material
---   ----           --------
72   longbow        yew [68]
73   plate armor    iron [79]
74   longsword      iron [79]
75   pike           wood [77]
85   lt. crossbow   wood [77]
40   hvy. crossbow  wood [77] & iron [79]

One unit of the input material may turned into one of the desired items each day. For example, make 72 2 would spend two days turning two yew [68] into two longbows [72].

4 Movement

Each location report indicates the adjacent locations and the time it would take to travel there:

    Routes leaving Great Forest:
       North, forest, to Great Forest [bc27], 8 days
       East, to Ocean [bd27], West Ocean, impassable
       South, forest, to Great Forest [bc25], 8 days
       West, forest, to Great Forest [bc28], 8 days
    Inner locations:
       Elgethon City [b2b3], port city, 1 day

The times shown are for walking. If a unit can ride (i.e., has a riding capacity greater or equal to its weight), then it will travel twice as fast. If a unit can fly (i.e., has flying capacity greater or equal to its weight), it will travel even faster (see ⟨undefined⟩ [fly], page ⟨undefined⟩).

Note that a direction may be listed as "impassable" to indicate that it cannot be travelled.

To travel at walking or riding place, use the ⟨undefined⟩ [move], page ⟨undefined⟩ command. To fly, use the ⟨undefined⟩ [fly], page ⟨undefined⟩ command. To sail a boat, use the ⟨undefined⟩ [sail], page ⟨undefined⟩ command.

4.1 Stacking

One unit may stack under another unit. Two or more units grouped in this way are referred to as a stack.

Stacks move together and fight together. Here is a stack of four units:

    Law Netexus [2020], accompanied by:
      Feasel the Wicked [1109]
      Drakkar the Trader [1752]
      Alion Krysaka [2785]

Law Netexus is the stack leader, the top-most unit in the stack.

Only one level of stack depth is shown, so all that can be determined from the location report is that Feasel, Drakkar and Alion are stacked somewhere beneath Law Netexus. The exact arrangement of stacking bonds is not shown.

Feasel might be stacked under Law Netexus, with Drakkar and Alion under Feasel. Or Feasel, Drakkar, and Alion may all be stacked directly beneath Law Netexus.

Generally, such internal arrangements are only important when the stack breaks up. If Drakkar is stacked beneath Feasel, he will stick with Feasel if Feasel drops out of the stack. But if Drakkar were stacked beneath Law Netexus, he would not follow Feasel if Feasel unstacked.

If Law Netexus issues a move order, the entire stack will move. If, however, Feasel issues a move order, he will first drop out of the stack before moving.

Similarly, if Law Netexus engages combat with the attack order, all nobles in the stack will fight together. If one member of the stack is attacked, the entire stack will respond in defense.

Multiple levels of internal stacking can be useful if one wants several stacks to join together for a while, but then split apart later into their old arrangements.

Ocean ships may not be stacked together. There is no way to cluster ships into a fleet.

4.1.1 Allowing stacking

By default, a unit may not stack with a noble belonging to another faction. A unit is also denied entry to a location controlled by another player if that location’s border is closed.

Players may allow units from other factions to stack with them or enter buildings or ships they control with the admit order. (See also ⟨undefined⟩ [Control of Locations], page ⟨undefined⟩).

4.2 Carrying capacity and Travel

Men and items are rated for how much they weigh, and how much they can carry, walking, riding or flying.

                    weight   walking    riding   flying
                    ------   -------    ------   ------
    noble             100      100        -         -
    peasant           100      100        -         -
    worker            100      100        -         -
    riding horse      300      300       150        -
    wild horse        300      self      self       -
    warmount          300      400       200        -
    ox               1000     1500       self       -
    winged horse      300      150       150       150
    lt. wagon   500     1000       500* -
    hvy. wagon   900     1800       900 -
    war wagon   900     1800       900* -
    * Light and war wagons cannot move at riding pace in mountains or swamps.

Note that the only men that will carry anything are peasants, workers and nobles. Military units (such as knights) refuse to demean themselves by acting as packhorses. Skirmishers are simply too unreliable and drunken to carry anything.

In order to ride, the total riding capacity must cover the weights of all the units that may not ride themselves.

Examples:

peasant + riding horse ride capacity is 150 - 100 = 50 walk capacity is 300 + 100 = 400

peasant + riding horse + wild horse ride capacity is 150 - 100 + 0 = 50 walk capacity is 100 + 300 + 0 = 400

The wild horse can walk or ride on its own, but will not carry anything.

peasant + riding horse + ox ride capacity is 150 - 100 = 50 walk capacity is 100 + 300 + 1500 = 1900

The ox may be driven alongside the horse, but will not carry anything when moving so quickly.

A stack will ride if there is enough riding capacity to carry all of the non-riders. Otherwise, the stack will walk. (See also ⟨undefined⟩ [Swamps], page ⟨undefined⟩.)

In order to fly, the total flying capacity for the stack must cover the weights of all units that can not fly themselves.

Stacks which are overloaded beyond their walking capacity will travel slower than normal. A stack loaded at 150% of its walking capacity will take 50% longer to traverse a route. Stacks overloaded to over 200% of walking capacity may not travel at all.

Weights and capacities are always considered for the stack as a whole. One unit may have all the men, and another unit may have all the horses. If they are stacked together, the distribution of items across units is irrelevant.

4.2.0.1 Swamps

Swamps, bogs and pits are so difficult to negotiate that they present special difficulties for the traveller.

No stack may move at riding speed into or inside a swamp province, or into a bog or pit. At best, a stack can move at walking pace.

If a stack contains animals, the difficult terrain will further slow travel by an extra day. A wagon is considered an animal because of its horses.

Knights,paladins and cavaliers receive a -50 penalty to both their attack and defense ratings when fighting on a ship or in a swamp province.

4.3 Ships

I must go down to the seas again, to the lonely sea and the sky, And all I ask is a tall ship and a star to steer her by, And the wheel’s kick and the wind’s song and the white sail’s shaking, And a gray mist on the sea’s face, and a gray dawn breaking.

Sea Fever, John Masefield

By studying Shipcraft [1000], a noble may learn to build and sail ships. Ships may be built to any desired configuration: wide-beamed, capacious cargo vessels, sleek and fast cutters, or heavily fortified warcraft.

4.3.1 Making ships

Building a ship takes several steps. First the shipbuilder lays down a hull, which determines the size and carrying capacity of the ship. Then he adds sails, rowing ports, a keel and fortifications, mixing and matching these features to create a ship with precisely the speed, capacity, and combat characteristics he desires.

Laying a hull requires a noble to have the Build Hull [1002] skill, a subskill of Shipcraft [1000]. To begin construction of a ship, the shipbuilder should unstack from beneath other nobles and issue the following build order:

build ship name hull-size The build ship order takes two arguments. The first is the name of the

ship to be built. The second is the size of the hull. Each unit of hull has a capacity of 2500 units of weight, and requires 10 wood and 100 worker-days to build. So the command:

    build ship "HMS Bounty" 3

would begin construction of a ship hull named "HMS Bounty" with a capacity of 7500 units weight, requiring 30 wood and 300 worker days to build.

There is no limit to the size of hull that may be built.

One-fifth of the lumber required will immediately be deducted from the shipbuilder’s inventory and put to use building the new ship. The shipbuilder and his workers will be placed inside the new ship.

At least three workers are needed to begin construction of a ship.

Until the ship is completed, it will be shown as ‘in progress’:

    Ships docked at port:
       HMS Pinafore [ih39], ship-in-progress, 28% completed, owner:
          Osswid the Constructor [5499], with five workers

The remaining construction materials are deducted from the builder’s inventory as work on the ship progresses. A second fifth of the lumber is required when the ship becomes 20% complete, a third fifth when the ship becomes 40% complete, etc. Construction halts if the builder runs out of materials

For example, a noble who wanted to build a ship hull of size 2 would need to know Build Hull [1002], have at least three workers, and start with at least 4 wood [77]. (16 more wood is required to bring the hull to completion).

To resume construction of a partially completed ship, first enter the ship, then issue build ship.

Once the laying of a hull has begun, the size of the hull cannot be changed. Ships may only be built in locations with a passable route to an ocean. Operating a ship requires 2 sailors [19] for every 20 units of hull. Addi-

tional sailors may be required to work the sails, row, etc.

4.3.2 Ship Features

A newly-completed hull is merely the skeleton for a ship. While it can be slowly rowed from place to place, to be of any real use it will need to have sails or rowing ports added and possibly a keel, fortifications and a ram. Each of these features is added with a Shipcraft [1000] subskill:

    Add Sails [1005]
    Add Fortification [1091]
    Add Rowing Ports [1006]
    Add Keel [1090]

There are also corresponding skills to remove features:

    Remove Sails [1096]
    Remove Fortification [1093]
    Remove Rowing Ports [1097]
    Remove Keel [1092]

4.3.2.1 Rowing Ports

Rowing ports are shuttered openings set into the hull of the ship and fitted with oarlocks. Alongside each rowing port is a bench on which the rower sits to pull his oar. With sufficient rowers, a ship can travel quickly regardless of wind conditions. In Olympia, sailors [19] double as oarsmen.

Each rowing port accommodates 4 oarsmen. (Each port is actually a matched pair of ports, one on either side of the hull, each seating a pair of oarsmen.) Each unit of hull may have a maximum of 3 ports, and each newly-constructed ship comes with one rowing port (not one per hull, but one total).

Adding rowing ports to a ship requires the Add Rowing Ports [1006] skill. Each additional port added to a ship requires 2 wood [77], 10 workers and takes 3 days. To add a rowing port to a ship, the shipbuilder should enter the ship and issue the order:

    use 1006

At the conclusion of the command, a new rowing port will be added to the ship.

To remove a rowing port from a ship, the shipbuild should enter the ship and issue the order:

    use 1097

Removing a rowing port requires 10 workers.

Rowing a ship at normal speed requires one fully-staffed rowing port per unit hull. For example, rowing a 3 unit hull ship at normal speeds requires that the ship have 3 rowing ports and at least 14 sailors (2 to man the hull and 4 each for the 3 rowing ports).

Each additional fully-staffed rowing port per unit hull cuts a ship’s rowing travel time by one day. For example, a 3 unit ship with 6 fully manned rowing ports would cut its travel time by one day, crossing a 6 day ocean province in just 5 days.

Conversely, if a ship has less than one fully-manned rowing port per unit hull, it will travel more slowly. For each unit hull which does not have a fullystaffed rowing port, one day will be added to the travel time. For example, a 3 hull ship with 3 rowing ports and 7 sailors would only be able to fully man 1 rowing port (2 sailors to man the hull, leaving 5 sailors to row), and would add 2 days to any travel time.

Finally, a ship with less than one fully-manned rowing port for each four units of hull cannot travel at all. For example, a 5 hull ship with 7 sailors on board would only be able to man 1 rowing port, and would not be able to row at all.

4.3.2.2 Sails

Sails permit a ship to use the wind for propulsion, cutting down greatly on the number of sailors required and sometimes providing the speediest of travel. With sufficient sail, a ship can cross Olympia’s wide oceans as fast as the wind can blow.

Each unit of hull may mount a maximum of 3 sails. Each sail requires one sailor to tend it, up to a maximum of 100 sailors. A sail uses up 250 units of a ship’s capacity.

Adding sails to a ship requires the Add Sails [1005] skill. Each addition sail added to a ship requires 1 wood [77] and 10 workers and takes 3 days. To add a sail to a ship, the shipbuilder should enter the ship and issue the order:

    use 1005

At the conclusion of the command, a new sail will be added to the ship. To remove a sail from a ship, the shipbuild should enter the ship and

issue the order:

    use 1096

Removing a sail requires 10 workers.

Sailing a ship at normal speed requires two fully-staffed sails per unit hull. For example, sailing a 3 unit hull ship at normal speeds requires that the ship have 6 sails and at least 8 sailors (2 to man the hull and 1 each for the 6 sails).

Each additional fully-staffed sail per unit hull cuts a ship’s travel time by two days. For example, a 3 unit ship with 9 fully manned sails would cut its travel time by two days, crossing a 6 day ocean province in just 4 days.

Conversely, if a ship has less than two full-staffed sails per unit hull, it will add two days of travel time. A ship with less than one sail per two units hull cannot move at all.

4.3.2.3 Keels

A keel is a longitudinal plate projecting downward from the center bottom of the boat. Made of heavy iron, it provides ballast and stability, and is essential equipment for ships that will be attempting ocean crossings. With sufficient keel, a ship will have a much-improved chance to survive the powerful storms of the open ocean.

Each unit of hull may mount one unit of keel. For example, a 3 unit hull could mount up to 3 units of keel.

Adding keel to a ship requires the Add Keel [1090] skill. Each unit of keel added to a ship requires 1 iron [79], 10 workers and takes 5 days. To add keel to a ship, the shipbuilder should enter the ship and issue the order:

    use 1090

At the conclusion of the command, a new unit of keel will be added to the ship.

To remove keel from a ship, the shipbuild should enter the ship and issue the order:

    use 1092

Removing keel requires 10 workers.

Due to the keel’s heavy weight, each unit of keel added to a hull reduces the capacity by 100 units weight.

Keel protects a ship against storm damage. Ships caught in a coastal storm have the following chance of sustaining damage:

    (15 - (units keel / unit hull) * 30)%

Thus a ship with at least one keel per two units of hull will (under normal circumstances) be immune to damage from coastal storms.

Ships caught in open ocean storms have the following chance of sustaining damage:

    (35 - (units keel / unit hull) * 30)%

Even a fully keeled ship (one unit of keel per unit of hull) still risks a 5% chance of damage in an open ocean storm.

The storm’s strength determines how much damage it will cause. A storm of strength "n" will do 1-n points of damage to a ship, although no storm will do more than 50 points damage to a ship in one blow. Multiple storms in a province act as one storm with strength equal to the sum of the strengths of all the storms. Only rain storms cause ship damage.

Keels also protect a ship from reef damage. Reef damage can occur when a ship sails along the coast, particularly if the coast is rocky (i.e., mountainous). There is a 1 in 75 chance per day that a ship is in a coast province that it will sustain 3 to 5 points of damage. This chance goes up to 1 in 50 to sustain 6 to 10 points of damage if the coast is rocky. Reef damage to a ship is reduced by (units keel / units hull)%. A fully keeled ship will not be damaged by reefs.

4.3.2.4 Ship Fortifications

Unlike merchantmen or coastal skiffs, warships are built to repel boarders and sustain damage from attack. While heavy and costly, these fortifications can mean the difference between victory and a watery grave.

Each unit of fortification added to ship uses 250 units of the ship’s carrying capacity. A fortified ship provides cover for

    (units fortification / units hull) * 10

men; each of these men receives a defensive bonus in combat of

    (units fortification / units hull) * 5

For example, a 3 unit hull carrying 7 units of fortification would provide a defensive bonus of +11 to the first 23 fighters on board the ship.

Adding fortification to a ship requires the Add Fortification [1091] skill. Each unit of fortification added to a ship requires 5 wood [77], 15 workers and takes 7 days. To add fortification to a ship, the shipbuilder should enter the ship and issue the order:

    use 1091

At the conclusion of the command, a new unit of fortification will be added to the ship.

To remove fortification from a ship, the shipbuilder should enter the ship and issue the order:

    use 1093

Removing fortification requires 15 workers.

4.3.3 Repairing a Ship

Ships may be damaged while sailing by storms, or by submerged rocks in coastal waters. Damaged ships may be repaired with the repair command. Repairing a damaged ship requires one unit of pitch [261].

The capacity of a damaged ship is a reduced in proportion to the amount of damage it has sustained.

4.3.4 How to sail

The king sits in Dumferline town, Drinking the blood-red wine: "O whar will I get a good sailor To sail this ship of mine?"

Sir Patrick Spens (traditional)

The Sailing [1001] subskill of Shipcraft [1000] is required to pilot a ship. Nobles with the Sailing skill can pilot ships using the ⟨undefined⟩ [sail], page ⟨undefined⟩ command.

A ship with both sails and rowing ports will always travel by the faster method, selecting rowing or sailing as appropriate. The ship need not have enough sailors to row and sail simultaneously, since that is never done.

Piloting a ship in deep water, out of sight of land, is dangerous without special skills. The captain must be able to navigate without landmarks, anticipate weather on the open sea, and so on. The Shipcraft [1000] subskill Navigate Deep Seas [1094] gives a captain these special skills.

A ship’s captain who attempts to pilot a ship in deep waters without this skill runs a 1/3 chance for each deep sea province he crosses of encountering some sort of problem:

    Chance  Problem
    ------  -------
     40%    3-10 days extra travel time.
     40%    Travel in a wrong direction.
     15%    Rough seas causing 1-20 points damage to ship.
      5%    Capsize!

Unlike land movement, which slows down when overloaded, an overloaded ship cannot be sailed at all.

4.3.5 Operating a ferry

Ferry me across the water,

Do boatman, do.

If you’ve a penny in your purse I will ferry you.

I have a penny in my purse,

And my eyes are blue;

So ferry me across the water, Do, boatman, do.

Step into my ferry-boat, Be they black or blue, And for the penny in your purse I will ferry you.

Boatman, Christina Giorgina Rosetti

There are several commands useful for operating a commercial ferry:

fee Sets the fee passengers will be charged

board Passengers use this to pay their fee and board a ferry

unload Unload passengers once the destination is reached

ferry Signals passengers waiting in port that they may board

The captain of a ferry uses the fee command to set the fee charged to passengers boarding his ship (see See ⟨undefined⟩ [fee], page ⟨undefined⟩). The fee can be expressed in terms of the number of nobles boarding, the number of men boarding, the weight boarding or any combination of the three.

For instance, if the captain wanted to charge 1/2 gold per unit weight of the passenger’s stack, he would issue ‘fee weight 500’. (a 500 gold fee for every 10000 weight).

Note that like any fee, the fee set for a ferry will not take effect until the end of the month.

The fee is a property of the ship, not the captain, but will be cleared when the last noble leaves the ship. The fee may also be cleared with the fee command. If no fee is set, then the ship is not considered to be operating as a ferry, and nobles are may not use the board order to enter the ship.

Passengers issue ‘board ship’ to board a ferry. The order will fail if the ship is not present, or if it is not operating as a ferry (the captain of the ship has no fee set). board will cause the noble to pay the captain the required boarding fee, then move the noble’s stack onto the ship.

The board command also prevents the ship from being overloaded. A stack that would overload the ship is turned away.

Note that the board command, unlike the move command, ignores the admit list of the ship’s captain.

4.3.6 Ferry synchronization

Suppose now that the captain has a ship, has set a fee, announced his service in the Game Newsletter and with post, and is now ready to ferry passengers. What should he tell them? How will the synchronization work?

Passengers should travel to the port the ferry will be arriving at and issue

‘wait ferry ship’.

When the ship arrives at port, the captain should order unload to eject his current load of passengers, then ferry to signal any passengers waiting in port that they may now board.

Note that passengers should not use wait ship to wait for the ferry. Otherwise, they will attempt to board as soon as the ship reaches the port. The captain’s unload order may not have executed at this point. In this case, board orders might fail because there won’t be any room to enter the ship until the existing load of passengers has disembarked.

Example:

Posted by Captain McCook [3402]:
"Captain McCook's ferry to Drassa departs each
 Sunsear on the 15th.  The fee is 1 gold/wt.
 Issue WAIT FERRY ht34 then BOARD ht34 for
 ferry service.   Arrr!"

Captain McCook’s orders:

> fee wieght 1000                 # 1 gold/wt. is our fee
> sail ...                # arrival at port
> unload                  # unload current passengers
> wait day 15             # wait until stated time of departure
> ferry                   # sound our horn
> sail ...                # on to the next port

Passengers wishing to travel on McCook’s ferry:

> wait ferry ht34             # [ht34] is McCook's ship
> board ht34                  # pay our gold and embark
> wait loc destination  # Captain McCook will unload us at the
                              # end of the journey, so wait un-
til we
                              # find ourselves in the desti-
nation city.

5 Battles

Combat in Olympia occurs when one stack attacks another stack.

20: Jean Malreaux I [j9n], Imperial Empire, with two workers, one riding
20: horse, arrived from the south.
20: Stack of gorgon [fp0c] attacks Jean Malreaux I [j9n]!
20:
20:    Stack of gorgon [fp0c], gorgons, number: 7
20:
20:    Jean Malreaux I [j9n], Imperial Empire, with two work-
ers, one riding
20:    horse
20:
20: Combat round: one
20:   Melee phase:
20:     Stack of gorgon's forces hit seven times.
20:     Jean Malreaux I's forces hit one time.
20:       Stack of gorgon [fp0c] loses one gorgon [288].
20:       Jean Malreaux I [j9n] loses two workers [11].
20: Combat round: two
20:   Melee phase:
20:     Stack of gorgon's forces miss!
20:     Jean Malreaux I's forces hit one time.
20:       Stack of gorgon [fp0c] loses one gorgon [288].
20: Combat round: three
20:   Melee phase:
20:     Stack of gorgon's forces miss!
20:     Jean Malreaux I's forces hit one time.
20:       Stack of gorgon [fp0c] loses one gorgon [288].
20: Combat round: four
20:   Melee phase:
20:     Stack of gorgon's forces miss!
20:     Jean Malreaux I's forces hit one time.
20:       Stack of gorgon [fp0c] loses one gorgon [288].
20: Combat round: five
20:   Melee phase:
20:     Stack of gorgon's forces miss!
20:     Jean Malreaux I's forces hit one time.
20:       Stack of gorgon [fp0c] loses one gorgon [288].
20: Combat round: six
20:   Melee phase:
20:     Stack of gorgon's forces miss!
20:     Jean Malreaux I's forces hit one time.
20:       Stack of gorgon [fp0c] loses one gorgon [288].

20: Combat round: seven
20:   Melee phase:
20:     Stack of gorgon's forces miss!
20:     Jean Malreaux I's forces miss!
20: Combat round: eight
20:   Melee phase:
20:     Stack of gorgon's forces miss!
20:     Jean Malreaux I's forces hit one time.
20:       Stack of gorgon [fp0c] was hit for one health point.
20:
20: Jean Malreaux I [j9n] is victorious!
20:
20:    Jean Malreaux I lost two workers.
20:
20:    Stack of gorgon lost six gorgons.
20:    Stack of gorgon [fp0c] was killed.
20:

A unit will be defended by any other nobles it is stacked with. Thus, a unit which is part of a stack cannot be attacked alone. No matter which unit is specified in the attack order, the entire stack containing the unit will respond.

Only the leader of a stack may initiate combat. If another unit in a stack wants to issue attack, it must first drop out of the stack.

Combat involves nobles, who possess strong heroic fighting abilities, and fighters who may accompany them. Fighters include soldiers, pikemen, lt. foot, knights, paladins, crossbowmen, archers, etc.

5.1 Resolution of battle

Each kind of fighter has an attack and defense rating:

                 attack    defense    missile
                +----------------------------
peasant         |   1         1
postulant       |   1         1
worker          |   1         1
sailor          |   1         1
skirmisher      |   3         3
soldier         |   5         5
ninja           |   5         5
pirate          |  10 (30)   10 (30)
pikeman         |   5        20 (*)
fanatic         |  10         5
lt. foot        |  30        30
hvy. foot       |  60        60
angel           |  25       100
cavalier        |  75 (25)   75 (25)
knight          |  90 (40)   90 (40)
paladin         | 180 (130)  180 (130)
lt. xbowman     |   1         1         15
hvy. xbowman    |   1         1         75
archer          |   1         1         35
horse archer    |   1        20         40
elite archer    |   1         1         45
(*) Pikemen double their defense against mounted troops (knights,
paladins, cavaliers, and centaurs).

• Numbers in parenthesis are for shipboard combat.

• Nobles have strong innate heroic abilities, and so even an unarmed, untrained noble is a formidable opponent on the battlefield.

• A noble is rated (attack=80, defense=80, missile=0).

• A pirate fights (30,30,0) on a ship, but only (10,10,0) on land.

• Knights,paladins and cavaliers receive a -50 penalty to both their attack and defense ratings when fighting on a ship or in a swamp province. The algorithm by which combat is resolved is as follows:

1. Combat is fought in rounds. Each round has three phases:

1. Special attacks phase. Special attacks include such things as dragon fire and battle spells. Unlike missile and melee attacks, some special attacks can hit any unit on the other side, including those in the back rows of the combat and nobles who would normally be protected by their men.

2. Missile attacks phase. All troops with missile weapons who are not in the front row of the battle attack. Missile attacks strike the front row of the opposing force.

3. Melee attacks phase. The front row of troops attacks the opposing front row.

After each phase, any men killed or wounded are removed from the battle (i.e., combat inside each phase is simultaneous).

2. At the beginning of a battle, the defenders get a free round of missile attacks.

3. Targets are chosen at random from amongst the applicable enemy troops. For melee and missile attacks, the target may be any of the men in the enemy’s front units. A ship or building housing the enemy also counts as a potential target. In other words, if there are five men in a tower, each man has 1/6 chance of being selected as a target, and the tower also has 1/6 chance of being selected. Each attacker selects his target independent of the other attackers, so it is possible that more than one attacker may select the same target. All attacks in a phase are considered simultaneous, so even if the first attacker kills the target, the remaining attacks against that target will not be redirected. For some special attacks, the target may be any man in the opposing force (not just those in the front row).

4. In melee combat, the chance that the attacker will score a hit against the target is:

A = attacker’s attack rating B = target’s defense rating

A / (A + B) For example:

A(attack=90) vs B(defense=45) A has a 2/3 chance of killing B A(attack=90) vs B(defense=90) A has a 1/2 chance of killing B

5. If the attacker scores a hit against a noble, the noble will receive a random wound of 1-100 health points. Note that there is a 1% chance that a perfectly healthy noble will be instantly killed, and a greater chance that a previously wounded noble will die (see ⟨undefined⟩ [Health], page ⟨undefined⟩). Wounded nobles do not continue fighting, even if their wounds are minor. If a hit is scored against a fighter (soldiers, pikemen, archers, etc.), the fighter is killed. A man successfully attacking a building or a ship will cause one point of damage to the structure. A catapult or battering ram attacking a building will cause 5-10 points of damage.

The chance to hit in missile combat is similar, except the attacker’s missile rating is used instead of his attack rating. Note that troops can only use missile attacks when in the back rows of the combat.

6. Combat rounds continue until a side breaks. A side breaks when its total offensive plus defensive value falls by 50% (but see see ⟨undefined⟩ [], page ⟨undefined⟩). For instance, a noble plus two pikemen has an offensive value of (80+80) + (5+20) + (5+20) = 210 with a break point of 105. A noble plus two knights has an offensive value of (80+80) + (45+45) + (45+45) = 340 with a break point of 170. Therefore, a noble with two pikemen will continue to fight if the pikemen are killed. However, a noble with two knights will be declared the loser if both knights are killed. The offensive value used is either the attack or the effective missile rating, whichever is higher.

7. All attacks in the missile and melee phases are "simultaneous." Although the battle printout will show one side attacking first, all units get an opportunity to attack before any units are removed from the battle, so the actual order in which attacks take place is irrelevant. The exception to this is during the special attacks phase. Some special attacks (particularly certain magic spells) have secondary effects, such as draining magicians of their aura. In these cases, order of attack does matter, so the order of attack in the special attacks phase is determined randomly.

8. At the end of the battle, the following things happen:

1. The losing stack stops all movement.

2. The winning stack takes prisoners (see ⟨undefined⟩ [Prisoners], page ⟨undefined⟩).

3. The nobles in the losing stack who have not been captured flee. They leave the stack, and if they are in a sublocation, move out to the province.

5.2 Front and rear

Units may issue the behind command to declare whether they will be in the front or the rear in battle. Rear units do not become targets for the enemy until all of the units in the rows in front of them have been killed. Only missile fighters, such as archers, may attack from the rear.

Note that rear units do not advance to fill in the front row until a successful attack by the other side. So if a force starts off with all its units in the rear (e.g., BEHIND 9) they will remain in the rear (possibly making missile attacks) until a hit from the other side forces some of them into the front row.

Regardless of whether they are in the front row or behind, nobles will not receive a hit until all the troops they personally command have been killed. The leader of each stack (the top-most unit in the stack) will be the last unit to receive hits, regardless of its behind status. That is, the leader is effectively "behind" all the other nobles.

The exception to all these rules are some special attacks, which hit nobles and men alike regardless of where they stand.

5.3 Limits to Missile Troops

Missile troops may only use missile attacks from the back rows of the battle. Missile troops in the front row will engage in melee combat using their attack rating.

Each missile trooper who makes an attack during the missile phase must be protected by a front row melee fighter. Therefore, the number of missile attacks taken by a side during the missile phase will never be more than the number of fighters that side has in its front row.

For example, if a stack has 25 lt. foot in the front row and 35 archers behind them, only 25 archers will be able to attack during the missile phase. The remaining 10 archers hold themselves in reserve; they do not advance to the front row.

If the location of the battle is windy, all missile ratings are halved. If it is raining, the missile rating for archers and elite archers is reduced to zero. The location of a battle is based on the defending stack. This may be different from the location of the attacking stack (e.g., when attacking from one province into another).

5.4 Limits to Melee Troops

Since there is a practical limit to how many combatants can attack a single defender, the number of melee attacks for a side is limited to a maximum of four times the number of defenders. For example, if a force of 120 men attacks a force of 20 men, the larger force will get only 80 attacks in each melee phase.

5.5 Allies

When a unit is attacked, it may call in allies to help defend itself.

In particular, when a unit is attacked, all other units in the same location who have placed that unit on their defend list will join on the defending unit’s side in the battle. In addition, if the unit being attacked is in a sublocation (such as a castle), allies in the outer location will also join in the defense.

However, since allied units are usually responding to the news of the attack from some distance away, they do not join the battle until the beginning of the second round. If the defender is defeated in the first round the battle will end immediately, before any allies have a chance to come to his aid.

For more on the defend list, see ⟨undefined⟩ [Declared attitudes], page ⟨undefined⟩.

5.6 Who Participates in Combat?

To be used effectively in battle, the Olympia fighting units (such as soldiers, archers, etc.) must be commanded. Any noble can possess a large number of fighting troops; not every noble can use them in battle. Similarly, any noble can be given fighting animals (such as dragons, minotaurs, etc.); whether they can control them under battle circumstances is a different question.

Young nobles receive basic training in small group combat. This permits them to control up to 10 men when attacking another unit. To control more men in a battle situation, a noble must use Control Men in Battle [1107], a subskill of Combat [1100]. Each use of this skill takes three days and increases the number of men a noble can control in battle by one.

If a noble owns more men than he can control, he will use the men with the highest total attack + missile + defense values, or in the case of a tie, the men with the higher attack + missile value. Any men he cannot command will not be involved in any way in combat. In particular, a noble does not "replenish" lost men from the men he cannot command.

If a noble owns fewer men than he can control, he will not “borrow” additional men from a stack mate.

Controlling men on defense is somewhat easier, since they have a natural motivation to save their lives. For this reason, a noble can always control 20 more men on defense than he can on the attack. An untrained noble can defend himself with 30 men.

Garrison commanders have the training to control 20 men on the attack and 40 men in defense.

Non-player units (such as elves) generally have no limit on the number of men (or beasts) they can control in battle.

Controlling fighting animals in battle is considerably more difficult, and untrained nobles cannot use any animals in battle — whether on attack or defense. To control animals in a battle situation, a noble must know Use Beasts in Battle [1195], a subskill of Combat [1100]. Each use of this skill takes seven days and increases the number of animals a noble can control in battle by one.

There is no bonus for the number of animals that can be controlled on defense.

5.7 Declared attitudes

Commands dealing with attitudes:

  • hostile
  • neutral
  • defend
  • default

Attitudes can be declared by or for either specific units, nations, or an entire faction. For instance, player [613] could declare an attitude for a nation (e.g., "Mandor"), a player [555], or a specific attitude for individual units within player 555’s faction.

Declaring a permission for a player works so long as the player’s units are not concealing their faction identity with Conceal faction [1205], a subskill of Stealth [1200].

A unit may have one of four combat attitudes to another unit:

hostile Attack on sight.

defend Defend other unit if attacked.

neutral Do nothing if other unit is attacked.

default Neutral to units in other factions; Defend units in the same faction unless either one is concealing its lord.

Every noble, and player faction entity, keeps three lists of units or other factions which have attitudes declared towards them. A unit is either on the hostile, defend, or neutral list. If a unit does not appear on any of the three lists, it has attitude default.

Example:

    nation          Imperial Empire         Mandor
    player          778                     816
                      hostile 816             hostile Imperial
    units           4205                    6499
                    4600, concealing nation 6530, concealing lord
                                            6599
                                              hostile monster

Player 778 has declared player 816 hostile. One of 816’s nobles is concealing its lord.

If 4205 or 4600 run into unit 6499, they will attack it on sight. However, since 6530 is hiding its affiliation with 816, it will not be attacked on sight.

If 6499 is attacked and both 6530 and 6599 are present, 6599 will aid in the defense, but 6530 will not, because that might give away its affiliation.

If player 816 wanted 6530 to defend the faction’s units anyway, either 816 or 6530 should issue the order ‘defend 816’. This would override the default attitude of units in the faction to one another.

Attitude toward units is considered before attitude toward the unit’s faction. Thus, one may declare a faction hostile, but exclude certain units within the faction by specifically declaring them neutral.

Player 816 has declared himself hostile towards the Imperial Empire nation. If unit 6499 encountered 4205, it will attack on sight. If it encounters 4600 it will not attack, since that unit is concealing its nation.

Unit 6599 has declared himself hostile towards monsters. A monster is a stack of monsters not controlled by another player. This does not include any noble stacks, even if those stacks are not controlled by a player (e.g., a noble who has deserted his faction). Essentially, "hostile monster" will cause a unit to attack all monster stacks.

A unit must be the top-most noble in its stack to aid in defense. If a unit joins a combat because of defend, it will bring its entire stack along, even if the other members of the stack have not declared a defend attitude.

Defenders only help when units are attacked, not when they initiate attacks. For example, if A has declared ‘defend B’, and B attacks C, A will not help B, even if B loses the battle.

Nobles declared defend to units which are guarding a province against pillaging will aid the guards if they are attacked, either explicitly with attack, or implicitly via ‘pillage 1’.

Units which joined a combat because of a defend declaration are shown with the qualification ‘ally’ in the combat report.

5.8 Combat Bonuses

Certain skills, terrains, items and fortifications can affect the course of combat. In general, combat bonuses are computed by adding in any fixed bonuses, and then multiplying by any percentage bonuses. For example, a noble with an attack factor of 80, a magic sword with an attack bonus of +10, and Foresee Defense prayer of +10% would have an effective of attack of (80+10)*1.1 or 99.

A unit can benefit from one fortification bonus, one terrain bonus, one item bonus, and any number of skill or other bonuses. For example, a unit in a castle in the mountains receives both the castle defensive bonus and the mountains defensive bonus.

5.8.1 Fortifications

Structures which may aid fighters in battle are rated for their defensive bonus.

    Castle Imperius [gx56], castle, defense 25

During battle, the fortification rating is added to the defense number for the men who fit inside the structure.

    structure               men protected           default bonus
    ---------               -------------           -------------
    Castle                  first 500     40
    Tower                   first 100     20
    Other structures        first 50      0

Attacking fighters may randomly select the structure instead of an enemy fighter. The attack is resolved in the same way as for two fighters, using the attacker’s attack rating and the structure’s defense rating. If the attacker is successful the structure’s defense rating will be lowered by one point.

Once the defense rating reaches zero, further hits will cause the building to become damaged. A fully damaged building (100% damage) will collapse, ejecting its occupants.

Siege engines always select the structure as a target.

    engine          attack  defense   missile  weight
    ------          ------  -------   -------  ------
    catapult          25      200       25      3000
    battering ram     30      250               3000
    siege tower        0      100               3000

These siege engines do 5-10 points of damage to the structure per hit. Note that catapults are missile weapons and may attack from the back rows. Catapults are not affected by moats.

Note that siege towers have no attack capability. Instead, siege towers permit attackers to easily scale defensive fortifications. Unlike catapults and battering rams, siege towers do not attack the structure. Instead, they reduce the defensive bonus of the structure. Each siege tower in the front row reduces the opposing defensive bonus by the following amount:

    engine          castle  tower  other
    ------          ------  -----  -----
    siege tower       1       5     10

For example, an army attacking a castle with 10 siege towers would reduce the defensive bonus of the castle from +25 to +15.

Although siege towers have no attack capability, they can still be chosen in the combat algorithm as an attack; in which case the attack fails.

Siege engines are not used in combat at sea. However, siege engines can be used against docked ships.

5.8.2 Terrain Effects

Some types of terrain favor the defender because they provide hiding places and obstacles for the attackers to overcome. Some types of terrain favor the attacker, because they slow the ability of the defender to escape.

The following chart shows the affect terrain has on the Defense Factor of the defending unit in combat:

    Terrain         Modifier
    -------         --------
    City              +25%
    Forest            +50%
    Mountains        +100%
    Swamp             -25%

So a noble attacking a pikeman in the mountains will find that the pikeman’s defense factor is doubled (from 20 to 40). Note that the bonus applies only to the stack being attacked, and not to the attacker. Thus the noble’s defense factor remains at 80.

A stack benefits only from its immediate terrain. A stack in a mountain city, for example, receives only the city bonus.

5.8.3 Item bonuses

Nobles may possess items which grant attack, defense or missile bonuses in combat. Only one item may be wielded for each category. If the noble possesses multiple items with bonuses in the same area, the item with the largest bonus will be chosen. See ⟨undefined⟩ [Appendix - Magical Artifacts], page ⟨undefined⟩ for more details.

Nobles automatically use items with combat bonuses in battle. No special orders are needed to wield them.

5.8.4 Skill Bonuses

The Combat Tactics [1131] skill, a subskill of Combat [1100], gives a noble and the men he directly commands a bonus to their attack factor.

The Defense Tactics [1132] skill, a subskill of Combat [1100], gives a noble and the men he directly commands a bonus to their defense factor.

5.8.5 Other Bonuses

Pikemen double their defensive value when defending against mounted troops (knights, paladins, cavaliers and centaurs).

Rulers (see ⟨undefined⟩ [Castles], page ⟨undefined⟩) can control additional men in combat due to their natural fame.

5.9 Combat Spells

During the special attacks phase, magicians on both sides of the battle may use combat spells. To use a combat spell, a noble must be in the rear ranks of the battle (i.e., "BEHIND"). Nobles in the front ranks are too busy defending themselves to cast spells.

5.9.1 Selection of Spells

During each special attacks phase, each noble will attempt to cast all the combat spells he knows, in the order in which he learned the spells. A noble who knows ten combat spells and has sufficient aura will cast all ten spells during the special attacks phase.

5.9.2 Round Two Rule

Combat spells cannot be cast until the second round of combat. There are two reasons for this rule.

First, this rule permits non-magical forces to capture a magician by overwhelming him before he can bring his potent magics to bear.

Second, this rules prevents attackers from draining a magician of his aura by sending a series of easily-defeated "nuisance" attacks against him.

5.9.3 Effects of Combat Spells

Many combat spells simply kill opposing units. This acts as a normal attack. Killed or wounded units remain on the field of battle (and may launch special attacks of their own) until the end of the special attacks phase.

Other combat spells have different effects, such as draining other magicians of their aura. These effects are immediate, and may affect a special attack later in the special attacks phase. For this reason, the order of attack (i.e., which side attacks first) during the special attacks phase is determined randomly.

5.10 Panic in Battle

When peasants, workers and skirmishers are involved in battle, they often panic and flee the battlefield when they see their mates being killed.

Panic occurs after every combat phase in which a unit takes losses. The chance for each peasant and worker to flee the field of battle is 50%; the chance for a skirmisher to flee is 33%. Panic occurs only when a similar unit under the same noble is lost.

If a noble is stacked with 10 peasants and 10 workers, and enemy missile fire kills 3 peasants, then at the conclusion of the missile phase, that noble’s peasants will be subject to panic. The workers (who were unscathed) are not affected, nor are any peasants stacked with other nobles.

5.11 Prisoners and Looting

The winner in combat will attempt to take prisoners. The chance that a given defeated unit will be taken prisoner is proportional to the sizes of the remaining forces.

(Taking prisoners in battle and claiming loot requires many soldiers to run after the fleeing enemy. Thus, the chance of success is based on numerical advantage rather than combat skill.)

    1:1     25% 2:1     50% 3:1     75%

If the winning side outnumbers the defeated force by 2:1, there is a 50% chance that a given defeated unit will be captured. Defeated units which are not captured retreat from battle. If they are occupying a building or are located in a city, they may flee into the outlying province.

The victor always has at least a 25% chance of taking a unit prisoner, but no better than a 75% chance.

The victor will claim the defender’s position in the location list if it is better, or will move into the defender’s structure, ejecting the losing force. (The attacker may specify a flag to the attack order to inhibit this behavior.)

Prisoners are stripped of their belongings by the victor. The stack leader of the winning force receives all of the loot from the battle.

Some of the men in the employ of the prisoner may also decide to join the victor’s forces. The chance of a man deserting an imprisoned noble to join his capturer is:

    5/(attack + defense)

That is, a common soldier [12], with an attack of 5 and a defense of 5, has a 5/10 or 50% chance to switch sides when captured. A paladin, on the other hand, has only a 5/(180+180), or 1.3% chance of joining with the victors.

Beasts (such as wolves or dragons) can never be captured in battle. Beasts are trained to obey a particular noble, and can never be forced to serve another.

Since prisoners are unable to report where they are and what they are seeing, they do not contribute to the turn report of their faction. The player’s turn report will show that that a unit is being held prisoner, but little else.

Prisoners will not execute any orders while they are in captivity. Queued orders will remain pending, but none will be processed.

Prisoners when spotted appear as stacked units, marked with the ‘prisoner’ string:

    Seen here:
       Kosar the Indefectible [2022], with six peasants, one archer,
       two soldiers, accompanied by:
          Alion Krysaka [2785], prisoner

Unstacking a prisoner sets them free. Kosar could free Alion by ordering

‘unstack 2785’.

Prisoners may be transferred between units with the give command. Kosar could transfer Alion to Osswid [xt8] by ordering:

    give 501 2785

5.12 Prisoner escapes

Prisoners are always on the lookout for ways to escape. Units holding prisoners can reduce their chances by remaining inside a structure, not transferring prisoners with give, and not traveling with prisoners.

Each week (four times each game turn), a prisoner being held by a unit which is outside of a building has a 2% chance of escaping. A prisoner being held by a unit which is inside a structure, such as a castle, tower, inn or ship, has a 1% chance of escaping.

Each time a prisoner is transferred with the give order, there is a 2% chance of a escape. Also, each time a unit holding a prisoner engages in travel which takes longer than one day, there is a 2% chance that the prisoner will be able to get free. Thus, short movement, such as entering or exiting a building, will not give the prisoners additional opportunities for escape, but traveling between provinces with prisoners will.

Prisoners inside building or sublocations will flee out into the surrounding location upon gaining their freedom. Escaped prisoners on ships will leap over the side and swim to a nearby shore.

5.13 Combat and Movement

A moving stack may be attacked from the location in which it began its move until the end of the day the move completes. (Note: a stack may receive a message that it has "arrived" in a location and still be attacked later that same day by a stack in the location where it began travel.) It may not be attacked from its destination until the move completes.

If an attack against a moving stack succeeds, the movement fails and the stack remains in its original location. If the attack fails, the movement completes.

5.14 Combat Examples

    Example 1: Missile Attacks and the Four to One Rule
    Party One
      2 cyclops (25,75)
    Party Two
      1 noble at behind 0 (80,80)
        10 pikemen (5,20)
      1 noble at behind 1 (80,80)
        5 lt. xbowmen (1,1,15)

The cyclopses attack the nobles. Before the first round, the defenders (the nobles) get a free phase of missile attacks. All five xbowmen get attacks, since there are 10 pikemen (and one noble) in the front row.

The first round opens with another phase of missile attacks from the xbowmen. In the melee phase, the two cyclops (assuming they haven’t been killed by the missile fire) attack, and the pikemen plus the noble get eight attacks. Only eight attacks are allowed because in melee one side can outnumber the other by at most four to one. Note that the four to one rule applies only during the melee phase. Whether the noble gets to attack is random; attackers and defenders are chosen randomly from those available. (However, each attacker can attack only once per round.)

    Example 2: "Behind"
    Party One
      1 noble at behind 0 (80,80)
        100 pikemen (5,20)
    Party Two
      1 noble at behind 0 (80,80)
      1 noble at behind 1 (80,80)
        5 lt. xbowmen (1,1,15)

During the initial missile phase and the missile phase of the first round, Party Two will take one missile attack, since the front row of Party Two only has one person behind which a missile troop can shelter. The other four xbowmen do nothing.

In the melee phase, Party One makes four attacks against the noble in the front row of Party Two. Even if the noble is killed, no attacks fall through to the second row of Party Two. Removal of dead units (and hence removal of "rows") occurs only between phases.

A party can lose at most three rows during a round of combat: one row each after the special attacks phase, the missile phase, and the melee phase.

6 Locations

6.1 Buildings

The following structures may be built by nobles:

    kind        effort    material   skill         where
    ----        ------    --------   -----         -----
    inn            300     75 wood    1400         province or city
    mine shaft     500        none    1703         mountain or mine shaft
    temple       1,000    50 stone    1400         anywhere
    tower        2,000   100 stone    1400         anywhere
    castle      10,000   500 stone    1400         province or city
    city     50,000  2000 stone    1490        province
                          500 wood

The term ‘anywhere’ means a province, city, or other sublocation. Buildings may not be built inside other buildings, with the exception of towers. Up to six towers may be built inside a castle.

Effort is in worker-days.

The builder must know the required skill, have at least three workers, and possess at least one-fifth of the necessary construction materials in his inventory.

To start building, unstack from beneath other nobles and issue one of the following build orders:

  • build inn "name of inn"
  • build shaft "name of mine"
  • build temple "name of temple"
  • build tower "name of tower"
  • build castle "name of castle"

For example, a noble who wanted to build a tower would need to know Construction [1400], have at least three workers [11], and possess at least 20 stone [78].

The construction materials will immediately be deducted from the builder’s inventory and put to use. The builder and his workers will be placed inside the new structure.

The remaining construction materials are paid as work on the structure progresses. A second fifth is required as the building becomes 20% complete, the third fifth at 40%, etc. Construction halts if the builder runs out of materials.

The building is completed when the required number of worker-days is invested in construction.

To resume construction of a partially completed building, first enter the structure, then issue the appropriate build order (such as build inn, build shaft, etc.)

6.1.1 The first tower you build is free

Since towers are essential to learning, the first tower a player builds requires no workers, no stone, and takes only one day. The player should choose an appropriate site for the tower, and issue a regular build tower order. Knowledge of Construction is not required to build the first tower.

6.1.2 Guild Halls

Guild halls are specially dedicated towers where the elite members of a skill group (e.g., [1000] Combat) come to discuss and learn the most complicated and expert skills. Guild halls are created by dedicating a tower using the guild command. Only guild members may enter a guild hall (but non-guild members can enter as guests when accompanying a guild member, i.e., stacked underneath a guild member).

A noble cannot dedicate a guild hall until he has met all the requirements to join the guild, i.e., he must have learned all the non-guild skills. He must also be alone in the tower when he dedicates it.

Guilds can only be created in a city, and only one guild of a particular type can be built in a city.

A noble must be in an appropriate guild hall to learn a guild skill (see ⟨undefined⟩ [Guild Skills], page ⟨undefined⟩). Some guilds also have other features. For example, the Trading guild hall has a market where special trade goods are bought and sold.

In some games, guild halls teach the base skill. To use a guild hall as a source of teaching, a noble will have to be escorted into the guild hall as part of a guild member’s stack. Note, however, that he still cannot learn guild skills (even though he is inside a guild hall) unless he becomes a guild member.

6.1.3 Inns

Inns must be built in provinces or cities. A good site for an inn is just outside a city, as the inn will benefit from the patronage of many travelers.

The restive environment inns provide help wounded nobles to heal. Nobles located in an inn have an 10% increased chance of fighting off infection each week.

Inns generate income each month from the visitors who stop for a meal and a pint of ale, or spend in the night in one of the inn’s rooms. (Nobles who enter are not directly charged. Other patrons are anonymous; the only indication of their presence is the income the inn generates.)

Inns generate between 50 and 75 gold per month. Looting and pillaging scares away customers, and so lowers the inn’s income. If more than one inn is built in a province, the profits are split between them.

Income is paid to the owner of the inn at the end of each month.

6.1.4 Castles

Castles are the foundation of land ownership. A castle is the symbol of strength that peasants recognize in a ruler. A noble backed by a castle may garrison land and collect taxes.

A province may contain only one castle. The castle must be built in the outer province or in a city, if the province has one. Castles may not be built inside other sublocations.

A castle alone is not sufficient to rule a province. A garrison must be stationed outside the castle in the province to protect it.

The castle owner gains status from the number of provinces under control. As a result, he can control additional men in combat. This benefit is received by all nobles in the castle owner’s faction:

    provinces       rank            control
    ---------       ----            -------
       1-3          lord              +4
       4-7          knight            +5
       8-11         baron             +6
      12-15         count             +7
      16-19         earl              +8
      20-22         marquess          +9
       23+          duke             +10
      region        king             +20 (The King only)

If a faction controls multiple castles, the nobles benefit only from the highest rank. If a noble controls a region of at least 25 provinces, then he receives the rank "King" and may control 20 additional soldiers. This King bonus applies only to the noble controlling the castle, and not to the other nobles in his faction.

Control of a province allows one to change its name or the name of any of its sublocations (⟨undefined⟩ [name], page ⟨undefined⟩), disband the garrison (⟨undefined⟩ [ungarrison], page ⟨undefined⟩), take items from the garrison, issue decrees to watch for certain units (⟨undefined⟩ [decree], page ⟨undefined⟩), have the garrison attack specified units on sight (⟨undefined⟩ [decree], page ⟨undefined⟩), control the borders (⟨undefined⟩ [decree], page ⟨undefined⟩), set taxes (⟨undefined⟩ [tax], page ⟨undefined⟩), and establish entrance fees (see ⟨undefined⟩ [Control of Locations], page ⟨undefined⟩).

Visitors to a province are informed of the castle to which the garrison is bound, and the owner of the castle:

    Province controlled by Amber Keep [cy09], castle, in Forest [cj12]
    Ruled by Erekosse [5210], baron
    ...
    Seen here:
      Garrison [780], garrison, on guard, with ten soldiers

The ruler of a province may grant certain privileges to other nobles by admitting them to his stack (⟨undefined⟩ [admit], page ⟨undefined⟩). Every noble on the ruler’s admit list (either directly, or because his faction or nation is on the admit list) is given the following privileges: he may take and give items to garrisons, he may stack with garrisons, he is not charged taxes or entrance fees, and he may name the province (⟨undefined⟩ [name], page ⟨undefined⟩). He is not permitted to turn borders on and off, set taxes or entrance fees, ungarrison, or issue decrees.

6.1.5 Mines and Mining

A mine is a deep shaft or tunnel which allows workers to extract valuable resources from the earth, such as iron and gold. Mine shafts must be started in mountain provinces; there may be at most one mine shaft in a location. A mine shaft is built with the "BUILD SHAFT" command. To build a mine shaft, a noble must know [1703] Build Mine Shaft. Building a mine requires at least 3 workers and 500 worker-days of effort.

The initial mine shaft built in a mountain goes down 100 feet. Upon completion, it will appear as a new location "down" from the province, e.g.,

    Routes leaving Landsdrop Mountains:
       North, mountain, to Trothguard [bo13], 10 days
       East, mountain, to Doriendor Corishev [bp14], 10 days
       South, to Ocean [bq13], Balatean Sea, impassable
       West, mountain, to Landsdrop Mountains [bp12], 10 days
       Down, mine shaft, to New mine shaft [g55], 0 days, depth 100 feet

To mine deeper, a noble must descend into this mine shaft and build another mine shaft there, which will descend an additional 100 feet. This may be repeated to tunnel to any desired depth.

6.1.5.1 Mining

Mines often contain precious resources, such as iron and gold. To obtain these resources, a noble must have a work crew and use the appropriate mining skill, e.g., [1701] Mine Iron.

As the depth of the shaft increases, the mix of resources obtainable changes. Iron is usually found nearest the surface, at depths of up to 1000 feet. Gold is found from 600 feet downward. Mithril is found below 1000 feet, and gate crystals are found only at the deepest levels.

A noble’s chance of finding a resource depends primarily upon his experience. Additional workers help slightly, but only the first 100 workers aid in finding ore. Of course, no amount of experience or additional workers will help find ore if there is none to be found, or if the level being mined is "played out".

When ore is discovered, the amount recovered depends upon the number of workers available to process the ore. Again, only the first 100 workers aid in processing the ore; more cannot fit into the close confines of the mine shaft.

At times nobles may hit rich lodes of ore that will increase (sometimes greatly) the amount of ore harvested, but such occurrences are fairly rare.

6.1.5.2 The Dangers of Mining

The deeper a mine becomes, the more frequently cave-ins or other accidents will occur. The chance of a cave-in occurring in an unimproved mine shaft is (depth/300)% per month. That is, a depth 300 mine shaft has a 1% chance of collapsing each month. A depth 1800 mine shaft has a 6% chance of collapsing each month.

When a mine shaft collapses, it destroys itself and all lower mine shafts. Any nobles in the collapsing shafts will evacuate, but will lose 1-20% of all belongings (including men and beasts) for each level of mine shaft they must evacuate. A noble in the depth 1500 mine shaft of a mine when a cave-in occurred on the depth 800 level would need to evacuate through 7 levels of collapsing mine shaft, losing 7-100% of his goods. Nobles themselves take 1-20 points of damage for each level they evacuate, so that in this case the noble would also suffer 7-140 points of damage.

The chance of a mine shaft collapsing can be reduced by adding wooden or iron shoring to the shaft. Wooden shoring reduces the chance of a mine shaft collapse by 2.3%. Iron shoring reduces the chance of a mine shaft collapse by 4.6%. Regardless of shoring, there is always a 0.2% chance per month that a shaft will collapse.

In addition to the danger of cave-ins, it has been reported that at great depths miners occasionally encounter various terrible monsters, most especially [310] balrogs.

6.2 Control of Locations

BOUNDARY, n: In political geography, an imaginary line between two nations, separating the imaginary rights of one from the imaginary rights of another.

The Devil’s Dictionary, Ambrose Bierce

Every location in Olympia can have an owner. The owner of a location controls who is allowed to enter and how much they must pay to enter. In the case of provinces and cities, the owner can also set various tax rates. This chapter describes how to control a location and what control permits you to do.

6.2.1 Control of Sublocations

A sublocation appears inside a province. Sublocations are shown in the turn report this way:

    Inner locations:
       Yew grove [em28], 1 day

Sublocations include inns, temples, guilds, and ships as well as bogs, caves, graveyards and other natural features. (The only places in Olympia that aren’t sublocations are provinces, cities and mines.)

The owner of a sublocation is the first noble in the sublocation. The owner of a sublocation controls who is permitted to enter, and what they will be charged (if anything) to enter.

6.2.1.1 Permission to Enter

The owner of a sublocation (see ⟨undefined⟩ [Control of Sublocations], page ⟨undefined⟩) can set the "border" of the sublocation to be open or closed.

If the border is open, everyone except those people on the owner’s hostile list (see ⟨undefined⟩ [hostile], page ⟨undefined⟩) are permitted to enter.

If the border is closed, no one except those nobles on the owner’s admit (see ⟨undefined⟩ [admit], page ⟨undefined⟩) list are permitted to enter – i.e., only those nobles who are permitted to stack with the owner are permitted to enter the sublocation. (Someone on both the hostile and admit lists will be turned away.)

Note that closing the border only stops travel over land. You can always fly or sail into a province or sublocation, even if the border is closed.

Use the ⟨undefined⟩ [border], page ⟨undefined⟩ command to open or close a location’s border.

Closed sublocations are noted in the turn report with "closed" in the sublocation description:

    Inner locations:
       Lair [ao9s], lair, 1 day
       Chateau Olympique [s85], castle, closed, defense 40, "with two bloody
       spots on the walls"
          Greater Tower of the Art [q45], tower, defense 20

Units in an interior location can always exit into the outer location, even if the outer location’s border is closed. In the above example, a unit in the Greater Tower of the Art [q45] would be able to exit to Chateau Olympique [s85], even though the castle’s borders are closed.

Whether a border is open or closed is a property of the location, not the owner. That is, if a noble closes a border and then leaves, causing a different noble to become the owner of the location, the border will still be closed for the new owner. However, a location with no owner always has an open border.

If a noble is refused admission to a location, he may still enter by attacking the location. If the attack is successful, the noble will enter the location and seize control (by becoming the owner of the location). (However, this does not permit nobles to enter temples or guilds they would not normally be able to enter.)

Control of a province is very similar to control of a sublocation, except that the owner of a province is the ruling noble (see ⟨undefined⟩ [Control of Provinces], page ⟨undefined⟩).

6.2.1.2 Entrance Fees

The owner of a location may set an "entrance fee" for the building. This fee is the cost in gold each noble will have to pay to enter the building. The entrance fee for a location appears in its description, e.g.,

    Inner locations:
      Lair [ao9s], lair, 1 day
      Chateau Olympique [s85], castle, entrance fee 100 gold per noble,
        closed, defense 40, "with two bloody spots on the walls"

The entrance fee for a location is set with the fee order (see See ⟨undefined⟩ [fee], page ⟨undefined⟩). Entrance fees can be charged per noble, per man, per weight, or any combination of those fees.

When a stack of nobles enters a location, the stack leader is responsible for paying the entrance fee for all the nobles in the stack, although he will borrow from other nobles in the stack to pay, if necessary. (But he will only borrow from nobles controlled by the same player – you cannot take money from another player this way.) A noble can control the maximum amount he will pay to enter a location with the maxpay command (see See ⟨undefined⟩ [maxpay], page ⟨undefined⟩).

The entrance fee is be paid to the owner of the location. A stack that cannot afford to enter a location will be refused permission to enter. A stack that attacks into a location does not have to pay the entrance fee.

6.2.2 Control of Provinces

Provinces are controlled by establishing a garrison in the province. Once a garrison has been established, the ruler of the province controls the borders to the province, the entrance fees, and any taxes in the province.

The ruler of the province is the first noble seated in the castle to which the garrison is pledged. If there is no noble in the castle, the province is unruled.

Control of a province allows one to change its name or the name of any of its sublocations (⟨undefined⟩ [name], page ⟨undefined⟩), disband the garrison (⟨undefined⟩ [ungarrison], page ⟨undefined⟩), take items from the garrison, issue decrees to watch for certain units (⟨undefined⟩ [decree], page ⟨undefined⟩), have the garrison attack specified units on sight (⟨undefined⟩ [decree], page ⟨undefined⟩), control the borders (⟨undefined⟩ [decree], page ⟨undefined⟩), set taxes (⟨undefined⟩ [tax], page ⟨undefined⟩), and establish entrance fees (see ⟨undefined⟩ [Control of Locations], page ⟨undefined⟩).

The ruler of a province may grant certain privileges to other nobles by admitting them to his stack (see ⟨undefined⟩ [admit], page ⟨undefined⟩). (Note that you do not admit nobles to the province; you admit them to the ruler’s stack.) Every noble on the ruler’s admit list (either directly, or because his faction or nation is on the admit list) is given the following privileges: he may take and give items to garrisons, he is not charged taxes or entrance fees, and he may name the province (⟨undefined⟩ [name], page ⟨undefined⟩). He is not permitted to turn borders on and off, set taxes or entrance fees, ungarrison, or issue decrees.

Nobles on the ruler’s admit list are considered rulers as well for the purposes of market taxes (⟨undefined⟩ [Sales Tax], page ⟨undefined⟩), etc.

6.2.2.1 Garrisons

Garrisons are groups of men who are stationed in provinces to protect them, control the borders, and collect taxes in the name of a castle. Garrisons must be created with the garrison order, and must be bound to a castle located in the same region.

For example, suppose that the region Lesser Atnos had 20 provinces. One of these provinces contains Amber Keep [cy09]. A garrison bound to Amber Keep could be stationed in each of the 20 provinces (including the province containing the castle itself).

Continuing the example, garrison units not in the Lesser Atnos region could not be bound to Amber Keep. The castle a garrison is bound to must be in the same region.

Garrisons can be bound to any castle in the region. If Lesser Atnos had two castles, some of the garrisons could be bound to one, and the rest to the second castle.

Ten soldiers are required to create a garrison. The garrison order must be issue at the outer level of a province; one can’t establish a garrison while inside a city, building or other sublocation.

Example:

    > garrison cy09
    Installed Garrison [780], garrison, on guard, with ten soldiers

Visitors to this province would see:

    Province controlled by Amber Keep [cy09], castle, in Forest [cj12]
    Ruled by Erekosse [5210], baron

The garrison pays the maintenance cost of its men from the gold in its inventory. Taxes collected from the province are placed into the garrison’s inventory, and so can be used to pay the garrison’s maintenance costs. For example, a garrison of ten soldiers would require 20 gold per month to support.

Since a garrison helps keep a province civilized and lawful, grateful peasants in the province contribute goods and services toward supporting the garrison. For every 20 peasants in a province, the cost of maintenance for the province garrison is reduced by 1 gold piece. The maximum reduction is 50 gold pieces. A garrison of ten soldiers in an empty province would require 20 gold per month in support; a garrison of ten soldiers in a province of 227 peasants would require only 9 gold per month in support.

Stacks can ⟨undefined⟩ [defend], page ⟨undefined⟩ a garrison just as any other stack.

6.2.2.2 Garrison reports

Garrisons do not provide full location reports to their owners. They do notice any resource depletion activity, such as timber cutting or mining, as well as any large or unusual parties which enter their province. This includes any stack of five units or more, any party of 20 or more men, and most monsters or wild beasts. Garrisons also report their inventory. Only the ruler of a location receives the garrison report.

    Garrison Report:
    garr where  men cost  tax rate castle ruler
    ---- -----  --- ----  --- ---- ------ ------
    ey7f  ba12   10   10 1000   10   d9l  n0z
    hj2l  ba13    6    6    0    0   d9l  n0z
    ey7f in Snowcap Mountains [ba12]
    -----------------------------------
        2,000  gold [1]                   0
           10  peasants [10]          1,000  cap 1,000 (1,1,0)
          195  wood [77]              5,850
    hj2l in Great Forest [ba13]
    -----------------------------------
            6  peasants [10]            600  cap 600 (1,1,0)

Garrisons do not monitor activity in hidden locations, even if the players who rule over the garrisons have discovered the hidden locations.

The ‘decree watch who’ order may be given by a ruler to instruct all garrisons to watch for a particular unit. This is useful for locating individuals who would otherwise go unnoticed by the garrisons.

6.2.2.3 Referring to garrisons

Since a province may only have one garrison, garrisons may be referred to without knowing their entity number. The keyword garrison will match the province’s garrison, if there is one.

Examples:

    give garrison 12 5
    attack garrison

6.2.2.4 Status

The number of provinces a noble controls determines his status or rank and allows all the nobles in that faction to control additional men in battle:

    provinces       rank            control
    ---------       ----            -------
       1-3          lord              +4
       4-7          knight            +5
       8-11         baron             +6
      12-15         count             +7
      16-19         earl              +8
      20-22         marquess          +9
       23+          duke             +10
      region        king             +20   (region must have at
                                            least 25 provinces)

In addition, if a noble has control over every province in a region, and the region contains at least 25 provinces, then the noble is given the rank of king.

6.2.2.5 Province Control

Border control for provinces is the same as border control for sublocations (see ⟨undefined⟩ [Permission to Enter], page ⟨undefined⟩), with some exceptions.

First, the owner of a province is the first noble in the castle to which the province’s garrison is pledged.

Second, the owner of a province also controls any city within the province. Third, in case of attack, the province and the city are defended by the

garrison, not by the first unit in the location.

Fourth, the garrison will not patrol the border (leaving it open) unless there is a noble seated in the controlling castle.

6.2.2.6 Province Entrance Fees

The owner of a province may set an entrance fee for the province or any city in the province (see ⟨undefined⟩ [Entrance Fees], page ⟨undefined⟩).

The entrance fees for the province and any city in the province are paid to the garrison.

The garrison will not collect entrance fees unless there is a noble seated in the controlling castle.

6.2.2.7 Taxes

A noble who controls a province may raise money by placing a general tax on the province and/or by placing a market tax on any city in the province.

6.2.2.8 Province Tax

A noble who rules a province (by virtue of being the owner of the castle which garrisoned the province) may tax the province. The tax on a province is a rate of 0 to 10%. Each month, gold equal to that percentage of the population is collected from the peasants.

For example, a province with 357 peasants taxed at 5% would collect 17 gold in taxes.

Taxes are collected out of the peasants’ collective wealth. Peasants in provinces without a city produce 1 gold per 20 peasants per month (5%). In provinces with cities, peasants produce 1 gold per 10 peasants per month (10%). If the peasants in a province do not have the gold to pay their taxes, they liquidate their property to get the gold, but this drives away one peasant for each gold needed.

For example, a province with 357 peasants taxed at 10% would collect 35 gold in taxes. If the peasants in the province had only 27 gold, raising the additional 12 gold would drive 12 peasants away, leaving the population at 345.

The tax rate is set with the tax order (see See ⟨undefined⟩ [tax], page ⟨undefined⟩).

A regular province with less than 100 peasants cannot be taxed. A province with a city cannot be taxed if it has less than 1000 peasants.

Money generated by taxing a province is placed into the province’s garrison. A garrison will collect the province tax only if there is a noble seated in the controlling castle.

6.2.2.9 Market Taxes

Markets are located in cities and in trading guilds (which are themselves located in cities). A noble who rules a market (by virtue of ruling the province in which the market is located) may tax the market. The tax on a market is a rate of 0 to 25%. For each sale that occurs in the market, that percentage of the sale is placed into the local garrison.

For example, if a market has a tax rate of 5% and a merchant sells 27 lumber for 13 gold apiece, then 5% of the total sale price (i.e., 5% of 351 gold, or 17 gold) will be placed into the local garrison and the seller will receive the remainder of the sale price.

The tax rate is set with the tax order. Note that new market tax rates do not go into effect until the end of the month.

7 Economics & Growth

7.1 Peasants

Ill fares the land, to hastening ills a prey, Where wealth accumulates, and men decay. Princes and lords may flourish or may fade,– A breath can make them, as a breath has made; But a bold peasantry, their country’s pride, When once destroy’d, can never be supplied.

The Traveller, Oliver Goldsmith

The basis of all government in Olympia is the peasant. The lowliest of citizens, forced to work long hours tilling barren land, conscripted into battles, driven willy-nilly across the countryside by wild magics, the peasant nonetheless is the basic unit upon which the Olympian society is built. The task of government is to nurture and grow the peasants upon which all else is built.

Each province contains some number of peasants. Provinces with less than 100 peasants are considered wilderness; provinces with 1000 or more peasants are usually cities. Left alone, the peasant population in a province will slowly grow over time. However, peasants are rarely left alone. Monsters prey upon them; nobles recruit them, pillage them, sell them opium – all which reduce their numbers.

In return for the protection of a local garrison of soldiers, peasants will permit a noble to tax them. But taxing too drives away the peasants, so the wise ruler must seek a balance between growing his peasant population and generating money by taxing them.

7.1.1 Population Growth

Each province contains a population of peasants. If the population exceeds 100 peasants it is shown in the province report:

    Barbary Coast [ap21], plain, in West Camaris, peasants: 127

The peasant population in each province grows at the end of each month. The rate of growth per month depends upon the type of province:

    Province       Growth rate Maximum population
    --------       ----------- ------------------
    city               3%      10,000
    plains             1%       1,000
    forest             1%       1,000
    mountain           1/2%       1,000
    swamp              1/2%       1,000
    other              0%                   500

If a province exceeds its maximum population, 10% of the excess peasants will die per turn.

To move peasants around, recruit them in one province and drop them in another.

The skill Encourage Growth [1390] can be used to increase the growth in a province by 2% (e.g., in plains, from 1% to 3%).

The month a province is pillaged, the peasant population does not grow.

7.1.2 Creating Wealth

Peasants create wealth. As they till the land, mine the soil, fish the shores, harvest the woods they increase their wealth, and that wealth flows outward to the nobles that nurture them. Or feed on them, as some would say.

In provinces without a city, peasants create 1 gold per 20 peasants per month. In provinces with cities, peasants create 1 gold per 10 peasants. This money remains with the peasants until they spend it or have it taken from them.

The first use of any money generated in a province is to pay taxes to the lawful ruler of the province. After that, the peasants may spend – or be parted from – their gold in many ways. They may donate money to a local temple, pay a magician to perform miracles, have their money stolen by thieves, lose it to pillagers, and etc.

Any gold unspent accumulates from month to month.

7.1.3 Taxing

The owner of a province may impose a tax on the province (see ⟨undefined⟩ [Province Tax], page ⟨undefined⟩). This tax is paid out of the peasants’ gold. If the tax is too high, the peasants will pay by liquidating their possessions, but this will drive away the peasants. A wise ruler must balance taxes and growth to create a sustainable economy.

7.1.4 Pillaging

Gangs of ten or more fighters may use the pillage command to seize loot from a province or a city. For every ten fighters pillaging a province, one peasant is killed each day and two gold seized. (Fractions are truncated.) Note that the peasants may have no gold to surrender to the pillagers if they have paid it all in taxes or spent it elsewhere.

To pillage, the fighters must have an attack factor of at least a skirmisher, and be controlled by the noble using the pillage command (i.e., the noble must be able to control them in battle). Beasts may be used to pillage if they are controlled by a noble. The noble himself does not count towards the number of pillagers, as they hold themselves aloft from such distasteful actions.

For example, a location with 215 peasants pillaged by a group of 27 men would kill 14 men (27/10 = 2, 2*7 = 14), and seize at most 28 gold (27/10 = 2, 2 * 2 * 7 = 28).

Pillagers must first defeat any units guarding the province, including the garrison, if one exists. There is also a 33% chance that a mob of 12-36 angry peasants will form to resist further pillaging attempts. Angry peasants have a combat rating of (attack-2, defense-1, missile-0). If a mob forms, there is a 33% chance that the mob will immediately attack the pillager in retaliation.

Only a stack leader may pillage.

A province with less than 100 peasants cannot be pillaged.

Pillaging a city is the same as pillaging the province the city is in.

Any month a province is pillaged, the peasant population does not grow.

7.1.5 Opium

Opium is produced in swamp regions, and consumed by markets in desert, plain, forest, and mountain provinces. All markets (except those in swamp regions) have some level of opium demand. However, this demand will not be visible in the market report at low levels.

Satisfying opium demand in a market will cause the next month’s demand to be higher. As peasants become addicted to opium, the increased demand will be shown in the location’s market report. If no opium is sold to a market, the demand will fall.

The money to buy opium is paid out of the peasants’ accumulated wealth. If there is insufficient wealth to purchase the needed opium, then one peasant will die for each additional gold needed, as addicts kill and rampage to find the gold to feed their habits.

7.2 Markets

The story is told / Of the power of gold And it’s lure on the unsuspecting. It glitters and shines / It badgers and whines And constantly needs protecting.

The Power of Gold, Dan Fogelberg/Tim Weisberg

Nobles issue buy and sell commands to indicate their desire to trade goods. Every trade must be between a buyer and a seller. Whenever a compatible buy and sell request are found, the trade will be executed.

Trades are only matched in cities, where merchants gather at the local bazaar to swap goods, and in Trading Guilds, where professional merchants meet to buy and sell trading goods. Trades are only made at the end of the day, so if you wish to trade you’ll have to stay in the city or the Trading Guild for at least one day.

(Note that nobles are always free to use the give command to exchange items, regardless of where they are. But buy and sell orders are only matched in cities.)

A buyer indicates what he wants to buy, how much of it he wants, and the most he is willing to pay. A seller indicates what item is being sold, how much of it is to be sold, and the minimum price he will accept.

Example: A noble who wants to buy five iron [79] at no more than 10 gold each issues:

    > buy 79 5 10
    Try to buy five iron [79] for 10 gold each.

A noble wants to sell five iron [79] at no less than 8 gold each:

    > sell 79 5 8

A trade occurs if there is a buyer willing to pay a seller’s minimum cost. In cases where there are multiple buyers and/or multiple sellers, an "auc-

tion" takes place. An auction works as follows:

The buyer willing to pay the most for a product buys as many as he can at a price just above that of the next highest bidder. If there is more than one seller, he buys first from the seller with the lowest asking price. When he is satisfied, the auction repeats with the next highest bidder. If there are no other bidders for an item, the buyer pays the seller’s asking price.

For example, suppose a market contains the following:

    Noble A selling 2 horses at 75 gold each.
    Noble B selling 2 horses at 80 gold each.
    Noble C buying 1 horse at 88 gold each.
    Noble D buying 3 horses at 100 gold each.
    Noble E selling 3 horses at 150 gold each.

Noble D would make the first purchase. He will buy horses at 89 gold each (just above the next highest’s bidder’s price of 88 gold each). He purchases first from A and then from B. Next, Noble C makes a purchase. There are no other bidders, so he buys a horse at the seller’s asking price. Noble E is not involved, since his asking price exceeds all the buyer’s prices.

    Noble D buys 2 horses at 89 gold each from Noble A.
    Noble D buys 1 horse at 89 gold from Noble B.
    Noble C buys 1 horse at 80 gold from Noble B.

Note that order in the market makes no difference. The first sale will go to the noble willing to pay the most.

Trades only happen at the end of the day. Nobles passing through a market without being there at the end of a day (e.g., moving in and out of sublocations such as towers) will not transact any trades.

If a buy or a sell order can’t be matched with pending trades from other units in the city, it will become a standing order and remain in effect until executed or canceled. To remove a pending trade, issue a trade with units of zero:

    > buy 79 0
    Cleared pending buy for iron [79].

Note that the buyer and seller can’t specify what noble they will deal with. They will trade with any unit that has a matching buy or sell order.

7.2.1 Market report

The location report for each city includes a market report listing pending trades.

    trade    who    price    qty   item
    -----    ---    -----    ---   ----
      buy   2019      100      1   peasant [10]
      buy   4274       74      3   iron [79]
     sell   3682       12     11   sailor [19]
     sell   2019       50      1   paladin [15]

A trade will not be listed unless it could be executed. For instance, a unit might issue an order to sell iron [79], even though the unit doesn’t possess any. (Perhaps the noble plans to get some iron later in the month, and wants the buy order to be in place when the iron arrives). This order will not be shown in the market report, because the seller doesn’t have any iron to sell.

Similarly, the buyer must have enough gold to buy at least one of the item. No buy order would be listed for a penniless unit that wanted to buy five iron at 10 gold each. If the unit later obtained 10 gold, enough to buy one of the five desired units of iron, the order would be listed in the market report for one iron, not five.

Some traders may work through middlemen to hide their identity. In such cases, the who field of the market report will not show their unit number, and their identity will not be revealed when the trade is executed.

Even sneakier traders may have their pending trades omitted from the market report entirely.

7.2.2 City purchasing

Some cities will issue buy or sell orders on their own behalf for certain goods. These may be produced by player units through the use of skills, or bought from other locations and transported to the buying cities.

All cities buy 1-10 wild horses a month at a base price of 15 gold.

Port cities purchase 100 fish per month at a base price of 2 gold.

Cities that are growing (have less than 9500 population) purchase wood and stone for construction. These purchases taper off as the city reaches its maximum size. Cities pay 13 gold for wood and 8 gold for stone. (Some cities may both purchase and sell wood.)

Large population cities (more than 9500 population) purchase small amounts (1-5) of lana bark, pretus bones, mallorn wood, yew, farrenstone, spiny root, and avinia leaf at various prices each month.

7.2.3 Resolution of trades

Buyers who are willing to pay the most trade first. They buy first from sellers who are asking the least. Ties (two buyers offering to buy at the same price, for instance) are resolved randomly.

7.2.4 Partial trades

Partial trades are executed, if possible. If a buyer wants 100 iron, but only 25 iron are available, he will buy 25. The pending buy order will be reduced so that 75 more will be bought whenever possible.

7.2.5 Sales Tax

The ruler of a city may place a tax on all sales transacted in the city’s market and the market in the city’s Trading Guild (if any). This tax is a percentage of the sale price of a good, and is taken from the amount paid by the buyer to the seller. For example, if a noble sells 10 iron at 20 gold each, and the city has a 10% tax, he will receive 180 gold, the other 10% of the 200 gold sale price being placed into the local garrison.

The tax on a market (if any) is shown in the market report.

Needless to say, all nobles who rule in the market’s province are exempt from sales tax. This includes all the nobles directly or indirectly on the ruler’s admit list (see ⟨undefined⟩ [Control of Provinces], page ⟨undefined⟩).

7.2.6 More market examples

A buyer who desires three iron at no more than five gold each orders:

    > buy 79 3 5
    Try to buy three iron [79] for 5 gold each.

Suppose the only seller of iron in this city had ordered:

    > sell 79 5 6
    Try to sell five iron [79] for 6 gold each.

The trade will not take place, because the seller’s price exceeds the buyer’s. Later, the buyer travels to a different city, where another seller has previously issued:

    > sell 79 10 4
    Try to sell ten iron [79] for 4 gold each.

At the end of the day, the trade will be matched:

    Bought three iron [79] for 12 gold.

Note that the trade takes place at the seller’s price (because there were no other buyers), and that the seller will still be offering seven iron at four gold each.

If the seller had only one unit of iron for sale, the trade would have executed, but the buyer would still be looking for two more iron to buy, at five gold each or better.

7.2.7 Trade goods

In addition to the common city markets, cities buy and sell trade goods (such as wool, rugs, old books, etc.) in a special market inside the local Trading Guild. Normally a city buys three trade goods and sells three trade goods.

Each month, the city produces the three goods it offers for sale and places them on sale, and places orders for the three goods it buys. The number of goods produced (or bought) depends upon the tradegood. For example, cities produce 50 wool per month, 10 rugs per month, and 1 old book per month.

Trade goods are initially offered (for sale or purchase) at their "base price". For example, the base price for wool is 19 gold, the base price for rugs is 125 gold, and the base price for old books is 1200 gold.

If a good is for sale and goes unsold, the price will be reduced. If the good is sold, it’s price will be raised. A good’s price will never be higher than approximately twice its base price, nor lower than approximately half its base price. There will never be more of a good for sale than the city produces in one year. A good that goes unsold for a time will be removed from the market and replaced with another good.

If a good is demanded and goes unfulfilled, the price will be raised. If the good is filled, it’s price will be reduced. A good’s price will never be higher than approximately twice its base price, nor lower than approximately half its base price. A good that goes unfulfilled for a time will be removed from the market and replaced with another good.

Trading skills exist to manipulate the market in trade goods in various ways.

Chapter 7: Economics & Growth 103

Base prices for trade goods can be found in ⟨undefined⟩ [Appendix - Reference Sheets], page ⟨undefined⟩.

8 Miscellaneous

8.1 Hints

Here are some starting hints for new players.

(1) Don’t try to be self-sufficient. Contact your fellow players and work with them. There is a huge advantage to playing as part of a "team".

(2) NPs are the most important resource in the game. Use them wisely. Most nobles can be contract loyalty. At the very least, wait until your initial contract loyalty wears off before you "oath" a noble. By then you might have lots of gold!

(3) Don’t spend too much of your starting gold before you establish some new sources of income. Here are the common ways to earn money in TAG:

Taxes/Tariffs: high income, but difficult to establish. Fishing: steady income, needs a port city. Inns: low income, ties up a noble. Pots/Baskets: low income, lots of competition, easy. Treasure Hunting: potentially high rewards, dangerous. Trade: high income, difficult to establish.

(4) The Age of Gods is not vanilla Olympia! Check the rules before you assume something works the same in both games. In particular, The Age of Gods is based on Olympia G1, not Olympia G2 or later.

8.2 Offenses against realism

• The training skills (such as Train peasant to worker) specify the time needed to train one man; time to train n men is linearly proportional. There are none of the economies of scale one would expect to find when teaching groups of men with a single instructor, and individual men can be trained in unreasonably short periods of time.

• Units in Olympia move by first waiting the duration of the journey in the starting location, then instantly jumping to the destination. One can see units long after they have departed. An interrupted movement order instantly brings units back to their starting location. Note that most interaction with units, such as giving items and stacking, as well as combat, is explicitly prevented after they have begun movement.

• Reports are received for every location visited by a player’s units up until the end of the month, even if the units have since left the viewed locations.

8.3 Credits

"Olympia: The Age of Gods" is based upon "Olympia". Much credit goes to Rich Skrenta for the development of that excellent game.

Several people helped with the design and offered suggestions about The Age of Gods. These include:

• Rich Skrenta

• Steve Chapin

• and Greg Lindahl

Appendix A Appendix - Reference Sheets

Training List

num kind  combat skill input man  input item  where

— ———– —— —– ————– —————- —— 11 worker (1,1) none [10] peasant 12 soldier (5,5) 1108 [36] skirmisher 13 archer (1,1,35) 1902 [12] soldier longbow [72] 14 knight (90,90) 1190 [37] hvy. foot warmount [53] 15 paladin (180,180) 1191 [14] knight guild 16 pikeman (5,20) 1100 [12] soldier pike [75] 19 sailor (1,1) 1001 [10] peasant 20 lt. foot (30,30) 1105 [12] soldier longsword [74] 21 lt. xbowman (1,1,15) 1100 [12] soldier lt. crossbow [85] 22 elite archer (1,1,45) 1902 [13] archer elvish arrow [39] 24 pirate (10,10) 1105 [19] sailor longsword [74] ship 28 fanatic (10,5) 2635 [12] soldier 29 ninja (5,5) 2733 [12] soldier 36 skirmisher (3,3) 1100 [10] peasant 37 hvy. foot (60,60) 1192 [20] lt. foot plate armor [73] 38 hvy. xbowman (1,1,75) 1100 [12] soldier hvy. crossbow [40] 41 horse archer (1,20,40) 1190 [13] archer riding horse [52] 43 cavalier (75,75) 1190 [12] soldier mithril [82],

warmount[53]

Item Information

item name                     mnt weight land ride fly cost beast?
---- ----                     --- ------ ---- ---- --- ---- ------
  1 gold                      0      0    0    0    0    0  N
 10 peasant (1,1,0)           1    100  100    0    0    0  N
 11 worker (1,1,0)            2    100  100    0    0    0  N
 12 soldier (5,5,0)           2    100    Y    0    0    0  N
 13 archer (1,1,35)           3    100    Y    0    0    0  N
 14 knight (90,90,0)          5    400    Y    Y    0    0  N
 15 paladin (180,180,0)       2    400    Y    Y    0    0  N
 16 pikeman (5,20,0)          3    100    Y    0    0    0  N
 18 ghost warrior             0      0    0    0    0    0  N
 19 sailor (1,1,0)            2    100  100    0    0    0  N
 20 lt. foot (30,30,0)        3    100    Y    0    0    0  N
 21 lt. xbowman (1,1,15)      2    100    Y    0    0    0  N
 22 elite archer (1,1,45)     4    100    Y    0    0    0  N
 23 angry peasant (2,1,0)     2    100  100    0    0    0  N
 24 pirate (10,10,0)          3    100    Y    0    0    0  N
 25 elf (50,50,100)           0    100  100    0    0    0  N
 26 spirit (50,50,0)          0    100  100    0    0    0  N
 27 postulant (1,1,0)         0    100  100    0    0    0  N
 28 fanatic (10,5,0)          0    100    Y    0    0    0  N
 29 ninja (5,5,0)             2    100    Y    0    0    0  N
 30 angel (25,100,0)          0    100  100    0  100    0  N
 31 undead (25,1,0)           0    100  100    0    0    0  N
 32 savage (1,1,0)            0    100  100    0    0    0  N
 33 skeleton (5,5,0)          0    100  100    0    0    0  N
 34 barbarian (2,1,0)         0    100  100    0    0    0  N
 36 skirmisher (3,3,0)        1    100    Y    0    0    0  N
 37 hvy. foot (60,60,0)       4    100    Y    0    0    0  N
 38 hvy. xbowman (1,1,75)     2    100    Y    0    0    0  N
 39 elvish arrow              0      1    0    0    0    0  N
 40 hvy. crossbow             0     20    0    0    0    0  N
 41 horse archer (1,20,40)    3    400    Y    Y    0    0  N
 43 cavalier (75,75,0)        4    400    Y    Y    0    0  N
 44 lt. wagon                 0    500 1000  500    0    0  Y
 45 hvy. wagon                0    900 1800  900    0    0  Y
 46 war wagon (+20 defense)   0    900 1800  900    0    0  Y
 51 wild horse                0    300    Y    Y    0    0  Y
 52 riding horse              0    300  300  150    0   25  Y
 53 warmount                  0    300  400  200    0    0  Y
 54 winged horse              0    300  150  150  150    0  Y
 55 nazgul (360,360,0)        0    300  150  150  150    0  Y
 56 zombie (15,15,0)          0    100  100    0    0    0  N
 57 wight (75,75,0)           0    100  100    0    0    0  N
 59 flotsam                   0     30    0    0    0    0  N
 60 battering ram (30,250,0)  0   3000    0    0    0    0  N
 61 catapult (25,200,25)      0   3000    0    0    0    0  N
 62 siege tower               0   3000    0    0    0    0  N
 63 ratspider venom           0      0    0    0    0    0  N
 64 lana bark                 0      1    0    0    0    0  N
 65 avinia leaf               0      1    0    0    0    0  N
 66 spiny root                0      1    0    0    0    0  N
 67 farrenstone               0      1    0    0    0    0  N
 68 yew                       0      3    0    0    0    0  N
 69 elfstone                  0      1    0    0    0    0  N
 70 mallorn wood              0     30    0    0    0    0  N
 71 pretus bones              0      1    0    0    0    0  N
 72 longbow                   0      3    0    0    0    0  N
 73 plate armor               0      5    0    0    0    0  N
 74 longsword                 0      5    0    0    0    0  N
 75 pike                      0      5    0    0    0    0  N
 76 ox                        0   1000 1500    Y    0  100  Y
 77 wood                      0     30    0    0    0   13  N
 78 stone                     0    100    0    0    0    0  N
 79 iron                      0     10    0    0    0   30  N
 80 leather                   0     30    0    0    0    0  N
 81 ratspider (5,5,0)         0      1    0    0    0    0  N
 82 mithril                   0      5    0    0    0    0  N
 83 gate crystal              0      1    0    0    0    0  N
 84 blank scroll              0      2    0    0    0    0  N
 85 lt. crossbow              0     10    0    0    0    0  N
 86 rug                       0    200    0    0    0  125  N
 87 fish                      0      2    0    0    0    2  N
 88 pepper                    0      5    0    0    0   10  N
 89 pipeweed                  0     10    0    0    0   27  N
 90 ale                       0     50    0    0    0   13  N
 91 wool                      0    200    0    0    0   19  N
 92 jewels                    0      3    0    0    0    0  N
 93 opium                     0      1    0    0    0    0  N
 94 woven basket              0      1    0    0    0    4  N
 95 clay pot                  0      5    0    0    0    7  N
 97 fish oil                  0     10    0    0    0    0  N
 98 drum                      0      2    0    0    0    0  N
 99 hide                      0     50    0    0    0  130  N
102 lead                      0    500    0    0    0   10  N
103 fine cloak                0     15    0    0    0  140  N
104 chocolate                 0     25    0    0    0  115  N
105 ivory                     0    100    0    0    0   40  N
106 cardamom                  0      5    0    0    0  205  N
107 honey                     0    100    0    0    0   20  N
108 ink                       0     50    0    0    0  160  N
109 licorice                  0     50    0    0    0  110  N
110 soap                      0     50    0    0    0   11  N
111 old book                  0     25    0    0    0 1200  N
112 jade idol                 0     10    0    0    0 1900  N
113 purple cloth              0    100    0    0    0   36  N
114 rose perfume              0      3    0    0    0 2150  N
115 silk                      0     50    0    0    0   35  N
116 incense                   0     10    0    0    0  185  N
117 ochre                     0     25    0    0    0  195  N
118 jeweled egg               0      2    0    0    0 2200  N
119 obsidian                  0    500    0    0    0   31  N
251 orange                    0    100    0    0    0   21  N
252 cinnabar                  0     50    0    0    0  165  N
253 myrrh                     0      4    0    0    0  210  N
254 saffron                   0      1    0    0    0 2250  N
255 dried fish                0      5    0    0    0    8  N
256 tallow                    0     50    0    0    0   35  N
257 candles                   0     25    0    0    0   85  N
258 wax                       0     50    0    0    0   45  N
259 sugar                     0    100    0    0    0    6  N
260 salt                      0    100    0    0    0    5  N
261 pitch                     0    100    0    0    0   50  N
262 linen                     0    100    0    0    0   18  N
263 beans                     0    100    0    0    0   14  N
264 walnuts                   0    100    0    0    0  145  N
265 flax                      0    100    0    0    0   12  N
266 flute                     0      5    0    0    0 1500  N
267 cassava                   0    100    0    0    0  170  N
268 plum wine                 0     50    0    0    0  130  N
269 vinegar                   0     50    0    0    0   80  N
270 tea                       0     25    0    0    0   15  N
271 centaur (30,30,30)        0    300  150  150    0    0  Y
272 minotaur (30,5,0)         0    500    Y    Y    0    0  Y
278 giant spider (150,100,0)  0    100    Y    Y    0    0  Y
279 rat (3,3,0)               0     20    Y    0    0    0  Y
280 lion (100,100,0)          0    300    Y    Y    0    0  Y
281 giant bird (300,250,0)    0    100    Y    Y    Y    0  Y
282 giant lizard (80,80,0)    0    100    Y    Y    0    0  Y
283 bandit (3,3,0)            0    100  100    0    0    0  N
284 chimera (250,250,0)       0    100    Y    Y    0    0  Y
285 harpie (80,120,0)         0    100    Y    Y    Y    0  Y
286 dragon (1500,1500,0)      0   5000    Y    Y 5000    0  Y
287 orc (20,15,0)             0    100  100    Y    0    0  N
288 gorgon (10,20,0)          0    100    Y    Y    0    0  Y
289 wolf (5,5,0)              0    100    Y    Y    0    0  Y
290 crystal orb               0      1    0    0    0    0  N
291 cyclops (25,75,0)         0    200  200    0    0    0  N
292 giant (75,25,0)           0    200  600    0    0    0  N
293 faery (1,1,1)             0     10   10    0    0    0  N
295 golden seagrass           0      1    0    0    0    0  N
296 firewort                  0      1    0    0    0    0  N
297 beastnip                  0      1    0    0    0    0  N
298 elf poppy                 0      1    0    0    0    0  N
299 ironwood                  0      1    0    0    0    0  N
300 king's fern               0      1    0    0    0    0  N
301 moon palm                 0      1    0    0    0    0  N
302 otter (10,10,0)           0     10   20    0    0    0  Y
303 mole (10,10,0)            0     10   20    0    0    0  Y
304 bull (10,10,0)            0     10   20    0    0    0  Y
305 eagle (10,10,0)           0     10   20    0    0    0  Y
306 monkey (10,10,0)          0     10   20    0    0    0  Y
307 hare (10,10,0)            0     10   20    0    0    0  Y
308 wardog (10,10,0)          0     10   20    0    0    0  Y
309 sand rat (10,10,0)        0     10   20    0    0    0  Y
310 balrog (1000,1000,0)      0   1000    Y    Y 1000    0  Y
311 dirt golem (125,125,0)    0    500 1000    0    0    0  N
312 flesh golems (250,250,0)  0    750 2000    0    0    0  N
313 iron golem (500,500,0)    0   2500 5000    0    0    0  N
314 lesser demon (250,250,0)  0    100  100    0    0    0  N
315 greater demon (500,500,0) 0    100  100    0    0    0  N
316 green rose                0      1    0    0    0    0  N
317 elf ear                   0      1    0    0    0    2  N
318 savage ear                0      1    0    0    0    2  N
319 nazgul tail               0      1    0    0    0   40  N
320 centaur hide              0     10    0    0    0    5  N
321 minotaur hide             0     15    0    0    0    5  N
322 spider eye                0      1    0    0    0    2  N
323 rat tail                  0      1    0    0    0    1  N
324 lion mane                 0      5    0    0    0    5  N
325 giant bird feather        0      1    0    0    0   10  N
326 lizard tail               0      5    0    0    0    5  N
327 bandit ear                0      1    0    0    0    1  N
328 chimera eye               0      1    0    0    0   10  N
329 harpie feather            0      1    0    0    0    5  N
330 dragon scale              0     50    0    0    0   40  N
331 orc scalp                 0      2    0    0    0    4  N
332 gorgon liver              0      2    0    0    0   10  N
333 wolf hide                 0      5    0    0    0    3  N
334 cyclops eye               0      1    0    0    0    8  N
335 giant's tongue            0     10    0    0    0   10  N
336 balrog horn               0      5    0    0    0  300  N

Cost is the base price an item receives when first offered for sale or purchase in a city or Trading Guild market.

Beast? indicates whether the item can be captured with the skill "Capture Beasts".

Combat Details

• A noble is rated (attack=80, defense=80, missile=0).

• Pirates fight (30,30,0) on a ship, but only (10,10,0) on land.

• Knights, paladins and cavaliers receive a -50 penalty to both their attack and defense ratings when fighting on a ship or in a swamp province.

    structure               men protected
    ---------               -------------
    Castle                  first 500
    Tower                   first 100
    Other structures        first 50
    Reduction in Defensive Bonus
    engine          castle  tower  other
    ------          ------  -----  -----
    siege tower       1       5     10

Command List

accept from-who item [qty]  pay to-who [amount] [have-left]
accept nation item [qty]  pillage [flag]

accept clear  players

admit who-or-what [‘ALL’] [units]

attack direction or target […] [flag]  post # of lines of following text
banner [unit] "message"                          press # of lines of text

behind number  practice skill
board ship  promote who

border where OPEN | CLOSE  quarry [number of stones] [days]
bribe who amount [flag]  quit
build structure "Name" [hulls] [max-days]  attack [target]
buy item qty price [have-left]  recruit [days]
catch [number of horses] [days]  repair [days]
claim item [number]  resend [turn]
collect item [number] [days]  rumor # of lines of text
contact who  sail direction or destination […]
decree decree  seek [who]
dedicate priest  sell item qty price [have-left]
default who  split bytes lines

defend who  stack noble

drop item qty [have-left]  stop
email new email address  study skill [study points flag]
execute [prisoner]  take who item [qty] [have-left]
exhume [body]  tax where [market | province | weight | men] rate

explore                                       teach skill

fee weight | men | nobles gold [where]  terrorize who severity

ferry                                         train kind [days]

fish [number of fish] [days]  ungarrison [garrison]

flag string  unload

fly direction or destination […]  unstack [who]
forget skill  use skill [arguments …]
form unit "name of new noble"                    email new email address for player list
format num  wait clear: waits for clear weather.
garrison castle  wait day n: waits until at least day N.
get who item [qty] [have-left]  wait ferry n: waits for ship N to sound its horn.
give to-who item [qty] [have-left]  wait flag f [n]: wait for noble N to signal flag F.
guard flag  wait fog: waits for fog.
guild skill school  wait gold n: wait until noble has N gold.
honor amount  wait item n q: wait until noble has Q of item N.
hostile who  wait loc n: waits for the noble to arrive at N.
identify item id  wait not: inverts sense of following condition.
make item [qty]  wait owner: waits until noble is first in the location.
message # of lines of text to-who  wait rain: waits for rain.
move direction or destination […]  wait ship n: waits until specified ship arrives.
maxpay [gold]  wait shiploc n: waits until ship is at location N.
name [unit] "new name for unit"                   wait stack n: waits until noble is in a stack with N.
nationlist                                    wait time n: waits for N days to elapse.
neutral who  wait top: waits until noble is the stack leader.
notab num  wait turn: n waits until at least turn N.
oath level  wait unit: n waits until N arrives.
password ["password"]  wait wind: waits for wind.

Appendix B Appendix - Mythology of Olympia

In the beginning was Sol. And Sol shone in the darkness.

And Sol was unhappy, for he was alone. He wept, and his tears collected into a cool blue ball that became Luna, called the Mother of All Gods.

Luna reflected Sol’s light and together the two were happy for many ages. And as will always be when two love one another, children were born.

At first, the children played in the cool blue light of their Mother and the bright yellow radiance of their father. But Sol realized that his children needed a place to grow and learn. So with Luna he created Olympia and placed his children there.

"Here you will learn to be gods, as your Mother and your Father are," said Sol to his children. "Here I command you to compete, to learn and to challenge; the greatest of you I shall call to join me in the Heavens."

And Sol filled Olympia with many creatures and wonders to delight and instruct his children. Foremost amongst all he created Man, and to Man alone he gave the power of speech and reason, that Man might be a companion to his sons and daughters.

For many ages Man lived in harmony with the children of Sol. Many children of Sol became great in wisdom and power and were called forth into the Heavens. The night sky blazed with a thousand of Sol’s children. And always Sol and Luna loved one another and sent new children to Olympia to learn to be Gods.

One of these children was named Nereid. Of all of the children of Sol and Luna, Nereid was the most lovely. It is said that when she walked flowers sprung up in her footsteps. She was also wise and kindly, and much loved by Man. The favorite of her father, she was soon called up into the Heavens, where she shone as the brightest star in the night sky.

As the ages passed, Sol loved Nereid more and more deeply. Always she moved closer to her father, so that soon Nereid alone shone in the daylight with her father. Blinded by his daughter’s beauty and wisdom, Sol conceived an illicit desire for his daughter. Although he hid his love, Sol’s heart was troubled. His visage became cloudy, and Man upon Olympia saw Sol reach out yearningly towards his daughter.

Finally Sol was overcome by his desire and cast himself upon his daughter. The heavens burned a sickly purple at what Sol had done, and the face of Luna became dark and swollen with sadness. Nereid herself was changed. Her visage, once clear and blue like her mother, now became the feeble orange of a guttering candle. She cast herself across the Heavens away from her father, hiding herself in shame, until soon she could be seen only on the darkest of nights.

Luna, too, turned away from Sol. Once Luna had hung low over Olympia, watching her children and protecting them from many dangers. Now she drew herself further away and turned her face away from Sol, so that for weeks at a time Luna could not be seen at all in the night sky.

Of the union of Sol and Nereid a son was born, called Tor. In the nights following Sol’s rape of his daughter, Tor could be seen rushing across the sky, tearing open a great wound in the Heavens in his haste to be with his father. But Sol refused to look upon Tor, and cast him down instead upon Olympia, where he landed with a great crash that shook the mountains and caused the seas to rise up in angry protest.

And so Tor came to Olympia. He had all the dazzling beauty of his mother Nereid, and all the great power of his blazing father. He did many great works as he walked upon Olympia, and soon many proclaimed him the greatest of all Sol’s children.

Then Tor began to teach those who followed him that Sol was wrong to place them on Olympia, and that all deserved to live in the Heavens as Gods. He whispered to Man that they should be as the sons of Sol, and he taught Men the secrets of magic, which Sol had forbidden. Through all this Sol did nothing. And many of the other sons and daughters of Sol, seeing this, joined with Tor.

At the height of his powers, Tor gathered his followers together and challenged even the power of death, opening gateways upon the face of Olympia into the realm of Hades, where the lost souls of fallen Men were said to wander. But even while Tor and his followers celebrated his great triumph, Sol grew angry at last.

He rose up in the sky and flashed with an angry, unnatural brilliance. The seas of Olympia boiled and the trees on the mountains burst into flame. Tor and his followers turned and fled into the deep recesses beneath Olympia, but many were killed by Sol’s anger.

And it seemed as if all would die, for the face of Olympia was aflame. But then Luna interceded to save her children. She drew a veil of cloud across Sol’s face and bathed the land in a healing rain. (This is why, in the eastern provinces of the Empire, Luna is called the Lady of the Veil.) When Luna next drew the veil back, Sol had retreated to his rightful place in the sky.

Since the Time of Fire, Sol has turned his back upon Olympia. No longer do the children of Sol come to Olympia; no longer do the learned and wise children of Sol rise up from Olympia to join him in the Heavens. Neither does Luna protect her children, though she still turns her face each month from Sol in shame and anger.

Of the children of Sol, eight survived the Time of Fire. These eight were called Eres, the God of Sea and Wind, Anteus, the God of Fire and Earth, Dol the Beastmaster, Timeid, Lady of Plants, Ham the Jester, Kireus the Smith, Halon Warmaster and Domingo the Hidden. Although they no longer show their visages upon Olympia for fear of angering their father, it is said that they sometimes wander the land in the guise of Man, guiding and helping their followers.

Of Tor and the rebel children of Sol less can be said. Some claim they still live under Olympia, in a kingdom they have created in the dark regions below. It is said that the orcs and other foul creatures who occasionally venture forth unto the face of Olympia from the realms below are the twisted children of the Men who followed Tor. Others claim that Tor and the other Dark Children are still worshipped in secret upon the face of Olympia, but no one has yet found a temple to one of those shamed children.

What is certain is that some men still seek greedily after the powers that Tor shared with Man. Wizards in lonely towers spend endless hours studying the tomes and artifacts from the time before the Rebellion hoping to unlock their secrets. Although some wizards start out as innocent scholars, most soon turn toward the dark side of power. Indeed, orcs and other foul creatures are sometimes seen in the service of the wizards, and other wizards surround themselves with undead and other fell beasts from beyond the grave.

Appendix C Appendix - History

All of Olympia is divided in two parts: the Imperial Empire and the Mandorian Confederacy.

In ages past, there was only the Imperial Empire, and peace reigned on the face of Olympia. Then a professional seeker of antiquities named Mandor discovered a library of books and artifacts in the basement of a ruined tower partway up a mountain. The books were in an ancient tongue, but after long study Mandor was able to translate them. Mandor had discovered the fabled lost library of Bastrestric ther Archymonaged, the greatest of the human mages that had fallen in the Time of Flames (see ⟨undefined⟩ [Appendix - Mythology of Olympia], page ⟨undefined⟩). Mandor began to study the books, and gathered a small group of like-minded seekers of knowledge. The lost art of Magic was refound.

It came to pass after some years that Mandor and his followers were discovered by priests of the Eight Churches. The priests denounced Mandor and his followers as disciples of Tor, and accused them of evil crimes against Man. Mandor and his followers protested that they were only innocent seekers of knowledge, but the priests refused to listen, and soon raised up an army and drove Mandor and his followers to the edge of the ocean.

Although harried by sword and bow, and driven like cattle, Mandor and his followers did not resist, only protesting again and again their innocence. But at the edge of the great sea, Mandor himself turned and stood against the army arrayed against him, crying out "See now what you have called upon yourself!" With that he unleashed great magics against the priests of the Eight Churches and their soldiers, such as had not been seen since the Time of Flames. Fireballs laid waste great sections of the army; lightning bolts rained down out of a clear sky upon the priests; and fallen men rose up from the dead to battle their brethren.

While Mandor stood against the armies, his followers gathered into ships and fled across the ocean and out of the known lands. And when they had gone, his magics exhausted, Mandor fell beneath the swords of the priests.

And in the Imperial Empire there was great rejoicing that the plague of Tor had again been cleansed from the land.

But the followers of Mandor sailed on across the ocean for long weeks. Many perished in the empty sea, but at last a few survivors beached upon the shores of a new land. There they built a city and a nation, naming it Mandoria after their fallen leader.

In time Mandoria grew in size to rival the Imperial Empire. In both nations memories grew dim. In the Imperial Empire, Mandor was little more than a bogeyman to scare children; in the Mandorian Confederacy their origins in the Imperial Empire were forgotten. Then daring explorers of both nations discovered the others, and long forgotten tales of horror were rekindled. In the Imperial Empire the priests of the Eight Churches preached against the horrors of Tor and the evils that magicians worked against their fellow men. In the Mandorian Confederacy, the streets were filled with stories of persecution and the bloodlust of the Empire.

And so the Age of Gods began, when religious wars raged across the face of Olympia, and man turned against man until the oceans themselves were red with blood.

Appendix D Appendix - Listing of all commands

accept

accept from-who item [qty] accept nation item [qty]

 accept clear
 • time: 0 days
 • priority: 0

The ACCEPT order must be used in order to receive items from other players. Gold is exempt from the ACCEPT check.

If ACCEPT is issued by a noble, it applies to items given to that noble only. If ACCEPT is issued by the player entity, it applies to every unit in the faction.

If from-who is 0, the item will be accepted from anyone. Otherwise fromwho should specify a noble, a faction, or a nation name. If item is 0, any item will be accepted. If qty is zero or is not present, any quantity of the item will be accepted.

If qty is specified, the given quantity must be less than or equal to the qty specified on the matching ACCEPT order. If the GIVE is successful, the given quantity will be deduced from the remaining qty value in the outstanding ACCEPT.

A GIVE which could be passed by multiple overlapping ACCEPT orders will choose the first which matches, first for the noble, then for the player entity. If an ACCEPT order specifies an item quantity limit, the quantity must be greater than or equal to the amount specified in the GIVE order. Multiple conflicting ACCEPT orders will not combined to satisfy a single GIVE.

"ACCEPT CLEAR" is used to clear all existing accepts. If used by a noble, it clears all accepts for the noble. If used by the player entity, it clears all the player-level accepts.

Examples:

         ACCEPT 0 0              # accept anything from anyone
         ACCEPT Mandor 0         # accept anything from Mandorians
         ACCEPT 0 10             # accept peasants from anyone
         ACCEPT 2950 0           # accept anything from noble 2950
         ACCEPT 2950 10          # accept peasants from noble 2950
         ACCEPT 2950 10 5        # accept up to 5 peasants from
                                 # noble 2950

         ACCEPT Pen 10 5         # accept up to 5 peasants from Pen
         ACCEPT 501 10           # accept peasants from any noble
                                 # in faction 501
         ACCEPT 501 10 5         # accept up to 5 peasants from
                                 # nobles in faction 501
         ACCEPT clear      # clear all existing accepts

The current status of what a noble will accept is shown in the turn report. The turn report first lists what the noble will specifically accept and then any accepts that he inherits from the player entity:

      Accepting:
          Accepting any amount of iron [76] from The Duchess [yy6].
          (Accepting any amount of anything from any member of Mandor
          Confederacy.)

In this case, the noble is specifically accepting iron from another player, and the entire faction is accepting anything from any member of the Mandor Confederacy.

admit

admit who-or-what [`ALL'] [nation | units]

 • time: 0 days
 • priority: 0

Allow units to stack with the specified noble, or to enter the given building or ship. By default, a unit may not stack with a noble belonging to another faction.

The first argument to admit may either be a noble, building or ship. If a noble is given, admit will specify who may stack with the unit. If a building or ship is given, admit will determine who may enter the structure.

There are three cases for the rest of admit’s arguments:

(1) If no further arguments are given, admit will clear the permissions for the given entity, and deny admission or stacking to any foreign unit:

    admit ce90                  # admit no one to ce90

(2) If a list of units, players, or nations is given, those units will be permitted to enter or stack with the given entity:

    admit ce90 778 2960 4240    # allow any member of faction 778,
                                # as well as 2960 and 4240, into ce90
    admit ce90 mandor 4481      # admit Mandorians and 4481

Note that if a faction is specified, units within the faction must not be concealing their lord in order to be recognized.

(3) If the keyword ‘ALL’ is used, then permission may be granted to every player’s units:

admit ce90 all                  # allow anyone into ce90

A list of units to exclude may follow ‘ALL’:

admit ce90 all 778 2960 4240    # allow anyone into ce90, except
                                # any member of faction 778, and
                                # units 2960 and 4240.
admit ce90 all Pen # Admit all, except Pen.

Summary examples:

admit ce90              # admit no one to ce90
admit ce90 all          # admit anyone to ce90
admit ce90 778          # admit any member of faction 778 to ce90
admit ce90 all 778      # admit anyone but a member of faction 778
admit ce90 2960 4240    # admit 2960 or 4240 to ce90
admit ce90 mandor       # admit anyone from Mandor
admit ce90 mandor 2960  # admit anyone from Mandor and 2960

If there are too many units to list on a line with a single admit order, more admit orders may be used:

admit ce90 2596 3921 3934 3999 4012 4045 4046 4256 4300

has the same meaning as:

admit ce90 2596 3921 3934 3999 4012
admit ce90 4045 4046 4256 4300

Don’t forget to give the entity again as the first argument on continuation admit orders! (ce90 in the above example)

Note that in TAG, unlike other versions of Olympia, the first admit order of a turn does not clear the admits for a unit! The only time admits for a unit are cleared is when admit is used with no arguments, e.g.,

        admit g83

clears the admit list for [g83]. Only admit permissions toward the top most noble of a stack are con-

sidered. Other units which do not have admit permission may accompany a noble who does.

Admit lists are kept for the faction as a whole, not for particular units. The admit order should be executed by the player entity.

attack

attack direction or target [...] [flag]

• time: varies
• priority: 3

Engage two stacks in combat. Only the top most noble in a stack may initiate combat.

target may be another unit, a ship, a building, or a location. Compass points may be abbreviated: ‘n’, ‘s’, ‘e’, ‘w’. Units may also attack out of a sublocation to the surrounding province or location.

A unit may attack any unit or location directly accessible from its current location: units in the location, sublocations, or adjacent locations. As a special case, a stack on a ship can attack another ship in the same province, but in general a stack cannot attack from a ship into an adjacent province (or sublocation). So to attack an island, you must first sail to the island and then attack.

Attacking a building or location is really an attack against the controlling unit in that location. The controlling unit for buildings and other sublocations is the first unit in the location. The controlling unit for a province is the garrison. Attacking an empty or unguarded location is the same as moving into the location (see ⟨undefined⟩ [move], page ⟨undefined⟩).

If flag is 1, the attacker will not attempt to move into the defender’s position. Note that in the case of multiple targets or directions, the flag must be the final argument.

For example, one would use the order ‘attack jg99’ to attack this castle:

    Inner locations:
       Castle [jg99], castle, defense 20, owner:
          Law Netexus [2020], with three peasants

Do not use ‘attack 2020’, as Law Netexus is not in the same location as nobles outside the castle. If successful, the attackers will move into the castle.

A unit may not attack another member of its own faction. Two units in the same stack may not engage in combat.

Prisoners may not be attacked. To attack the stack holding a prisoner, issue an attack order for another member of the stack (such as the top unit in the stack).

Multiple targets or directions may be given as arguments to the attack order. If the first fails, the next will be tried, until a target is found, or attack runs out of arguments. Note that only one target will be attacked with each attack order.

The attack order takes one day plus any necessary travel time. Attacking a stack in the same location takes one day. Attacking a location that is 8 days away (i.e., attacking the garrison guarding an adjacent province) takes 9 days – 8 days travel time plus one day for the attack itself. An attack against an empty or unguarded location also requires one day plus travel time, even though it will not result in a battle.

A moving stack does not leave its current location until the day it completes the move. This means that the stack can be attacked until its move completes, e.g., on an eight day move the stack could be attacked for eight days after it begins the move.

For details on the resolution of battles, see ⟨undefined⟩ [Battles], page ⟨undefined⟩.

banner

banner [unit] "message"

 • time: 0 days
 • priority: 1

Provide a short message describing something about the unit. The message will appear after the unit in the turn report.

If a unit number is specified, the given unit’s banner will be changed. This is useful for setting banners on entities which don’t execute orders directly, such as ships and castles. Banners may not be set for units controlled by another player.

Example:

    > banner "carrying a gold standard"
    > banner 2019 "carrying a gold standard"
    Seen here:
       Fighters of Pelenth [2019], "carrying a gold standard"

The message should be limited to 50 characters or it will be rejected.

behind

 behind number
 • time: 0 days
 • priority: 1

Indicate how close to the front lines the unit should stand in combat. Number should be between 0 and 9. 0 is the front, 9 the rear.

Only missile fighters such as archers may attack from the rear. Lt. foot, knights, etc. may not. When all of the units in a row have been killed or taken prisoner, the next row will advance (skipping any empty rows).

board

 board ship
 • time: 0 days
 • priority: 2

Pay a fee to board a ship operating as a ferry. If a maximum fee has been set with maxpay, no more than that amount of gold will be paid as a boarding fee (see See ⟨undefined⟩ [maxpay], page ⟨undefined⟩).

The board order requires that the ship be operating as a ferry. In other words, the ship’s captain must have set a boarding fee with the fee order.

The board order will fail if the ship is already overloaded, or if the ship would be overloaded if the stack boarded the ship.

Note that the board command, unlike the move command, ignores the admit list of the ship’s captain.

Passengers waiting to board a ferry should use wait ferry. Ferry captains should sound their horn before departure with the ferry order.

border

 border where OPEN | CLOSE
 • time: 0 days
 • priority: 0

Open or close a border.

For a small location, such as a building, the noble issuing the border command must be the owner of the building.

For a province or a city, the noble issuing the border command must be in the same faction as the ruler of the province.

For more information, see See ⟨undefined⟩ [Permission to Enter], page ⟨undefined⟩.

breed

breed first second

 • time: 7 days
 • priority: 3

The same as USE 2233. See ⟨undefined⟩ [Breed beasts [2233]], page ⟨undefined⟩.

bribe

bribe who amount [flag]

 • time: 7 days
 • priority: 3

Offer a noble gold in an attempt to win their service. The Persuasion subskill Bribe noble [1301] is required to use the bribe order. The amount must be at least 250 gold, or the receiving noble will not consider it a serious offer.

Nobles with oathbound loyalty ignore bribes. They may accept the gift, but they will not renounce loyalty to their lord.

Nobles with contract loyalty will accept the bribe if it exceeds their current loyalty rating. For example, a noble of loyalty contract-750 would ignore any bribe of less than 751 gold.

Nobles with fear loyalty will accept any bribe of at least 250 gold.

Note that "refugee" nobles may have any sort of loyalty (oath, contract, or fear).

If the bribe is successful, the noble will renounce loyalty to his current lord, and swear fealty to the bribing noble. Nobles who are successfully bribed will have contract loyalty of half the amount of the bribe over their old loyalty. For example, a noble with contract loyalty of 570 who was bribed with 1100 gold would switch loyalties and have a new loyalty of contract-265 ((1100-570)/2). Nobles with fear loyalty before the bribe are treated as if they had contract-250 loyalty.

If flag is 1, and the bribe attempt is successful, the bribed noble will stack with the briber.

Bribes will be refused by nobles who have already switched lords during the month because of bribery. For example, suppose Osswid successfully bribes Feasel, and Feasel swears loyalty to Osswid. Later that turn, if Procrustes attempts to bribe Feasel, Feasel will refuse to accept the bribe.

Examples:

    bribe 5499 750      # Offer [5499] 750 gold to switch factions.
                        # If successful, [5499] will not stack with the
                        # briber.
    bribe 3409 300 1    # Offer [3409] 300 gold to switch factions.
                        # If successful, [3409] will stack with the briber.

build

build structure "Name" [max-days] [identifier] build structure [max-days] build ship "Name" hulls [max-days] [identifier]

 build ship [max-days]
 • time: varies
 • priority: 3

Build the indicated structure or ship. All construction tasks require at least three workers to begin. The noble should acquire the required quantity of building material, unstack from beneath other nobles, and issue one of the following build commands:

    build inn "name of inn"
    build shaft "name of mine"
    build temple "name of temple"
    build tower "name of tower"
    build castle "name of castle"

build ship "name of ship" number of hulls Note that you are *required* to provide a name when starting a building,

and a name and a number of hulls when starting a ship. If you do not provide a name and/or a number of hulls, the build command will likely fail.

Any of these commands can be followed by the maximum number of days to spend building, or the maximum number of days and an identifier (see below).

For example, to start building a castle, one might issue:

    build castle "Castle Blood"

or

    build castle "Castle Blood" 14

indicating that the noble should build for 14 days, or

    build castle "Castle Blood" 14 aj4

indicating that the noble should build for 14 days and the id "aj4" should be used.

If you wish to specify an identifier for the building, but not a maximum number of days to work, you should use zero for the maximum number of days, e.g.,

    build castle "Castle Blood" 0 aj4

Do not simply leave out the maximum number of days:

    build castle "Castle Blood" aj4

This will be misinterpreted by the game, with unpredictable results.

A new structure or ship will be created, and the builder will be placed inside. Construction proceeds until completed unless maximum days has been specified.

kind        effort    material   skill         where
----        ------    --------   -----         -----
inn            300     75 wood    1400         province or city
mine shaft     500        none    1703         mountain or mine shaft
temple       1,000    50 stone    1400         anywhere
tower        2,000   100 stone    1400         anywhere
castle      10,000   500 stone    1400         province or city
city     50,000  2000 stone    1490        province
                     500 wood
ship hull      100    10 wood     1002         port or ocean province

Effort is in worker-days. Construction [1400] is required to erect buildings. Build Hull [1002], a

subskill of Shipcraft [1000], is required to build ships. Mines may be built with Build Mine Shaft [1703].

To resume construction on a ship or building, enter the uncompleted structure and use one of the following commands:

build inn
build shaft
build temple
build tower
build castle
build ship

Note that when resuming work you do not specify a name (or the number of hulls for a ship). Any of these commands can also be followed by a maximum number of days to work.

Do not include the name of the structure when resuming work (or the number of hulls). This will be misinterpreted by the game with unpredictable results.

The first tower a player builds requires no workers, no stone, and only takes a day to build.

If the player specifies an identifier for a structure, it must be drawn from the list of identifiers for new nobles listed at the top of the player’s turn report:

The next five nobles formed will be:  af8 8e8 ae8 a8e a88

Note that this identifier follows the maximum days (and possibly the number of hulls), so these must be specified as well. For example, to build a tower using one of your assigned identifiers and to work on it until completed, you would use the command:

build tower "My Tower" 0 ae8

Here, maximum days is specified as zero as a "placeholder" so that the game correctly understand that you intend "ae8" to be the identifier for the new tower.

For more details, see ⟨undefined⟩ [Buildings], page ⟨undefined⟩ and ⟨undefined⟩ [Ships], page ⟨undefined⟩.

buy

buy item qty price [have-left]

 • time: 0 days
 • priority: 1

Buy quantity of the given item for no more than price each.

If quantity is zero, any pending buy order for the indicated item will be canceled.

Only one pending buy order may be active for each item at a time. If two buy orders are issued for the same item, and the first has not yet completed, the second buy order will replace the first.

Buys and sells are only matched in cities, where markets may be found to execute trades. However, buy and sell may be issued anywhere.

Trades only occur at the end of the day. Therefore, you must stay in a location for at least one day to execute a trade. If you step into a market, issue a buy or sell, and then leave immediately, no trade will occur. To be certain to stay at least one day, you may issue a ⟨undefined⟩ [wait], page ⟨undefined⟩ command.

Example:

    > buy 79 3 5
    Try to buy three iron [79] for 5 gold each.

If there is a seller of iron in the city, selling for 5 gold or less, the trade will be executed:

    1: Bought three iron [79] for 15 gold.

Otherwise, the buy will be placed on the noble’s pending trades list:

    Pending trades:
    trade    price    qty   item
    -----    -----    ---   ----
    buy        5      3   iron [79]

At the end of the turn, the buy order will appear in the city’s market report:

    Market report:
    trade    who    price    qty   item
    -----    ---    -----    ---   ----
      buy   2019        5      3   iron [79]

A trade will match between a buyer and a seller if the buyer’s price is greater than or equal to the seller’s price. The actual price that the trade will be made at is the seller’s price.

Have-left is an amount of gold that the noble will not spend on this trade. For instance, if the unit issued buy 79 2 10 5, this would mean

‘Buy two iron [79] for 10 gold each, but not if I'll have less than

5 gold left over.’ If the noble had 15 gold, only one of the iron would be bought.

catch

catch [number of horses] [days]

 • time: as given
 • priority: 3

Attempt to catch any wild horses found roaming the plains or in a pasture. Each day the noble will have a 25% chance of catching a wild horse, if any can be found.

Each plains or pasture location contains a population of horses. Plains provinces can support 100 wild horses, while pastures, being smaller, can support only 20 horses. At the end of each month, the horse population breeds new horses. In plains, one new horse is bred each month. In pastures, where forage is better, three new horses are born each month.

If the population of horses in any location drops to zero, it will take some time for the horse population to recover. A noble can add wild horses to a location by dropping them from his inventory in that location.

A unit must know the Catch wild horses [1930] subskill of Ranger [1900] in order to catch horses.

Examples:

    catch        # catch as many horses as possible
    catch 1      # stop after catching 1 wild horse
    catch 0 10   # try to catch horses for 10 days
    catch 1 10   # stop after catching 1 wild horse, or 10 days,
                 # whichever comes first

In all cases, horse-catching ceases when no more horses can be found roaming the plains. catch is an alias for ‘collect 51’.

claim

claim item [number]

 • time: 0 days
 • priority: 1

New players are given a few things to get them started. Rather than load them with possessions when they first set foot in the world, slowing their travel and make them tempting targets for bandits, they may obtain these items at any time with the claim command.

New players are given:

    5,000 gold [1]
       50 wood [77]

Any unit in the player’s faction may claim the items.

    > claim 77
    Claimed 50 wood.
    > claim 1 1000
    Claimed 1,000 gold.

collect

collect item [number] [days]

 • time: as given
 • priority: 3

Catch, harvest, collect or produce the given item from the unit’s current location, if possible. Collection proceeds until the desired quantity of the item has been obtained, or the given number of days has passed, whichever comes first.

If neither number or days is given, collect will continue until the location’s supply of the item is fully depleted.

Collection ceases when no more of the resource is available in the location. Thus, ‘collect 87 10’ may obtain less than 10 fish [87], if the current ocean province’s supply of fish runs out.

Examples:

    collect 10         # Recruit as many peasants [10] as possible
    collect 68 2       # Harvest two yew trees
    collect 87 0 10    # Fish for 10 days
    collect 87 25 10   # Fish until 25 fish [87] have been caught, or
                       # for 10 days, whichever comes first

Many collection commands have aliases so they can be more easily remembered. For example, the above orders could be written:

    recruit yew 2 fish 0 10 fish 25 10

Most collection tasks require the unit to possess a particular skill, and to be in a likely location to obtain the desired item.

    num  item        who     where            qty  skill  chance  alias
    ---  ----        ---     -----            ---  -----  ------
    -----
     10  peasant     noble   city, province^{1} 10   none   100%   recruit
     68  yew         worker  yew groves         5   1603   100%   yew
     77  wood        worker  forest            30   1602   100%   wood
     78  stone       worker  mountains         50   1402   100%   quarry
     78  stone       worker  rocky hills       10   1402   100%   quarry
     78  stone       worker  desert            10   1402   100%   quarry
     87  fish        sailor  ocean             50   1004    50%   fish
     51  wild horse  noble   plains            ^{2} 1930    25%   catch

Other rare items which may be collected are listed in Collect rare foliage [1604], a subskill of Forestry [1600], and Collect rare elements [1503], a subskill of Alchemy [1500].

Explanation of table columns:

num item The resource to be collected. For example, Fish [87].

who Some tasks may only be performed by a noble, while others must be done by men, such as workers or sailors. For instance, recruiting requires the attention of a noble. Additional men are of no benefit.

Fishing, however, requires sailors. The noble himself does not participate. The more sailors available for fishing, the quicker the fish will be caught.

where qty Resources are replenished at the end of each month. Each ocean

province will have 50 fish that may be caught. Mountains yield up to 50 stone per turn.

skill Skills are required for most collection tasks (except recruiting):

activity        subskill                  category
--------        --------                  --------
recruiting      none                      none
cutting yew     Harvest yew [1603]        Forestry [1600]
cutting timber  Harvest lumber [1602]     Forestry [1600]
quarrying       Stone quarrying [1402]    Construc-
tion [1400]
fishing         Fishing [1004]            Shipcraft [1000]
horse catching  Catch wild horses [1930]  Ranger [1900]

chance The rate of collection is certain for most tasks. Cutting timber yields one wood [77] per day per worker, as long as the forest has a supply of wood. However, some tasks have an element of randomness. Horsecatching, for example, is a difficult, uncertain task. Each day there is only a 25% chance that a noble will catch a wild horse. The chance is for the unit as a whole, not for individual men engaged in the task. Thus, a unit of sailors will collect fish about half of the days they spend fishing. The other half, no fish will be caught.

alias Common collection tasks have an alias so they are easier to remember.

task            alias
----            -----
collect 10      recruit
collect 68      yew
collect 77      wood
collect 78      quarry
collect 87      fish
collect 51      catch

For example, the orders ‘collect 51 0 10’ and ‘catch 0 10’ are equivalent.

Note that if no arguments are supplied to the collect order (or one of its aliases), it may never terminate. For instance, suppose a unit with one sailor fishes alone in an ocean province:

fish    -or-    collect 87

Since no arguments were given for the fish order, fishing will continue as long as the ocean has a supply of fish. The sailor will produce, on average, a fish every two days, or 15 fish per month. Since this is less than the monthly supply of fish in an ocean province, the sailor will never exhaust the supply, and thus is likely to continue fishing forever, until interrupted with a stop order.

contact

 contact who
 • time: 0 days
 • priority: 0

Used by nobles hidden through stealth, magic, or because of fog to inform other units of their presence.

For example, if Osswid [5501] is employing stealth skills to hide, he will not be visible to other nobles. Orders issued by other units referencing [5501] will fail.

However, Osswid may desire to allow certain units to interact with him while he is hidden. If this is the case, he should issue contact orders for those units to be informed of his presence.

The contact order may also be specify a player number. In this case, every unit in the specified player’s faction will be aware of the hidden unit’s presence.

Contact information is maintained until the hidden unit moves. Once the unit has moved, it must issue new contact orders if other units are to continue to be able to interact with it.

The contact order is silent; it does not provide a message to the contacted player or faction.

Units belonging to the same player are always aware of each other’s presence. They do not need to use contact.

Hidden units are not shown in the location report, even if their presence has been reveal through contact or seek.

Related orders: ⟨undefined⟩ [seek], page ⟨undefined⟩

decree

 decree decree
 • time: 0 days
 • priority: 0

Issue a decree to all garrisons that the noble controls. There is no way to issue selective decrees to individual garrisons; all garrisons under control receive the same decree.

The decree may be one of the following: watch who

Instruct the garrisons to watch for the given unit. Each garrison may accept up to three decree watch order per month. Additional watch decrees are ignored.

hostile who Instruct the garrisons to attack the specified unit on sight. Each garrison will accept up to three decree hostile orders per month. Additional hostile decrees are ignored. Unlike nobles and players, who can declare themselves hostile to nations or monsters, garrisons can only be declared hostile against specific units.

dedicate

 dedicate priest
 • time: 7 days
 • priority: 3

By use of this command, a noble dedicates himself to a priest (and, by extension, the priest’s god).

Once dedicated, a noble may not change his dedication until the priest is dead and laid to rest. Priests may not use this command; they are already dedicated to their god. Mages also may not use this command, since the gods will not accept mages as followers.

The priest and the dedicating noble must be in the same location for the entire time of the dedication ceremony.

default

 default who
 • time: 0 days
 • priority: 0

Clear any hostile, neutral or defend attitude for the unit or faction. May be issued by either a noble or the player entity.

who may be either a unit id, a nation name, or "monster".

See ⟨undefined⟩ [Declared attitudes], page ⟨undefined⟩ for more information about declared attitudes.

Related orders: ⟨undefined⟩ [defend], page ⟨undefined⟩ ⟨undefined⟩ [neutral], page ⟨undefined⟩ ⟨undefined⟩ [hostile], page ⟨undefined⟩

defend

 defend who
 • time: 0 days
 • priority: 0

Declare a unit or player to have attitude defend towards another unit, nation, or player who. When a unit is attacked, other units which have defend attitude to the unit will join the combat and aid with defense.

To declare an attitude toward a nation, use the nation’s name (or any non-ambiguous first part of the name) in the command:

    defend Mandor

Case is ignored.

When a combat occurs, the attitude of the top-most noble in each stack in the location is checked. If any unit is declared defend to the victim of the attack, the unit’s stack will join the combat on the victim’s side.

who may be either a unit id or a nation name.

See ⟨undefined⟩ [Declared attitudes], page ⟨undefined⟩ for more information about declared attitudes.

Related orders: ⟨undefined⟩ [default], page ⟨undefined⟩ ⟨undefined⟩ [neutral], page ⟨undefined⟩ ⟨undefined⟩ [hostile], page ⟨undefined⟩

drop

drop item qty [have-left]

 • time: 0 days
 • priority: 1

Discard the specified quantity of item. If have-left is specified, at least this many will not be discarded. If quantity is zero, the entire inventory of the given item will be discarded.

For example, if Feasel has 100 gold [1]:

    drop 1 50          #  Throw away 50 gold
    drop 1 0           #  Throw away 100 gold  (0 means all)
    drop 1 100         #  Throw away 100 gold
    drop 1 100 25      #  Throw away 75 gold   (keep 25)
    drop 1 0 25        #  Throw away 75 gold   (keep 25)

Common items including gold, lumber, swords, etc. are permanently lost when discarded. It is assumed that they are quickly scavenged by others.

Unique items, such as magical artifacts, scrolls, potions and rare books, do not become permanently lost. Such items will be found through exploration of the province in which they were lost. Unique items lost at sea will wash ashore on a nearby island or coast, where they may also be found through exploration.

The drop order is also be used to release men such as peasants, workers, soldiers, etc. from service:

    drop 10 3               # Release 3 peasants from service

When used on a ship, drop will throw items overboard.

email

email new email address

Change the address to which turn results should be sent to the given address. Must appear after the begin order.

To change your email address, send in

    begin 999
    email new@address.com   # give your new address
    end

As a security measure, the confirmation will be sent to both the new and old addresses.

If you want to send your turn report to two or more addresses, use both addresses in the email command, separated with a comma:

    begin 999
    email new@address.com,other@address.com
    end

execute

 execute [prisoner]
 • time: 0 days
 • priority: 1

Put the given prisoner to death. If no prisoner is specified, all of the noble’s prisoners will be executed.

explore

 explore
 • time: 7 days
 • priority: 3

Explore the current location for hidden features or routes. If there is something to find, there is a 33% chance with each attempt that the explorer will find it.

    1: > explore
    1: Explore location for hidden features.
      ...
    7: Exploration uncovers no new features.

Exploration in a province may also turn up rare items which have been lost.

If the exploring noble is on a ship on the ocean, the surrounding ocean province will be searched for hidden islands or sea routes.

Priests of Domingo have a 50% chance of success.

fee

 fee weight | men | nobles gold [where]
 • time: 0 days
 • priority: 1

Set a fee required to board a ship or enter a location.

The first argument to fee is the word "weight", "men" or "nobles". This selects whether the fee is per 1000 weight entering, per 100 men entering, or per noble. The second argument is the cost:

    > fee nobles 5  # charge each noble 5 gold to enter

Fees may be charged in combination by using the fee command several times:

    > fee nobles 5  # charge each noble 5 gold to enter a17.
    > fee men 10    # Also charge 10 gold for each 100 men they have.

For small locations such as ships, towers, yew groves, and etc., the fee can only be set by the owner of the location.

For provinces and cities, the fee can only be set by the ruler of the province or city. The ruler must have garrisoned the province and be the owner of the controlling castle. (For convenience, any noble in the ruler’s faction is permitted to issue the fee command.)

To set the fee for a province or city, use the optional third argument to specify where you are setting the fee:

    > fee nobles 10 a17  # charge each noble 5 gold to enter a17.

Fees for provinces and cities are collected only when a garrison is present and a noble is sitting in the controlling castle.

New fees do not take affect until the end of the month. (This prevents an unscrupulous player from springing a "surprise" fee on you.) Fees appear in the turn reports. Note that if you enter a location sight unseen, you might be charged an unexpected fee!

To remove the fee from a location, set the fee to zero. When a "small" location is vacated, any fee set in that location will be removed. Fees in provinces and cities persist until removed.

ferry

 ferry
 • time: 0 days
 • priority: 1

Sound a ship’s horn, signaling units in port waiting with wait ferry that they may begin boarding the ferry. ferry may only be used by the captain of a ship.

Units waiting to board a ferry should use wait ferry instead of wait ship to allow existing passengers on the ferry time to unload. If those trying to board do so before existing passengers have unloaded, the board order may fail because of ship weight limits.

fish

fish [number of fish] [days]

 • time: as given
 • priority: 3

A unit must be on a ship on the ocean in order to fish. Each day, the unit has a 50% chance of encountering fish. If fish are found, each sailor will catch one fish that day, as long as fish stocks in the ocean province last. Ocean provinces generate 50 catchable fish per month.

A unit must know the Fishing [1004] subskill of Shipcraft [1000] in order to fish, and must have at least one sailor.

Only sailors may be used for fishing. Pirates will not fish.

Examples:

fish         # obtain as many fish as possible
fish 25      # stop after getting 25 fish
fish 0 5     # fish for five days
fish 25 5    # stop after five days, or 25 fish, whichever
             # comes first

In all cases, fishing will cease as soon as the stocks in the ocean province have been exhausted. fish is an alias for ‘collect 87’.

flag

flag string
• time: 0 days
• priority: 1

Signal nobles waiting with a wait flag order. The string is an arbitrary sequence of nobles or numbers, but it should match the flag specified in the

wait flag order.

For example, a noble may wait for a particular signal from Osswid [5499]:

> wait flag yellow 5499

This will wait for Osswid [5499] to raise a flag saying ‘yellow’. Osswid would issue:

> flag yellow

Which would cause the wait order to complete. The wait flag order defaults to watching for flags from units in the

player’s faction. It may also watch for flags from units in another faction, a flag from a specific unit, or a flag from any unit in the game:

wait flag yellow        # wait for 'flag yellow' from a unit
                        # in our faction
wait flag yellow 5499   # wait for 5499 to order 'flag yellow'
wait flag yellow 918    # wait for any unit belonging to player
                        # 918 to order 'flag yellow'.
wait flag yellow 0      # wait for any unit in the game to give
                        # the yellow flag

Nobles do not need to be in the same location to flag. Flags are remembered for the duration of the month. If a wait flag is

issued for a flag which has already been raised, the wait will immediately finish (it will take zero time).

However, flags which have been raised are not remembered between turns, so a flag raised on the 30th of month 5 will not satisfy a wait flag on the 1st of month 6.

fly

fly direction or destination [...]

 • time: varies
 • priority: 2

Fly in the indicated direction, or to the specified location. Refer to the move order for details about movement.

Nobles must have some means of flight to be able to use fly. For instance, a sufficient number of winged horses must be available to carry the entire stack and its belongings. If the weight of the stack exceeds the flight weight capacity, the fly order will fail.

Flying traverses any link in a maximum of three days.

    rated time      time to fly
     of route         across
    ----------      -----------
    0 days          0 days
    1 day           1 day
    3 days          3 days
    4 days          3 days
    8 days          3 days

Flying animals, such as winged horses, become tired when flying over the ocean, as they have no place to land and rest. Up to fifteen days may be spent flying over the ocean. At the end of the fifteenth day, the exhausted animals will be unable to continue, and the entire stack will plunge into the sea, killing all of the stack’s nobles.

Any arrival at a land or ship location will allow the flying creatures to rest and regain their strength.

Multiple destinations or destinations may be given as arguments to fly. If the first fails, the next will be tried, until a passable route is found, or fly runs out of arguments.

For example, ‘fly n e w’ would first try to fly north. If travel north was not possible for any reason, east would next be tried, then west. In this way exploring units can attempt to deal with the case of running into a hole in the map or other impassable province.

forget

 forget skill
 • time: 0 days
 • priority: 1

Forget all knowledge of a skill. If a category skill is specified, all subskills within the category will also be forgotten. You *never* recover any NPs (or gold) you might have spent to learn a skill when you forget it.

In some cases, effects from learned skills persist even after you’ve forgotten the skill. For example, if you use [1102] Fight to the Death to set your breakpoint to 0%, that breakpoint may remain even after you forget that skill.

form

form unit "name of new noble"

 • time: 7 days
 • priority: 3

Spend a Noble Point to obtain a new noble for the player’s faction. Nobles must be formed in a city. Note that if you’re in a tower (or in a castle, a guild, a temple, etc.) in a city, form will fail. You must be directly in the city.

The unit parameter should be selected from the line near the top of the turn report listing new noble numbers:

    The next five nobles formed will be:  5717 3215 4902 4489 5628

If 0 is given for unit, the first number in the list will be chosen. However, if more than one form is to be issued during the month, it can be uncertain which will execute first. By specifying the new unit number, the player will know the correct entity number when submitting orders for the new noble.

Any noble in the player’s faction may issue form. The newly obtained noble will appear and stack beneath the forming noble. The new noble is given the name specified as the second argument to form.

Orders for new units may be submitted in the turn they are formed. As soon as the new unit comes into existence, it will begin to execute its commands.

New units are unequipped and untrained, and join with loyalty contract-500.

Example:

    1: > form 3215 "Feasel the Wicked"
    1: Attempt to form a new noble.
    7: Feasel the Wicked [3215] stacks beneath us.

More than five units may not be formed in a turn.

format

 format [CLEAR] [HTML] [TEXT] [RAW] [TAGS] [ALT] [PATH http://...]

The format command is used to set the format of the turn report.

Four basic turn report formats are available:

The RAW format produces "raw" output. Raw output is largely unformatted and contains both the embedded tags and all of the text output. The raw format is useful mainly to players who want to post-process their turns themselves, e.g., to produce their own version of HTML output. For details on the raw format, see ⟨undefined⟩ [Appendix - Output Formats], page ⟨undefined⟩.

The TEXT format produces a text-only turn report.

The HTML format produces a turn report suitable for viewing with a Web browser.

Specifying the TAGS format embeds Age of God tags into your turn report. These tags can be used to automatically process the turn report, e.g., to extract location information for a map database. For details on the format of tags, see ⟨undefined⟩ [Appendix - Output Formats], page ⟨undefined⟩.

You may select the format of your turn report by using the format command followed by one or more of the HTML, TEXT, or RAW keywords. You may set multiple formats (e.g., HTML and RAW). In this case, you will be sent multiple turn reports, one in each format you’ve selected. The turn reports will contain a keyword in the subject line of the email indicating the format (so that you can automatically sort them).

To reset your format choices, use "FORMAT CLEAR". This will clear all formats. (If you do not select a format, you will receive only a text format turn report.)

Some players find the location activity logs less useful than others, and prefer them to be banished to the end of the turn report. The command "FORMAT ALT" will do this. This alternate format is also cleared by the "FORMAT CLEAR" command.

0 Noble reports listed first, followed by location reports including activity logs.

1 Location reports without activity logs, followed by noble reports, followed by location activity logs.

Finally, you may use the format command to set an initial path for the external links in your HTML format turn report.

In the HTML turn report format, any noble, location or item that does not appear in your turn report is linked to an HTML page of the same name. For example, a noble "John Hawkins [ag7]" will be linked to "http:ag7.html". If you create a web page "ag7.html" in the directory where you keep your turn reports, then clicking on the name "John Hawkins [ag7]" will link you to that web page (where you could, for example, store information about him).

However, you might want that link to point to some other location. For instance, you might be part of an alliance that keeps all information on a central server. In that case, you can use the "FORMAT PATH" command to make the link point to that server. For example, if you give the command "FORMAT PATH http://home.me.org/tag/", then "John Hawkins [ag7]" will be linked to "http://home.me.org/tag/ag7.html".

To clear the path, use "FORMAT PATH".

garrison

 garrison castle
 • time: 1 day
 • priority: 3

Claim a province in the name of a castle by leaving a garrison of soldiers to defend it.

The castle must be completed in order to bind a garrison to it.

There may not be any existing garrison in the location. Only provinces may be garrisoned (the noble should be at the outermost level, not inside a sublocation or building).

The noble must have ten soldiers [12] to create a garrison. The soldiers will be removed from the noble’s inventory and placed in the new garrison.

See ⟨undefined⟩ [Garrisons], page ⟨undefined⟩ for details about how garrisons work.

get

get who item [qty] [have-left]

 • time: 0 days
 • priority: 1

Get the given quantity of item from the specified target.

Nobles may get items from other units in their faction, or from garrisons which they control. get may not be used to take items from another faction’s units; the other faction’s units must issue give orders.

Specifying zero for the quantity will cause the unit to attempt to take all of the target’s inventory of the item. If have-left is given, at least that many will not be taken.

For example, if Feasel [5499] had 100 gold[1]:

    get 5499 1 50          #  Get 50 gold from Feasel
    get 5499 1 0           #  Get all 100 gold from Feasel
    get 5499 1 100         #  Get all 100 gold from Feasel
    get 5499 1 100 25      #  Get 75 gold from Feasel
    get 5499 1 0 25        #  Get 75 gold from Feasel

Prisoners may also be transferred with the get order.

    get 5499 4569          #  Transfer prisoner [4569] from [5499]

When taking men from a garrison, the garrison must be left with at least ten men, or the get will fail. For example, if a garrison had five peasants and six soldiers, either a peasant or a soldier could be taken. No men can be taken from a garrison with less than ten men.

give

give to-who item [qty] [have-left]

 • time: 0 days
 • priority: 1

Give quantity of an item in inventory to another noble. Specifying zero for quantity will cause the unit to give all of its inventory of the item. An amount specified by have-left will be retained.

For example, if Feasel has 100 gold [1]:

    give 5499 1 50          #  Give 50 gold to [5499]
    give 5499 1 0           #  Give all 100 gold to [5499]
    give 5499 1 100         #  Give all 100 gold to [5499]
    give 5499 1 100 25      #  Give 75 gold to [5499]
    give 5499 1 0 25        #  Give 75 gold to [5499]

Prisoners may also be transferred with the give order.

    give 5499 4569          #  Give prisoner [4569] to [5499]

If you give something to a noble from another faction, that noble must use the accept command or the gift will be refused. See ⟨undefined⟩ [accept], page ⟨undefined⟩.

guard

 guard flag
 • time: 0 days
 • priority: 1

Protect a city or province from looting and pillaging. ‘guard 1’ will cause the unit to suppress pillage attempts by other nobles. The guarding unit must be the top unit in its stack.

Any number of units in a location may be on guard. The message ‘on guard’ will appear next to guard units in the location report.

A noble attempting to pillage the city or province may specify a flag which causes them to attack the guards. In this case, all of the stacks on guard will respond to the attack. However, allies of the guard units will not respond. Only guard stacks will be involved in combat to suppress looting and pillaging.

Specifying a flag of 0 turns off guarding. Any movement by the guarding noble will also clear the guard flag. For example, if a unit issued ‘guard 1’ in a city, then moved out of the city, the guard flag would be cleared, and the unit would no longer be on guard.

The guarding units must be inside they location they are to guard. Units in a castle in a city will not protect the city. Units guarding a city will not protect the surrounding province.

Note that a guarding unit will not come to the defense of a garrison if it is directly attacked. To help defend a garrison, use the ⟨undefined⟩ [defend], page ⟨undefined⟩ command.

guild

guild skill school

 • time: 7 days
 • priority: 3
 • cost: 250 gold

From inside a tower, convert it to a guild hall for the indicated skill school (see ⟨undefined⟩ [Guild Halls], page ⟨undefined⟩). To use this command the noble must be the only occupant of the tower.

To create a guild hall, a noble must either (1) already be a member of the guild for the indicated skill school, or (2) be ready to join that guild (i.e., must know all the non-guild skills of that skill school). Creating a guild hall causes the creator to join the appropriate guild.

Guilds must be created in cities, and only one guild of a type is allowed in a city.

honor

honor amount

 • time: 0 days
 • priority: 3

Pay a unit for its service. Units paid will take gold out of their own inventory and use it for their own purposes.

Nobles with oathbound loyalty will refuse to be paid, and reject the gold. Nobles with fear loyalty will be converted to contract loyalty.

The amount of money paid accumulates in the noble’s contract loyalty rating. For instance, a fear-10 noble honored with 500 gold would have loyalty contract-500. A contract-250 noble honored with 500 gold would become contract-750.

hostile

 hostile who
 • time: 0 days
 • priority: 0

Declare unit, nation or player who hostile to the unit or player issuing the order. Units declared hostile will be attacked on sight. Players may also specify "hostile monster" which will cause all monsters to be attacked on sight.

To declare an attitude toward a nation, use the nation’s name (or any non-ambiguous first part of the name) in the command:

    hostile Mandor

Case is ignored.

If the player entity issues hostile, any unit in the player’s faction will attack the hostile unit on sight. If an individual noble issues hostile, the noble will attack the hostile unit on sight.

Units will only initiate automatic attacks if they are in perfect health. Hostile units which are being held prisoner will not be attacked. Automatic attacks take place at priority 3 (the same as ⟨undefined⟩ [attack], page ⟨undefined⟩).

An entire faction may be declared hostile by specifying another player entity instead of an individual unit. Any unit not concealing its affiliation with the faction declared hostile would be attacked on sight.

Examples:

    hostile 4240            # attack 4240 on sight
    hostile 778             # attack any member of faction 778 on sight

See ⟨undefined⟩ [Declared attitudes], page ⟨undefined⟩ for more information about declared attitudes.

Related orders: ⟨undefined⟩ [default], page ⟨undefined⟩ ⟨undefined⟩ [defend], page ⟨undefined⟩ ⟨undefined⟩ [neutral], page ⟨undefined⟩

identify

 identify item id
 • time: 0 days
 • priority: 3

The identify command causes a noble to closely examine an item for the tiny, cryptic runes that indicate the purpose of the ancient magical artifacts that can sometimes be found in Olympia (see ⟨undefined⟩ [Appendix - Magical Artifacts], page ⟨undefined⟩). If the runes can be found and read, the identify command will display the item’s enchantment, e.g.,

    8: > identify atq
    8: You carefully read the runes on this artifact and identify it as:
    8: Reduced Maintenance, 9%.

Some magical artifacts will have no enchantment, while others will be missing runes, and so cannot be identified, e.g.,

    8: > identify cz0
    8: You carefully read the runes on this artifact and identify it as:
    8: no enchantment.
    8: > identify atq
    8: You are unable to identify that item.

Note that identify does not change the name of the artifact. You can change the name of the artifact using the ⟨undefined⟩ [name], page ⟨undefined⟩ command, if you’d like. Keep in mind that someone scrying your noble will see that new name!

make

make item [qty]

• time: as given
• priority: 3

Make the indicated item. qty defaults to as many as possible. Each of the following items requires one day each to produce:

num   item           skill    material
---   ----           -----    --------
40   hvy. crossbow  1106     wood [77] & iron [79]
72   longbow        1106     yew [68]
73   plate armor    1106     iron [79]
74   longsword      1106     iron [79]
75   pike           1106     wood [77]
80   leather        none     ox [76]
84   blank scroll   1500     lana bark [64]
85   crossbow       1106     wood [77]
94   woven basket   none     none
95   clay pot       none     none

Knowledge of Weaponsmithing [1106], a subskill of Combat [1100], is required to make weapons and armor. Alchemy [1500] must be known to produce blank scrolls from lana bark. No skills or input materials are needed to make woven baskets or clay pots.

Examples:

To make 5 longbows from 5 yew (a kind of wood):

make 72 5               # takes 5 days

To make 1 plate armor from 1 iron:

make 73 1               # takes 1 day

To make 5 woven baskets:

make 94 5               # takes 5 days

Although making pots and baskets is considered menial labor, some desperate nobles have resorted to selling them in cities to earn gold. Each city in Olympia either buys woven baskets or clay pots.

message

message # of lines of text to-who

 • time: 1 day
 • priority: 3

Send a message to the target named in the second parameter. The target may be any valid entity, such as a noble, a player, or a location (including provinces, cities, ships, buildings, or other sublocations).

The number of lines in the message should be given as the first argument to message. The text of the message should follow the message order.

For example:

    message 2 2909
      Aye, ye not be safe in this world for long, Oleg.
      I be remembering ya, and yer past fould deeds.  Arrr!
    move ev18
    recruit

(The text of the message is indented here to make the orders easier to read. The order scanner will ignore leading spaces on each line.)

This would send the message to noble [2909]. The recipient would see the message in his turn report:

    Received a message from Captain McCook [2104]:
       "Aye, ye not be safe for long in this world, Oleg.
    I be remembering ya, and yer past foul deeds.  Arrr!"

Take care to carefully count the lines of text in the message. If the wrong number of lines is given, your noble may end up sending his orders to the recipient instead of executing them.

Each line of message text must be less than 60 characters long, or the message will be rejected.

move

move direction or destination [...]

 • time: varies
 • priority: 2

Travel in the indicated direction, or to the specified location. If a location is specified, there must be a direct route to it from the noble’s current location.

Compass points may be abbreviated: ‘n’, ‘s’, ‘e’, ‘w’. Units may also move out of a sublocation to the surrounding province or location, or move in to a sublocation from the surrounding province or location. Note that move in is ambiguous if there are two sublocations (e.g., a sacred grove and a temple). In that case you should move to the id of the sublocation you want, so that you don’t end up someplace unexpected (e.g., in the sacred grove when you expected to be in the temple).

Units may also move up and move down to go up and down. Currently this is only useful when moving up and down through the levels of a mine.

Visible routes leaving each province are listed in the turn report.

For example:

    Routes leaving Plain [cw04]:
       North, to Ocean [cv04], South Sea, 1 day
       East, to Plain [cw05], 7 days
       South, to Plain [cx04], 7 days
       West, to Plain [cw03], 7 days

The following would be valid move commands from the example province:

    move e  -or-  move cw05  -or-  move east
    move s  -or-  move cx04  -or-  move south
    move w  -or-  move cw03  -or-  move west

Note that move cannot be used to enter the ocean region [cv04]. The noble would need to issue the sail order as captain of a ship to move north into the South Sea.

Route distances are rated in the number of days it would normally take a lightly loaded noble, walking, to traverse them. Actual travel times may differ from the rated times.

Nobles overloaded with belongings will make slower progress. Horses usually speed progress, but over especially rough or treacherous terrain, they may actually slow travel because they must be led and managed.

Inventory is considered for the stack as a whole. For instance, if one noble possesses enough horses for every member of the stack, the stack may ride. Similarly, if one noble is badly overloaded, but other nobles have extra carrying capacity, they will assist their stackmate. Distribution of goods between units within a stack is irrelevant to movement.

The units stacked beneath a moving noble will move with the stack leader. If the unit issuing move is stacked beneath another noble, he will first automatically unstack before movement begins.

For example:

    Seen here:
       Osswid the Constructor [5499], accompanied by:
          Feasel the Wicked [502]

If Osswid issues a move order, Feasel will go with Osswid. If Feasel issues a move order, however, Feasel would first unstack from Osswid before beginning travel.

Multiple destinations or directions may be given as arguments to the move order. If the first fails, the next will be tried, until a passable route is found, or move runs out of arguments.

For example, ‘move n e w’ would first try to move north. If travel north was not possible for any reason, east would next be tried, then west. In this way exploring units can attempt to deal with the case of running into an impassable province or a coast.

Some locations may charge noble to enter. See ⟨undefined⟩ [Entrance Fees], page ⟨undefined⟩.

The size of the entrance to some locations is limited, and only stacks of that size or smaller may enter (or attack) that location. For example:

    Inner locations:
       Lair [ad7n], lair, 1 day, entrance size: 4, owner:
          Harpie [jn7t], harpies, number: 12

Lair [ad7n] can only be entered – or attacked – by stacks of size four or less. (If you attack with more than that number in your stack, the attack will immediately fail.) Notice that this does not restrict the size of stacks inside the location, only the number that may enter at one time.

Some locations may not permit a noble to enter. For example, a province guarded by a garrison may close its border, refusing admission to nobles not on the ruler’s admission list. In this case, the move order will fail in zero time.

To enter a closed location, a noble should use the attack order. It is sometimes convenient to use move and attack in combination, to ensure entry to a location. For example, if Feasel gave the commands:

    move cw05 attack cw05

he would first try to move into cw05. If that succeeds, the subsequent attack command will fail (since Feasel will already be in cw05). If the move fails because the border is closed, the attack will happen. If the move fails for some other reason (such as the location not existing) then the attack will fail as well.

A moving stack does not leave its current location until the day it completes the move. This means that the stack can be attacked until its move completes, e.g., on an eight day move the stack could be attacked for eight days after it begins the move.

maxpay

 maxpay [gold]
 • time: 0 days
 • priority: 1

Sets the maximum amount a unit will pay to enter a location (see See ⟨undefined⟩ [fee], page ⟨undefined⟩). If set to zero, a unit will refuse to enter any location charging a fee.

The default is a maxpay of zero (pay nothing).

name

name [unit] "new name for unit"

 • time: 0 days
 • priority: 1

Change the name of an item or unit. If a unit number is specified, the given unit’s name will be changed. This is useful for setting the name on entities which don’t execute orders directly, such as ships and castles. Names may not be set for units controlled by another player.

Names of nobles should be no longer than 35 characters. Names for other entities, such as towers, ships and items should be no longer than 25 characters. Longer names will be rejected.

Note that the name must be enclosed in quotes or it will not be recognized. Examples:

    > name "Feasel the Wicked"
    > name cx53 "Castle Chardia"
neutral

 neutral who
 • time: 0 days
 • priority: 0

Declare unit, nation or player who to have attitude neutral to the player or faction issuing the order. Units declared neutral will not be aided if they are attacked. Players may also specify "neutral monster".

To declare an attitude toward a nation, use the nation’s name (or any non-ambiguous first part of the name) in the command:

    neutral Mandor

Case is ignored.

Units in the same faction will defend each other if they are attacked, unless either unit is concealing its lord. This behavior may be suppressed with neutral. A unit may declare other units in its faction neutral. Or, the entire faction may declare itself neutral. In this case, members of the faction who were attacked would not aid each other, unless they were stacked together.

See ⟨undefined⟩ [Declared attitudes], page ⟨undefined⟩ for more information about declared attitudes.

Related orders: ⟨undefined⟩ [default], page ⟨undefined⟩ ⟨undefined⟩ [defend], page ⟨undefined⟩ ⟨undefined⟩ [hostile], page ⟨undefined⟩

notab

 notab num

Ordinarily, tab characters may appear in turn reports. Some players read their reports on computer equipment which cannot correctly display tab characters. If this is the case, issue the command:

    notab 1

and tabs will be converted to spaces for your report.

oath

oath level

 • time: 1 day
 • priority: 3

Have a noble take an oath of loyalty. Either 1 or 2 should be specified for level. One or two NPs will be deducted from the player’s faction, and the specified noble’s loyalty will become oath-1 or oath-2.

Noble Points’s spent on oathbound loyalty accumulate. Thus, oath-1 may be upgraded to oath-2 with the ‘oath 1’ order.

password

 password ["password"]

Set or clear a password for the player. If no argument is given, any existing password will be cleared. Otherwise, the specified password will be set.

Examples:

    begin 663
    password "xyzzy9"
    end

This would set a password ‘xyzzy9’ for the faction. Further orders will require the password to be specified after the player number on the begin line:

    begin 663 "xyzzy9"
    password                # clear the password
    end
pay

pay to-who [amount] [have-left]

 • time: 0 days
 • priority: 1

Pay amount of gold to indicated entity, keeping at least have-left funds available. The pay command is shorthand for the equivalent give command to transfer gold from one unit to another.

The amount of gold to be paid will default to all of the unit’s gold if not specified or if zero.

For example, if Feasel has 100 gold:

    pay 5499 100            # pay 100 to [5499]
    pay 5499 0              # pay 100 to [5499]
    pay 5499 0 25           # pay 75 to [5499]
    pay 5499 100 25         # pay 75 to [5499]
pillage

 pillage [flag]
 • time: 7 days
 • priority: 3

Loot and pillage a province or city. If flag is not given, pillaging will not take place if the province is being protected by guards. If flag is 1, the looters will first attack the guards, then proceed with pillaging if the guards are defeated.

At least 10 fighters must be in the pillaging stack. Pillaging may be done with beasts or men, but the noble directing the pillaging must be able to control the pillagers in battle. Only the top most noble in a stack may initiate pillaging; only the men he directly controls are involved in the pillaging. Other nobles in his stack will continue with whatever actions they are doing.

Pillaging kills one peasant for every two fighters pillaging until there are 100 or less peasants left in a province. It then becomes more difficult to find the peasants, and five fighters are required to find and kill each peasant. Each peasant kill yields 2 gold. For example, a location with 112 peasants pillaged by a group of 218 men would kill 51 peasants and yield (at most) 102 gold. If the peasants do not have the gold to pay off the pillages (e.g., they have paid it in taxes) then the pillagers may collect less than the maximum gold (or none at all).

Pillaging can be dangerous. There is a 33% chance that a mob of 12-36 angry peasants will form to resist further pillaging attempts. Angry peasants have a combat rating of (attack-2, defense-1, missile-0). If a mob forms, there is a 33% chance that the mob will immediately attack the pillager in retaliation.

players

 players

Ask the order scanner to send a list of the players in the game.

post

post # of lines of following text

 • time: 1 day
 • priority: 3

Post a public message in the location. All nobles in the location will see the text of the message in the location description section of their turn reports. Signs remain visible for one turn.

The number of lines in the post should be given as an argument to post, followed by the text of the sign. For example, here is a post command followed by a move and a recruit:

    post 3
      All ye beware!  These lands are the property of
      noble Osswid.  Those who trespass here without
      his permission will come to a GRAVE end!
    move ev18
    recruit

(The text of the post is indented here to make the orders easier to read. The order scanner will ignore leading spaces on each line.)

Take care to carefully count the lines of text in the post. If the wrong number of lines is given, your noble may end up posting his orders instead of executing them.

Each line of posted text must be less than 60 characters long, or the post will be rejected.

press

press # of lines of text

 • time: 0 days
 • priority: 1

Submit a signed message to the Game Newsletter.

The press order may be given by a noble or by the player entity. The identity of the submitting faction will appear in the newsletter. beneath the entry.

The number of lines in the entry should be given as an argument to press, followed by the text of the entry. For example, here is a press item submitted by Osswid [xt8].

    PRESS 3
      All ye beware!  Osswid hereby claims the entire continent
      of Provinia as his rightful domain.  Those who trespass here
      without his permission will come to a GRAVE end!

Take care to carefully count the lines of text in the post. If the wrong number of lines is given, your noble may end up posting his orders instead of executing them.

Alternatively, one may specify 0 for the lines of text argument, and terminate the message with the end keyword:

    PRESS 0
      All ye beware!  Osswid hereby claims the entire continent
      of Provinia as his rightful domain.  Those who trespass here
      without his permission will come to a GRAVE end!
    END

Take great care when using the 0 lines/end keyword feature. If you forget the end, or mistype it, none of the orders following the press command will execute. Instead, they will be published in the newsletter. This is obviously very bad, so if you intend to use the end keyword to avoid having to count lines, please be careful.

Each line of posted text must be less than 78 characters long, or the post will be rejected.

Editorial guidelines for submissions to the Game Newsletter:

1. No profanity.

2. No personal insults to other humans (in-character roleplaying insults are fine).

In some cases, the Times pays money for submissions. If so, it will be noted in the Game description.

practice

 practice skill
 • time: varies
 • priority: 3
 • cost: varies

Some skills may be improved with practice. See ⟨undefined⟩ [Appendix - Skill Dictionary], page ⟨undefined⟩ for the skills that can be improved with practice. Skills that can be practiced list the number of days of continuous practice required to raise the skill one experience level and the cost of practicing the skill.

There is no credit for "partial practice" if practice is interrupted with the STOP command.

promote

 promote who
 • time: 0 days
 • priority: 1

Advance the given unit so that it appears before the noble issuing promote in the location list.

The noble to be promoted should be stacked immediately beneath the promoting noble, or appear later in the location list at the same level.

Examples:

Before:

 Osswid the Constructor [5499], accompanied by:
    Feasel the Wicked [502]

Osswid issues ‘promote 502’:

 Feasel the Wicked [502]
 Osswid the Constructor [5499]

Before:

 Osswid the Constructor [5499], accompanied by:
    Feasel the Wicked [502]
 Candide [1269]

Osswid issues ‘promote 1269’:

 Candide [1269]
 Osswid the Constructor [5499], accompanied by:
    Feasel the Wicked [502]

A common mistake is to think that promote can be used to swap the stack leader and someone in the stack. For example, if you have

 Osswid the Constructor [5499], accompanied by:
    Feasel the Wicked [502]

and Osswid issues ‘promote 502’, you do NOT get this:

 Feasel the Wicked [502] accompanied by:
    Osswid the Constructor [5499]

instead you get two separate stacks, with the Feasel stack ahead of the Osswid stack in the location report:

 Feasel the Wicked [502]
 Osswid the Constructor [5499]

To swap stack leaders, you need some combination of ⟨undefined⟩ [stack], page ⟨undefined⟩ and ⟨undefined⟩ [unstack], page ⟨undefined⟩ commands.

quarry

quarry [number of stones] [days]

• time: as given
• priority: 3

Building stone is used in the construction of temples, castles and towers. Up to 50 stone may be removed from mountain provinces each month. Ten stone may be taken from desert provinces or rocky hills each month.

A unit must have the Stone quarrying [1402] subskill of Construction [1400] in order to cut building stone. At least one worker is also required.

Each worker will cut and prepare one building stone [78] per day.

Examples:

quarry         # cut as much building stone as possible
quarry 25      # stop after cutting 25 stone
quarry 0 5     # quarry stone for five days
quarry 25 5    # stop after five days, or 25 stone, whichever
               # comes first

In all cases, quarrying will cease as soon as the supply of ready stone in the province has been exhausted.

quit

quit

Quit Olympia. This command must be issued by the player entity. No turn report is mailed for the turn a quit order is processed.

raze

raze [target] [days]

• time: varies
• priority: 3

Tear down a building. The command completes when the structure has been completely destroyed, or when days have elapsed. Note that you must specify target to specify days.

A building may be razed from the inside or the outside. If razed from the outside, a building must be unoccupied. The argument target, if present, is the building to be razed.

If razed from the inside, the razing unit must be the owner of the structure (the top-most unit inside the building). The noble will safely leave the building before it collapses.

Sturdier structures require more men to do a given amount of damage:

                 man-days per
structure       point of damage
---------       ---------------
ship                 1
temple               2
inn                  2
tower                2
mine                 2
castle               3

For example, two men can cause 1% of damage to a tower in a day.

Only men in the noble’s inventory will help with destruction. The noble is counted when the number of men are computed. The noble does not have to be the stack leader to raze; it’s possible for two nobles in the same stack to be razing different buildings at the same time.

recruit

 recruit [days]
 • time: as given
 • priority: 3

Recruit one peasant each day for the given number of days. If days is not specified, recruiting will obtain as many peasants as possible.

Peasants may be recruited in any province with more than 100 peasants or any city or city province with more than 1000 peasants. Recruiting can not be done in other sublocations (such as rocky hills, forest glades, etc.)

Up to 5% of a province’s population may be recruited each month. One peasant is recruited each day the recruit command is used.

The recruit order will stop or fail when no further peasants are available to be recruited.

For example, one player could recruit for up to 10 days per month in a mountain province with 200 peasants, obtaining one peasant on each of the 10 days.

If two players in the province each issued recruit 10, both would recruit five peasants (each player would get a peasant a day until all 10 had been recruited).

Recruit is an alias for ‘collect 10’.

repair

 repair [days]
 • time: as given
 • priority: 3

Repair the damaged ship or building that the noble is inside. Requires at least one worker [11].

                    worker-days per
    structure       point of damage
    ---------       ---------------
    ship                 1
    temple               2
    inn                  2
    tower                2
    mine                 2
    castle               3

Work proceeds until the structure has been fully repaired, or the given number of days has passed. Multiple units may work at repair simultaneously.

For example, a tower with 50% damage would require 100 worker-days to fully repair. A ship with 50% damage would require 50 worker-days. A castle, 150 worker-days.

Each use of repair on a ship requires one unit of pitch [261]. One unit of pitch is sufficient for a ship repair job, no matter how many days the repair takes. Pitch may be purchased in any port city.

resend

 resend [turn]

Ask the order scanner to resend a turn report. The turn argument defaults to the turn most recently run.

Because of their large size, turn reports other than the most recent may not be kept by the order scanner. If you request a report and the scanner responds that the report cannot be found, then the reports for that turn were deleted to reclaim disk space.

rumor

rumor # of lines of text

 • time: 0 days
 • priority: 1

Submit an unsigned message to the Game Newsletter.

The rumor order may be given by a noble or by the player entity. The identity of the submitting entity will not appear with the message; rumor posts are anonymous.

The number of lines in the entry should be given as an argument to rumor, followed by the text of the entry. For example: here is a rumor someone submitted.

    RUMOR 3
    As whispered from ear to ear in the great hall of the Chateau Olympique:
    Q:  Why do so many people take an instant dislike to Oleg the Loudmouth?
    A:  Saves time.

Take care to carefully count the lines of text in the post. If the wrong number of lines is given, your noble may end up posting his orders instead of executing them.

Alternatively, one may specify 0 for the lines of text argument, and terminate the message with the END keyword:

    RUMOR 0
    As whispered from ear to ear in the great hall of the Chateau Olympique:
    Q:  Why do so many people take an instant dislike to Oleg the Loudmouth?
    A:  Saves time.
    END

Take great care when using the 0 lines/END keyword feature. If you forget the END, or mistype it, none of the orders following the rumor command will execute. Instead, they will be published in the newsletter. This is obviously very bad, so if you intend to use the END keyword to avoid having to count lines, please be careful.

Each line of posted text must be less than 78 characters long, or the post will be rejected.

Editorial guidelines for submissions to the newsletter:

1. No profanity.

2. No personal insults to other humans (in-character roleplaying insults are fine).

In some cases, the Times pays money for submissions. If so, it will be noted in the Game description.

sail

sail direction or destination [...]

 • time: varies
 • priority: 4

Pilot a ship in the given direction, or to the specified location. If a location is specified, there must be a direct route to it from the current location of the ship.

Compass points may be abbreviated: ‘n’, ‘s’, ‘e’, ‘w’.

The Sailing [1001] subskill of Shipcraft [1000] is required to use the sail order.

The noble issuing the sail order must be the ship’s captain, i.e. the noble must be the first unit listed in the ‘Seen here’ report of the ship.

The sail order will fail if the ship is overloaded. Use drop to throw items overboard if necessary.

Multiple destinations or directions may be given as arguments to sail. If the first fails, the next will be tried, until a passable route is found, or sail runs out of arguments.

For example, ‘sail n e w’ would first try to sail north. If travel north was not possible for any reason, east would next be tried, then west. In this way exploring units can attempt to deal with the case of running into an impassable ocean route, or a coast.

seek

 seek [who]
 • time: 7 days
 • priority: 3

Seek for a hidden noble. Nobles may be hidden because of fog lying over a province, or through the application of stealth or magical skills.

Hidden nobles can not be interacted with by a unit unless they contact the unit first, or the unit successfully finds them with seek. The seek order will return immediately if the noble is visible, or has already been found with a previous seek order.

If a noble is specified, there is a 50% chance that the noble will be found during each of the seven days the order is being executed.

This chance is reduced by 5% for each experience level of [1290] Conceal Self the hidden noble possesses (provided, of course, that he is using the skill to conceal himself). For example, a noble with experience level 4 in [1290] Conceal Self will be spotted only 30% (50-4*5) of the time.

The chance to spot a hidden noble is increased by 5% for each experience level of [1901] Spot Hidden Noble possessed by the noble performing the seek. For example, a noble with experience level 3 in [1901] Spot Hidden Noble would have an additional 15% chance to spot a hidden noble.

If the who argument is not given, seek will seek for any hidden nobles present. In this case there will be one-half the normal chance of finding a hidden noble during each of the seven days.

If a unit found with seek moves, the contact information is lost, and the unit again becomes inaccessible.

Hidden units are not shown in the location report, even if their presence has been revealed through contact or seek.

Related orders: ⟨undefined⟩ [contact], page ⟨undefined⟩

sell

sell item qty price [have-left]

 • time: 0 days
 • priority: 1

Sell quantity of the given item for max-price each.

If quantity is zero, any pending sell order for the indicated item will be canceled.

Only one pending sell order may be active for each item at a time. If two sell orders are issued for the same item, and the first has not yet completed, the second sell order will replace the first.

Buys and sells are only matched in cities, where markets may be found to execute trades. However, buy and sell may be issued anywhere.

Trades only occur at the end of the day. Therefore, you must stay in a location for at least one day to execute a trade. If you step into a market, issue a buy or sell, and then leave immediately, no trade will occur. To be certain to stay at least one day, you may issue a ⟨undefined⟩ [wait], page ⟨undefined⟩ command.

Example:

    > sell 79 3 5
    Try to sell three iron [79] for 5 gold each.

If there is a buyer of iron in the city, buying for 5 gold or more, the trade will be executed:

    1: Sold three iron [79] for 5 gold.

Otherwise, the sell will be placed on the noble’s pending trades list:

    Pending trades:
    trade    price    qty   item
    -----    -----    ---   ----
     sell        5      3   iron [79]

At the end of the turn, the sell order will appear in the city’s market report:

    Market report:
    trade    who    price    qty   item
    -----    ---    -----    ---   ----
     sell   2019        5      3   iron [79]

A trade will match between a buyer and a seller if the buyer’s price is greater than or equal to the seller’s price. The actual price that the trade will be made at is the seller’s price.

Have-left is an amount of the item that noble will keep and not sell. For instance, if the unit issued ‘sell 79 10 15 5’, this would mean "Sell ten iron for 15 gold each, but keep 5 iron for myself." If there noble only have ten iron, this order would put five up them up for sale.

split

split bytes lines

 • time: 0 days
 • priority: 1

Split large turn reports before mailing them.

Some mail networks and online services have restrictions on how large an email message may be. Some typical limits which may be encountered are 5,000 lines, 50k, and 64k. Since Olympia turn reports can be arbitrarily large, players with such restrictions on their incoming email may have trouble getting their reports. Reports may be truncated, or not delivered at all.

The split order asks Olympia to divide turn reports into several pieces before mailing them. The lines parameter, if nonzero, gives a maximum number of lines to allow in a turn report section. The bytes parameter, if nonzero, gives a maximum number of bytes to allow in a turn report section. Reports will be split at the first of these limits reached.

Note that headers added by the operating system add to the size of a message after it has been mailed. Leaving 2000 bytes or 20 lines leeway before your desired limit is a good idea.

For example, suppose your incoming messages are truncated at 50k bytes. Issue the order:

    split 0 48000

This will cause messages to be split after the first 48,000 bytes. Lines are ignored for the purposes of splitting.

If your messages are rejected if they are larger than 64k or 5,000 lines, whichever is less, issue the order:

    split 4980 62000 # note 20 line / 2000 byte margin

To turn off report splitting, issue the order:

    split 0 0

Reports will be split on line boundaries (lines will not be split in the middle). "[part n of total]" will be added to the Subject: line of each section. For example,

    Subject: Olympia turn 1 report [part 1 of 3]

The SPLIT order is parsed by the order scanner, so it takes effect immediately, rather than when the next turn runs. Suppose you received a turn report which was truncated because of a mail size limit of 32k. You could request a resend of your turn report, split into sections, with the following message to the order scanner:

    begin <your player id>
    split 0 30000 # split at 32k, with 2k leeway
    resend # resend my last report to me
stack

 stack noble
 • time: 0 days
 • priority: 1

Stack beneath the specified noble. If the noble is already stacked beneath another unit, the noble will first unstack before stacking with the new noble.

Example: Two nobles, Feasel and Osswid, are in the same location:

    Seen here:
       Osswid [5499]
       Feasel the Wicked [502]

If Feasel issued the order ‘stack 5499’, he would stack beneath Osswid:

    Seen here:
       Osswid [5499], accompanied by:
          Feasel the Wicked [502]

Nobles may also stack beneath garrisons and controlled units.

Any attempt to stack will fail if you do not have "admit" permissions (⟨undefined⟩ [admit], page ⟨undefined⟩).

stop

 stop

Interrupt a command left running at the end of a turn. Must be the first order queued for a unit.

For example, suppose on day 28 a unit begins an explore command. Explore takes seven days, so as of the end of the turn, it will still have four days left before it completes. The order template for the unit will list the executing command:

    unit 5499  # Osswid
       # > explore (executing for four more days)

The player may interrupt this command with the stop order, and begin unit 5499 executing a new set of orders (say, beginning with a move) on day 1 of the next turn:

    unit 5499  # Osswid
       # > explore (executing for four more days)
       stop
       move
       ...

(Remember that ‘#’ begins a comment in the orders, from the ‘#’ to the end of the line).

The stop order will interrupt the executing explore before day 1 begins on the next turn. Some commands, such as collect, will give partial results; the unit will have whatever it has already collected. Other orders, such as explore, give no partial benefit, and the three days already invested in exploration will be wasted.

If a noble executing a move order is interrupted with stop, the noble will return to the original location, as if no travel had taken place.

study

study skill [study points flag]

 • time: 7 days
 • priority: 3

Study the given skill for one week.

If the skill has been studied before, learning will continue. Otherwise, the noble must meet the requirements necessary to begin learning the skill.

To begin study of a skill requires the following:

• Payment of 100 gold to begin study.

• Payment of Noble Points to learn advanced skills

• For category skills, a source of instruction. − A noble teaching the skill. − A city teaching the skill. − A book of instruction in the skill. − Study points.

To use your study points (see ⟨undefined⟩ [Study Points], page ⟨undefined⟩) to learn a skill, you must provide the study point flag after the skill to be studied. Just provide a "1" to indicate that you wish to use study points:

                    study 1000 1

This prevents you from inadvertantly spending your study points (e.g., when you thought you were in a tower with a teacher, but he never showed up).

Note that multiple study orders during a turn yield diminishing results: − First week: 7/7 − Second week: 4/7 − Third week: 2/7 − Fourth week: 1/7

See ⟨undefined⟩ [Skills], page ⟨undefined⟩ for more detailed information about study.

take

take who item [qty] [have-left]

 • time: 0 days
 • priority: 1

Same as get. See ⟨undefined⟩ [get], page ⟨undefined⟩ for a description.

tax

 tax where [market | province | weight | men] rate
 • time: 0 days
 • priority: 1

The tax command is used to set the tax rate in a province, market or city. For all forms of the tax command, the first argument to the command is the location to be taxed, the second argument is one of the keywords "market", "weight", "men" or "province", and the third argument is the tax rate. In all cases you must be the ruler of a location to impose a tax. (When taxing a province or a market, any noble in the ruler’s faction may issue the tax command.)

For markets, the tax rate may be from 0 to 25 percent. That percentage will be deducted from the sale price of all goods sold in the market and placed in the city garrison. To tax a city market, give the city id as the first argument to the tax command. To tax a market in a trader’s guild, give the id of the guild as the first argument to the tax command. The tax rate for a market will not change until the end of the month.

For provinces, the tax rate may be from 0 to 10 percent. Each month, gold equal to that percentage of the population is collected from the peasants. The gold produced will be placed into the local garrison.

City rulers may also impose a tax on the men and/or goods entering a city.

The command tax <city> men <rate> imposes a tax on the men entering <city>. The rate is the number of gold that must be paid for each 100 men entering the city. Fractional amounts are truncated. For example, a noble with 213 men in his stack entering a city with a man tax rate of 10 gold per 100 men would incur a tax of 21 gold.

The command tax <city> weight <rate> imposes a tax on the goods entering <city>. The rate is the number of gold that must be paid for each 1000 weight of goods entering the city. The weight of goods includes everything a stack carries except the weight of the men and beasts themselves. Fractional amounts are truncated. For example, a stack with 11714 weight in goods entering a city with a weight tax rate of 25 gold per 1000 weight would incur a tax of 292 gold.

teach

teach skill [days]

 • time: 7 days default
 • priority: 4

The teach command is used to provide instruction in skill to other nobles. The first argument to this command is the skill to teach. The second, optional, argument is the number of days to teach. This defaults to 7 days.

Teaching must be done in a tower. Only one class can take place at a time in a tower. The teacher is the first noble in the tower issuing a "teach" command. The first five nobles in the tower who would benefit from studying that subject are his students.

For example, suppose that on a particular day, the following nobles are in a tower executing the indicated commands:

       Newark Academy [z38], tower, defense 20, owner:
          Fenris [re7]                      (study 1001)  *
          Kosar the Indefectible [aw0]      (study 1193)
          Captain McCook [ea0]              (study 1000)   *
          Oleg [o1eg]                       (teach 1000)   +
          Dr. Pangloss [ac6y]               (study 1005)  *
          Law Netexus [au0]                 (study 1000)   *

          Alion Krysaka [e1y]               (study 1100)
          Dr. Pain [em28]                   (teach 1100)
          S.K. Renta [ep76]                 (study 1000)   *
          Armweak [em78] (study 1000)

The plus in this example indicates the teacher, the asterisks indicate students who benefit from his teaching.

The only teacher in this tower is Oleg. Dr. Pain is not considered a teacher, because he is not the first teacher in the tower. Fenris benefits from Oleg’s teaching because [1001] is a subskill of [1000] and he is one of the first five students. Kosar does not benefit, because 1193 is not a subskill of [1000]. Likewise Alion does not benefit because he is studying [1100]. Note that Armweak does not benefit because, although he is studying the proper subject, he is not one of the first five students in the tower.

You cannot teach [3500] Basic Religion or [3300] Advanced sorcery.

terrorize

terrorize who severity

 • time: 7 days
 • priority: 3

Terror may be used to convert prisoners to a faction, or to reinforce fear loyalty among the faction’s own units. The noble to be terrorized must be stacked beneath the terrorizing unit.

When used on one’s own units, terrorize will establish or increase fear loyalty. Nobles of contract loyalty will be converted to fear. The severity amount will be added to the noble’s fear rating. For instance, a fear-4 noble terrorized at severity 10 will become fear-14.

The severity indicates how much the noble will be abused to gain fear. The noble will suffer a severity hit to health as a result. Thus, there is a danger that the noble will fall ill with increasing amounts of terror. See ⟨undefined⟩ [Health], page ⟨undefined⟩ for details about illness.

Prisoners may also be terrorized. Prisoners of contract or fear loyalty lord have severity percent chance of swearing fealty to the terrorizing unit. Oathbound units can not be induced to convert through terror.

Note that "refugee" nobles may have any sort of loyalty (oath, contract, or fear).

Thus, the more severe the terrorization, the greater the chance that a prisoner will swear to the terrorizer, but also the greater the victim’s wounds. Prisoners have been known to die as a result of over-zealous conversion attempts.

Nobles belonging to other player factions may not be terrorized unless they are being held prisoner.

Units of fear loyalty may not use the terrorize order.

train

train kind [number]

• time: variablen
• priority: 3

Attempt to produce the given kind of men through training. One day is required to train each man.

See ⟨undefined⟩ [Training], page ⟨undefined⟩ for more information.

num  kind            skill  input man       input item        where
---  -----------     -----  --------------  ----------------
------
11  worker           none  peasant [10]
36  skirmisher       1100  peasant [10]
12  soldier          1108  skirmisher [36]
28  fanatic       2635  soldier [12]
29  ninja       2733  soldier [12]
13  archer           1902  soldier [12]    longbow [72]
14  knight           1190  hvy. foot [37]  warmount [53]
15  paladin       1191  knight [14]         guild
16  pikeman          1100  soldier [12]    pike [75]
19  sailor           1001  peasant [10]
20  lt. foot         1105  soldier [12]    longsword [74]
37  hvy. foot        1192  lt. foot [20]   plate armor [73]
21  lt. xbowman      1100  soldier [12]    lt. crossbow [85]
38  hvy. xbowman     1100  soldier [12]    hvy. crossbow [40]
22  elite archer     1902  archer [13]     elvish arrow [39]
24  pirate           1105  sailor [19]     longsword [74]    ship
41  horse archer     1190  archer [13]     riding horse [52]
43  cavalier       1190  soldier [12]    mithril [82], warmount[53]

Examples: To train 5 peasants into 5 workers:

train 11 5              # takes 5 days

To train 5 peasants into 5 sailors (requires knowledge of Sailing [1001]):

train 19 5              # takes 5 days

To train 3 soldiers into 3 lt. foot (requires knowledge of Train Swordsmen [1105], and a longsword [74] for each soldier to be trained):

train 20 3              # takes 3 days

ungarrison

ungarrison [garrison]
• time: 1 day
• priority: 3

Disband the garrison in the current province. The noble removing the garrison must be a member of a faction which has rulership over the province.

If a specific garrison is given as an argument to ungarrison, only that garrison will be disbanded. This can be useful as a safety check, in case a failed move order might cause the wrong garrison unit to be disbanded.

The garrison’s men will be transferred to the noble issuing ungarrison.

Examples:

   ungarrison 5611         # remove Garrison [5611] from this province
   ungarrison              # remove any garrison from this province
unload

unload
• time: 0 days
• priority: 3

Unload all passengers from a ship. May only be used by the captain of a ship. Every noble not in the captain’s stack will disembark from the ship.

Passengers may only be unloaded while the ship is docked in a port. They will not jump off a ship at sea.

unstack

unstack [who]
• time: 0 days
• priority: 1

When used without an argument, unstack removes the noble from being stacked beneath another noble:

Before:

   Osswid the Constructor [5499], accompanied by:
      Feasel the Wicked [502]
      Captain McCook [2019], with five workers

502: > unstack

After:

Osswid the Constructor [5499]
Feasel the Wicked [502], accompanied by:
   Captain McCook [2019], with five workers

A stack leader may give an argument to unstack, specifying a unit to be dropped from the stack:

Before:

Osswid the Constructor [5499], accompanied by:
   Feasel the Wicked [502]
   Captain McCook [2019], with five workers
5499: > unstack 502

After:

Osswid the Constructor [5499]
Feasel the Wicked [502], accompanied by:
   Captain McCook [2019], with five workers

Note that in this example [2019] is stacked beneath [502], not [5499]. That is why Captain McCook remains with Feasel when Feasel unstacks. See ⟨undefined⟩ [Stacking], page ⟨undefined⟩ for more details on hierarchical stacks.

A special case of unstack is invoked if the noble specifies itself as the argument. In this case, the noble will drop out of the stack alone, leaving behind both those stack above as well as below:

Before:

Osswid the Constructor [5499], accompanied by:
   Feasel the Wicked [502], accompanied by:
      Captain McCook [2019], with five workers
502: > unstack 502

After:

Osswid the Constructor [5499], accompanied by:
   Captain McCook [2019], with five workers
Feasel the Wicked [502]

The effect on the stack in this case is the same as if Feasel had died or been captured in battle.

More examples: Before:

Osswid the Constructor [5499], accompanied by:
   Feasel the Wicked [502]

After Feasel unstacks:

Osswid the Constructor [5499]
Feasel the Wicked [502]

If other nobles were stacked beneath Feasel, they would remain with him. Before:

 Osswid the Constructor [5499], accompanied by:
    Feasel the Wicked [502], accompanied by:
       Candide [1269]

If Candide were stacked beneath Feasel, and Feasel unstacked, then:

 Osswid the Constructor [5499]
 Feasel the Wicked [502], accompanied by:
    Candide [1269]

On the other hand, if Candide were stacked beneath Osswid, then:

 Osswid the Constructor [5499], accompanied by:
    Candide [1269]
 Feasel the Wicked [502]
use

use skill [arguments ...]

• time: varies
• priority: 3

Make use of a skill or spell that the noble knows. Sub-skills often have a function which is invoked through use.

Category skills do not have direct use functions. For example, one could use the Sailing [1001] skill to sail a ship:

 use 1001 west

But no use function exists for the category skill, Shipcraft [1000].

email

email new email address for player list

Set a different email address to appear in the player list than the one the turn reports are sent to.

For example:

 email "skrenta@shadow.com"
 vis_email "skrenta@pbm.com"

Would cause turn reports to be sent to skrenta@shadow.com, while the player list shows skrenta@pbm.com as player 501’s email address:

 501    House of Osswid
        Rich Skrenta <skrenta@pbm.com>

wait

wait conditions

 • time: varies
 • priority: 1

Wait for one of the the specified conditions to occur. The first condition which is satisfied will cause the wait command to terminate. If one of the specified wait conditions is true when the wait order is issued, no wait will take place, and the order will complete in zero time.

The following conditions may be waited on:

alone Wait until the noble is alone in a location, meaning that his stack is the only stack in the location. (See ⟨undefined⟩ [guild], page ⟨undefined⟩.)

clear Wait until it is not foggy, windy or rainy.

day n Wait until the day of the month is n, or later. For example, "wait day 15" waits until it is the 15th of the month or later.

turn n Wait until it is turn n, or later.

ferry n Wait for the ship n to sound its horn with the ferry order. Units waiting to board a ferry should use wait ferry instead of wait ship. This will allow passengers on the ship time to disembark. When passenger unloaded is complete, the captain will sound the ship’s horn, and those wishing to board may proceed. Otherwise, those waiting to board may try to get on the ship before the current load of passengers has left, causing board orders to fail because of ship weight limits.

flag f [n] Wait until noble n, or a noble belonging to player n, signals with flag f. n defaults to the player issuing the flag order. If n is zero, wait until any noble signals with flag f. See ⟨undefined⟩ [flag], page ⟨undefined⟩ for more details

fog Wait until it is foggy in the noble’s province.

gold n Wait until the noble has at least n gold

item n q Wait until the noble has at least quantity q of item n

loc n Wait until arrival at the given location.

not Invert the sense of the following condition.

owner Wait until the noble is the first noble in the location list.

rain Wait until it is raining in the noble’s province. ship n Wait until the specified ship arrives in port.

shiploc n Many players have been bitten by using wait loc while they’re on a ship to wait for the ship’s arrival at a particular port. Don’t do this! wait loc waits for the noble to arrive at a particular location. If the noble is on a ship, the ship is the location the noble is in, not the ship’s location. Use wait shiploc to wait for the ship’s arrival at a particular place.

stack n Wait until the noble and unit n are members of the same stack. Position within the stack is not important; The condition is true whenever n and the noble share a common stack leader.

time n Wait for n days to elapse

top Wait until the unit becomes a stack leader. This could happen if the units’s current stack leader unstacked or promoted the unit.

unit n Wait until the noble and the specified unit are in the same location

wind Wait until it is windy in the noble’s province.

teacher n Wait until a source of specific instruction for the specified skill is available. This can be a book or a teacher. Note that if you wait for a subskill (e.g., 3401), this wait condition will not be satisfied by someone teaching the skill school (e.g., 3400).

For example:

    1: > wait time 1
    1: Wait finished: one day has passed.
    2: > wait time 10 day 5
    5: Wait finished: today is day 5.

The first condition to become true will end the wait. This can be used to specify the maximum amount of time that the noble will wait for something.

For instance, suppose we want to wait for [5499] to show up, but we don’t want to wait past day 15.

    > wait unit 5499 day 15

This order will finish on day 15 or as soon as 5499 shows up, which ever comes first. If 5499 is already here when the order is issued, no wait will take place. Likewise, if it is already past day 15, no wait will take place.

Keep in mind that the order priority mechanism eliminates the need to use waits in many cases. Consider synchronization in the following example:

Feasel [502] arrives at Osswid’s [5499] location. Feasel gives Osswid some gold, and then leaves. Osswid departs after receiving the gold.

    unit 502   # Feasel
      move ...
      pay 5499 100
      move ...
    unit 5499  # Osswid
      wait unit 502
      move ...

The orders above will work; no additional waits are needed. After the wait unit becomes true, both 502’s pay and 5499’s move are ready to execute. Since pay has a lower priority than move, pay will happen first.

This technique also works with stack and give, since they also have a lower priority than move.

The NOT keyword inverts the sense of the condition which follows. For example:

    > wait not rain time 7

Will wait until it is not raining, or for seven days, whichever comes first.

    > wait not rain not wind

Will wait until it is not raining or not windy.

wood

wood [number of wood] [days]

 • time: as given
 • priority: 3

The wood command is an abbreviation for "collect 77 number of wood days". See ⟨undefined⟩ [collect], page ⟨undefined⟩.

yew

yew [number of yew] [days]

 • time: as given
 • priority: 3

The yew command is an abbreviation for "collect 68 number of yew days". See ⟨undefined⟩ [collect], page ⟨undefined⟩.

Appendix E Appendix - Skill Dictionary

This appendix lists all the available skills in Olympia: The Age of Gods.

Shipcraft [1000]

1001   Sailing                            two weeks
1002   Build Hull                         two weeks
1095   Add ram to ship                    four weeks
1004   Fishing                            two weeks
1005   Add Sails to Ship                  two weeks
1006   Add Rowing Ports to Ship           two weeks
1096   Remove Sails from Ship             two weeks
1097   Remove Rowing Ports from Ship      two weeks
1098   Remove Ram from Ship               two weeks
1090   Add Keel to Ship                   two weeks
1091   Add Fortification to Ship          two weeks
1092   Remove Keel from Ship              two weeks
1093   Remove Fortification from Ship     two weeks
1094   Navigate Deep Seas                 two weeks

Combat [1100]

1102   Fight to the death                 four weeks
1194   Construct catapult                 two weeks
1105   Train Swordsmen                     two weeks
1106   Weaponsmithing                     two weeks
1107   Control Men in Battle              four weeks
1108   Instill Combat Discipline          two weeks
1195   Use beasts in battle               four weeks
1131   Combat Tactics                     four weeks
1132   Defense Tactics                    four weeks
1190   Train Mounted                      two weeks
1191   Train Paladin                      two weeks
1192   Train Armored Troops               two weeks

Stealth [1200]

1201   Petty thievery                     two weeks
1202   Determine inventory of noble   two weeks
1203   Determine skills of noble      two weeks
1204   Determine noble's lord         two weeks
1205   Conceal faction                    four weeks
1230   Learn of richest nearby noble      two weeks
1231   Torture prisoner                   four weeks
1232   Conceal identity of trader         two weeks
1233   Sneak into structure               four weeks
1290   Conceal self                       four weeks

1291   Conceal Nation                     four weeks
1292   Assassinate Noble                  two weeks

Persuasion [1300]

1301   Bribe noble                        two weeks
1302   Raise peasant mob                  two weeks
1303   Rally peasant mob                  four weeks
1304   Train angry peasants               two weeks
1330   Persuade oathbound noble           four weeks
1331   Incite mob violence                four weeks
1332   Summon wild men                    four weeks
1333   Persuade wild men to remain        four weeks
1390   Encourage Population Growth        two weeks

Construction [1400]

1401   Construct siege tower              two weeks
1402   Stone quarrying                    two weeks
1490   Build City                         two weeks
1491   Fortify Castle                     three weeks
1492   Strengthen Castle Walls            three weeks
1493   Add Moat to Castle                 three weeks

Alchemy [1500]

1501   Brew healing potion                four weeks
1502   Brew potion of death               four weeks
1503   Collect rare elements              two weeks
1530   Extract venom from ratspider       four weeks
1531   Make potion of slavery             four weeks
1532   Record skill on scroll             four weeks
1533   Turn lead into gold                four weeks
1590   Make potion of weightlessness      four weeks
1591   Make potion of fiery death         four weeks

Forestry [1600]

1601   Construct battering ram            two weeks
1602   Harvest lumber                     two weeks
1603   Harvest yew                        two weeks
1605   Harvest opium                      two weeks
1604   Collect rare foliage               two weeks
1630   Harvest mallorn wood               two weeks
1631   Improve opium production           two weeks

Mining [1700]

1701   Mine iron                          two weeks
1702   Mine gold                          two weeks
1703   Build Mine Shaft                   three weeks
1730   Mine mithril                       two weeks

1731   Mine gate crystal                  two weeks
1790   Add Wooden Shoring                 three weeks
1791   Add Iron Shoring                   three weeks

Trading [1800]

1801   Build Wagons                       two weeks
1802   Increase Demand                    two weeks
1803   Decrease Demand                    two weeks
1804   Increase Supply                    two weeks
1805   Decrease Supply                    two weeks
1830   Smuggle Goods                      two weeks
1831   Smuggle Men                        two weeks
1832   Avoid Taxes                        two weeks
1890   Hide Money                         two weeks
1891   Hide Item                          two weeks

Ranger [1900]

1901   Spot Hidden Noble                  two weeks
1902   Archery                            four weeks
1903   Find Food in Wilderness            two weeks
1930   Catch wild horses                  two weeks
1931   Train wild horse to riding horse   two weeks
1932   Train wild horse to warmount       four weeks
1990   Protect Noble                      two weeks

Church of Eres [2000]

2030   Fierce wind                        four weeks
2031   Dissipate storm                    four weeks
2032   Renew storm strength               four weeks
2033   Storm fury                         four weeks
2034   Seize control of storm             four weeks
2035   Fog of death                       four weeks
2036   Summon rain                        four weeks
2037   Summon wind                        four weeks
2038   Summon fog                         four weeks
2039   Direct storm                       four weeks
2040   Summon Water Elemental             four weeks

Church of Anteus [2100]

2130   Find Mountain Trail                four weeks
2131   Obscure Mountain Trails            four weeks
2132   Improve Mining                     four weeks
2133   Conceal Mine                       four weeks
2134   Protect Mine                       four weeks
2135   Bless Building                     four weeks
2136   Weaken Building                    four weeks
2137   Boulder Trap                       four weeks

Church of Dol [2200]

2231   Bird spy                           four weeks
2232   Capture beasts           four weeks
2233   Breed beasts                       four weeks
2234   Detect Beasts                      four weeks
2235   Snake Trap                         four weeks

Church of Timeid [2300]

2330   Find Forest Trail                  four weeks
2331   Obscure Forest Trails              four weeks
2332   Improve Forestry                   four weeks
2333   Reveal Forest                      four weeks
2334   Create Dense Thicket               four weeks
2335   Create Deadfall                    four weeks
2336   Recruit Elves                      four weeks

Church of Ham [2400]

2430   Enchant City Guards                four weeks
2431   Urchin Spy                         four weeks
2432   Draw Crowds                        four weeks
2433   Arrange Mugging                    four weeks
2434   Receive vision                     four weeks
2435   Create Holy Instruction            four weeks
2436   Shroud location from magical scry  two weeks
2437   Protection from Aura Blast   four weeks

Church of Kireus [2500]

2530   Improve Quarrying                  four weeks
2531   Improve Smithing                   four weeks
2532   Edge of Kireus                     four weeks
2533   Create Mithril                     four weeks
2534   Set Quicksand Trap                 four weeks

Church of Halon [2600]

2630   Calm Angry Peasants                four weeks
2631   Improve Charisma                   four weeks
2632   Mesmerize Crowd                    four weeks
2633   Improve Taxes                      four weeks
2634   Guard Loyalty                      four weeks
2635   Instill Fanaticism                 four weeks
2636   Create Holy Instruction            four weeks

Church of Domingo [2700]

2730   Find Hidden Location               four weeks
2731   Conceal Wilderness Location        four weeks
2732   Mists of Concealment               four weeks
2733   Train Ninjas                       four weeks

Magic [2800]

2801   Meditate                           two weeks
2802   Perform common tasks for gold      two weeks
2803   Appear common                      two weeks
2804   View current aura level of others  four weeks
2830   Speed casting of next spell        four weeks
2831   Reveal abilities of another mage   four weeks
2832   Tap health for aura                four weeks
2833   Modern magic script                four weeks
2834   Shroud abilities from scry         four weeks
2835   Detect ability scry                four weeks
2836   Dispel ability shroud              four weeks
2837   Advanced meditation                four weeks
2838   Hinder meditation                  four weeks
2839   Save speeded cast                  four weeks
2840   Lightning bolt                     four weeks
2841   Create Dirt Golem                  two weeks

Scrying [2900]

2901   Scry location                      four weeks
2930   Shroud location from magical scry  two weeks
2931   Dispel location shroud             four weeks
2932   Ciphered writing of Areth-Pirn     four weeks
2933   Create magical barrier             four weeks
2934   Remove magical barrier             four weeks
2935   Locate noble                       four weeks
2936   Detect location scry               four weeks
2937   Farcasting                         four weeks
2938   Save farcast state                 four weeks
2939   Banish undead                      four weeks
2940   Foresee Defense                    four weeks

Gatecraft [3000]

3001   Detect gates                       two weeks
3002   Jump through gate                  two weeks
3030   Teleport                           two weeks
3031   Seal gate                          two weeks
3032   Unseal gate with key               two weeks
3033   Notify if gate unsealed            two weeks
3034   Forcefully unseal gate             four weeks
3035   Reveal gate key                    four weeks
3036   Notify of gate jumps               two weeks
3039   Drain Mana                         four weeks
3037   Language of the Ancients           four weeks

3038   Reverse jump through gate          four weeks

Artifact construction [3100]

3101   Detect Artifact   two weeks
3102   Reveal Artifacts   two weeks
3103   Destroy artifact                   two weeks
3104   Deep Identify   four weeks
3130   Forge auraculum                    four weeks
3131   Mutate Artifact           two weeks
3132   Conceal Artifacts   two weeks
3133   Obscure Artifact           two weeks
3134   Forge magical weapon               four weeks
3135   Forge magical armor                four weeks
3136   Forge magical bow                  four weeks
3137   Remove Obscurity   two weeks
3138   Arcane symbols                     four weeks
3139   Forge palantir                     four weeks

Necromancy [3200]

3201   Summon ghost warriors              four weeks
3202   Raise undead                       four weeks
3203   Aura blast                         four weeks
3230   Runes of Evil                      four weeks
3231   Summon demon lord                  four weeks
3232   Renew demon bond                   four weeks
3233   Banish demon lord                  four weeks
3234   Eating of the dead                 four weeks
3235   Absorb aura blast                  four weeks
3236   Raise soldiers                     four weeks
3237   Create Flesh Golem                 three weeks
3238   Transcend death                    four weeks, 1 NP req'd

Advanced sorcery [3300]

3330   Trance                             four weeks
3331   Teleport items                     four weeks
3332   Fireball                           four weeks
3333   Create Iron Golem                  four weeks

Heroism [3400]

3401   Swordplay   two weeks
3402   Defense   two weeks
3403   Personal Fight to the Death   two weeks
3480   Avoid Wounds   two weeks
3481   Avoid Illness   two weeks
3482   Improved Recovery   two weeks

3483   Survive fatal wound   two weeks
3484   Forced March   two weeks
3485   Extra Melee Attacks   two weeks
3486   Extra Missile Attacks   two weeks
3487   Acute Senses   two weeks
3488   Improved Explore   two weeks
3489   Uncanny Accuracy   two weeks
3490   Blinding Speed   two weeks

Basic Religion [3500]

3501   Heal                               two weeks
3502   Last Rites                         two weeks
3503   Create Holy Symbol                 two weeks
3504   Dedicate Temple                    two weeks
3505   Devotional Prayer                  two weeks
3506   Bless Follower                     two weeks
3507   Gather Holy Plant                  two weeks
3508   Create Holy Instruction            two weeks
3509   Proselytise                        two weeks
3530   Resurrect Dead Noble               four weeks
3531   Protection from Aura Blast         two weeks
3532   Hinder Meditation                  two weeks
3533   Vision of Holy Terrain             two weeks

, Appendix - Skill Dictionary, Appendix - Skill Dictionary,]

E.1 Shipcraft [1000]

Time to learn: 7 days.

All skills concerning ocean travel fall under this category. Shipcraft encompasses the building and repair of ships, training of sailors, and navigation at sea.

The following teachable skills may be studied directly once Shipcraft is known:

num    skill                              time to learn
---    -----                              -------------
1001   Sailing                            two weeks
1002   Build Hull                         two weeks
1004   Fishing                            two weeks
1005   Add Sails to Ship                  two weeks
1006   Add Rowing Ports to Ship           two weeks

The following guild skills may be studied by guildmembers:

num    skill                              time to learn
---    -----                              -------------
1090   Add Keel to Ship                   two weeks
1091   Add Fortification to Ship          two weeks
1092   Remove Keel from Ship              two weeks
1093   Remove Fortification from Ship     two weeks
1094   Navigate Deep Seas                 two weeks
1095   Add ram to ship                    four weeks
1096   Remove Sails from Ship             two weeks
1097   Remove Rowing Ports from Ship      two weeks
1098   Remove ram from ship               four weeks

, Shipcraft [1000], Shipcraft [1000],]

E.1.1 Sailing [1001]

Usage: USE 1001 <direction or location> Or: SAIL <direction or location> Requires: noble must be ship’s captain Time: varies with route taken Time to learn: 14 days. Time to use: Variable.

Pilot a ship in the given direction, or to the specified location. If a location is specified, there must be a direct route to it from the current location of the ship.

Compass points may be abbreviated: N, S, E, W.

The noble issuing the Pilot ship order must occupy the first position inside the ship (the noble must be the captain of the ship).

The captain must have a crew (sailors or pirates) to assist in the handling of the ship. To sail, one sailor is required for each sail. To row, each rowing port can accommodate up to 4 rowers.

Overloaded ships do not sail well. They may take longer than expected to travel across routes, and are almost certain to capsize and sink in a storm. Captains should be careful not to leave port in a vessel carrying too much cargo.

For convenience, one may use the SAIL order rather than "USE 1001". Knowledge of Pilot ship [1001] is still required, however.

, Sailing [1001], Shipcraft [1000],]

E.1.2 Build Hull [1002]

Usage: BUILD ship "Name of ship" size Requires: at least three workers [11]

10 wood [77] per unit size 100 worker-days per unit size

Time to learn: 14 days.

To begin construction of a ship, the shipbuilder should unstack from beneath other nobles and issue a BUILD order. One-fifth of the lumber is required to start construction. The rest is deducted as progress on the ship is made (the second fifth at 20% completion, the third fifth at 40% completion, etc.)

The shipbuilder and his workers will be placed inside the new ship.

As soon as the required number of worker-days has been invested in construction, the ship will be christened and declared seaworthy.

To resume construction of a partially completed ship, first enter the ship, then issue "build ship".

Ships may only be built in locations with a passable route to an ocean, such as a port city.

, Build Hull [1002], Shipcraft [1000],]

E.1.3 Navigate Deep Seas [1094]

Usage: Automatic as needed Time to learn: 14 days. Time to use: Variable.

This skill permits a noble to captain a ship in deep seas (out of sight of land) without danger of getting lost, delayed, capsized, etc. Knowledge of Sailing [1001] is required to learn this skill.

, Navigate Deep Seas [1094], Shipcraft [1000],]

E.1.4 Add ram to ship [1095]

Usage: USE 1095 Time to learn: 28 days. Time to use: 10 days. Requires: three workers [11] (not consumed) Requires: three wood [77] Requires: one iron [79]

Requires: three workers [11] 3 wood [77] 1 iron [79]

Time: 10 days

A ship fitted with a ram is a most effective ocean fighting vessel. The ram allows a ship galley to crush the sides of enemy ships.

To use this skill, the noble must enter the ship and issue USE 1095. The ship will be fitted with an iron-tipped ram. The noble need not be in control of the ship in order to perform the modification.

, Add ram to ship [1095], Shipcraft [1000],]

E.1.5 Fishing [1004]

Usage: FISH [number of fish] [days] Requires: at least one sailor Time to learn: 14 days.

Unit must be on a ship on the ocean in order to fish. Each day spent fishing there is a 50% chance that the ship will come across a school of fish. If fish are found, each sailor will produce one fish that day. Provinces generate 50 fish per month which may be obtained by fishing.

Examples:

fish # obtain as many fish as possible fish 25 # stop after getting 25 fish fish 0 5 # fish for five days fish 25 5 # stop after five days, or 25 fish, whichever # comes first

In all cases, fishing will cease as soon as the stocks in the ocean province have been exhausted.

, Fishing [1004], Shipcraft [1000],]

E.1.6 Add Sails to Ship [1005]

Usage: USE 1005 Time to learn: 14 days. Time to use: 3 days. Requires: ten workers [11] (not consumed) Requires: one wood [77]

This skill adds one sail to a ship.

Each unit of hull may mount a maximum of 3 sails. Each sail requires one sailor to tend it, up to a maximum of 100 sailors. A sail uses up 250 units of a ship’s capacity.

Sailing a ship at normal speed requires two fully-staffed sails per unit hull. For example, sailing a 3 unit hull ship at normal speeds requires that the ship have 6 sails and at least 8 sailors (2 to man the hull and 1 each for the 6 sails).

Each additional fully-staffed sail per unit hull cuts a ship’s travel time by two days. For example, a 3 unit ship with 9 fully manned sails would cut its travel time by two days, crossing a 6 day ocean province in just 4 days.

Conversely, if a ship has less than two full-staffed sails per unit hull, it will add two days of travel time. A ship with less than one sail per two units hull cannot move at all.

, Add Sails to Ship [1005], Shipcraft [1000],]

E.1.7 Add Fortification to Ship [1091]

Usage: USE 1091 Time to learn: 14 days. Time to use: 7 days. Requires: ten workers [11] (not consumed) Requires: five wood [77]

This skill adds one fortification to a ship.

Each unit of fortification added to ship uses 250 units of the ship’s carrying capacity. A fortified ship provides cover for

(units fortification / units hull) * 10

men; each of these men receives a defensive bonus in combat of (units fortification / units hull) * 5 For example, a 3 unit hull carrying 7 units of fortification would provide

a defensive bonus of +11 to the first 22 fighters on board the ship.

, Add Fortification to Ship [1091], Shipcraft [1000],]

E.1.8 Add Rowing Ports to Ship [1006]

Usage: USE 1006 Time to learn: 14 days. Time to use: 3 days. Requires: ten workers [11] (not consumed) Requires: two wood [77]

This skill adds one rowing port to a ship.

Each rowing port accommodates 4 oarsmen. Each unit of hull may have a maximum of 3 ports, and each newly-constructed hull comes with one rowing port.

Rowing a ship at normal speed requires one fully-staffed rowing port per unit hull. For example, rowing a 3 unit hull ship at normal speeds requires that the ship have 3 rowing ports and at least 14 sailors (2 to man the hull and 4 each for the 3 rowing ports).

Each additional fully-staffed rowing port per unit hull cuts a ship’s rowing travel time by one day. For example, a 3 unit ship with 6 fully manned rowing ports would cut its travel time by one day, crossing a 6 day ocean province in just 5 days.

Conversely, if a ship has less than one full-staffed rowing port per unit hull, it will add one day of travel time. A ship with less than one oarsman per unit hull cannot move at all.

, Add Rowing Ports to Ship [1006], Shipcraft [1000],]

E.1.9 Add Keel to Ship [1090]

Usage: USE 1090 Time to learn: 14 days. Time to use: 5 days. Requires: ten workers [11] (not consumed) Requires: one iron [79]

This skill adds one unit of keel to a ship.

Each unit of hull may mount one unit of keel. For example, a 3 unit hull could mount up to 3 units of keel.

Due to the keel’s heavy weight, each unit of keel added to a hull reduces the capacity by 100 units weight.

Appendix E: Appendix - Skill Dictionary 186

Keel protects a ship against storm damage. Ships caught in a coastal storm have the following chance of sustaining damage:

(15 - (units keel / unit hull) * 30)% Thus a ship with at least two units of keel for each three units of hull will

(under normal circumstances) be immune to damage from coastal storms.

Ships caught in open ocean storms have the following chance of sustaining damage:

(35 - (units keel / unit hull) * 30)% Keels also protect a ship from reef damage. Reef damage to a ship is

reduced by (units keel / units hull)%. A fully keeled ship will not be damaged by reefs.

, Add Keel to Ship [1090], Shipcraft [1000],]

E.1.10 Remove Sails from Ship [1096]

Usage: USE 1096 Time to learn: 14 days. Time to use: 3 days. Requires: ten workers [11] (not consumed)

This skill removes one sail from a ship. , Remove Sails from Ship [1096], Shipcraft [1000],]

E.1.11 Remove Fortification from Ship [1093]

Usage: USE 1093 Time to learn: 14 days. Time to use: 7 days. Requires: 15 workers [11] (not consumed)

This skill removes one fortification from a ship. , Remove Fortification from Ship [1093], Shipcraft [1000],]

E.1.12 Remove Rowing Ports from Ship [1097]

Usage: USE 1097 Time to learn: 14 days. Time to use: 3 days. Requires: ten workers [11] (not consumed)

This skill removes one rowing port from a ship.

, Remove Rowing Ports from Ship [1097], Shipcraft [1000],]

E.1.13 Remove Keel from Ship [1092]

Usage: USE 1092 Time to learn: 14 days. Time to use: 5 days. Requires: ten workers [11] (not consumed)

This skill removes one unit of keel from a ship.

, Shipcraft [1000],]

E.1.14 Remove Ram from Ship [1098]

Usage: USE 1098 Time to learn: 14 days. Time to use: 7 days.

This skill removes the ram from a ship.

, Shipcraft [1000], Appendix - Skill Dictionary,]

E.2 Combat [1100]

Time to learn: 7 days.

The student of combat learns the training of soldiers and how to lead them well in battle. With study and experience, the battle lord will find strategies to be used in difficult combat situations.

The following teachable skills may be studied directly once Combat is known:

num    skill                              time to learn
---    -----                              -------------
1102   Fight to the death                 four weeks
1105   Train Swordsmen                     two weeks
1106   Weaponsmithing                     two weeks
1107   Control Men in Battle              four weeks
1108   Instill Combat Discipline          two weeks

The following unteachable skills may be studied directly once Combat is known:

num    skill                              time to learn
---    -----                              -------------
1131   Combat Tactics                     four weeks
1132   Defense Tactics                    four weeks

The following guild skills may be studied by guildmembers:

num    skill                              time to learn
---    -----                              -------------
1190   Train Mounted                      two weeks
1191   Train Paladin                      two weeks
1192   Train Armored Troops               two weeks
1194   Construct catapult                 two weeks
1195   Use beasts in battle               four weeks

, Combat [1100], Combat [1100],]

E.2.1 Fight to the death [1102]

Usage: USE 1102 <break-point> Time: 0 days Time to learn: 28 days.

Normally a side will break when its offensive plus defensive total falls by 50% or more. Knowledge of 1102 allows a leader to override this default break point and set it to any value between 0 and 100.

A break point of zero means that a side will not break, and will fight until the last man has been killed. A break point of 100 means the side will break immediately after the first volley of attacks is launched.

REPEAT: If you do "USE 1102 100" you will give up immediately! To "fight to the death", do "USE 1102 0".

Default setting: USE 1102 50 Fight to the death: USE 1102 0 Break at 30: USE 1102 30

Note that "break at 30%" means that the side will break when its offensive plus defensive total falls to 30% of its original value.

During combat, only the stack leader’s setting for 1102 is checked. 1102 works both for attacking and defending stacks.

Note that prisoners may still be taken from a side fighting to the death, as wounded nobles may be too weak to resist being taken captive.

, Fight to the death [1102], Combat [1100],]

E.2.2 Use beasts in battle [1195]

Usage: Automatic Time to learn: 28 days. Time to practice: 3+ days. Cost to practice: 10 gold.

This skill grants a noble the ability to control animals in battle.

Each PRACTICE of this skill increases the number of animals the noble can use in battle by one. After one use, a noble is able to control one animal. After three uses, the noble will be able to use three animals, etc.

If a noble owns more animals than he can control, he will use the animals with the highest total attack + defense values, or in the case of a tie, the animals with the higher attack value.

Only the animals directly possessed by the noble will aid combat. Animals owned by units which do not possess knowledge of 1195 will not be used in battle.

The following animals may be used in combat: nazgul [55] (360,360,0) centaur [271] (30,30,30) minotaur [272] (30,5,0) giant spider [278] (150,100,0) rat [279] (3,3,0) lion [280] (100,100,0) giant bird [281] (300,250,0) giant lizard [282] (80,80,0) chimera [284] (250,250,0) harpie [285] (80,120,0) dragon [286] (1500,1500,0) gorgon [288] (10,20,0) wolf [289] (5,5,0) otter [302] (10,10,0) mole [303] (10,10,0) bull [304] (10,10,0) eagle [305] (10,10,0) monkey [306] (10,10,0) hare [307] (10,10,0) wardog [308] (10,10,0) sand rat [309] (10,10,0) balrog [310] (1000,1000,0)

With additional 10 beasts a noble can use, the time to practice this skill increases by one day, i.e., the first ten uses of this skill require three days each, the second ten four days each, and so on.

, Use beasts in battle [1195], Combat [1100],]

E.2.3 Construct catapult [1194]

Usage: USE 1194 Time to learn: 14 days. Time to use: 14 days. Requires: thirty wood [77] Requires: three leather [80] Requires: thirty workers [11] (not consumed) Produces: catapult [61]

The engineer directs the construction of a wheeled wooden frame containing a mechanism to supply stress to the end of a long wooden pole. The pole contains a basket which may be loaded with stones or other missiles to be hurled against the enemy’s walls, or over them. A winch to lower the pole and a release mechanism finish the catapult.

Catapults are sometimes loaded with the bodies of the enemy’s dead and hurled over their walls, to dishearten the besieged and spread disease amongst them.

Catapults weigh 3000 units. Leather may be made from oxen (see ⟨undefined⟩ [make], page ⟨unde-

fined⟩).

, Construct catapult [1194], Combat [1100],]

E.2.4 Train Swordsmen [1105]

Time to learn: 14 days.

Knowledge of this skill allows the noble to train lt. foot and pirates via the TRAIN order.

, Train Swordsmen [1105], Combat [1100],]

E.2.5 Weaponsmithing [1106]

Usage: MAKE <item> <days> Time: as given Time to learn: 14 days.

Weapons and armor are required for the training of some kinds of fighters. Archers require longbows, for instance, hvy. foot require plate armor, etc.

Weapons and armor are made with the MAKE command. The following weapons and armor may be produced:

num item material

— —- ——–

72 longbow yew [68]

73 plate armor iron [79]

74 longsword iron [79]

75 pike wood [77]

85 lt. crossbow wood [77]

40 hvy. crossbow wood [77] & iron [79] One unit of the input material may be turned into one of the desired

items each day. For example, MAKE 72 2 would spend two days turning two yew [68] into two longbows [72].

See the MAKE command for further information. , Weaponsmithing [1106], Combat [1100],]

E.2.6 Control Men in Battle [1107]

Usage: Automatic Time: 3 days Time to learn: 28 days. Time to practice: 3+ days. Cost to practice: 3 gold.

This skill grants a noble the ability to control men in battle.

Each PRACTICE of this skill increases the number of men the noble can use in battle by one. Initially, a noble can control ten men in battle. With each additional level of skill, a noble is able to control one additional man. With level three experience, the noble will be able to use thirteen men, etc.

If a noble owns more men than he can control, he will use the men with the highest total attack + defense values, or in the case of a tie, the men with the higher attack value.

Only the men directly possessed by the noble will aid in combat. If a noble owns fewer men than he can control, he will not “borrow” additional men from a stack mate.

With additional 10 men a noble can control, the time to practice this skill increases by one day, i.e., the first ten uses of this skill require three days each, the second ten four days each, and so on.

, Control Men in Battle [1107], Combat [1100],]

E.2.7 Combat Tactics [1131]

Usage: USE 1131 Time to learn: 28 days. Time to practice: 7 days.

By practicing this skill, a noble studies various combat tactics that will increase the attack effectiveness of the men he commands in combat.

Upon learning this skill, a noble has an attack multiplier of 1.0. With each PRACTICE, the multiplier is increased by 0.02, to a maximum of 2.0. In combat, the attack rating of all units under this noble’s direct command will be multiplied by the attack multiplier. This bonus applies in all phases of combat (i.e., melee, missile and special attacks).

For example, a lt. foot with an attack rating of 15 commanded by a noble with an attack multiplier of 1.56 would attack with an effective attack rating of 23.40.

Attack ratings are rounded to the nearest hundredth.

, Combat Tactics [1131], Combat [1100],]

E.2.8 Defense Tactics [1132]

Usage: Automatic Time to learn: 28 days. Time to practice: 7 days.

By practicing this skill, a noble studies various combat tactics that will increase the defense effectiveness of the men he commands in combat.

Upon learning this skill, a noble has an defense multiplier of 1.0. With each PRACTICE, the multiplier is increased by 0.02, to a maximum of 2.0. In combat, the defense rating of all units under this noble’s direct command will be multiplied by the defense multiplier.

For example, a pikeman with an defense rating of 20 commanded by a noble with an defense multiplier of 1.56 would defend with an effective defense rating of 31.20.

Defense ratings are rounded to the nearest hundredth.

, Defense Tactics [1132], Combat [1100],]

E.2.9 Instill Combat Discipline [1108]

Usage: Automatic when needed Time to learn: 14 days.

A noble with this skill has the ability to instill combat discipline in raw military recruits. Combat discipline – the ability to maintain one’s calm and follow orders in the chaos of battle – is the primary skill distinguishing peasants, skirmishers and other such troops from the disciplined military troops.

This skill is required to train skirmishers [36] into soldiers [12].

, Instill Combat Discipline [1108], Combat [1100],]

E.2.10 Train Armored Troops [1192]

Usage: Automatic when needed Time to learn: 14 days.

This skill permits a noble to train troops to fight in heavy armor. A soldier accustomed to fighting without encumberment must learn new skills to deal with the special problems of fighting while encased in heavy metal armor.

This skill is required to train lt. foot [20] to hvy. foot [37]. One unit of plate armor [73] is required for each hvy. foot trained.

, Train Armored Troops [1192], Combat [1100],]

E.2.11 Train Mounted [1190]

Usage: Automatic when needed Time to learn: 14 days.

This skill permits a noble to train troops to fight while on horseback. A soldier accustomed to fighting on foot must learn new skills to deal with the special problems of fighting while staying on the back of a massive warmount.

This skill is required to train hvy. foot [37] to knights [14], to train archers [13] to horse archers [41], and to train soldiers [12] to cavaliers [43]. One warmount [53] is required for each knight trained, one riding horse [52] is required for each horse archer trained, and one mithril [83] and one warmount [53] is required for each cavalier trained.

, Combat [1100],]

E.2.12 Train Paladin [1191]

Usage: Automatic when needed Time to learn: 14 days.

This skill permits a noble to train a knight to a paladin. By instilling in the knight extraordinary virtues of courage and purity, the noble creates one of the most powerful fighting forces on Olympia.

Paladins are extraordinarily devoted to the noble who created them, and cannot be transferred to another noble.

, Combat [1100], Appendix - Skill Dictionary,]

E.3 Stealth [1200]

Time to learn: 14 days.

Stealth includes the arts of theft, spying, assassination, kidnapping, and other covert activities. Most stealth skills will not make use of soldiers or other followers, instead relying on the sneakiness of a lone noble.

The following teachable skills may be studied directly once Stealth is known:

num    skill                              time to learn
---    -----                              -------------
1201   Petty thievery                     two weeks
1202   Determine inventory of noble   two weeks
1203   Determine skills of noble      two weeks
1204   Determine noble's lord         two weeks
1205   Conceal faction                    four weeks

The following unteachable skills may be studied directly once Stealth is known:

num    skill                              time to learn
---    -----                              -------------
1230   Learn of richest nearby noble      two weeks
1231   Torture prisoner                   four weeks
1232   Conceal identity of trader         two weeks
1233   Sneak into structure               four weeks

The following guild skills may be studied by guildmembers:

num    skill                              time to learn
---    -----                              -------------
1290   Conceal self                       four weeks
1291   Conceal Nation                     four weeks
1292   Assassinate Noble                  two weeks

, Stealth [1200], Stealth [1200],]

E.3.1 Conceal Nation [1291]

Usage: USE 1291 nation-name Time to learn: 28 days. Time to use: 7 days.

This skill causes a noble to appear to be from a different nation. The differences between the nations are so striking that the noble must use extensive make-up, prosthetic devices and other artifices to conceal his true identity. For this reason, the effect of this skill lasts for only two weeks and then must be renewed. While it is being renewed, the noble’s true nation will be evident.

The noble selects the nation he wishes to imitate by giving a unique first portion of the nation’s name. For example, if the noble wanted to appear to be from the "Imperial Empire" he would use the command:

USE 1291 Imperial

or even

USE 1291 I

so long as there were no other nations in the game whose name started with "I".

The noble may only impersonate existing nations. That is, he may not simply make up a new nation name.

, Conceal Nation [1291], Stealth [1200],]

E.3.2 Assassinate Noble [1292]

Usage: USE 1292 target Time to learn: 14 days. Time to use: 7 days. Time to practice: 28 days. Cost to practice: 50 gold.

This skill is used to assassinate a noble. If successful, the attack will instantly kill the target noble. Assassinations can only take place in cities, where the assassin can hide in the crowds, or in castles, where the assassin can make use of secret passages and the like to sneak up on the victim.

The chance of success is dependent upon several factors: luck, the skill of the assassin, and the skill of any bodyguards.

The base chance of success is randomly -25% to 25%.

Each level of experience the assassin has in 1292 increases the base chance of success by 5%.

Each level of experience in the skill [1990] Protect Noble possessed by any noble stacked with the target noble reduces the chance of success by 5%.

As an example, suppose an assassin with 13 levels of experience in 1292 attempted to assassinate a noble in a stack that held two bodyguards with 3 and 6 levels respectively of [1990] Protect Noble. If the base chance was randomly determined to be 7%, then the overall chance of success would be 7 + (13*5) - (9*5) = 27%.

The maximum chance of success for assassination is 75%.

Assassination takes 7 days. Most of this time is spent in preparation. The target noble must be present only on the 7th day.

Assassination attempts are always visible to everyone in the location, but the identity of the assassin is not revealed.

If the assassination fails, or the slain target has bodyguards, there is a 25% chance that the assassin will be captured after the assassination.

, Assassinate Noble [1292], Stealth [1200],]

E.3.3 Petty thievery [1201]

Usage: USE 1201 Time: seven days Time to learn: 14 days. Time to use: 7 days.

The skill of the common burglar, 1201 allows a thief to work a city or civilized province for gold. The thief may pick pockets, steal from merchants in the market, enter homes in search of loot, or deceive townsfolk into parting with their gold with confidence games.

The thief must be alone in order to thieve. It may not be stacked with any other nobles, or possess any visible followers, such as peasants, soldiers, workers, etc. If the province or city where the thief is working is garrisoned, there is a possibility he will be captured by the garrison.

How much a thief will earn, and his likelihood of being captured, depends upon the experience level of the thief:

    Description      Amount         Chance of Capture
    -----------      ------         -----------------
    apprentice        7 gold              5%
    journeyman       14 gold              3%
    adept            28 gold              1%
    master           49 gold              -
    grand master     70 gold              -

Of course, thieving depends upon the peasants having something to steal. Any money a thief takes from a province comes out of the peasants’ accumulated wealth. If they have no gold, the thief will come away empty-handed.

Thieving has a detrimental effect on the populace, and will drive 7-21 peasants away with each use.

A province with less than 100 peasants cannot be thieved. Thieving a city is the same as thieving the province the city is in.

, Petty thievery [1201], Stealth [1200],]

E.3.4 Determine inventory of character [1202]

Usage: USE 1202 <character> Time: seven days Time to learn: 14 days. Time to use: 7 days.

Sneakily learn the contents of the specified character’s inventory. The spy must be in the same location as the target.

, Determine inventory of noble [1202], Stealth [1200],]

E.3.5 Determine skills of character [1203]

Usage: USE 1203 <character> Time: seven days Time to learn: 14 days. Time to use: 7 days.

Sneakily learn what skills the specified character possesses by crafty observation and spying.

The spy must be in the same location as the target.

, Determine skills of character [1203], Stealth [1200],]

E.3.6 Determine character’s lord [1204]

Usage: USE 1204 <character> Time: seven days Time to learn: 14 days. Time to use: 7 days.

Reveals what player the specified character is sworn to, unless the target character is concealing its identity. The spy must be in the same location as the target.

, Determine character’s lord [1204], Stealth [1200],]

E.3.7 Learn of richest nearby noble [1230]

Usage: USE 1230 Time: seven days Time to learn: 14 days. Time to use: 7 days.

Must be used in an inn. The character will chat with the local patrons of the inn, inquiring as to the recent comings and goings of other nearby characters, paying special attention to clues which would indicate the presence of wealthy units.

Use of 1230 will reveal the identity of the richest noble within three provinces of the inn. Any character with less than 500 gold is not considered wealthy, and is ignored by 1230.

, Learn of richest nearby noble [1230], Stealth [1200],]

E.3.8 Torture prisoner [1231]

Usage: TORTURE <prisoner> Time: seven days Time to learn: 28 days. Time to use: 7 days.

Torture is the practice of inflicting pain on a captive prisoner in order to extract information against the victim’s will.

Torture will attempt to learn what faction the prisoner belongs to. Each torture attempt inflicts a 50 point health loss on the prisoner.

The chance of success depends on the prisoner’s loyalty:

loyalty chance ——- —— oath-2 0% oath-1 10% contract 50% fear 100%

, Torture prisoner [1231], Stealth [1200],]

E.3.9 Conceal identity of trader [1232]

Usage: BUY <item> <qty> <price> <have-left> 1

SELL <item> <qty> <price> <have-left> 1 Time to learn: 14 days.

This skill allows a trader to conceal his identity when dealing with a market. By working through middlemen and employing shady contacts, the trader will not reveal his identity to the other party to his transaction, or to the market at large.

To use this skill, the trader should supply a "1" flag as the fifth argument to the BUY or SELL command. Note that The first four arguments must be specified; use "0" for have-left if it is to be left as the default.

For example, to conceal one’s identity when selling 10 pots for 2 gold each, issue the order:

    sell 95 10 2 0 1

If the conceal trade flag is supplied, but the character does not know 1232, the BUY or SELL order will fail.

, Conceal identity of trader [1232], Stealth [1200],]

E.3.10 Sneak into structure [1233]

Usage: SNEAK [building or ship] Time: three days Time to learn: 28 days. Time to use: 3 days.

The character will sneak into the specified building or ship. Once inside, other occupants of the structure will see the character in their location reports, unless the unit employs additional methods of concealment. Occupants will not see the character enter, however.

To exit the structure by the same means, use SNEAK with no arguments.

The unit must be alone in order to sneak into and out of buildings or ships. It may not be stacked with any other characters, or possess any followers, such as peasants, soldiers, workers, etc.

, Sneak into structure [1233], Stealth [1200],]

E.3.11 Conceal faction [1205]

Usage: automatic when necessary Time to learn: 28 days.

Knowledge of 1205 causes a character to keep its faction affiliation secret. Other units attempting to learn the what player the unit belongs with the skill Determine character’s lord [1204] will fail.

Permissions set for the player of a character concealing its faction affiliation will not include the character. For example, suppose unit [5499] is concealing the fact that it belongs to player [613]. Permissions declared for player [613] (such as ADMIT, DEFEND, etc.) will not include the sneaky unit [5499], since [5499] will not admit to an affiliation with [613].

Furthermore, such units will not DEFEND units from their own faction by default, since this might reveal their identity.

, Stealth [1200],]

E.3.12 Conceal self [1290]

Usage: USE 1290 <flag> Time to learn: 28 days. Time to use: 3 days. Time to practice: 7 days. Cost to practice: 35 gold.

This skill allows a character to conceal itself from view. When activiated, other characters will not see the unit in their location reports, and will not see the unit come and go. The unit will also be invisible to garrisons, and so may enter a province where it would be turned away if visible. Note that the unit is still prevented from sneaking into a city or smaller sublocation (e.g., a tower).

The unit must be alone in order to conceal itself. It may not be stacked with any other characters, or possess any visible followers, such as peasants, soldiers, workers, etc. It may also not be the first character inside a sublocation, building or ship. The owners of these small locations are always visible.

Hiding takes three days to accomplish. However, turning off concealment takes zero time.

Use of the skill will set whether or not the unit is trying to hide from view:

    use 1290 1 # Hide from view, takes three days

    use 1290 0 # Don't conceal self from view, zero days

Note that attempts to gain entrance to buildings or ships with the MOVE order will subject to the unit to the ADMIT system. A failure to gain entrance with MOVE will give the unit away to the structure’s owner.

Concealment does not hide noticeable actions such as attacking another unit, digging up a body, etc. Other than movement, actions which would generate a message in the location log will still generate a message.

Hidden nobles may be spotted by other nobles using the SEEK command. Nobles using SEEK to look for a particular noble have a base chance of 50% per day of finding the noble. Nobles using SEEK without a particular target have a 25% chance per day of spotting any hidden noble.

For each level of skill a noble has in 1290, the chance to discover that noble through SEEK is reduced by 5% if the noble is specifically being sought, and by 2.5% if the search is general. The chance to spot a noble is improved if the seeker has the skill [1901] Spot Hidden Noble.

, Stealth [1200], Appendix - Skill Dictionary,]

E.4 Persuasion [1300]

Time to learn: 14 days.

The skill of smooth-tongued oratators, Persuasion involves the manipulation of others through eloquent verbal dialog. From subverting a noble’s loyalty to rousing peasant mobs with firey speeches, Persuasion can a powerful tool or weapon.

The following teachable skills may be studied directly once Persuasion is known:

    num    skill                              time to learn
    ---    -----                              -------------
    1301   Bribe noble                        two weeks
    1302   Raise peasant mob                  two weeks
    1303   Rally peasant mob                  four weeks
    1304   Train angry peasants               two weeks

The following unteachable skills may be studied directly once Persuasion is known:

    num    skill                              time to learn
    ---    -----                              -------------
    1330   Persuade oathbound noble           four weeks
    1331   Incite mob violence                four weeks
    1332   Summon wild men                    four weeks
    1333   Persuade wild men to remain        four weeks

The following guild skills may be studied by guildmembers:

    num    skill                              time to learn
    ---    -----                              -------------
    1390   Encourage Population Growth        two weeks

, Persuasion [1300], Persuasion [1300],]

E.4.1 Bribe noble [1301]

Usage: BRIBE <target> <gold> [flag] Time: seven days Time to learn: 14 days. Time to use: 7 days.

Offer a character gold in an attempt to win their service. The amount must be at least 250 gold, or the receiving noble will not consider it a serious offer.

Nobles with oathbound loyalty ignore bribes. They may accept the gift, but they will not renounce loyalty to their lord.

Nobles with contract loyalty will accept the bribe if it exceeds their current loyalty rating. For example, a noble of loyalty contract-750 would ignore any bribe of less than 751 gold.

Nobles with fear loyalty, or nobles not sworn to any lord, will accept any bribe of at least 250 gold.

Nobles that accept a bribe will renounce loyalty to their current lord, and swear fealty to the bribing character. Characters who are successfully bribed will have contract loyalty of half the amount of the bribe over their old loyalty. For example, a noble with contract loyalty of 570 who was bribed with 1100 gold would switch loyalties and have a new loyalty of contract-265 ((1100-570)/2). Nobles with fear loyalty before the bribe are treated as if they had contract-250 loyalty.

If flag is 1, and the bribe attempt is successful, the bribed character will stack with the briber.

Bribes will be refused by characters who have already switched lords during the month because of bribery. For example, suppose Osswid successfully bribes Feasel, and Feasel swears loyalty to Osswid. Later that turn, if Procrustes attempts to bribe Feasel, Feasel will refuse to accept the bribe.

Examples:

bribe 5499 750 # Offer [5499] 750 gold to switch factions. # If successful,

[5499] will not stack with the # briber.

bribe 3409 300 1 # Offer [3409] 300 gold to switch factions. # If successful, [3409] will stack with the briber.

, Bribe noble [1301], Persuasion [1300],]

E.4.2 Persuade oathbound noble [1330]

Usage: USE 1330 <target> [flag] Time: seven days Time to learn: 28 days. Time to use: 7 days.

Attempt to persuade an oathbound noble to defect. There is only a chance of success when used on oath-1 nobles. Oath-2 nobles will not leave their lord under any circumstances.

1330 costs 25 gold per use, to pay for gifts, fine meals, handsome clothing, and other props useful in the persuasion attempt.

Each seven day use yields a 10% chance of success.

If flag is 1, and the persuasion attempt is successful, the target character will stack with the persuading character.

, Persuade oathbound noble [1330], Persuasion [1300],]

E.4.3 Raise peasant mob [1302]

Time to learn: 14 days. Time to use: 7 days.

Usage: RAISE Time: seven days

The noble gives an impassioned speech, rousing peasants and causing them to gather into a crowd. New peasant crowds will set their GUARD flag to protect their homes from pillaging.

Peasant mobs may be formed in provinces and cities. Only one crowd may be raised per location. Only when the raised mob is later killed, or if the peasants return to their homes, may another mob be raised there.

, Raise peasant mob [1302], Persuasion [1300],]

E.4.4 Rally peasant mob [1303]

Usage: RALLY <mob> Time: seven days Time to learn: 28 days. Time to use: 7 days.

Deliver a fiery speech to an existing crowd of peasants, causing them to stack with the noble. They will remain with the noble for three months. The mob must be specified as an argument to RALLY.

Use of 1303 on a mob already stacked with the noble will renew the peasants’ enthusiasm, and add an additional three months to the time they will remain stacked, to a maximum of five months.

Mobs stacked with the noble may not be given any orders, but will participate in any combat the noble encounters.

A maximum of three peasant mobs will stack with a noble.

, Rally peasant mob [1303], Persuasion [1300],]

E.4.5 Incite mob violence [1331]

Time to learn: 28 days.

Usage: INCITE <mob> <target> Time: seven days

Speak out against the target in such as a way as to inspire public hatred and ire. The target will be made to seem a malicious influence responsibile for all of the common man’s troubles.

An existing crowd of peasants must be specified as the first argument, and the target of their violence as the second. The mob may not already be stacked under another noble who has rallied it.

There is a 50% chance that the mob will be roused to attack the target. There is also a 33% chance that patrons in any local inns will learn that

the noble is trying to incite a mob to violence against the target. , Incite mob violence [1331], Persuasion [1300],]

E.4.6 Summon wild men [1332]

Usage: USE 1332

MAKE 98 1 Time: one day Time to learn: 28 days. Time to use: 1 day.

Wild men, or savages, inhabit remote parts of the world. Savages communicate by beating their drums to indicate calls-to-arms, battle challenges, or victory celebrations.

Use of 1332 will attempt to summon wild men with drumbeats, and persuade them to stack with the noble. Savages will remain with the noble for 3 months, during which time they will participate in any battles the noble fights.

To summon wild men, the noble must first use the MAKE 98 1 order to produce a drum [98]. A drum requires 1 mallorn wood [70] to produce. MAKE 98 1 requires knowledge of 1332.

The drum may be used to summon wild men with the USE 1332 order. The noble should be in a remote province, preferably surrounded by areas uninhabited by other characters. The noble should allow time for the savage men to hear and follow the drumbeats. Any wild men summoned to the province will stack with the noble.

For example:

make 98 1 # make a drum use 1332 # use the drum to summon the savages wait time 15 # give savages time to arrive and # stack with us

The number of savage units in the world is limited. If all of the available savages have already been summoned, none will come when the drum is used.

At most three savage units will stack with a noble. If a noble attempts to summone more savages while already stacked with three savage units, the summoning will fail.

, Summon wild men [1332], Persuasion [1300],]

E.4.7 Persuade wild men to remain [1333]

Usage: USE 1333 <savage> Time: seven days Time to learn: 28 days. Time to use: 7 days.

Usage of 1333 will induce a group of wild men to remain stacked with the noble that summoned them. An application of 1333 will add two months to the duration the wild men will stay with the noble, to a maximum of four months.

The savage unit to be persuaded to remain must be specified as an argument to USE 1333.

, Persuade wild men to remain [1333], Persuasion [1300],]

E.4.8 Train angry peasants [1304]

Time to learn: 14 days. Usage: TRAIN 23 [days]

TRAIN 10 [days] Time: as given

The noble employs persuasive means to rouse mild-tempered peasants into a state of anger. Angry peasants fight with twice the attack value of regular peasants. However, in their excited state they are untrainable to any other kind of man.

Examples:

    train 23 # train all peasants to angry peasants
    train 23 5 # train five peasants into angry peasants
    train 10 # calm all angry peasants back into peasants

Angry peasants require a monthly maintenance cost of 2 gold each.

, Persuasion [1300],]

E.4.9 Encourage Population Growth [1390]

Usage: USE 1390 Time to learn: 14 days. Time to use: 7 days.

By using this skill, a glib noble convinces peasants to settle down in a province and make their futures there. By convincing the local peasants that the province is a wonderful place, he raises the growth rate of the population in the province by 2%.

This effect lasts one month and is not cumulative.

, Persuasion [1300], Appendix - Skill Dictionary,]

E.5 Construction [1400]

Time to learn: 7 days.

The domain of engineers, builders and architects, Construction concerns the erection of inns, castles, city walls, and other fortifications.

The following teachable skills may be studied directly once Construction is known:

num    skill                              time to learn
---    -----                              -------------
1401   Construct siege tower              two weeks
1402   Stone quarrying                    two weeks

The following guild skills may be studied by guildmembers:

num    skill                              time to learn
---    -----                              -------------
1490   Build City                         two weeks
1491   Fortify Castle                     three weeks
1492   Strengthen Castle Walls            three weeks
1493   Add Moat to Castle                 three weeks
1494   Widen Entrance   two weeks

, Construction [1400], Construction [1400],]

E.5.1 Fortify Castle [1491]

Usage: USE 1491 [days] Time to learn: 21 days. Time to use: Variable.

This skill permits a noble to add fortification to an existing castle, raising its defense rating. To use this skill, a noble should be inside the castle to be improved and have at least 1 worker.

Improving the fortification of a castle requires [78] stone. The amount required is the current defense of the castle plus one. For example, to improve a castle from a defense rating of 40 to 41 requires 41 stone.

The effort required is 5 times the current defense rating of the castle in worker-days. For example, to improve a castle from a defense rating of 40 to 41 requires 40*5 = 200 worker-days.

If the "days" argument is present, the noble will work on adding fortification for that many days. Otherwise, he will work until the job is completed.

A damaged castle cannot be improved until it has been fully repaired. , Fortify Castle [1491], Construction [1400],]

E.5.2 Strengthen Castle Walls [1492]

Usage: USE 1492 Time to learn: 21 days. Time to use: Variable.

This skill permits a noble to strengthen the walls of an existing castle. Strengthening the walls of a castle adds to the damage the castle must take before it is destroyed. Initially, a castle takes 100 hits to destroy. Each time the walls are strengthened, 5 hits are added to the castle. To use this skill, a noble should be inside the castle to be improved and have at least 1 worker.

Strengthening the walls of a castle requires [78] stone. The amount required is the current hit rating of the castle. For example, to improve a castle from a hit rating of 110 to 115 requires 110 stone.

The effort required is 5 times the current hit rating of the castle in workerdays. For example, to improve a castle from a defense rating of 110 to 115 requires 110*5 = 550 worker-days.

If the "days" argument is present, the noble will work on adding fortification for that many days. Otherwise, he will work until the job is completed.

A damaged castle cannot be improved until it has been fully repaired.

, Strengthen Castle Walls [1492], Construction [1400],]

E.5.3 Add Moat to Castle [1493]

Usage: USE 1493 Time to learn: 21 days. Time to use: Variable.

This skill permits a noble to add a moat to an existing castle. A moat prevents attacking missile troops to approach the walls of the castle close enough to be effective against troops inside the castle. A castle with a moat adds 100 to the defense rating against missile weapons of all the troops it protects. This bonus does not apply to missile siege weapons attacking the castle, since they are stronger and not bothered by the added distance. To use this skill, a noble should be inside the castle to be improved and have at least 1 worker.

Building a moat requires 200 [78] stone, and 4000 worker-days of effort.

A damaged castle cannot be improved until it has been fully repaired.

, Add Moat to Castle [1493], Construction [1400],]

E.5.4 Construct siege tower [1401]

Usage: USE 1401 Time to learn: 14 days. Time to use: 14 days. Requires: thirty wood [77] Requires: thirty workers [11] (not consumed) Produces: siege tower [62]

The engineer directs thirty workers to build a wooden tower. The tower contains several ladders and a platform from which soldiers may attack.

In battle, the tower is wheeled up against an enemy’s walls and used by attacking soldiers to match the height of the defenders. The attackers may then climb from the tower onto the walls.

Siege towers weigh 3000 units.

, Construct siege tower [1401], Construction [1400],]

E.5.5 Stone quarrying [1402]

Usage: QUARRY [number of stones] [days] Requires: at least one worker Time to learn: 14 days. Time to use: Variable.

Building stone is used in the construction of temples, castles and towers. Up to 50 stone may be removed from mountain provinces or rocky hills each month, or up to ten stone from desert provinces.

Each worker will cut and prepare one building stone [78] per day.

Examples:

quarry # cut as much building stone as possible quarry 25 # stop after cutting 25 stone quarry 0 5 # quarry stone for five days quarry 25 5 # stop after five days, or 25 stone, whichever

# comes first

In all cases, quarrying will cease as soon as the supply of ready stone in the province has been exhausted.

, Stone quarrying [1402], Construction [1400],]

E.5.6 Build City [1490]

Usage: BUILD city "Name of City" Time to learn: 14 days. Time to use: Variable.

One of the greatest construction projects a noble can undertake is to build a new city. By laying the city walls, grading streets, creating rudimentary sewers and perhaps port facilities, a noble prepares the way for a thriving new metropolitan center.

But building a new city is not a trivial effort. It requires 2000 stone, 500 wood, 50,000 worker-days of effort. It is said that the fastest a city has ever been built was when the Emperor Dato built the city of Taejon in 57 days.

The conditions must also be right for the creation of a new city. The province must contain at least 1000 peasants and adjacent provinces must contain at least 100 peasants. In addition, the nearest existing city must be at least 5 provinces away.

, Construction [1400],]

E.5.7 Widen Entrance [1494]

Usage: USE 1494 <sublocation> Time to learn: 14 days. Time to use: 1 week

This skill permits a Construction guildmaster to widen the entrance to a sublocation. Some sublocations limit the number of men that can enter at one time. For example:

    Inner locations:
       Lair [ad7n], lair, 1 day, entrance size: 4, owner:
          Harpie [jn7t], harpies, number: 12

Lair [ad7n] can only be entered – or attacked – by stacks of size four or less.

For every ten workers the guildmaster possesses when using this skill, he widens the entrance to the sublocation by 1 man. For example, one use of this skill on [ad7n] with 40 workers would widen the entrance to 8 men at a time.

, Construction [1400], Appendix - Skill Dictionary,]

E.6 Alchemy [1500]

Time to learn: 14 days.

Called petty magician, druggist, and sometimes fool, the alchemist seeks to brew useful potions from elements and rare extracts. A method to transmute matter into gold is eagerly sought, but remains unknown.

The following teachable skills may be studied directly once Alchemy is known:

num    skill                              time to learn
---    -----                              -------------
1501   Brew healing potion                four weeks
1502   Brew potion of death               four weeks
1503   Collect rare elements              two weeks

The following unteachable skills may be studied directly once Alchemy is known:

num    skill                              time to learn
---    -----                              -------------
1530   Extract venom from ratspider       four weeks
1531   Make potion of slavery             four weeks
1532   Record skill on scroll             four weeks
1533   Turn lead into gold                four weeks

The following guild skills may be studied by guildmembers:

num    skill                              time to learn
---    -----                              -------------
1590   Make potion of weightlessness      four weeks
1591   Make potion of fiery death         four weeks

, Alchemy [1500], Alchemy [1500],]

E.6.1 Extract venom from ratspider [1530]

Usage: USE 1530 Produces: one ratspider venom [63] Time to learn: 28 days. Time to use: 7 days. Requires: one ratspider [81] Produces: ratspider venom [63]

The alchemist performs a delicate operation to extract the highly poisonous ratspider venom.

The top of the ratspider’s head is removed by inserting a sharp knife above the head joint and twisting. Two glands, from which the ratspider venom is later extracted, are removed from behind the mandible.

Risk is minimized by taking extreme care when handling the ratspider. The ratspider is killed in the process.

, Extract venom from ratspider [1530], Alchemy [1500],]

E.6.2 Make potion of slavery [1531]

Usage: USE 1531 Time to learn: 28 days. Time to use: 7 days. Requires: one ratspider venom [63]

The alchemist brews a potion which, when ingested by a character, causes that character to become a servant of the alchemist. The unfortunate noble will swear fealty to the alchemist and promise to do his bidding.

The rare and highly poisonous ratspider venom is used in the preparation of this potion. Upon completion, the new potion will be added to the alchemist’s inventory. The potion will, of course, be unlabeled.

A character wishing to drink the potion should specify it to the USE order. For instance, to drink potion [601], issue the order: USE 601.

, Make potion of slavery [1531], Alchemy [1500],]

E.6.3 Brew healing potion [1501]

Usage: USE 1501 Time to learn: 28 days. Time to use: 7 days. Requires: one avinia leaf [65]

The alchemist brews a potion of healing. Any character ingesting the potion will be immediately healed of any wounds or other maladies.

Upon completion, the potion will be added to the alchemist’s inventory. The potion’s label will not reveal its purpose.

A character wishing to drink the potion should specify it to the USE order. For instance, to drink potion [602], issue the order: USE 602.

, Brew healing potion [1501], Alchemy [1500],]

E.6.4 Brew potion of death [1502]

Usage: USE 1502 Time to learn: 28 days. Time to use: 10 days. Requires: one lana bark [64]

The alchemist brews a potion which will cause instant death to any character which drinks it.

Upon completion, the potion will be added to the alchemist’s inventory. The potion’s label will not reveal its sinister purpose.

A character wishing to drink the potion should specify it to the USE order. For instance, to drink potion [603], issue the order: USE 603.

, Brew potion of death [1502], Alchemy [1500],]

E.6.5 Make potion of weightlessness [1590]

Usage: USE 1590 Time to learn: 28 days. Time to use: 7 days. Requires: one pretus bones [71]

The alchemist brews a potion which will make any noble "weightless" for 7 days.

While under the effect of this potion, the noble will have his total weight (his weight, plus the weight of everything in his inventory) reduced by 500 units. For example, a noble using a potion of weightlessness could carry 5 stone [78] by himself without being overburdened.

Upon completion, the potion will be added to the alchemist’s inventory. The potion’s label will not reveal its purpose.

A character wishing to drink the potion should specify it to the USE order. For instance, to drink potion [603], issue the order: USE 603.

, Make potion of weightlessness [1590], Alchemy [1500],]

E.6.6 Make potion of fiery death [1591]

Usage: USE 1591 Time to learn: 28 days. Time to use: 14 days. Requires: one spiny root [66]

The alchemist brews a potion which, when hurled in battle, will create a small explosion. Upon completion, the potion will be added to the alchemist’s inventory. The potion’s label will not reveal its purpose.

Beginning in the second round of combat, an alchemist who knows this skill and carries a potion of fiery death will hurl it into the front rank of the opposing forces, where it will explode. The attack occurs during the special attacks phase, and the alchemist must not be in the front rank of the battle. Each exploding potion acts as a 30 strength attack against 5 fighters. There is a 5% chance that a potion will be a "dud" and not explode at all. Note that the alchemist hurls one potion per round until he has exhausted his supply of potions.

Other nobles may carry potions of Fiery Death but cannot use them in battle.

A character who mistakenly drinks the potion (e.g., through the USE command), will suffer 10-20 points damage from the fiery concoction.

, Make potion of fiery death [1591], Alchemy [1500],]

E.6.7 Record skill on scroll [1532]

Usage: USE 1532 <skill> [# days instruction] [book] Requires: blank scrolls [84], 100 gold Time to learn: 28 days. Time to use: 7 days.

Create a book for the given skill, or add instruction to an existing book. The book can be used as a source of instruction. Books may be created only for category skills and teachable subskills.

If the days of instruction argument is missing, a default of 7 days is assumed. One scroll is required for every 7 days of instruction being added.

There is a cost of 100 gold to begin a new book. This covers the cost of the binding, covers, etc.

The skill must be known in order to be written. The character will still know the skill after writing it.

This skill requires one day for every page written.

The optional third argument is the id number of an existing book in the noble’s possession. If this is a book for the given skill, then additional pages will be added to the book. In this case, there is no 100 gold fee to begin a new book.

1532 cannot be used to write books of any magic or religion skills. Examples: USE 1532 7 Creates a new book with 7 days of instruction in 1532. USE 1532 3 j3k2 Adds three days instruction to book [j3k2], provided that book teaches

1532.

, Record skill on scroll [1532], Alchemy [1500],]

E.6.8 Collect rare elements [1503]

Usage: COLLECT <item> [number] [days] Time: as given or necessary Time to learn: 14 days.

The noble will seek rare elements to collect. See the COLLECT order for details on the syntax of the arguments and an explanation of the columns in the following table.

The followings kinds of rare elements may be collected with this skill:

num item  who where  qty skill chance
— —-  — —–  — —– ——

67 farrenstone noble caves 2 1503 25% 71 pretus bones noble sand pits 1 1503 25% 81 ratspider noble ruins 1 1503 1%

, Alchemy [1500],]

E.6.9 Turn lead into gold [1533]

Usage: USE 1533 [amount] Time: one week Requires: one farrenstone [67] Time to learn: 28 days. Time to use: 7 days.

Using obscure chemical processes, the Alchemist may turn lead into gold. One unit of farrenstone is required as a catalyst for the conversion.

Up to 20 units of lead may be converted with each use of 1533. Amount defaults to 20 or all of the lead in the Alchemist’s inventory, whichever is less. Each unit of lead yields 10 gold after conversion.

, Alchemy [1500], Appendix - Skill Dictionary,]

E.7 Forestry [1600]

Time to learn: 7 days.

Forestry is skill at managing forests and harvesting their products, such as timber, yew, and rare foliage.

The following teachable skills may be studied directly once Forestry is known:

num    skill                              time to learn
---    -----                              -------------
1601   Construct battering ram            two weeks
1602   Harvest lumber                     two weeks
1603   Harvest yew                        two weeks
1604   Collect rare foliage               two weeks
1605   Harvest opium                      two weeks

The following unteachable skills may be studied directly once Forestry is known:

num    skill                              time to learn
---    -----                              -------------
1630   Harvest mallorn wood               two weeks
1631   Improve opium production           two weeks

, Forestry [1600], Forestry [1600],]

E.7.1 Construct battering ram [1601]

Usage: USE 1601 Time to learn: 14 days. Time to use: 14 days. Requires: thirty wood [77] Requires: five iron [79] Requires: thirty workers [11] (not consumed) Produces: battering ram [60]

The craftsman constructs a battering ram by selecting a large, heavy tree trunk, sharpening the end somewhat and plating it with iron, and then suspending it in a movable wooden frame such that it may be swung back and forth against walls or gates.

The frame receives an iron plated roof to protect the operators from boiling oil or pitch, rocks, or other missile weapons hurled down upon them by the enemy.

Battering rams weigh 3000 units.

, Construct battering ram [1601], Forestry [1600],]

E.7.2 Harvest lumber [1602]

Usage: WOOD [number] [days] Time: as given or necessary Requires: at least one worker [11] Time to learn: 14 days. Time to use: Variable.

Workers enter the forest and select trees which are suitable for felling and sawing. Lumber is used in the construction of ships, buildings and some weapons. Timber is also used to support the roofs of mine shafts.

Must be used in a forest province. Forests generate 30 wood [77] each month, and no more than 30 wood per month can be harvested from a province.

Examples:

wood # obtain as much wood as possible

wood 25 # stop after getting 25 wood wood 0 5 # cut timber for five days wood 25 5 # stop after five days, or 25 wood, whichever

# comes first

In all cases, timber cutting will cease as soon as the forest’s supply of wood for the month has been exhausted.

WOOD is an alias for COLLECT 77.

, Harvest lumber [1602], Forestry [1600],]

E.7.3 Harvest yew [1603]

Usage: YEW [number] [days] Times: as given or necessary Requires: at least one worker [11] Time to learn: 14 days. Time to use: Variable.

Workers enter a yew grove and select yew trees which are suitable for collection. Yew is used in the making of longbows.

Must be used in a yew grove. Yew groves generate 5 yew [68] each month. Examples:

yew # obtain as much yew as possible

yew 1 # stop after getting 1 wood yew 0 4 # cut yew for four days

In all cases, yew harvesting will cease as soon as the grove’s supply of yew for the month has been exhausted.

YEW is an alias for COLLECT 68. , Harvest yew [1603], Forestry [1600],]

E.7.4 Collect rare foliage [1604]

Usage: COLLECT <item> [number] [days] Time: as given or necessary Time to learn: 14 days.

The noble will seek rare foliage to collect. See the COLLECT order for details on the syntax of the arguments and an explanation of the columns in the following table.

The followings kinds of rare foliage may be collected with this skill:

num item  who where  skill chance
— —-  — —–  —– ——

64 lana bark noble circle of trees 1604 50% 65 avinia leaf noble mallorn grove 1604 20% 66 spiny root noble bogs, pits 1604 25% 66 spiny root noble swamps 1604 25%

Each location contains a limited supply of foliage. At the end of each month, new foliage grows. The following table shows the maximum amount of each kind of foliage in a location and the rate at which it grows.

num item max grow

— —- — —- 64 lana bark 30 3 65 avinia leaf 4 1 66 spiny root 2 2 (bogs, pits) 66 spiny root 1 1 (swamp)

If the foliage in any location drops to zero, it will take some time for the foliage to recover. A noble can "re-plant" a location by dropping the item from his inventory into that location.

, Collect rare foliage [1604], Forestry [1600],]

E.7.5 Harvest mallorn wood [1630]

Usage: MALLORN [number] [days] Time: as given or necessary Time to learn: 14 days.

Mallorn groves produce two units of mallorn wood [70] each month. A maximum of two units may be harvested each month.

Because of the careful selection required, and the small quantity of mallorn which becomes ready to cut each month, a noble works at this task alone. No workers are needed for mallorn wood cutting.

Each day spent using MALLORN gives a 20% chance of obtaining one unit of mallorn wood [70].

Examples:

mallorn # obtain as much mallorn wood as possible mallorn 2 # stop after getting 2 mallorn wood mallorn 0 5 # cut mallorn wood for five days mallorn 2 5 # stop after five days, or 2 mallorn wood, whichever

# comes first

In all cases, mallorn cutting will cease as soon as the groves’s supply for the month has been exhausted.

MALLORN is an alias for COLLECT 70. , Harvest mallorn wood [1630], Forestry [1600],]

E.7.6 Harvest opium [1605]

Usage: OPIUM [number] [days] Time: as given or necessary Time to learn: 14 days.

Poppy fields produce 25 harvestable opium [93] each month. A noble using the OPIUM command will collect one unit of opium each day.

Examples:

opium # obtain as much opium as possible

opium 25 # stop after getting 25 opium opium 0 5 # harvest opium for five days opium 25 5 # stop after five days, or 25 opium, whichever

# comes first

In all cases, opium harvesting will cease as soon as the field’s supply for the month has been exhausted.

OPIUM is an alias for COLLECT 93.

, Forestry [1600],]

E.7.7 Improve opium production [1631]

Usage: USE 1631 Time: seven days Time to learn: 14 days. Time to use: 7 days.

Use of 1631 causes the noble to pay special attention to the development of a poppy field. Through appropriate fertilization and agressive weed and pest control, the opium production of the poppy field will be doubled in the following month, from 25 to 50 units of opium.

, Forestry [1600], Appendix - Skill Dictionary,]

E.8 Mining [1700]

Time to learn: 7 days.

Mining is the skill of sinking deep shafts into mountainous provinces for the purpose of extracting valuable elements. Mining skills require a mine shaft to be constructed first with the BUILD SHAFT order.

The likelihood of death because of cave-ins or poisonous gases seeping into the shafts tends to make mining dangerous work.

The following teachable skills may be studied directly once Mining is known:

    num    skill                              time to learn
    ---    -----                              -------------
    1701   Mine iron                          two weeks
    1702   Mine gold                          two weeks
    1703   Build Mine Shaft                   three weeks

The following unteachable skills may be studied directly once Mining is known:

    num    skill                              time to learn
    ---    -----                              -------------
    1730   Mine mithril                       two weeks
    1731   Mine gate crystal                  two weeks

The following guild skills may be studied by guildmembers:

    num    skill                              time to learn
    ---    -----                              -------------
    1790   Add Wooden Shoring                 three weeks
    1791   Add Iron Shoring                   three weeks

, Mining [1700], Mining [1700],]

E.8.1 Build Mine Shaft [1703]

Usage: build shaft ["Name of shaft"] Time to learn: 21 days. Time to use: Variable.

This skill permits a character to build a mine shaft.

A mine is a deep shaft or tunnel which allows workers to extract valuable resources from the earth, such as iron and gold. Mine shafts must be built in mountain provinces or in another mine shaft; there may be at most one mine shaft in a location. A mine shaft is built with the "build shaft" command. To build a mine shaft, a noble must know [1703] Deepen Mine Shaft. Building a mine requires at least 3 workers and 500 worker-days of effort.

When a mine shaft is completed, it will be "down" from the location in which it was built. Each mine shaft descends 100 feet, so that the initial mine shaft in a location goes down 100 feet; a mine shaft built in that mine shaft descends to 200 feet, and so on. E.g.,

    Province controlled by Revelstone [d53], castle, in Westron Moun-
    tains [bo14]
    Ruled by Lord Variol [dw7r], duke
    Routes leaving Landsdrop Mountains:
       North, mountain, to Trothguard [bo13], 10 days
       East, mountain, to Doriendor Corishev [bp14], 10 days
       South, to Ocean [bq13], Balatean Sea, impassable
       West, mountain, to Landsdrop Mountains [bp12], 10 days
       Down, mine shaft, to New mine shaft [g55], 0 days, depth 100 feet

NOTE that an incomplete mine shaft is not "down" from the location in which it was built! You should use "move out" to exit from an incomplete mine shaft, not "move up"!

, Build Mine Shaft [1703], Mining [1700],]

E.8.2 Add Wooden Shoring [1790]

Time to learn: 21 days. Time to use: 7 days. Requires: ten workers [11] (not consumed) Requires: 25 wood [77]

Use of this skill adds wooden shoring to a mine shaft. Normally the chance each month of a cave-in in a mine shaft is (depth/300)%. Wooden shoring reduces this chance by 2.3%.

, Add Wooden Shoring [1790], Mining [1700],]

E.8.3 Add Iron Shoring [1791]

Time to learn: 21 days. Time to use: 14 days. Requires: ten workers [11] (not consumed) Requires: ten iron [79]

Use of this skill adds iron shoring to a mine shaft. Normally the chance each month of a cave-in in a mine shaft is (depth/300)%. Iron shoring reduces this chance by 4.6%.

, Add Iron Shoring [1791], Mining [1700],]

E.8.4 Mine iron [1701]

Usage: USE 1701 [days] Requires: at least ten workers Time to learn: 14 days. Time to use: Variable. Requires: ten workers [11] (not consumed)

Iron is widely used in the making of weapons and armor. It is a required component of many of the weaponsmithing skill uses, and is also used for applying a metal tip to ship rams and battering rams. An iron roof provides a protective cover for battering rams.

The character must be in a mine shaft to USE 1701. The quantity of iron obtained will vary with the depth of the shaft, and whether or not the supply of iron at that level has been mined out. Iron is commonly found at relatively shallow depths.

If [days] is not specified, the character will mine for 7 days.

, Mine iron [1701], Mining [1700],]

E.8.5 Mine gold [1702]

Usage: USE 1702 [days] Requires: at least ten workers Time to learn: 14 days. Time to use: Variable. Requires: ten workers [11] (not consumed)

Gold is widely accepted as a currency in Olympia. Its scarcity lends gold a high value, but it is not so rare that it cannot be used in the everyday exchange of goods. It has a high value, but weighs little, and thus is easy to carry. Gold is also readily assayable to guarantee its value.

The character must be in a mine shaft to USE 1702. The quantity of gold obtained will vary with the depth of the shaft, and whether or not the supply of gold at that level has been mined out.

If [days] is not specified, the character will mine for 7 days.

, Mine gold [1702], Mining [1700],]

E.8.6 Mine mithril [1730]

Usage: USE 1730 [days] Requires: at least ten workers Time to learn: 14 days. Time to use: Variable. Requires: ten workers [11] (not consumed)

Mithril is a magical metal used in the by the ancient elf-lords to construct their bright, fearsome weapons. Today mithril is the component of some rare construction skills.

The character must be in a mine shaft to USE 1730. The quantity of mithril obtained will vary with the depth of the shaft, and whether or not the supply of mithril at that level has been mined out. Mithril is usually found only in very deep mines. As the dangers of poison gasses or collapse of the mine shaft increase with depth, mithril mining tends to be risky.

If [days] is not specified, the character will mine for 7 days.

, Mining [1700],]

E.8.7 Mine gate crystal [1731]

Usage: USE 1731 [days] Requires: at least ten workers Time to learn: 14 days. Time to use: Variable. Requires: ten workers [11] (not consumed)

Gate crystals are rare stones found deep in the earth. Formed at the same time as the world, they bind up great magical powers and are much coveted by magicians.

The character must be in a mine shaft to USE 1731. The quantity of gate crystals obtained will vary with the depth of the shaft, and whether or not the supply of gate crystals at that level has been mined out. Gate crystals are usually found only in very deep mines. As the dangers of poison gasses or collapse of the mine shaft increase with depth, gate crystal mining tends to be risky.

If [days] is not specified, the character will mine for 7 days.

, Mining [1700], Appendix - Skill Dictionary,]

E.9 Trading [1800]

Time to learn: 7 days.

The student of Trading learns to manipulate markets, smuggle, hide money, and all the other tricks and tactics that make a great merchant. While not as prestigious as some, the merchant has the satisfaction of travel and a life made luxurious through great wealth.

The following teachable skills may be studied directly once Trading is known:

    num    skill                              time to learn
    ---    -----                              -------------
    1801   Build Wagons                       two weeks
    1802   Increase Demand                    two weeks
    1803   Decrease Demand                    two weeks
    1804   Increase Supply                    two weeks
    1805   Decrease Supply                    two weeks

The following unteachable skills may be studied directly once Trading is known:

    num    skill                              time to learn
    ---    -----                              -------------
    1830   Smuggle Goods                      two weeks
    1831   Smuggle Men                        two weeks
    1832   Avoid Taxes                        two weeks

The following guild skills may be studied by guildmembers:

    num    skill                              time to learn
    ---    -----                              -------------
    1890   Hide Money                         two weeks
    1891   Hide Item                          two weeks

, Trading [1800], Trading [1800],]

E.9.1 Smuggle Goods [1830]

Usage: USE 1830 <flag> Time to learn: 14 days.

Some places charge each noble who enters a fee for the goods they carry. The amount of the fee is noted in the place’s description:

Imperial City [aa0m], port city, safe haven, entrace fee 25 gold per 1000 weight, entrance fee 10 gold per 100 men, 1 day

In this case, the Imperial City is charging entering nobles 25 gold for every 1000 weight in goods they carry. Goods are any non-mobile items.

Using of this skill, a merchant attempts to avoid this fee by means of hiding goods, bribing the gatekeepers, finding hidden routes into the city, and so on. If the merchant is successful, he will reduce the fee he pays on goods by 50-100%. If he is unsuccessful, there is a 25% he’ll be caught smuggling. In that case, he will pay the full fee plus a fine of 62.5-162.5%. If he cannot pay the entire fine, the gatekeepers will take as much of the fine as possible and turn the merchant away. In any case, if he is caught smuggling the merchant will be detained and questioned for three days.

The chance of success at smuggling goods is:

(50 + # of previous successful goods smuggling attempts)%

To smuggle goods, the merchant must be the stack leader. He will attempt to smuggle into the city all the goods in the stack, not just those in his own inventory. If he fails, the fine (like the fee) is charged against the entire stack. If the merchant cannot completely pay the fine, then other members of the stack will be charged until the fine is paid.

To begin smuggling, the merchant should use this skill with a <flag> of 1. From that time on, the merchant will attempt to smuggle goods every time he enters a city that charges a tariff on goods. To stop smuggling, a merchant should use this skill with a <flag> of 0.

, Smuggle Goods [1830], Trading [1800],]

E.9.2 Smuggle Men [1831]

Usage: USE 1831 <flag> Time to learn: 14 days.

Some places charge each noble who enters a fee for the men in their employ. The amount of the fee is noted in the city description:

Imperial City [aa0m], port city, safe haven, weight fee 25 gold per 1000 weight, entrance fee 10 gold per 100 men, 1 day

In this case, the Imperial City is charging entering nobles 10 gold for every 100 men they employ. In this case, men are any fighting creatures (including beasts).

Using of this skill, a merchant attempts to avoid this fee by means of disguising men, bribing the gatekeepers, finding hidden routes into the city, and so on. If the merchant is successful, he will reduce the fee he pays on smuggled men by 50-100%. If he is unsuccessful, there is a 25% he’ll be caught smuggling. In that case, he will pay the full fee plus a fine of 62.5-162.5%. If he cannot pay the entire fine, the gatekeepers will take as much of the fine as possible and turn the merchant away. In any case, if he is caught smuggling the merchant will be detained and questioned for three days.

The chance of success at smuggling men is:

(50 + # of previous successful man smuggling attempts)%

To smuggle men, the merchant must be the stack leader. He will attempt to smuggle into the city all the men in the stack, not just those in his own inventory. If he fails, the fine (like the fees) is charged against the entire stack. If the merchant cannot completely pay the fine, then other members of the stack will be charged until the fine is paid.

Some places charge a fee for each noble who enters, e.g.,

Imperial City [aa0m], port city, safe haven, entrance fee 1 gold per noble, 1 day

This skill is also used to smuggle nobles, in exactly the same way it is used to smuggle men.

To begin smuggling, the merchant should use this skill with a <flag> of 1. From that time on, the merchant will attempt to smuggle goods every time he enters a city that charges a fee on goods. To stop smuggling, a merchant should use this skill with a <flag> of 0.

, Smuggle Men [1831], Trading [1800],]

E.9.3 Avoid Taxes [1832]

Usage: Automatic when needed. Time to learn: 14 days.

A trader who knows this skill will pay half taxes when he sells an item in a market by selling some of the goods through the grey market, by mis-stating the quantity of goods to the tax collectors, and other shady means.

, Avoid Taxes [1832], Trading [1800],]

E.9.4 Build Wagons [1801]

Usage: MAKE 44 or MAKE 45 or MAKE 46 Time to learn: 14 days. Time to use: 7 or 10 or 14 days. Requires: two or four riding horses [52] Requires: five or ten wood [77] Requires: zero or one iron [79] Produces: wagon [44]

Wagons are used to transport large loads of goods across land routes. There are three types of wagons:

Type           Weight  Ride Capacity  Walk Capacity
----           ------  -------------  -------------
[44] light       500       500           1000
[45] heavy      1100       900           1800
[46] war        1100       900          1800

Wagons travel at riding pace until they exceed their ride capacity. Light and war wagons slow down to walking pace in mountains. All wagons slow to walking pace in swamps.

The three types of wagons have different build requirements:

Type        Build Time      Wood   Riding Horses   Iron
----        ----------     ------  -------------  ------
[44] light       7            5          2          0
[45] heavy      10           10          4          1
[46] war        14           10          4          1

In battle, the war wagon provides a +25 defense bonus for the up to 20 men, but only if the stack is not already inside a fortification.

, Build Wagons [1801], Trading [1800],]

E.9.5 Increase Demand [1802]

Usage: USE 1802 <good> Time to learn: 14 days. Time to use: 7 days.

By use of this skill, a canny merchant can increase the demand for a trade good in any market. By advertising, talking up a product in inns and bars, and similar means, the merchant will whip the buying public into frenzy, driving the price of the trade good up from 6 to 8 percent. The price increase persists until something changes it.

This skill must be used in the market where <good> is being traded. The price for a good cannot be driven higher than twice the base price of the good.

This skill can only be used on items sold or bought by the market itself, i.e., it cannot be used on nobles, and it only works if the good is currently being traded. For example, consider this interaction:

    8: > use 1802 94
    8: That good is not traded here.

[94] is the identifier for woven baskets. Why did this skill use fail? Although a city may be buying baskets on Day 1, by Day 8 someone may have sold the city all the baskets it requires. Consequently, the attempt to influence the market for baskets fails, because no baskets are being traded. (The city typically adds a new demand after the turn is over.)

, Increase Demand [1802], Trading [1800],]

E.9.6 Decrease Demand [1803]

Usage: USE 1803 <good> Time to learn: 14 days. Time to use: 7 days.

By use of this skill, a canny merchant can decrease the demand for a trade good in a market. By intimating that a product is inferior, had bad sideeffects, and similar means, the merchant will discourage the buying public, driving the price of the trade good down from 6 to 8 percent. The price decrease persists until something changes it.

This skill must be used in the market where <good> is being traded. The price for a good cannot be driven lower than half the base price of the good.

This skill can only be used on items sold or bought by the market itself, i.e., it cannot be used on nobles, and it only works if the good is currently being traded. For example, consider this interaction:

    8: > use 1803 94
    8: That good is not traded here.

[94] is the identifier for woven baskets. Why did this skill use fail? Although a city may be buying baskets on Day 1, by Day 8 someone may have sold the city all the baskets it requires. Consequently, the attempt to influence the market for baskets fails, because no baskets are being traded. (The city typically adds a new demand after the turn is over.)

, Decrease Demand [1803], Trading [1800],]

E.9.7 Increase Supply [1804]

Usage: USE 1804 <good> Time to learn: 14 days. Time to use: 7 days.

By use of this skill, a merchant can increase the supply of a trade good in a market. By encouraging more artisans to produce the product, by finding new supplies of the good, and similar means, the merchant will increase the number available for trade by 15-33% of the normal monthly supply.

Note that the monthly supply of a good is likely to be a small number. If you see 80 rare books for sale in a market, that’s probably an entire year’s production. That means that only 10 books are produced monthly (the Olympian year has 8 months), so this skill will increase the number for sale by only 1-3, not 12-24 as you might expect.

Also note that this skill does not change the monthly supply of a good, only the amount currently available in the market.

This skill must be used in the market where <good> is being traded. There can never be more goods for trade than the yearly limit.

This skill can only be used on items sold or bought by the market itself, i.e., it cannot be used on nobles, and it only works if the good is currently being traded. For example, consider this interaction:

    8: > use 1804 94
    8: That good is not traded here.

[94] is the identifier for woven baskets. Why did this skill use fail? Although a city may be buying baskets on Day 1, by Day 8 someone may have sold the city all the baskets it requires. Consequently, the attempt to influence the market for baskets fails, because no baskets are being traded. (The city typically adds a new demand after the turn is over.)

, Increase Supply [1804], Trading [1800],]

E.9.8 Decrease Supply [1805]

Usage: USE 1805 <good> Time to learn: 14 days. Time to use: 7 days.

By use of this skill, a merchant can decrease the supply of a trade good in a market. By encouraging artisans to produce other goods, by closing down supplies of the good, and similar means, the merchant will decrease the number available for trade by 15-33% of the normal monthly supply.

Note that the monthly supply of a good is likely to be a small number. If you see 80 rare books for sale in a market, that’s probably an entire year’s production. That means that only 10 books are produced monthly (the Olympian year has 8 months), so this skill will decrease the number for sale by only 1-3, not 12-24 as you might expect.

Also note that this skill does not change the monthly supply of a good, only the amount currently available in the market.

This skill must be used in the market where <good> is being traded. This skill can only be used on items sold or bought by the market itself,

i.e., it cannot be used on nobles, and it only works if the good is currently being traded. For example, consider this interaction:

    8: > use 1805 94
    8: That good is not traded here.

[94] is the identifier for woven baskets. Why did this skill use fail? Although a city may be buying baskets on Day 1, by Day 8 someone may have sold the city all the baskets it requires. Consequently, the attempt to influence the market for baskets fails, because no baskets are being traded. (The city typically adds a new demand after the turn is over.)

, Decrease Supply [1805], Trading [1800],]

E.9.9 Hide Money [1890]

Usage: USE 1890 <amount> Time to learn: 14 days. Time to use: 3 days.

Many men are greedy to possess the wealth that merchants enjoy, and sometimes use dishonorable methods such as thievery and banditry to separate a merchant from his cash. To avoid this, merchants have developed a wide range of techniques to hide their cash from would-be robbers. False bottoms in storage containers, hidden pockets, magical charms and other such tactics are all well-known to the brotherhood of merchants.

By use of this skill, a merchant hides <amount> money about his person and premises. If the amount given is less than the amount currently hidden, money is removed from hiding. To remove all money from hiding: "USE 1890 0".

Money in hiding will not be captured in battle, nor can it be stolen by thieves. There is a cumulative 1% chance per month that any money in hiding will be lost through forgetfulness, accident or other mishap.

, Trading [1800],]

E.9.10 Hide Item [1891]

Usage: USE 1891 <item> Time to learn: 14 days. Time to use: 3 days.

Many men are greedy to possess the wealth that merchants enjoy, and sometimes use dishonorable methods such as thievery and banditry to separate a merchant from his valuable possessions. To avoid this, merchants have developed a wide range of techniques to hide their possessions from wouldbe robbers. False bottoms in storage containers, hidden pockets, magical charms and other such tactics are all well-known to the brotherhood of merchants.

By use of this skill, a merchant hides a single item somewhere on his person. Amazingly enough, this item can be as big as a dragon (large items are concealed with charms of illusion), but regardless of the size, only a single item may be concealed. To remove an item from hiding: "USE 1891 0".

The item in hiding will not be captured in battle, cannot be stolen by thieves, nor can it be detected by magical means. There is a cumulative 1% chance per month that any item in hiding will be lost through forgetfulness, accident or other mishap.

, Trading [1800], Appendix - Skill Dictionary,]

E.10 Ranger [1900]

A ranger learns skills useful for living and fighting in the wilderness, and for spotting and neutralizing stealthy opponents. Time to learn: 7 days.

The following teachable skills may be studied directly once Ranger is known:

num    skill                              time to learn
---    -----                              -------------
1901   Spot Hidden Noble                  two weeks
1902   Archery                            four weeks
1903   Find Food in Wilderness            two weeks

The following unteachable skills may be studied directly once Ranger is known:

num    skill                              time to learn
---    -----                              -------------
1930   Catch wild horses                  two weeks
1931   Train wild horse to riding horse   two weeks
1932   Train wild horse to warmount       four weeks

The following guild skills may be studied by guildmembers:

num    skill                              time to learn
---    -----                              -------------
1990   Protect Noble                      two weeks

, Ranger [1900], Ranger [1900],]

E.10.1 Spot Hidden Noble [1901]

This skill is used to spot hidden nobles (i.e., a noble using the skill [1290] Conceal Self). Time to learn: 14 days. Time to use: Variable. Time to practice: 7 days. Cost to practice: 21 gold.

For each level of skill a noble has in 1901, he increases his chance to spot a hidden noble when using SEEK. If SEEK is used to find a particular noble, the chance is increased by 5%. If a general SEEK is performed, the chance is increased by 2.5%.

, Spot Hidden Noble [1901], Ranger [1900],]

E.10.2 Protect Noble [1990]

Time to learn: 14 days. Usage: Automatic Time to practice: 7 days. Cost to practice: 7 gold.

This skill is used to guard nobles against attacks from assassins (i.e., a noble using the skill [1292] Assassinate Noble). For each level of skill a noble has in 1990, he reduces the chances of any assassination attempt against any noble in his stack by 5%. This skill is used automatically when an attack occurs; the bodyguard does not have to USE this skill.

The protection of multiple bodyguards is cumulative.

, Protect Noble [1990], Ranger [1900],]

E.10.3 Find Food in Wilderness [1903]

Time to learn: 14 days. Time to use: 7 days.

With this skill, a ranger finds food in the wilderness to help feed his followers. This skill may only be used in wilderness provinces (i.e., with less than 100 peasants).

Each use finds food for a number of men equal to ten men plus the experience level of the ranger in 1903. For example, a ranger with experience level 2 in 1903 would find food for 12 men.

In mountain terrain, the amount of food found is halved. In desert, the amount of food found is quartered.

Food reduces the monthly maintenance costs of all units by 1 gp. Excess food provides no further reduction. To receive the food reduction, a unit must be in the ranger’s inventory.

For example, suppose a ranger with 10 peasants 12 soldiers and experience level 8 in 1903 used 1903 twice in a wilderness forest province. The ranger would find food for 36 men. Having only 22 men in his command, their maintenance costs for the month would be reduced by 22 gp. The excess 14 food would go to waste. Nobles and men stacked with the ranger would not benefit either.

, Find Food in Wilderness [1903], Ranger [1900],]

E.10.4 Catch wild horses [1930]

Usage: CATCH [number of horses] [days] Time: as given or necessary Time to learn: 14 days. Time to use: Variable.

Attempt to catch any wild horses found roaming the plains or pastures. Each day the noble will have a 25% chance of catching a wild horse, if any can be found.

Examples:

       catch        # catch as many horses as possible
       catch 1      # stop after catching 1 wild horse
       catch 0 10   # try to catch horses for 10 days
       catch 1 10   # stop after catching 1 wild horse, or 10 days, whichever
                    # comes first

Plains provinces produce one horse per month, up to a maximum of 100 horses in the province. Pastures produce three horses per month, up to a maximum of twenty horses in the pasture. In all cases, horse-catching ceases when no more horses can be found roaming in the location.

, Catch wild horses [1930], Ranger [1900],]

E.10.5 Train wild horse to riding horse [1931]

Usage: MAKE 52 Requires: one wild horse [51] Produces: one riding horse [52] Time to learn: 14 days. Time to use: 5 days.

Wild horses are first broken to saddle, then trained to respect and obey whatever riders they may serve. Once broken, riding horses become too timid and docile to be of any use in battle.

, Train wild horse to riding horse [1931], Ranger [1900],]

E.10.6 Train wild horse to warmount [1932]

Usage: MAKE 53 Requires: one wild horse [51] Produces: one warmount [53] Time to learn: 28 days. Time to use: 7 days.

Unlike riding horses, war horses must not be broken. The war horse should be trained without destroying any of its spirit. A warmount is also be taught much more than a simple riding animal, as it must know how to defend itself and its rider and to remain calm in the most frightening situations.

, Ranger [1900],]

E.10.7 Archery [1902]

Usage: USE 1902 Time: seven days Time to learn: 28 days. Time to use: 7 days.

Learning 1902 gives the noble a +50 missile attack rating.

Use of 1902 will cause the noble to spend time and effort practicing archery further. Every seven-day use will increase the noble’s missile rating by one point.

Knowledge of Archery also allows the noble to train archers and elite archers, via the TRAIN order.

, Ranger [1900], Appendix - Skill Dictionary,]

E.11 Basic Religion [3500]

When a noble learns a religious skill and becomes a priest, he is also automatically granted knowledge of the skill school Basic Religion. There is no noble point or gold charge for learning this skill. The prayers in this skill are common to all the religions.

The following teachable skills may be studied directly once Basic Religion is known:

num    skill                              time to learn
---    -----                              -------------
3501   Heal                   two weeks
3502   Last Rites                   two weeks
3503   Create Holy Symbol                 two weeks
3504   Dedicate Temple                   two weeks
3505   Devotional Prayer                  two weeks
3506   Bless Follower                   two weeks
3507   Gather Holy Plant                  two weeks
3508   Create Holy Instruction            two weeks
3509   Proselytise                   two weeks

The following unteachable skills may be studied directly once Basic Religion is known:

num    skill                              time to learn
---    -----                              -------------
3530   Resurrect Dead Noble               four weeks
3531   Protection from Aura Blast         two weeks
3532   Hinder Meditation                  two weeks
3533   Vision of Holy Terrain             two weeks

, Basic Religion [3500], Basic Religion [3500],]

E.11.1 Heal [3501]

Usage: USE 3501 <unit> Time to learn: 14 days. Time to use: 7 days. Piety to use: 3. Required to use: Holy plant

This prayer causes the designated noble to be miraculously healed of any sickness. The noble must be a member of the Priest’s religion. The noble and the priest must be in the same location.

, Heal [3501], Basic Religion [3500],]

E.11.2 Last rites [3502]

Usage: USE 3502 <dead noble> Time to learn: 14 days. Time to use: 7 days. Piety to use: 5. Required to use: Holy symbol

Commend a dead noble to his final resting place. This prayer prevents the noble from ever being raised as undead or resurrected.

To use this prayer, the priest must be in a temple of the priest’s religion. If the dead noble is one of the priests’ followers he may be laid to rest without the body. Otherwise, the body is required.

, Last rites [3502], Basic Religion [3500],]

E.11.3 Create Holy Symbol [3503]

Usage: USE 3503 Required Item: 100 gold Time to learn: 14 days. NPs to learn: 1. Time to use: 14 days. Piety to use: 5.

By melting down a large quantity of gold and praying over it continuously for two weeks, a Priest may create a holy symbol consecrated to the Priest’s diety. The holy symbol increases the faith of the Priest who possesses it and is necessary for many prayers. This prayer must be used in a Temple.

A priest may create a holy symbol for another priest of the same religion, and a priest who finds or captures an appropriate holy symbol may use it. Since this skill costs 1 NP to learn, and is needed only to create a holy symbol, a noble might be wise to instead buy a holy symbol from a priest who has already learned this skill.

, Create Holy Symbol [3503], Basic Religion [3500],]

E.11.4 Dedicate Temple [3504]

Usage: USE 3504 Requires: 500 gold Time to learn: 14 days. Time to use: 7 days. Piety to use: 5.

The Dedicate Temple prayer dedicates a new temple to the priest’s god. In a week long ceremony, the priest prays himself into a religious ecstasy and then blesses the structure, inscribes it with various holy symbols, and decorates it with icons. The decorations and other required implements cost 500 gold.

A temple may only be dedicated if it has never been dedicated before (i.e., it is newly built) and if it is in the proper holy terrain. The priest must be the building owner (i.e., the topmost noble in the structure) during the dedication ceremony. There may be only one temple (of any particular religion) in any location.

Once dedicated, a temple may only be entered by stacks headed by an appropriate priest, although they may still be attacked or razed from outside.

, Dedicate Temple [3504], Basic Religion [3500],]

E.11.5 Devotional Prayer [3505]

Usage: USE 3505 Time to learn: 14 days. Time to use: 7 days.

By use of this prayer, a Priest demonstrates his faith to his Deity. Devotional prayer is the cornerstone of every religion. By use of devotional prayer a Priest gains holy power (piety). Pray raises the current Piety of the praying Priest by two.

To pray, the Priest must be in a Temple dedicated to his Deity.

See also ⟨undefined⟩ [Hinder meditation [2838]], page ⟨undefined⟩ and ⟨undefined⟩ [Hinder meditation [3532]], page ⟨undefined⟩.

, Devotional Prayer [3505], Basic Religion [3500],]

E.11.6 Bless follower [3506]

Usage: USE 3506 <unit> Time to learn: 14 days. Time to use: 1 day. Piety to use: 1.

This prayer calls a blessing down upon the designated noble. This noble must be a member of the Priest’s religion. If so, the noble and all the men he commands in battle will receive a defense factor bonus of +10. This effect lasts 30 days.

Priests cannot bless magicians, undead, or any noble possessing undead. , Bless follower [3506], Basic Religion [3500],]

E.11.7 Gather Holy Plant [3507]

Usage: USE 3507 Time to learn: 14 days. Time to use: 7 days. Piety to use: 1.

By means of this prayer, a Priest gathers his holy plant from his current location. A Priest may only gather his holy plant in terrain that is blessed to his religion. The Priest has a 50% chance to find 1 plant.

, Gather Holy Plant [3507], Basic Religion [3500],]

E.11.8 Create Holy Instruction [3508]

Usage: USE 3508 <skill> [# days instruction] [book] Requires: blank scrolls [84], 100 gold Required piety: two points per page Time to learn: 14 days. Time to use: <variable>

Create a book for the given skill, or add days of instruction to an existing book. The book can be used as a source of instruction. Books may be created only for category skills and teachable subskills.

If the number of number of days argument is missing, a default of 7 days is assumed.

There is a cost of 100 gold to begin a new book. This covers the cost of the binding, covers, etc.

This skill requires one page for every seven days instruction written, and consumes two points of piety for every page written. It takes one day to write one page.

The skill must be known in order to be written. The character will still know the skill after writing it.

The optional third argument is the id number of an existing book in the noble’s possession. If this is a book for the given skill, then additional pages will be added to the book. In this case, there is no 100 gold fee to begin a new book.

Only religious prayers may be scribed with this skill.

Examples:

USE 3508 3508 7

Creates a new book with 7 days instruction in 3508. (A new book is created because the optional third argument, the book to add pages to, is not given.)

USE 3508 2041 3 j3k2

Adds three days instruction to book [j3k2], provided that book teaches 2041.

, Create Holy Instruction [3508], Basic Religion [3500],]

E.11.9 Proselytise [3509]

Usage: USE 3509 <number> Time to learn: 14 days. Time to use: Variable. Piety to use: 3.

Attempt to raise postulants from amongst the peasantry. Note that Proselytizing only works where there is already support for a deity: 5 postulants per month may be raised in provinces with a temple to the Priest’s religion.

Postulants have the same statistics a peasants, but have no maintenance costs.

This prayer expends 3 piety regardless of whether any postulants are raised.

, Proselytise [3509], Basic Religion [3500],]

E.11.10 Banish undead [3510]

Usage: USE 3510 <target> [piety] Time to learn: 14 days. Time to use: 7 days.

Banish all undead in the inventory of <target>. Requires one piety point for every undead to be banished. The prayer will banish as many undead as possible, limited only by the amount specified when the prayer is used and by the caster’s piety. No piety is used, nor time expended, if the target has no undead.

, Banish undead [3510], Basic Religion [3500],]

E.11.11 Resurrect dead noble [3530]

Usage: USE 3530 <body> Time to learn: 28 days. Time to use: 30 days. Piety to use: 10 (50). Required to use: Holy symbol, holy plant

Raise a dead noble back to life. To use this prayer, the noble to be resurrected must be a follower of the priest and be in a temple of the priest’s religion. (A priest may be resurrected by any other priest of the same religion.)

If successful, the resurrected noble will be brought back to life and stack beneath the priest. The resurrected noble’s loyalties will remain the same as they were at the time of the noble’s death.

If the priest possesses the body he is resurrecting, then this prayer uses 10 piety. If the priest is not in possession of the body, then 50 piety is required.

, Resurrect dead noble [3530], Basic Religion [3500],]

E.11.12 Protection from Aura Blast [3531]

Usage: Automatic when needed. Time to learn: 14 days.

A priest knowing this spell is protected from aura blasts. In the case he is aura blasted, the aura will be harmlessly diverted away from the priest.

, Protection from Aura Blast [3531], Basic Religion [3500],]

E.11.13 Hinder meditation [3532]

Usage: USE 3532 <target> Time to learn: 14 days. Time to use: 7 days. Piety to use: 3.

Clouds the target’s mind with disturbing images and incoherent thoughts. Whether a mage or a priest, the target will have difficulty achieving the proper mental state during his next attempt at meditation or devotional prayer.

The next meditate spell issued by the target may fail, depending upon the current piety of the prayer:

piety chance of failure —- —————– 1 50% 2 75% 3 90%

The prayer begins with a piety of 3 and loses one point of piety per month. Although multiple priests (and magicians) may attempt to hinder a noble’s meditation, no more than three points of piety/aura may accumulate on a target.

There is a 50% chance that the target will have a disturbing dream that will warn him of his situation.

See also (⟨undefined⟩ [Hinder meditation [2838]], page ⟨undefined⟩).

, Basic Religion [3500],]

E.11.14 Vision of Holy Terrain [3533]

Usage: USE 3533 <province> Time to learn: 14 days. Time to use: 7 days. Piety to use: 3.

By use of this prayer, a priest is given a vision of a distant province. The province viewed must be or contain the priest’s holy terrain, e.g., a priest of Eres can only receive visions of shore provinces, while a priest of Domingo can only receive visions of provinces which contain a hidden location. If a priest tries to vision a province that is not his holy terrain, the prayer will fail, but will still take seven days.

As with all prayers, holy visions cannot cross open water.

This prayer may be blocked by ⟨undefined⟩ [Shroud location from magical scry [2930]], page ⟨undefined⟩ and detected by ⟨undefined⟩ [Detect location scry [2936]], page ⟨undefined⟩.

, Basic Religion [3500], Appendix - Skill Dictionary,]

E.12 Church of Eres [2000]

Time to learn: 21 days. NPs to learn: 1.

Eres is the God of Sea and Wind. He appears to his followers as a tall, strongly built man of advancing years, with a wild mane of white hair. Priests of Eres typically build their temples near the sea and often serve on sailing ships. It is said that a ship with a Priest of Eres on it will never capsize or be lost in a storm. Certainly it is true that Priests of Eres have saved many a ship with their powerful prayers to control the weather.

Temples to Eres must be built in ocean provinces (on islands), or in coastal provinces next to the ocean. Attempts to consecrate temples elsewhere will fail.

The following unteachable skills may be studied directly once Church of Eres is known:

    num    skill                              time to learn
    ---    -----                              -------------
    2030   Fierce wind                        four weeks
    2031   Dissipate storm                    four weeks
    2032   Renew storm strength               four weeks
    2033   Storm fury                         four weeks
    2034   Seize control of storm             four weeks
    2035   Fog of death                       four weeks
    2036   Summon rain                        four weeks
    2037   Summon wind                        four weeks
    2038   Summon fog                         four weeks
    2039   Direct storm                       four weeks
    2040   Summon Water Elemental             four weeks

Note that upon learning Church of Eres [2000], a noble will also automatically be given knowledge of Basic Religion [3500], and can immediately begin learning those prayers.

, Church of Eres [2000], Church of Eres [2000],]

E.12.1 Fierce wind [2030]

Usage: USE 2030 <storm> <target> [piety] Required Item: holy symbol Piety: 1 Time to learn: 28 days. Time to use: 7 days. Piety to use: 1. Required to use: Holy symbol

Use a wind storm to damage a building or a ship. The target will be buffeted with fierce winds, causing an amount of damage to the target equal to the amount of the wind storm’s piety spent in the attack.

For example, "USE 2030 <storm> ce90 5" would cause <storm> to spend five of its remaining piety points against target [ce90], causing five points of damage.

If piety is not specified, the storm will expend all of its remaining energy in the attack. The mage does not need to be in the same location as the storm.

, Fierce wind [2030], Church of Eres [2000],]

E.12.2 Summon rain [2036]

Usage: USE 2036 <piety> ["Name of storm"] Required piety: three or more points (variable) Time to learn: 28 days. Time to use: 7 days. Required to use: Holy symbol, holy plant

Summon a rain weather entity. Rain storms may be summoned from forest or ocean provinces. Once a location has produced a rain storm, another rain storm may not be summoned from that place until the first storm has dissipated.

The storm will have a strength of twice the amount of piety invested in its summoning. Storms lose one point of strength at the end of each month. When a storm’s strength reaches zero, it dissipates.

While the storm exists, the summoner will get a location report for the storm’s province.

, Summon rain [2036], Church of Eres [2000],]

E.12.3 Summon wind [2037]

Usage: USE 2037 <piety> ["Name of storm"] Time to learn: 28 days. Time to use: 7 days. Required to use: Holy symbol, holy plant Required piety: three or more points (variable)

Summon a wind weather entity. Wind may be summoned from plains, ocean, mountain or desert provinces. Once a location has produced a wind entity, another wind may not be summoned from that place until the first has dissipated.

The wind entity will have a strength of twice the amount of piety invested in its summoning. Wind loses one point of strength at the end of each month. When a wind entity’s strength reaches zero, it dissipates.

While the wind exists, the summoner will get a location report the province that it is in.

, Summon wind [2037], Church of Eres [2000],]

E.12.4 Summon fog [2038]

Usage: USE 2038 <piety> ["Name of storm"] Required piety: three or more points (variable) Time to learn: 28 days. Time to use: 7 days. Required to use: Holy symbol, holy plant

Summon a fog. Fog may be summoned from forest, swamp and ocean provinces. Once a location has produced a fog, another fog may not be summoned from that place until the first has dissipated.

The fog will have a strength of twice the amount of piety invested in its summoning. Storms lose one point of strength at the end of each month. When a fog’s strength reaches zero, it dissipates.

While the fog exists, the summoner will get a location report for the province that it is in.

Character sightings will not appear in foggy province’s location report. Characters in sublocations of the province will still be seen, however. Consequently, automatic HOSTILE attacks do not function in provinces when fog is present.

Also, all sublocations of a foggy province are treated as if they are hidden. Thus, if a sublocation is already known to a character, the character will continue to see the sublocation. However, locations which have not been visited by a character will have to be found with EXPLORE before they can be entered (or until the fog passes.)

, Summon fog [2038], Church of Eres [2000],]

E.12.5 Direct storm [2039]

Usage: USE 2039 <storm> <direction or destination> Required aura: 0 Time to learn: 28 days. Time to use: 1 day. Required to use: Holy symbol

Direct an existing storm which the priest has previously summoned to move to an adjacent province at the end of the month. Storms only move once per month, and may only exist at the province level (they may not move into sublocations or buildings.)

The priest need not be in the same location as the storm in order to direct it.

At the end of the month, the storm will move to its new province, and the priest will receive a report for the storm’s new location.

, Direct storm [2039], Church of Eres [2000],]

E.12.6 Dissipate storm [2031]

Usage: USE 2031 <storm> Time to learn: 28 days. Time to use: 7 days. Required to use: Holy symbol

Dissipate an existing storm. The storm must be owned by the priest, and be in the same province as the caster. Half of the remaining strength of the storm will be added to the priest’s current piety rating.

, Dissipate storm [2031], Church of Eres [2000],]

E.12.7 Renew storm strength [2032]

Usage: USE 2032 <storm> <piety> Required piety: as given (variable) Time to learn: 28 days. Time to use: 3 days. Required to use: Holy symbol

Invest piety in an existing storm. The storm must be in the same province as the caster. Each point of the priest’s piety used will add two points of piety to the storm’s strength.

, Renew storm strength [2032], Church of Eres [2000],]

E.12.8 Storm fury [2033]

Usage: USE 2033 <storm> <target> [piety] Required piety: none Time to learn: 28 days. Time to use: 7 days. Required to use: Holy symbol, holy plant

Use a rain storm to deliver a lightning bolt against a target in the province. Piety is the amount of storm strength to deliver in the blast. If piety is not specified, the storm will expend all of its remaining engery in the bolt. The priest does not need to be in the same location as the storm.

The lightning bolt may be delivered against any building or character directly in the storm’s province. It will do an amount of damage equal to the strength of the bolt.

Priests of Eres are immune to lightning bolt attacks; they also protect any structures they are in.

, Storm fury [2033], Church of Eres [2000],]

E.12.9 Seize control of storm [2034]

Usage: USE 2034 <storm> Time to learn: 28 days. Time to use: 7 days. Piety to use: 5. Required to use: Holy symbol, holy plant

Seize control of a storm from another priest. The storm must be in the caster’s province.

, Seize control of storm [2034], Church of Eres [2000],]

E.12.10 Fog of death [2035]

Usage: USE 2035 <storm> <target> [aura] Time to learn: 28 days. Time to use: 7 days. Required to use: Holy symbol, holy plant

Use a fog to kill an enemy target’s men and beasts. Only units whose attack factor + defense factor + missile factor is less than 25 are vulnerable. One unit is killed for each point of the storm’s piety used. If piety is not given, all of the remaining strength of the storm will be used.

The caster does not need to be in the same location as the fog to cast this spell.

, Church of Eres [2000],]

E.12.11 Summon Water Elemental [2040]

Usage: USE 2040 Time to learn: 28 days. Time to use: 7 days. Piety to use: 7. Required to use: Holy symbol, holy plant

This spell summons a water elemental into a ocean or coastal province. The water elemental will remain in the province for up to one year.

If a stack or ship enters the province during that time, the water elemental will attack and attempt to drown as many as 10 men (or beasts). Each man (or beast) attacked by the water elemental has a 50% chance of being washed drowned.

The water elemental’s targets are selected exactly as in combat. Due to their superior strength, nobles cannot be drowned. If attacked, they will instead suffer a 1-100 point wound.

After attacking once, the water elemental will disperse.

Water elementals completely ignore stacks which have a priest of Eres.n , Church of Eres [2000], Appendix - Skill Dictionary,]

E.13 Church of Anteus [2100]

Time to learn: 21 days. NPs to learn: 1.

Anteus is the God of Fire and Earth. He appears to his followers as a short, stout man with a bulbous nose, a long cap and he is often seen carrying a pickaxe or hammer. Priests of Anteus build their temples in the mountains and are often seen working with miners. It is said that dwarves are the special minions of Anteus, and that a party which contains a Priest of Anteus will never be attacked by dwarves and will have free passage through dwarven tunnels.

Temples to Anteus must be in mountain provinces. Attempts to consecrate temples elsewhere will fail.

The following unteachable skills may be studied directly once Church of Anteus is known:

    num    skill                              time to learn
    ---    -----                              -------------
    2130   Find Mountain Trail                four weeks
    2131   Obscure Mountain Trails            four weeks
    2132   Improve Mining                     four weeks
    2133   Conceal Mine                       four weeks
    2134   Protect Mine                       four weeks
    2135   Bless Building                     four weeks
    2136   Weaken Building                    four weeks
    2137   Boulder Trap                       four weeks

Note that upon learning Church of Anteus [2100], a noble will also automatically be given knowledge of Basic Religion [3500], and can immediately begin learning those prayers.

, Church of Anteus [2100], Church of Anteus [2100],]

E.13.1 Find Mountain Trail [2130]

Usage: USE 2130 Duration: 1 day Time to learn: 28 days. Time to use: 1 day. Piety to use: 1. Required to use: Holy symbol

By means of this prayer, the Priest finds a faster trail into a mountain province, speeding the movement of the stack he is in as it travels. A stack under the influence of this prayer will travel into a mountain province in half the normally required time (e.g., under normal conditions, in 5 days instead of 10).

The Priest must be the stack leader for this prayer to be effective. The stack must move immediately after the casting of the spell for the prayer to be effective. The prayer has no effect on flying stacks.

, Find Mountain Trail [2130], Church of Anteus [2100],]

E.13.2 Obscure Mountain Trails [2131]

Usage: USE 2131 Required Item: holy symbol Piety: 5 Duration: 30 days Time to learn: 28 days. Time to use: 7 days. Piety to use: 5.

By use of this prayer, a Priest of Anteus obscures the mountain trails leading into a mountain province. Units moving into the province will have their movement time doubled. This effect will last for 30 days.

This prayer will not affect units moving out of the province, or other Priests of Anteus.

, Obscure Mountain Trails [2131], Church of Anteus [2100],]

E.13.3 Improve Mining [2132]

Usage: USE 2132 <mine shaft> Required Item: holy plant, holy symbol Piety: 3 Duration: 30 days. Time to learn: 28 days. Time to use: 7 days. Piety to use: 3.

By means of this prayer, the priest improves the efficiency of the mining operations in a particular mine shaft. While the spell lasts, production in the mine shaft will be improved by 50%. Multiple uses of this prayer are not cumulative.

To use this prayer, the priest must be in the mine shaft to be blessed.

, Improve Mining [2132], Church of Anteus [2100],]

E.13.4 Conceal Mine [2133]

Usage: USE 2133 <mine shaft> Required Item: holy symbol, holy plant Piety: 5 Time to learn: 28 days. Time to use: 30 days. Piety to use: 5.

By use of this prayer, a priest permanently conceals the entrance to a mine shaft. The priest’s faction will have knowledge of the mine’s location, but other factions must located it through exploration.

To use this prayer, the priest must be in the mine shaft to be concealed.

, Conceal Mine [2133], Church of Anteus [2100],]

E.13.5 Protect Mine [2134]

Usage: USE 2134 <mine> Required Item: holy symbol Piety: 3 Duration: 30 days. Time to learn: 28 days. Time to use: 7 days. Piety to use: 3.

By means of this prayer, the priest protects a mine against cave-ins and other disasters. Multiple uses of this prayer are not cumulative.

To use this prayer, the priest must be in the mine shaft to be protected.

, Protect Mine [2134], Church of Anteus [2100],]

E.13.6 Bless Building [2135]

Usage: USE 2135 <building> Required Item: holy symbol Piety: 3 Duration: 30 days Time to learn: 28 days. Time to use: 7 days. Piety to use: 3.

Priests of Anteus may use this prayer to increase the strength of a fortification. For the following month, the blessed fortification will have extra resistance to damage. Attacks against the fortification will have only a 50% chance to do damage. This prayer also provides complete protection against the prayer Weaken Fortification.

Multiple uses of this prayer are not cumulative.

To use this prayer, the priest must be in the same province as the building to be protected.

, Bless Building [2135], Church of Anteus [2100],]

E.13.7 Weaken Building [2136]

Usage: USE 2136 <building> <piety> Required Item: holy symbol, holy plant Piety: Variable Time to learn: 28 days. Time to use: 3 days.

By this prayer, the priest of Anteus causes the walls of a fortification to weaken and crumble. This prayer does 3 points of damage to a fortification for each point of piety used. If no piety is specified, the default piety of 5 points (causing 15 points damage to the structure) will be used.

This prayer will not affect buildings that have been blessed by a Priest of Anteus.

To use this prayer, the priest must be inside or in the same location as the building to be weakened.

, Church of Anteus [2100],]

E.13.8 Boulder Trap [2137]

Usage: USE 2137 Time to learn: 28 days. Time to use: 7 days. Piety to use: 7.

This spell creates a boulder trap in a mountain province. The trap will remain in the province for up to one year. If a stack enters the province during that time, the trap will collapse and kill as many as 10 men (or beasts). Each man (or beast) hit by a boulder has a 50% chance of being crushed.

The trap’s targets are selected exactly as in combat. Due to their superior footspeed, nobles will not be automatically killed by the trap. If hit, they will instead suffer a 1-100 point wound.

After attacking one stack, the trap will be dispersed.

Priest of Anteus easily spot and avoid boulder traps. Wild beasts moving through the province also avoid boulder traps.

, Church of Anteus [2100], Appendix - Skill Dictionary,]

E.14 Church of Dol [2200]

Time to learn: 21 days. NPs to learn: 1.

Dol is the God of the Animals, called the Beastmaster. He appears to his followers a tall, well-built man with the head of a bull. Priests of Dol build their temples in the plains, and are often seen working with beasts. It is said that the Priests of Dol will not be attacked by any wild beast.

Temples to Dol must be in plains provinces. Attempts to consecrate temples elsewhere will fail.

The following unteachable skills may be studied directly once Church of Dol is known:

    num    skill                              time to learn
    ---    -----                              -------------
    2231   Bird spy                           four weeks
    2232   Capture beasts           four weeks
    2233   Breed beasts                       four weeks
    2234   Detect Beasts                      four weeks
    2235   Snake Trap                         four weeks

Note that upon learning Church of Dol [2200], a noble will also automatically be given knowledge of Basic Religion [3500], and can immediately begin learning those prayers.

, Church of Dol [2200], Church of Dol [2200],]

E.14.1 Detect Beasts [2234]

Usage: USE 2234 Time to learn: 28 days. Time to use: 3 days. Piety to use: 2. Required to use: Holy symbol

This prayer detects all wild beasts within one province of the Priest. This includes wild beasts that are under a players control (such as orc units) but does not detect trained beasts in a unit’s inventory. This prayer does not detect wild horses, or beasts that are in hiding.

, Detect Beasts [2234], Church of Dol [2200],]

E.14.2 Snake Trap [2235]

Usage: USE 2235 Time to learn: 28 days. Time to use: 7 days. Piety to use: 7. Required to use: Holy symbol, holy plant

This spell creates a nest of poisonous snakes in a plains province. The snakes will remain in the province for up to one year. If a stack enters the province during that time, the snakes will attack the stack and attempt to bite as many as 10 men (or beasts). Each man (or beast) bitten has a 50% chance of dying from the poison.

The snakes’ targets are selected exactly as in combat. Due to their superior health, nobles will not be automatically killed by the poison. If attacked, they will instead suffer a 1-100 point wound.

After attacking one stack, the snakes will disperse.

Snakes completely ignore stacks which contain a priest of Dol. Wild beasts moving through the province also avoid snake traps.

, Snake Trap [2235], Church of Dol [2200],]

E.14.3 Bird spy [2231]

Usage: USE 2231 <direction or destination> Time to learn: 28 days. Time to use: 3 days. Piety to use: 1. Required to use: Holy symbol

The priest employs the use of a bird to fly into an nearby location and return with a report. A direction may be given, as with the MOVE order, or a destination may be specified. The destination may be any sublocation in the same province, or any neighboring location.

For example, a bird could fly from inside a castle in a province to spy on a tower in a city in the same province. The bird could spy on a neighboring province, but not enter sublocations in the neighboring province.

, Bird spy [2231], Church of Dol [2200],]

E.14.4 Capture beasts [2232]

Usage: USE 2232 <stack> Piety: variable Time to learn: 28 days. Time to use: 7 days. Required to use: Holy symbol, holy plant

The wild and fantastic beasts that populate the Olympia wilderness are completely inimical to man. Only by the holy power of Dol can they be captured, trained, and put to use. By means of this prayer, the priest of Dol mesmerizes a wild beast and causes the beast to join the priest.

The piety required to perform this amazing feat depends upon the power and ferocity of the beast to be captured. 1 point of piety is required, plus 1 point for each 50 points of the beast’s attack and defense factors (rounded). To capture a beast with an attack factor of 30 and a defense factor of 30 would require 2 piety. To capture a beast with an attack factor of 50 and a defense factor of 30 would cost 3 piety. If a priest lacks the piety to capture a beast, the prayer will fail after seven days. This prayer will capture one beast at a time.

The target of this skill is a stack. If the stack contains several types of beasts, the beast captured will be selected at random. Beasts may be captured from any stack; this prayer may be used to lure beasts from another noble’s inventory.

, Church of Dol [2200],]

E.14.5 Breed beasts [2233]

Usage: BREED <item> <item> Piety: 3 or 0 (see below) Time to learn: 28 days. Time to use: 7 days. Required to use: Holy symbol

Attempt to breed two animals together to produce an offspring. Not all animals produce successful offspring. Occasionally, cross-breeding different beasts will produce strange results.

For example:

To breed two wild horses together:

BREED 51 51

To attempt a cross-breed of a horse and an ox:

BREED 51 76

Wild horses [51], riding horses [52] and warmounts [53] may be used interchangably with BREED.

When two like normal animals are bred together (e.g., BREED 52 52 to breed two riding horses), no piety is required. Natural instincts, along with the guidance of the priest, is all that is needed to consumate the union. Such unions are always successful.

The following animals are considered normal:

[52] riding horse

[53] warmount

[76] ox

Breeding dissimilar animals (e.g., BREED 52 76 to cross a horse and an ox) or breeding fantastic animals (e.g., BREED 286 286 to breed two dragons) is a different matter. Such unions require the calming effect of Dol to have any chance of success. This use costs the priest 3 piety and requires a holy symbol.

But even under the power of Dol, cross-breeding and breeding fantastic beasts is a difficult matter. Some cross-breeds are infertile and can produce no offspring. Even if the cross-breed is fertile (or if two fantastic creatures are being bred together), there is only a 75% chance of an offspring. Further, there is a small chance that one of the breeding parents will be killed in the attempt.

, Church of Dol [2200], Appendix - Skill Dictionary,]

E.15 Church of Timeid [2300]

Time to learn: 21 days. NPs to learn: 1.

Timeid is the Goddess of Plants. She appears to her followers as a tall, lithe woman dressed in a silky green dress. Priests of Timeid build their temples in the woods, and live solitary lives among the trees. It is said that the Priests of Timeid are beloved of all woodland creatures, and most especially by the elves.

Temples to Timeid must be in forest provinces. Attempts to consecrate temples elsewhere will fail.

The following unteachable skills may be studied directly once Church of Timeid is known:

    num    skill                              time to learn
    ---    -----                              -------------
    2330   Find Forest Trail                  four weeks
    2331   Obscure Forest Trails              four weeks
    2332   Improve Forestry                   four weeks
    2333   Reveal Forest                      four weeks
    2334   Create Dense Thicket               four weeks
    2335   Create Deadfall                    four weeks
    2336   Recruit Elves                      four weeks

Note that upon learning Church of Timeid [2300], a noble will also automatically be given knowledge of Basic Religion [3500], and can immediately begin learning those prayers.

, Church of Timeid [2300], Church of Timeid [2300],]

E.15.1 Find Forest Trail [2330]

Usage: USE 2330 Duration: 1 day Time to learn: 28 days. Time to use: 1 day. Piety to use: 1. Required to use: Holy symbol

By means of this prayer, the Priest finds a faster trail into a forest province, speeding the movement of the stack he is in as it travels. A stack under the influence of this prayer will travel into a forest province in half the normally required time (e.g., under normal conditions, in 4 days instead of 8).

The Priest must be the stack leader for this prayer to be effective. The stack must move immediately after the casting of the spell for the prayer to be effective. The prayer has no effect on flying stacks.

, Find Forest Trail [2330], Church of Timeid [2300],]

E.15.2 Obscure Forest Trails [2331]

Usage: USE 2331 Duration: 30 days Time to learn: 28 days. Time to use: 7 days. Piety to use: 5. Required to use: Holy symbol

By use of this prayer, a Priest of Timeid obscures the forest trails leading into a forest province. Units moving into the province will have their movement time doubled. This effect will last for 30 days.

This prayer will not affect units moving out of the province, or other Priests of Timeid.

, Obscure Forest Trails [2331], Church of Timeid [2300],]

E.15.3 Improve Forestry [2332]

Usage: USE 2332 target Duration: 30 days. Time to learn: 28 days. Time to use: 7 days. Piety to use: 3. Required to use: Holy symbol, holy plant

By means of this prayer, the priest improves the efficiency of a logging operation. While the spell lasts, the rate at which wood will collected at the logging operation will be improved by 50%. There will also be 50% more wood available for collection. Multiple uses of this prayer are not cumulative.

This prayer improves the production of all types of wood, including [77] wood, [68] yew, and [70] mallorn wood. The target of this prayer is the province, yew grove, or mallorn grove to be blessed. If no target is given, this prayer blesses the priest’s current location.

To use this prayer, the priest must be in the same province as the logging operation to be blessed.

, Improve Forestry [2332], Church of Timeid [2300],]

E.15.4 Reveal Forest [2333]

Usage: USE 2333 Time to learn: 28 days. Time to use: 7 days. Piety to use: 2. Required to use: Holy symbol

This prayer reveals to the priest any hidden sublocations or exits in the current province. This prayer works only in forest provinces.

, Reveal Forest [2333], Church of Timeid [2300],]

E.15.5 Create Dense Thicket [2334]

Usage: USE 2334 <fortification> Duration: 30 days Time to learn: 28 days. Time to use: 7 days. Piety to use: 7. Required to use: Holy symbol, holy plant

Using this prayer, the priest causes a dense thicket of thorns to grow over a fortification. This thicket will last thirty days and will add 25 to the fortification’s defense factor.

Damage to the fortification from attack will always occur first to the thicket of thorns.

Multiple uses of this spell are not cumulative.

, Create Dense Thicket [2334], Church of Timeid [2300],]

E.15.6 Create Deadfall [2335]

Usage: USE 2335 Duration: 300 days Time to learn: 28 days. Time to use: 7 days. Piety to use: 7. Required to use: Holy symbol, holy plant

This spell creates a deadfall trap in a forest province. The trap will remain in the province for up to one year. If a stack enters the province during that time, the trap will collapse and kill as many as 10 men (or beasts). Each man (or beast) hit by a deadfall has a 50% chance of being crushed.

The trap’s targets are selected exactly as in combat. Due to their superior footspeed, nobles will not be automatically killed by the trap. If hit, they will instead suffer a 1-100 point wound.

After attacking one stack, the trap will be dispersed.

Priests of Timeid easily spot and avoid deadfall traps. Wild beasts also avoid deadfalls.

, Church of Timeid [2300],]

E.15.7 Recruit Elves [2336]

Usage: USE 2336 Duration: variable Time to learn: 28 days. Time to use: 7 days. Piety to use: 3. Required to use: Holy symbol

By praying for a week in the forest, the priest of Timeid who uses this spell entices a small number of elves to come out of hiding and join him. Elves being flighty creatures, some number of the elves stacked with a priest of Timeid will leave at the end of each month.

Each use of this spell entices 5-10 elves to join the priest. This prayer works only in the woods. At the end of each month, there is a chance for each elf to wander off. If the priest ends the month in a forest province, then each elf has a 25% chance to wander off. If the priest ends the month in any other type of province, each elf has a 50% chance to wander off.

, Church of Timeid [2300], Appendix - Skill Dictionary,]

E.16 Church of Ham [2400]

Time to learn: 21 days. NPs to learn: 1.

Ham is the God of Laughter and Devilment. He appears as a laughing young man dressed in the gay colorful garb of a court jester. Priests of Ham traditionally build their Temples in cities, where they serve as centers of entertainment. It is said that the followers of Ham never lose heart, and that no act goes unseen by Ham’s keen-eyed Priests.

Temples to Ham must be in cities. Attempts to consecrate temples elsewhere will fail.

The following unteachable skills may be studied directly once Church of Ham is known:

    num    skill                              time to learn
    ---    -----                              -------------
    2430   Enchant City Guards                four weeks
    2431   Urchin Spy                         four weeks
    2432   Draw Crowds                        four weeks
    2433   Arrange Mugging                    four weeks
    2434   Receive vision                     four weeks
    2436   Shroud location from magical scry  two weeks

Note that upon learning Church of Ham [2400], a noble will also automatically be given knowledge of Basic Religion [3500], and can immediately begin learning those prayers.

, Church of Ham [2400], Church of Ham [2400],]

E.16.1 Receive vision [2434]

Usage: USE 2434 <target> Time to learn: 28 days. Time to use: 10 days. Piety to use: 3.

Receive a vision of a character, place or unique item. A character may only see a vision for a particular target once. Each attempt to receive a vision has a 50% chance of success.

As with all prayers, Receive Vision will fail if the target is across an ocean from the priest. This prayer cannot be blocked by ⟨undefined⟩ [Shroud location from magical scry [2930]], page ⟨undefined⟩ but will be detected by ⟨undefined⟩ [Detect location scry [2936]], page ⟨undefined⟩.

, Receive vision [2434], Church of Ham [2400],]

E.16.2 Enchant City Guards [2430]

Usage: USE 2430 <target> Time to learn: 28 days. Time to use: 7 days. Piety to use: 4. Required to use: Holy symbol, holy plant

This prayer increases the gullibility of the city guard and increases the persuasiviness of the priest of Ham, allowing the priest of Ham to suggest to the City Guard that a particular unit was involved in a crime. This will result in the given unit being detained for questioning for 7 days. During that time the unit will be unable to execute commands (as if it were a prisoner).

While being detained, the unit can still make auto-attacks against units on it’s attack list, defend itself against attacks against, and participate as an ally in combat.

Both the priest of Ham and the target unit must be in the same city for the duration of the prayer. Other priests of Ham are not affected by this prayer.

, Enchant City Guards [2430], Church of Ham [2400],]

E.16.3 Urchin Spy [2431]

Usage: USE 2431 <target> Cost: $50 Time to learn: 28 days. Time to use: 1 day. Piety to use: 3. Required to use: Holy symbol

This prayer increases the persuasiveness of the priest of Ham, allowing the priest of Ham to suggest that an urchin follow a particular unit around the city for a week and report to the priest anything interesting he discovers. The priest also gives the urchin $50 to use for bribing shopkeepers, city guards, etc.

During the 7 days that the urchin spies on the target, the Priest will be informed of every action the target attempts, although not whether the action succeeds or fails. Both the target and the priest must be in the city in which the prayer was said for the urchin to report.

Both the priest of Ham and the target unit must be in the same city for the duration of the prayer. Other priests of Ham are not affected by this prayer.

Multiple uses of this prayer result in multiple urchins spying on the same target. They do not increase the duration of the spying.

, Urchin Spy [2431], Church of Ham [2400],]

E.16.4 Draw Crowds [2432]

Usage: USE 2432 Time to learn: 28 days. Time to use: 7 days. Piety to use: 3. Required to use: Holy symbol, holy plant

By use of this prayer, the priest increases the size of the crowds in a city. The bigger crowds will remain for 30 days. This prayer is only effective in a city.

While the bigger crowds are in effect, recruiting of peasants and postulants is increased by 50%. Units using [1201] Petty Thievery will also gain 50% more income.

, Draw Crowds [2432], Church of Ham [2400],]

E.16.5 Arrange Mugging [2433]

Usage: USE 2433 <victim> Time to learn: 28 days. Time to use: 7 days. Piety to use: 7. Required to use: Holy symbol, holy plant

Using this prayer, the priest arranges for muggers to attack the hapless victim. The muggers will rob the victim of some of his gold, and possibly rough him up.

The victim will lose 20-50% of his gold. Note that the muggers keep the gold for themselves. There is a 25% chance that the victim will take 5-15 points of damage in the attack, but the muggers are careful never to kill anyone.

Both the priest of Ham and the target unit must be in the same city for the duration of the prayer. Other priests of Ham are not affected by this prayer.

, Church of Ham [2400],]

E.16.6 Shroud location from magical scry [2436]

Usage: USE 2436 [piety] Time to learn: 14 days. Time to use: 3 days.

Creates a shroud over the priest’s current province. The shroud hinders attempts to view the location by magical scrying or holy visions (except for orbs, and ⟨undefined⟩ [Receive vision [2434]], page ⟨undefined⟩, which succeed even if a shroud exists).

The shroud is invested with twice the piety used in its creation. A mage or priest attempting to view the location from a distance with a scry spell or holy vision must expend a greater amount of aura (piety) than the shroud possesses to succeed.

The shroud loses one point of piety at the end of each month. Successive uses of 2436 are additive. For example, USE 9448 1 followed by USE 2436 3 will result in a location shroud with an piety of eight points.

The default piety to use is one point.

, Church of Ham [2400], Appendix - Skill Dictionary,]

E.17 Church of Kireus [2500]

Time to learn: 21 days. NPs to learn: 1.

Kireus is the patron of smiths. He appears as a thick-set man dressed in blacksmith’s garb, with bushy red hair and a beard singed by smithing fires. Priests of Kireus traditionally build their Temples in swamps where they have easy access to the coal they need for their forges, and where fires do not burn out of hand. Priests of Kireus often associate with engineers and weapon smiths. It is said that no priest of Kireus can be harmed by a blade and no wall built of stone can bar them.

Temples to Kireus must be in swamps. Attempts to consecrate temples elsewhere will fail.

The following unteachable skills may be studied directly once Church of Kireus is known:

   num    skill                              time to learn
   ---    -----                              -------------
   2530   Improve Quarrying                  four weeks
   2531   Improve Smithing                   four weeks
   2532   Edge of Kireus                     four weeks
   2533   Create Mithril                     four weeks
   2534   Set Quicksand Trap                 four weeks

Note that upon learning Church of Kireus [2500], a noble will also automatically be given knowledge of Basic Religion [3500], and can immediately begin learning those prayers.

, Church of Kireus [2500], Church of Kireus [2500],]

E.17.1 Improve Quarrying [2530]

Usage: USE 2530 Duration: 30 days. Time to learn: 28 days. Time to use: 7 days. Piety to use: 3. Required to use: Holy symbol, holy plant

By means of this prayer, the priest improves the efficiency of collecting stone in the priest’s current province. While the spell lasts, the rate at which stone will collected in the province will be improved by 100%. There will also be 100% more stone available for collection. Multiple uses of this prayer are not cumulative.

, Improve Quarrying [2530], Church of Kireus [2500],]

E.17.2 Improve Smithing [2531]

Usage: USE 2531 target Duration: 30 days Time to learn: 28 days. Time to use: 7 days. Piety to use: 3. Required to use: Holy symbol, holy plant

By means of this prayer, the priest improves the efficiency of a smith. A smith blessed by this prayer can produce two longswords per day or two plate armors per day, instead of the usual one per day. The smith must still have the skill [1106] Weaponsmithing, and must have the ingredients for the second item.

The smith and the priest must be in the same location at the time the prayer is cast.

Multiple uses of this prayer are not cumulative.

, Improve Smithing [2531], Church of Kireus [2500],]

E.17.3 Edge of Kireus [2532]

Usage: USE 2532 Duration: 7 days Time to learn: 28 days. Time to use: 7 days. Piety to use: 5. Required to use: Holy symbol

By use of this prayer, the priest of Kireus calls down a blessing upon all the edged weapons borne by the troops stacked with him. This blessing causes the edges of the weapons to glow, and magically maintains their sharpness in battle.

The Edge of Kireus increases the attack factor of all troops bearing edged weapons stacked with the priest by 25. The following troops bear edged weapons:

[14] knight [15] paladin [20] lt. foot [24] pirate [37] heavy foot [43] cavalier This prayer lasts 7 days. Multiple uses of this prayer are not cumulative. , Edge of Kireus [2532], Church of Kireus [2500],]

E.17.4 Create Mithril [2533]

Usage: USE 2533 Also Required: 10 units iron, 100 units gold Time to learn: 28 days. Time to use: 20 days. Piety to use: 10. Required to use: Holy symbol

By use of this prayer, the priest of Kireus creates one unit of the magical metal mithril. The rarest of metals, mithril is prized by wizards and weaponmakers. Each unit of mithril created consumes ten units of iron and one hundred units of gold from the priest’s inventory.

, Church of Kireus [2500],]

E.17.5 Set Quicksand Trap [2534]

Usage: USE 2534 Time to learn: 28 days. Time to use: 7 days. Piety to use: 7.

This spell creates a quicksand trap in a swamp province. The trap will remain in the province for up to one year. If a stack enters the province during that time, the quicksand will engulf as many as 10 men (or beasts). Each man (or beast) who falls into the quicksand has a 50% chance of being drowned.

The trap’s targets are selected exactly as in combat. Due to their quick wits, nobles will not be automatically killed by the trap. If hit, they will instead suffer a 1-100 point wound.

After attacking one stack, the trap will be dispersed.

Priest of Kireus easily spot and avoid quicksand traps. Wild beasts also avoid quicksand traps.

, Church of Kireus [2500], Appendix - Skill Dictionary,]

E.18 Church of Halon [2600]

Time to learn: 21 days. NPs to learn: 1.

Halon was the leader of the Faithful Eight when they rejected Tor’s suggestion to overthrow Sol. He appears a tall, strikingly handsome man with glossy black hair and penetrating grey eyes. Priests of Halon build their temples where men congregate, and are often found as advisors and confidants of great leaders. It is said that Halon’s touch upon a priest makes him a natural commander of men.

Temples to Halon must be within a city. Attempts to consecrate temples elsewhere will fail.

The following unteachable skills may be studied directly once Church of Halon is known:

   num    skill                              time to learn
   ---    -----                              -------------
   2630   Calm Angry Peasants                four weeks
   2631   Improve Charisma                   four weeks
   2632   Mesmerize Crowd                    four weeks
   2633   Improve Taxes                      four weeks
   2634   Guard Loyalty                      four weeks
   2635   Instill Fanaticism                 four weeks

Note that upon learning Church of Halon [2600], a noble will also automatically be given knowledge of Basic Religion [3500], and can immediately begin learning those prayers.

, Church of Halon [2600], Church of Halon [2600],]

E.18.1 Calm Angry Peasants [2630]

Usage: USE 2630 Time to learn: 28 days. Time to use: 7 days. Piety to use: 2. Required to use: Holy symbol

The priest of Halon first prays and then gives an impassioned speech urging angry peasants to calm down and go back to work. At the end of 7 days, all angry peasant mobs in the priest’s current location (including those stacked with nobles) will disappear. Angry peasants in the inventory of nobles will also disappear.

Angry peasants in the possession or stacked immediately under a priest of Halon are not affected.

, Calm Angry Peasants [2630], Church of Halon [2600],]

E.18.2 Improve Charisma [2631]

Usage: USE 2631 <unit> Time to learn: 28 days. Time to use: 7 days. Piety to use: 3. Required to use: Holy symbol, holy plant

By means of this prayer, the priest of Halon magically enhances a unit’s charisma for 30 days. The unit’s improved charisma will permit it to command double the number of men in combat it can normally command. The unit being blessed must be a follower of Halon.

The unit being blessed and the priest of Halon must be in the same location for the duration of the prayer. Multiple uses of this prayer are not cumulative. This prayer does not increase the number of beasts a unit can control in combat.

, Improve Charisma [2631], Church of Halon [2600],]

E.18.3 Mesmerize Crowd [2632]

Usage: USE 2632 Time to learn: 28 days. Time to use: 1 day. Piety to use: 5. Required to use: Holy symbol

Using this prayer, a priest of Halon mesmerizes a crowd for seven days with a hypnotic speech. All nobles in the same location as the priest have a 35% chance of standing dazed for seven days. During this time they will not accomplish any orders, but will still auto-execute attacks and participate in combat as allies.

This prayer does not affect priests of Halon. The priest giving the speech is also occupied for 7 days following the casting of the prayer.

, Mesmerize Crowd [2632], Church of Halon [2600],]

E.18.4 Improve Taxes [2633]

Usage: USE 2633 Time to learn: 28 days. Time to use: 7 days. Piety to use: 3. Required to use: Holy symbol

By use of magical leadership, the priest of Halon convinces the local inhabitants to contribute higher taxes without complaint. For the month in which the prayer was used, the local province will increase its taxation by 100%. Multiple uses of this prayer are not cumulative.

, Improve Taxes [2633], Church of Halon [2600],]

E.18.5 Guard Loyalty [2634]

Usage: USE 2634 Time to learn: 28 days. Time to use: 1 day. Piety to use: 2. Required to use: Holy symbol

For thirty days, this prayer protects a unit from all attempts to change its loyalty. This prayer also protects a unit with contract, fear or summoned loyalty from deserting when their loyalty reaches zero.

The priest and the target unit must be in the same location when the prayer is pronounced.

, Church of Halon [2600],]

E.18.6 Instill Fanaticism [2635]

Usage: USE 2635 [num] Required Item: up to 10 [12] soldiers Time to learn: 28 days. Time to use: 20 days. Piety to use: 6. Required to use: Holy symbol, holy plant

During this prayer, the Priest of Halon will preach to his soldiers about the importance of loyalty and the holy nature of their leaders. The result will be to convert the soldiers into fanatics [28]. Fanatics have an AF of 10 and a DF of 5.

If the optional argument [num] specifies the number of soldiers to convert. The default (and maximum number to convert) is 10.

Fanatics have no maintenance cost but cannot be transfered to other nobles.

, Church of Halon [2600], Appendix - Skill Dictionary,]

E.19 Church of Domingo [2700]

Time to learn: 21 days. NPs to learn: 1.

Of all the Faithful Eight, least is known about Domingo. The ancient texts from the time of the Rebellion do not mention him, and there is no record of his typical appearance. Some say that he is one of the children of Sol who followed Tor, returned now to Olympia to spy upon the man and the Faithful. Because of this, Priests of Domingo build their temples in secret places and are cautious about announcing their faith. It is said that the priests of Domingo know all the secret places in the world.

Priests of Domingo see through fogs and mists. Priests of Domingo also get a 20% bonus to find hidden locations through EXPLORE.

Temples to Domingo must be in a hidden location. Attempts to consecrate temples elsewhere will fail.

The following unteachable skills may be studied directly once Church of Domingo is known:

    num    skill                              time to learn
    ---    -----                              -------------
    2730   Find Hidden Location               four weeks
    2731   Conceal Wilderness Location        four weeks
    2732   Mists of Concealment               four weeks
    2733   Train Ninjas                       four weeks

Note that upon learning Church of Domingo [2700], a noble will also automatically be given knowledge of Basic Religion [3500], and can immediately begin learning those prayers.

, Church of Domingo [2700], Church of Domingo [2700],]

E.19.1 Find Hidden Location [2730]

Usage: USE 2730 Time to learn: 28 days. Time to use: 7 days. Piety to use: 2. Required to use: Holy symbol

Domingo mystically reveals to his priest any hidden locations in the priest’s current location.

, Find Hidden Location [2730], Church of Domingo [2700],]

E.19.2 Conceal Wilderness Location [2731]

Usage: USE 2731 [location] Time to learn: 28 days. Time to use: 30 days. Piety to use: 10. Required to use: Holy symbol, holy plant

By use of this prayer, the priest permanently conceals the entrance to a sublocation. The sublocation must be in a wilderness province, although it will remain hidden should province later become civilized. The priest must be in the sublocation or surrounding province during the casting of the prayer.

, Conceal Wilderness Location [2731], Church of Domingo [2700],]

E.19.3 Mists of Concealment [2732]

Usage: USE 2732 Time to learn: 28 days. Time to use: 7 days. Piety to use: 5. Required to use: Holy symbol, holy plant

By use of this prayer, the priest of Domingo calls up a concealing mist in the province he occupies. The mist will last till the end of the following month.

The mists do not affect priests of Domingo.

, Church of Domingo [2700],]

E.19.4 Train Ninjas [2733]

Usage: USE 2733 [num] Required Item: at most 10 soldiers [12] Time to learn: 28 days. Time to use: 20 days. Piety to use: 6. Required to use: Holy symbol, holy plant

During the 20 days of this prayer, the priest of Domingo will teach to the soldiers he commands temple secrets of concealment and train them harshly. The result will be to convert the soldiers into [28] ninjas. Ninjas have the same attributes and costs as [12] soldiers, but with the ability of concealment.

Because of their skills of concealment, ninjas do not appear in a unit’s description, and they do not prevent him from using skills which require him to be alone, e.g., [1290] Conceal Self.

Note that a priest while a priest may only create 10 ninjas at a time, he may use this prayer multiple times to create more ninjas.

, Church of Domingo [2700], Appendix - Skill Dictionary,]

E.20 Magic [2800]

Time to learn: 14 days. NPs to learn: 1.

No apprentice magician can rightfully don the robes of his order without some knowledge of Magic [2800]. Magic contains many spells useful on their own, as well as in conjunction with spells from the other schools of sorcery.

The following teachable skills may be studied directly once Magic is known:

num    skill                              time to learn
---    -----                              -------------
2801   Meditate                           two weeks
2802   Perform common tasks for gold      two weeks
2803   Appear common                      two weeks
2804   View current aura level of others  four weeks

The following unteachable skills may be studied directly once Magic is known:

num    skill                              time to learn
---    -----                              -------------
2830   Speed casting of next spell        four weeks
2831   Reveal abilities of another mage   four weeks
2832   Tap health for aura                four weeks
2833   Modern magic script                four weeks
2834   Shroud abilities from scry         four weeks
2835   Detect ability scry                four weeks
2836   Dispel ability shroud              four weeks
2837   Advanced meditation                four weeks
2838   Hinder meditation                  four weeks
2839   Save speeded cast                  four weeks
2840   Lightning bolt                     four weeks
2841   Create Dirt Golem                  two weeks

, Magic [2800], Magic [2800],]

E.20.1 Meditate [2801]

Usage: USE 2801 Time to learn: 14 days. Time to use: 7 days.

The magician enters a restful trance which regenerates body, mind and current aura level. The magician must be alone to meditate, which means that (1) he cannot be stacked with anyone, and (2) he cannot have any men or beasts in his inventory (except for ninjas, angels, and golems).

Current aura level will rise by four points if the magician is the sole occupant of a tower and two points otherwise. If the magician possesses an auraculum, he will gain an additional two points.

Meditate can only be used when the magician’s current aura level is below his maximum effective aura.

See also ⟨undefined⟩ [Hinder meditation [2838]], page ⟨undefined⟩ and ⟨undefined⟩ [Hinder meditation [3532]], page ⟨undefined⟩.

, Meditate [2801], Magic [2800],]

E.20.2 Perform common tasks for gold [2802]

Usage: USE 2802 Time to learn: 14 days. Time to use: 7 days.

Even powerful magicians sometimes require gold to further their plans. This skill allows the magician to perform trivial or common tasks for the local population of a city or province to earn gold.

Examples of common tasks are curing runny noses, concocting love potions, dowsing for water, or performing conjuring tricks.

The amount earned depends upon the experience level of the magician:

    Description      Amount
    -----------      ------
    apprentice       35 gold
    journeyman       56 gold
    adept            70 gold
    master           84 gold
    grand master     91 gold

Of course, this skill depends upon the peasants having money to spend. Any money paid to the magician comes out of the peasants’ accumulated wealth. If they have no gold, the magician will come away empty-handed. The skill ends as soon as the peasants run out of money, e.g., if they only have money for four days of the magician’s services, the skill will end after four days.

The services the magician provides are genuinely valuable, and every day the magician is providing paid services to the peasants, 1-5 new peasants will be attracted to the province.

This skill cannot be used in provinces with less than 100 peasants. Using this skill in a city is the same as using it in the province the city is in.

In provinces which contain a temple, the peasants will not associate with a magician, so this skill will fail.

, Perform common tasks for gold [2802], Magic [2800],]

E.20.3 Appear common [2803]

Usage: USE 2803 <aura> Time to learn: 14 days. Time to use: 1 day.

The caster expends aura to conceal the fact that he is a magician. Nobles who possess magical powers are identified to others in the turn reports with a tag:

max aura tag ——– — 5-10 mage 11-16 wizard 17+ sorcerer

Use of 2803 allows the tag to be suppressed for a number of turns equal to the amount of aura invested in the cast.

Uses of 2803 are addititive. A mage with 4 turns of concealment remaining who ordered "use 2803 3" would have a total of seven turns of concealment.

, Appear common [2803], Magic [2800],]

E.20.4 View current aura level of others [2804]

Usage: USE 2804 [aura] Time to learn: 28 days. Time to use: 7 days.

The specified aura will be invested to learn the current aura levels of other mages in the location. The default amount of aura to use is one point.

The current aura level of other mages present in the same location as the caster will be shown, unless they are shrouding their abilities. If a mage is employing an ability shroud, the aura invested in USE 2804 must exceed the shroud’s aura, or that mage’s current aura level won’t be shown.

For instance, USE 2804 1 will show the current aura for mages that are not shrouding their abilities. USE 2804 3 will show the current aura for all mages except those shrouding their abilities with three or more points of aura.

, View current aura level of others [2804], Magic [2800],]

E.20.5 Speed casting of next spell [2830]

Usage: USE 2830 <amount> Time to learn: 28 days. Time to use: 4 days.

The mage trades aura for time. The time necessary to cast the next spell will be reduced by the amount of aura invested in the cast of [2830], up to a maximum of 3 aura.

For example, a mage who issued USE 2830 3 would lose three current aura points. Whatever spell the mage casts next will take three days less than it normally would.

Uses of 2830 are additive, but no spell may be sped by more than three days. All spells take a minimum of one day to cast.

, Speed casting of next spell [2830], Magic [2800],]

E.20.6 Reveal abilities of another mage [2831]

Usage: USE 2831 <target> <magical skill category> [aura] Time to learn: 28 days. Time to use: 7 days.

Used against another magician, this spell will attempt to reveal all spells known by the other mage within the given magical skill category.

If the target is shrouding his abilities, the aura invested in USE 2831 must exceed the shroud’s aura, or the spell will fail. For instance, USE 2831 4 (costing four aura points) is required to overcome a three-aura ability shroud.

, Reveal abilities of another mage [2831], Magic [2800],]

E.20.7 Tap health for aura [2832]

Usage: USE 2832 <amount of aura> Time to learn: 28 days. Time to use: 7 days.

The mage will dip into his own health to generate current aura. Within 60 days of using this spell, the magician will suffer a wound equal to five times the amount of aura generated.

For example, USE 2832 10 would generate 10 points of current aura, and cause the mage to receive a 50 point wound sometime within the next 60 days.

Use of this spell is dangerous, as any wound runs the risk of infection, and death could result. See the section on Health in the rules for more information about wounds and healing.

, Tap health for aura [2832], Magic [2800],]

E.20.8 Lightning bolt [2840]

Usage: Automatic in battle Time to learn: 28 days. Time to use: Automatic in combat Aura to use: 10. This is a combat spell.

Lightning bolt is a combat spell. During the special attacks phase of the combat round, the magician sends a lightning bolt against the toughest defender in the front row of the opposing force. The lightning bolt acts as a 500 point attack against that defender.

Note that nobles protected by troops are not defenders, and so will not be targeted by this spell until all their troops have been eliminated.

, Lightning bolt [2840], Magic [2800],]

E.20.9 Modern magic script [2833]

Usage: USE 2833 <skill> [# days instruction] [book] Requires: blank scrolls [84], 100 gold Required aura: two points per page Time to learn: 28 days. Time to use: 7 days.

Create a book for the given skill, or add days of instruction to an existing book. The book can be used as a source of instruction. Books may be created only for category skills and teachable subskills.

If the number of number of days argument is missing, a default of 7 days is assumed.

There is a cost of 100 gold to begin a new book. This covers the cost of the binding, covers, etc.

This skills consumes two points of aura for every day of instruction written.

The skill must be known in order to be written. The character will still know the skill after writing it.

This skill requires one day for every day of instruction written. The optional third argument is the id number of an existing book in the

noble’s possession. If this is a book for the given skill, then additional pages will be added to the book. In this case, there is no 100 gold fee to begin a new book.

Only Magic [2800] spells may be scribed with this skill. Examples: USE 2833 7 Creates a new book with 7 pages of instruction in 2833. USE 2833 3 j3k2 Adds three days of instruction to book [j3k2], provided that book teaches

2833.

, Modern magic script [2833], Magic [2800],]

E.20.10 Shroud abilities from scry [2834]

Usage: USE 2834 [aura] Time to learn: 28 days. Time to use: 3 days.

Attempt to prevent other mages from successfully scrying the caster to learn magical abilities. Creates an ability shroud with the given amount of aura. A greater amount of aura must be spent by another mage to pierce the shroud.

The default amount of aura is one point.

The magical shroud loses one point of aura at the end of each month. Successive uses of 2834 are additive; USE 9169 1 followed by USE 2834 3 will result in an ability shroud with an aura of four points.

, Shroud abilities from scry [2834], Magic [2800],]

E.20.11 Detect ability scry [2835]

Usage: USE 2835 Time to learn: 28 days. Time to use: 7 days. Required aura: 1 (to practice)

Knowledge of 2835 will inform the mage of attempts made by other magicians to scry his abilities. No USE is required in order for scry attempts to be detected. The skill is automatically employed whenever the mage is the target of an ability scry spell.

USE of the skill will cause the magician to practice scry detection. Practice should improve the experience rating with this skill over time, and cause the mage to become more adept at discovering scry attempts and identifying their source.

Aura is only consumed by this spell when it is used for practice. No aura is depleted when a scry is detected.

, Detect ability scry [2835], Magic [2800],]

E.20.12 Dispel ability shroud [2836]

Usage: USE 2836 <target> Time to learn: 28 days. Time to use: 3 days. Required aura: 3

Dispels any magical shroud inhibiting scrying of the specified character’s abilities.

, Dispel ability shroud [2836], Magic [2800],]

E.20.13 Advanced meditation [2837]

Usage: USE 2837 Time to learn: 28 days. Time to use: 28 days.

Advanced meditation is a technique by which a magician can raise his maximum aura. By meditating alone and undisturbed in a tower for one month, the magician can raise his maximum aura by one point.

See also ⟨undefined⟩ [Hinder meditation [2838]], page ⟨undefined⟩ and ⟨undefined⟩ [Hinder meditation [3532]], page ⟨undefined⟩.

, Advanced meditation [2837], Magic [2800],]

E.20.14 Hinder meditation [2838]

Usage: USE 2838 <target> [aura] Time to learn: 28 days. Time to use: 10 days.

Clouds the target’s mind with disturbing images and incoherent thoughts. The target, whether mage or priest, will have difficulty achieving the proper mental state during his next attempt at meditation or devotional prayer.

The next meditate spell issued by the target may fail. The more aura invested in 2838, the greater the chance the target mage will fail:

aura chance of failure —- —————– 1 50% 2 75% 3 90%

If the target does not meditate, the spell will lose one point of aura at the end of each month until it has dissipated. Casts by multiple mages accumulate on the target, although the spell may not accumulate more than three points of aura.

For example, if two mages each cast USE 2838 501 2, character [xt8] will have a 90% chance of failure in his next attempt to meditate. (One aura point has been wasted, since only three may accumulate on the target). At the end of the month, the strength of the 2838 spell against [xt8] will diminish to two points, and [xt8] will have a 50% chance of failure during his next meditation.

A mage may cast multiple 2838 spells; they accumulate on the targets, not on the casting magician.

The default amount of aura is one point.

There is a 50% chance that the target will have a disturbing dream that will warn him of his situation.

See also (⟨undefined⟩ [Hinder meditation [3532]], page ⟨undefined⟩).

, Hinder meditation [2838], Magic [2800],]

E.20.15 Save speeded cast [2839]

Usage: USE 2839 Time to learn: 28 days. Time to use: 7 days. Required aura: 3

The state where the mage’s next cast will be speeded may be saved with this spell. A potion will be created which, upon use, will revive the accelerated cast state.

For example, a mage establishes that his next spell cast will be speeded by three days. He then issues USE 2839 to preserve this state. Magical potion [8765] is created. He later issues USE 8765 to drink the potion. He will gain a three day speedup for his next cast.

The mage will no longer have a speeded cast state after use of this spell. , Magic [2800],]

E.20.16 Create Dirt Golem [2841]

Usage: USE 2841 Time to learn: 14 days. Time to use: 14 days. Aura to use: 10.

By use of this spell, the magician creates a magical, animated being out of common dirt. The dirt golem weighs 500 units, can carry 1000 units of weight and fights at (125,125).

Golems will only serve their creator and cannot be transferred between nobles. Golems are not counted when determining if a noble is "alone" (e.g., ⟨undefined⟩ [Meditate [2801]], page ⟨undefined⟩). Golems are commanded in battle as if they were men (i.e., they are not "beasts").

Because dirt is base substance, dirt golems last only about six months before suddenly crumbling into pieces.

, Magic [2800], Appendix - Skill Dictionary,]

E.21 Scrying [2900]

Time to learn: 21 days. NPs to learn: 1.

Scrying concerns the gathering of information by means of magical spells. Simple spells may allow the mage to learn information about other characters nearby, or to view distant regions.

Advanced study delves into spells wrought long ago in by the men of Areth-Pirn as tools to rule their land, including the erection of magical barriers to travel, as well as additional spells of far-seeing.

The following teachable skills may be studied directly once Scrying is known:

num    skill                              time to learn
---    -----                              -------------
2901   Scry location                      four weeks

The following unteachable skills may be studied directly once Scrying is known:

num    skill                              time to learn
---    -----                              -------------
2930   Shroud location from magical scry  two weeks
2931   Dispel location shroud             four weeks
2932   Ciphered writing of Areth-Pirn     four weeks
2933   Create magical barrier             four weeks
2934   Remove magical barrier             four weeks
2935   Locate character                   four weeks
2936   Detect location scry               four weeks
2937   Farcasting                         four weeks
2938   Save farcast state                 four weeks
2939   Banish undead                      four weeks
2340   Foresee Defense                    four weeks

, Scrying [2900], Scrying [2900],]

E.21.1 Foresee Defense [2940]

Usage: Automatic in battle Time to learn: 28 days. Time to use: Automatic in combat Aura to use: 10. This is a combat spell.

Forsee Defense is a combat spell. During the special attacks phase of the combat round, the scryer looks a short distance into the future to predict the defense the other side will use in the upcoming round, and then quickly passes word of what he has seen to his troops. This gives the troops personally commanded by the scryer a 10% bonus during the attack phase.

, Foresee Defense [2940], Scrying [2900],]

E.21.2 Scry location [2901]

Usage: USE 2901 <location> [aura] Time to learn: 28 days. Time to use: 7 days.

Grants the mage a vision of the specified place, no matter how far away he might be. However, like all magical spells, Scry location cannot cross open water.

If the target location is magically shrouded from scry attempts, sufficient aura must be used for 2901 to pierce the shroud. For example, a location shroud of three aura would require USE 2901 <location> 4 to overcome (costing four aura to cast).

The default aura to use is one point.

, Scry location [2901], Scrying [2900],]

E.21.3 Shroud location from magical scry [2930]

Usage: USE 2930 [aura] Time to learn: 14 days. Time to use: 3 days.

Creates a magical shroud over the mage’s current province. The shroud hinders attempts to view the location by magical scrying or holy visions (except for orbs, and ⟨undefined⟩ [Receive vision [2434]], page ⟨undefined⟩, which succeed even if a shroud exists).

The shroud is invested with twice the aura used in its creation. A mage or priest attempting to view the location from a distance with a scry spell or holy vision must expend a greater amount of aura (piety) than the shroud possesses to succeed.

The shroud loses one point of aura at the end of each month. Successive uses of 2930 are additive. For example, USE 9142 1 followed by USE 9142 3 will result in a location shroud with an aura of eight points.

The default aura to use is one point.

, Shroud location from magical scry [2930], Scrying [2900],]

E.21.4 Dispel location shroud [2931]

Usage: USE 2931 <location> Time to learn: 28 days. Time to use: 3 days. Required aura: 3

Dispels any magical shroud inhibiting scrying of the specified location. , Dispel location shroud [2931], Scrying [2900],]

E.21.5 Ciphered writing of Areth-Pirn [2932]

Usage: USE 2932 <skill> [# days instruction] [book] Requires: blank scrolls [84], 100 gold Required aura: two points per page Time to learn: 28 days. Time to use: 7 days.

Create a book for the given skill, or add days of instruction to an existing book. The book can be used as a source of instruction. Books may be created only for category skills and teachable subskills.

If the number of number of days argument is missing, a default of 7 days is assumed.

There is a cost of 100 gold to begin a new book. This covers the cost of the binding, covers, etc.

This skills consumes two points of aura for every day of instruction written.

The skill must be known in order to be written. The character will still know the skill after writing it.

This skill requires one day for every day of instruction written. The optional third argument is the id number of an existing book in the

noble’s possession. If this is a book for the given skill, then additional pages will be added to the book. In this case, there is no 100 gold fee to begin a new book.

Only Scrying [2900] spells may be scribed with this skill. Examples: USE 2932 7 Creates a new book with 7 pages of instruction in 2932.

USE 2932 3 j3k2

Adds three days of instruction to book [j3k2], provided that book teaches 2932.

, Ciphered writing of Areth-Pirn [2932], Scrying [2900],]

E.21.6 Create magical barrier [2933]

Usage: USE 2933 [aura] Time to learn: 28 days. Time to use: 10 days.

Erects a magical barrier which prevents anyone from entering the location. The barrier will lose one point of strength at the end of each turn, and will dissipate when it has no strength remaining. The maximum strength for a magical barrier is 4 (lasting 4 months). The cost of casting this spell is the square of the strength, e.g., a strength 4 barrier costs 16 aura to cast.

Aura defaults to one.

Barriers cannot be cast on provinces.

When a magical barrier dissipates, lingering effects of the spell prevent another barrier from being cast on the same location for four months.

, Create magical barrier [2933], Scrying [2900],]

E.21.7 Remove magical barrier [2934]

Usage: USE 2934 <direction> [aura] Time to learn: 28 days. Time to use: 7 days.

Attempt to remove a magical barrier from the location in the given direction. If direction "0" is given, a barrier over the current location will be removed.

The chance that the barrier will be successfully removed is proportional to the aura invested in the attempt:

aura chance of success —- —————– 1 10% 2 25% 3 50% 4 75%

When a magical barrier is removed, lingering effects of the spell prevent another barrier from being cast on the same location for four months.

, Remove magical barrier [2934], Scrying [2900],]

E.21.8 Locate character [2935]

Usage: USE 2935 <character> [aura] Time to learn: 28 days. Time to use: 10 days.

Attempt to locate the specified character. If successful, the mage will learn the location in which the character is currently located. However, like all magical spells, Locate character cannot cross open water.

The more aura expended, the greater the chance that the character will be located.

aura chance of success —- —————– 1 50% 2 75% 3 90%

, Locate character [2935], Scrying [2900],]

E.21.9 Detect location scry [2936]

Usage: USE 2936 Time to learn: 28 days. Time to use: 7 days. Aura required: 1 (for practice)

Knowledge of 2936 will inform the mage of attempts made to scry or vision the mage’s current location. No USE is required in order for attempts to be detected. The skill is used automatically whenever the mage’s location is the target of a scry or vision spell.

USE of the skill will cause the magician to practice scry detection. Practice should improve the experience rating with this skill over time, and cause the mage to become more adept at discovering scry attempts and identifying their source.

Aura is only consumed by this spell when it is used for practice. No aura is depleted when a scry is detected.

, Detect location scry [2936], Scrying [2900],]

E.21.10 Farcasting [2937]

Usage: USE 2937 <location> Time to learn: 28 days. Time to use: 7 days.

If successful, the location check for the next spell cast attempted will be made from the given location instead of the location the mage is in.

The amount of aura consumed is equal to the distance in provinces from the mage’s location to the target location. For instance, to project a cast into an adjoining province requires two aura. To project a cast between any of the sublocations within a province requires one aura.

The projected cast state will persist until the mage’s next use of magic. The mage may move before casting the next spell; movement will not alter the target location for spell projection.

Unless the mage is projecting to the same location he is in, a magical shroud over the target location will cause this spell to fail.

Not all spells may be farcasted. The following are the spells that may be farcasted:

num    skill
----   -----
2804   View current aura level of others
3102   Reveal artifacts
3132   Conceal artifacts
3001   Detect gates
3031   Seal gate
3032   Unseal gate with key
3033   Notify if gate unsealed
3034   Forcefully unseal gate
3035   Reveal gate key
3036   Notify of gate jumps
2831   Reveal abilities of another mage
2930   Shroud location
2933   Create magical barrier
2934   Remove magical barrier
2937   Farcasting
2836   Dispel ability shroud
2838   Hinder meditation
3233   Banish demon lord
3203   Aura blast

, Farcasting [2937], Scrying [2900],]

E.21.11 Save farcast state [2938]

Usage: USE 2938 Time to learn: 28 days. Time to use: 7 days.

The state where a mage’s next cast will be projected to a different location may be saved with this spell. A potion will be created which, upon use, will revive the projected cast state.

For example, a mage establishes that his next spell cast will be based from Drassa [cx59]. He then issues USE 2938 to preserve this state. Magical potion [8764] is created. He later issues USE 8764 to drink the potion. His next spell will be based from Drassa.

The mage will no longer have a projected cast state after use of this spell. , Scrying [2900],]

E.21.12 Banish undead [2939]

Usage: USE 2939 <target> [aura] Time to learn: 28 days. Time to use: 7 days.

Banish all undead in the inventory of <target>. Requires one aura point for every undead to be banished. The spell will banish as many undead as possible, limited only by the indicated aura and the caster’s current aura. No aura is used, nor any time expended, if the target has no undead.

, Scrying [2900], Appendix - Skill Dictionary,]

E.22 Gatecraft [3000]

Time to learn: 21 days. NPs to learn: 1.

Long ago the Ancients, a race of wise men who ruled the land with the aid of powerful magic, constructed a network of teleportation gateways to connect distant cities. As the reach of their magical portals grew, so grew the reach of their armies and knowledge. At their height the rule of the Ancients extended to every corner of the known world, and beyond.

With the passing of the Ancients, the system of gates fell into disuse and decay, and now retains only a hint of its former majesty. Gatecraft, once practiced by the most common among the Ancients, has since been forgotten by nearly all men.

The school of Gatecraft retains the only knowledge of the use of these mystical portals. Because knowledge of Gates can be a powerful tool or weapon, Gatecraft spells are jealously guarded even within the school. Some aspects of Gatecraft, such as the secret of their construction, have been completely lost.

The following teachable skills may be studied directly once Gatecraft is known:

num    skill                              time to learn
---    -----                              -------------
3001   Detect gates                       two weeks
3002   Jump through gate                  two weeks

The following unteachable skills may be studied directly once Gatecraft is known:

num    skill                              time to learn
---    -----                              -------------
3030   Teleport                           two weeks
3031   Seal gate                          two weeks
3032   Unseal gate with key               two weeks
3033   Notify if gate unsealed            two weeks
3034   Forcefully unseal gate             four weeks
3035   Reveal gate key                    four weeks
3036   Notify of gate jumps               two weeks
3037   Language of the Ancients           four weeks
3038   Reverse jump through gate          four weeks
3039   Drain Mana                         four weeks

, Gatecraft [3000], Gatecraft [3000],]

E.22.1 Teleport [3030]

Usage: USE 3030 <location> Time to learn: 14 days. Time to use: 1 day.

This skill allows a mage to teleport to the given location without a gate. A gate crystal is required to fuel the jump. The gate crystal is consumed by the jump.

One point of aura is required for each 250 units of weight in the caster’s stack to be transported.

, Teleport [3030], Gatecraft [3000],]

E.22.2 Detect gates [3001]

Usage: USE 3001 Time to learn: 14 days. Time to use: 7 days. Required aura: one point

Identify any gates at caster’s location, or (if the mage is Farcasting) at the farcasted location.

If there are no gates at the casting location, Detect gates [3001] will indicate how far away the nearest gate is.

, Detect gates [3001], Gatecraft [3000],]

E.22.3 Jump through gate [3002]

Usage: USE 3002 <gate> [key] Time to learn: 14 days. Time to use: 1 day. Required aura: variable

The caster will teleport through the specified gate. The gate must be present in the current location. If the gate is sealed with a key, the key must be supplied for the jump to succeed.

One point of aura is required for each 250 units of weight in the caster’s stack to be transported.

, Jump through gate [3002], Gatecraft [3000],]

E.22.4 Seal gate [3031]

Usage: USE 3031 <gate> <key> Time to learn: 14 days. Time to use: 7 days. Required aura: six points

Seal the gate with the given key. The gate must be present in the current location, and must not already be sealed.

The key is a numeric password which will be required for further access to the gate for jumps or other manipulations. Only those who supply the key number will be granted access. The number should be between 1 and 999.

For example, if Gate [8192] is sealed with USE 3031 8192 123 then the key "123" must be supplied as a parameter to the jump spell, or the jump will fail.

The spell fails immediately if the caster does not have sufficient aura. In all other cases, the spell requires the full casting time and consumes any Farcast, even if the spell fails (e.g., if the gate does not exist).

, Seal gate [3031], Gatecraft [3000],]

E.22.5 Unseal gate with key [3032]

Usage: USE 3032 <gate> <key> Time to learn: 14 days. Time to use: 7 days. Required aura: three points

Remove the seal from a gate by supplying the key. The key will no longer be required for jumps through the gate.

For example, if Gate [8192] were sealed with key "123", the spell USE 3032 8192 123 would unseal the gate. Jumps through the gate could proceed without supplying the key.

The spell fails immediately if the caster does not have sufficient aura or does not supply a key. In all other cases, the spell requires the full casting time, aura and consumes any Farcast, even if the spell fails (e.g., if the gate does not exist).

, Unseal gate with key [3032], Gatecraft [3000],]

E.22.6 Notify if gate unsealed [3033]

Usage: USE 3033 <gate> <key> Time to learn: 14 days. Time to use: 7 days. Required aura: five points

Inform the sealer of the gate if the seal is broken, and of the identity of the mage who removes the seal. The gate must be present in the current location to apply this spell. However, notification will occur wherever the caster happens to be.

The spell fails immediately if the caster does not have sufficient aura or does not supply a key. In all other cases, the spell requires the full casting time, aura and consumes any Farcast, even if the spell fails (e.g., if the gate does not exist).

, Notify if gate unsealed [3033], Gatecraft [3000],]

E.22.7 Forcefully unseal gate [3034]

Usage: USE 3034 <gate> Time to learn: 28 days. Time to use: 7 days. Required aura: eight points

Forcefully remove a seal from a gate without specifying the key. The key will no longer be required to perform jumps through the gate. The gate must be present in the current location.

The spell fails immediately if the caster does not have sufficient aura. In all other cases, the spell requires the full casting time, aura and consumes any Farcast, even if the spell fails (e.g., if the gate does not exist).

, Forcefully unseal gate [3034], Gatecraft [3000],]

E.22.8 Reveal gate key [3035]

Usage: USE 3035 <gate> Time to learn: 28 days. Time to use: 7 days. Required aura: ten points

Reveal the numeric key necessary to make use of a sealed gate.

The spell fails immediately if the caster does not have sufficient aura. In all other cases, the spell requires the full casting time, aura and consumes any Farcast, even if the spell fails (e.g., if the gate does not exist).

, Reveal gate key [3035], Gatecraft [3000],]

E.22.9 Notify of gate jumps [3036]

Usage: USE 3036 <gate> Time to learn: 14 days. Time to use: 7 days. Required aura: six points

Notify the caster of the identity of any mages jumping through the specified gate. The gate must be present in the current location to apply this spell. However, the caster will be notified if the gate is used, wherever he may be at the time.

Only one [3036] spell may be active on a gate at a time. Casting this spell on a gate already notifying another mage of jumps will displace the previous mage’s spell.

The spell fails immediately if the caster does not have sufficient aura. In all other cases, the spell requires the full casting time, aura and consumes any Farcast, even if the spell fails (e.g., if the gate does not exist).

, Notify of gate jumps [3036], Gatecraft [3000],]

E.22.10 Drain Mana [3039]

Usage: Automatic in battle Time to learn: 28 days. Time to use: Automatic in combat Aura to use: 10. This is a combat spell.

Drain Mana is a combat spell. During the special attacks phase of the combat round, if the gatecrafter detects magicians or priests in the opposing forces, he opens a chaos gate on the other side of battlefield. This gate immediately drains all magicians and priests on that side of the battle of up to 10 aura (or piety) and then destroys itself. The gatecrafter himself is able to siphon off 1/10 of the drained aura and use it to replenish his own aura.

About 25% of the time, the gatecrafter will mistakenly detect magicians or priests in the opposing forces, and use the spell even though it will be ineffective. The gatecrafter may also mistake undead for magicians.

, Drain Mana [3039], Gatecraft [3000],]

E.22.11 Language of the Ancients [3037]

Usage: USE 3037 <skill> [# days instruction] [book] Requires: blank scrolls [84], 100 gold Required aura: two points per page Time to learn: 28 days. Time to use: 7 days.

Create a book for the given skill, or add days of instruction to an existing book. The book can be used as a source of instruction. Books may be created only for category skills and teachable subskills.

If the number of number of days argument is missing, a default of 7 days is assumed.

There is a cost of 100 gold to begin a new book. This covers the cost of the binding, covers, etc.

This skills consumes two points of aura for every day of instruction written.

The skill must be known in order to be written. The character will still know the skill after writing it.

This skill requires one day for every day of instruction written. The optional third argument is the id number of an existing book in the

noble’s possession. If this is a book for the given skill, then additional pages will be added to the book. In this case, there is no 100 gold fee to begin a new book.

Only Gatecraft [3000] spells may be scribed with this skill. Examples: USE 3037 7 Creates a new book with 7 pages of instruction in 3037. USE 3037 3 j3k2 Adds three days of instruction to book [j3k2], provided that book teaches

3037.

, Gatecraft [3000],]

E.22.12 Reverse jump through gate [3038]

Usage: USE 3038 <gate> [key] Time to learn: 28 days. Time to use: 1 day. Required aura: variable

Gates were long thought to be one way portals through space. Most mages believed that returning to the starting location required a second gate or a conventional journey.

However, it is possible to jump back to a gate from the gate’s destination. Reversing the direction of teleportation is difficult, so the jump requires twice as much aura as a forward jump.

As with a conventional jump, the mage must know the gate’s entity number as well as the key, if necessary.

, Gatecraft [3000], Appendix - Skill Dictionary,]

E.23 Artifact construction [3100]

Time to learn: 28 days. NPs to learn: 1.

The domain of high level wizards, artifact construction concerns the creation and manipulation of items which may focus and store aura, contain spells, or have other magical properties.

The following teachable skills may be studied directly once Artifact construction is known:

num    skill                              time to learn
---    -----                              -------------
3101   Detect Artifact   two weeks
3102   Reveal Artifacts   two weeks
3103   Destroy artifact                   two weeks
3104   Deep Identify                   four weeks

The following unteachable skills may be studied directly once Artifact construction is known:

num    skill                              time to learn
---    -----                              -------------
3130   Forge auraculum                    four weeks
3131   Mutate Artifact           two weeks
3132   Conceal Artifacts   two weeks
3133   Obscure Artifact           two weeks
3134   Forge magical weapon               four weeks
3135   Forge magical armor                four weeks
3136   Forge magical bow                  four weeks
3137   Remove Obscurity   two weeks
3138   Arcane symbols                     four weeks
3139   Forge palantir                     four weeks

, Artifact construction [3100], Artifact construction [3100],]

E.23.1 Forge auraculum [3130]

Usage: USE 3130 <aura-level> "Name" Time to learn: 28 days. Time to use: 14 days. Requires: 100 gold [1] Requires: five mithril [82]

Forge a magical auraculum, giving it the specified name.

An auraculum is a magical artifact which concentrates and amplifies a mage’s power. A permanent investment of maximum aura is made when constructing the auraculum. The artifact doubles the invested aura. However, since a portion of the mage’s pneuma has been bound in the auraculum, it is important to safeguard it from being lost or stolen.

In addition to depleting the caster’s current aura level, the specified aura amount will be permanently removed from the magician’s maximum aura rating. The new object, while possessed by its maker, will augment the mage’s maximum aura rating by twice the amount of aura invested in its creation.

Possession of an auraculum doubles the amount of aura a mage naturally recovers each month, from two points to four.

This spell cannot be learned until the magician has learned all the other schools of magic, i.e., [2800] Magic, [2900] Scrying, [3000] Gatecraft, [3100] Artifact construction and [3200] Necromancy.

Creation of an auraculum also grants knowledge of the Advanced sorcery magical skill category. Advanced sorcery may not be learned in any other way. Learning Advanced Sorcery is exempt from the normal NP costs of learning a skill.

Example:

       A mage of current aura level 10, maximum aura 10 who issues
       USE 3130 5 "Ring of power" will be reduced to current aura level 5,
       maximum aura level 5.  However, the new artifact will add 10 points
       to the maximum aura rating of its creator, but only while it is in
       his possession.
           Current aura:   10
           Maximum aura:   10
        1: > use 3130 5 "Ring of power"
        1: Attempt to forge an auraculum.
           ...
       14: Created Ring of power [js80].
       14: Learned Advanced sorcery [166].
           Current aura:   5
           Maximum aura:   15 (5+10)

Characters other than the auraculum’s creator do not receive an aura bonus from possession of the object.

Since the mage has bound a portion of his aura into the artifact, its destruction would deal him an instant, fatal blow. Thus, the creator of such an object should use all means to prevent it from being lost or stolen. However, magical items of such power are hard to destroy. An auraculum may only be destroyed in the province in which it was created.

This spell may only be used once. A magician may only create one auraculum during his lifetime.

The creator of the artifact may change its name later with the NAME order. The object must be possessed in order to rename it.

> name js80 "Wooden staff" Ring of power [js80] will now be known as Wooden staff [js80].

, Forge auraculum [3130], Artifact construction [3100],]

E.23.2 Detect Artifact [3101]

Usage: USE 3101 Time to learn: 14 days. Time to use: 7 days. Required aura: two points.

This spell informs the user the distance to the nearest artifact. Artifacts which have been concealed with Conceal Artifacts [3132] will not be detected.

, Detect Artifact [3101], Artifact construction [3100],]

E.23.3 Reveal Artifacts [3102]

Usage: USE 3102 <target> Time to learn: 14 days. Time to use: 7 days. Required aura: four points.

Reveals any artifacts the target is carrying (but does not identify what they are). Artifacts which have been concealed with Conceal Artifacts [3132] will not be detected.

, Reveal Artifacts [3102], Artifact construction [3100],]

E.23.4 Mutate Artifact [3131]

Usage: USE 3131 <artifact> Time to learn: 14 days. Time to use: 30 days. Required aura: 8 points

This spell turns an artifact in the possession of the caster into another, random artifact. The following kinds of artifacts cannot be mutated: auraculums, palantirs, and combat artifacts. An artifact that mutates into one of these types will no longer be mutable.

Note that the mutated artifact will have a different identifier than the original artifact.

, Mutate Artifact [3131], Artifact construction [3100],]

E.23.5 Conceal Artifacts [3132]

Usage: USE 3132 <target> Time to learn: 14 days. Time to use: 7 days. Required aura: 4 points. Duration: 30 days.

Conceals all the artifacts held by <target> from detection. Caster must be in the same location as <target>. Note that the spell is not cast upon an artifact, but upon a noble, and conceals all the artifacts he is carrying.

, Conceal Artifacts [3132], Artifact construction [3100],]

E.23.6 Obscure Artifact [3133]

Usage: USE 3133 <artifact> Time to learn: 14 days. Time to use: 14 days. Required aura: eight points.

Permanently obscures the tiny runes which identify the function of an artifact, so that any attempts to identify it will fail. This obscurity may be removed with Remove Obscurity [3137]. Note that this spell is cast upon the artifact, and the caster must be in possession of the artifact at the time of casting.

, Obscure Artifact [3133], Artifact construction [3100],]

E.23.7 Forge magical weapon [3134]

Usage: USE 3134 <aura> "Name of item" Time to learn: 28 days. Time to use: 7 days. Requires: 1 mithril [82]

Create a magical weapon. The weapon will grant its owner an attack bonus equal to 5 * the amount of aura invested in its creation.

Up to 20 aura may be invested in a magical weapon. Example:

 use 3134 5 "Sword of Death"
 > created Sword of Death [ha53].
 Inventory:
            1  Sword of Death [ha53]           0  +25 attack

, Forge magical weapon [3134], Artifact construction [3100],]

E.23.8 Forge magical armor [3135]

Usage: USE 3135 <aura> "Name of item" Time to learn: 28 days. Time to use: 7 days. Requires: 1 mithril [82]

Create magical armor. The armor will grant its owner a defense bonus equal to 5 * the amount of aura invested in its creation.

Up to 20 aura may be invested in magical armor. Example:

 use 3135 5 "Shield of Achilles"
 > created Shield of Achilles [ha53].
 Inventory:
            1  Shield of Achilles [ha53]           0  +25 defense

, Forge magical armor [3135], Artifact construction [3100],]

E.23.9 Forge magical bow [3136]

Usage: USE 3136 <aura> "Name of item" Time to learn: 28 days. Time to use: 7 days. Requires: 1 mallorn wood [70]

Create a magical bow. The bow will grant its owner a missile bonus equal to 5 * the amount of aura invested in its creation.

Up to 20 aura may be invested in a magical bow. Example:

 use 3136 5 "Hand of Legolas"
 > created Hand of Legolas [ha53].
 Inventory:
            1  Hand of Legolas [ha53]           0  +25 missile

, Forge magical bow [3136], Artifact construction [3100],]

E.23.10 Remove Obscurity [3137]

Usage: USE 3137 <artifact> Time to learn: 14 days. Time to use: 7 days. Required aura: four points

Removes one Obscure Artifact [3133] cast upon an artifact. The artifact must be in the caster’s inventory. If an artifact has been obscured multiple times, it will require multiple removals.

E.23.11 Arcane symbols [3138]

Usage: USE 3138 <skill> [# days instruction] [book] Requires: blank scrolls [84], 100 gold Required aura: two points per page Time to learn: 28 days. Time to use: 7 days.

Create a book for the given skill, or add days of instruction to an existing book. The book can be used as a source of instruction. Books may be created only for category skills and teachable subskills.

If the number of number of days argument is missing, a default of 7 days is assumed.

There is a cost of 100 gold to begin a new book. This covers the cost of the binding, covers, etc.

This skills consumes two points of aura for every day of instruction written.

The skill must be known in order to be written. The character will still know the skill after writing it.

This skill requires one day for every day of instruction written.

The optional third argument is the id number of an existing book in the noble’s possession. If this is a book for the given skill, then additional pages will be added to the book. In this case, there is no 100 gold fee to begin a new book.

Only Artifact construction [3100] spells may be scribed with this skill. Examples: USE 3138 7 Creates a new book with 7 pages of instruction in 3138. USE 3138 3 j3k2 Adds three days of instruction to book [j3k2], provided that book teaches

3138.

, Remove Obscurity [3137], Artifact construction [3100],]

E.23.12 Destroy artifact [3103]

Usage: USE 3103 <artifact> Time to learn: 14 days. Time to use: 7 days. Required aura: two points.

Attempt to destroy a magical artifact. The artifact must be in the caster’s possession. The casting mage will capture 1-20 points of aura from the destroyed artifact, i.e., his current aura will increase by 1-20 points. However, the caster will never regain more points than originally invested in the artifact.

, Destroy artifact [3103], Artifact construction [3100],]

E.23.13 Deep Identify [3104]

Usage: USE 3104 <artifact> Time to learn: 28 days. Time to use: 7 days. Required aura: four points.

This spell identifies artifacts that have lost their runes and cannot be identified with the "identify" command. This spell can also identify the maker and power of an auraculum. However, this spell cannot penetrate the Obscure Artifact spell.

, Artifact construction [3100],]

E.23.14 Forge palantir [3139]

Usage: USE 3139 Time to learn: 28 days. Time to use: 10 days. Requires: one mithril [82] Requires: one gate crystal [83] Required aura: 8

A palantir, or stone of far-seeing, is a crystal orb used to view images from a great distance. Once created, the palantir may be used to view a distant location:

use <palantir> <location>

For example, if a USE 3139 creates Palantir [6289], then one may view City of Drassa [cx61] with the palantir by issuing the order:

use 6289 cx61

Use of the palantir takes seven days. Any character may use the palantir; no magical abilities are required. A palantir has between 3 and 9 charges when created. Each use consumes one charge, and when the charges are used up, the palantir will no longer function.

, Artifact construction [3100], Appendix - Skill Dictionary,]

E.24 Necromancy [3200]

Time to learn: 28 days. NPs to learn: 1.

The way of the necromancer leads to fresh graves, the pits of Hades, and unspeakable horrors found only in dark corners of men’s souls.

Power is gained through animation and enslavement of the dead, both followers as well as nobles. A necromancer looks upon a bloody battlefield strewn with the dead as a potential army waiting to be trained.

Advanced necromancers wield fierce powers, manipulating evil Runes to effect their will upon the world, and making pacts with demons to obtain power in return for souls.

With great power also comes great danger, and necromancy is the most dangerous of the magical arts. Only those willing to risk death at the hands of civilized men, or worse fates meted out by demons sent from the underworld, should even consider its study.

The following teachable skills may be studied directly once Necromancy is known:

num    skill                              time to learn
---    -----                              -------------
3201   Summon ghost warriors              four weeks
3202   Raise undead                       four weeks
3203   Aura blast                         four weeks

The following unteachable skills may be studied directly once Necromancy is known:

num    skill                              time to learn
---    -----                              -------------
3230   Runes of Evil                      four weeks
3231   Summon demon lord                  four weeks
3232   Renew demon bond                   four weeks
3233   Banish demon lord                  four weeks
3234   Eating of the dead                 four weeks
3235   Absorb aura blast                  four weeks
3236   Raise soldiers                     four weeks
3237   Create Flesh Golem                 three weeks
3238   Transcend death                    four weeks, 1 NP req'd

, Necromancy [3200], Necromancy [3200],]

E.24.1 Transcend death [3238]

Usage: automatic when needed Time to learn: 28 days. NPs to learn: 1.

Transcend death allows a mortal to pass through Hades unmolested by the foul and dangerous spirits who live there. The necromancer must be alone, however. He cannot be stacked with any other nobles, or be accompanied by any followers.

This skill also insulates the necromancer against the spirit-draining properties of Hades.

When a necromancer with knowledge of 3238 dies, the body vanishes, and the mage is recreated somewhere in Hades, albeit lacking any of his worldly possessions.

, Transcend death [3238], Necromancy [3200],]

E.24.2 Raise soldiers [3236]

Usage: Automatic in battle Time to learn: 28 days. Time to use: Automatic in combat Aura to use: 10. This is a combat spell.

Raise soldiers is a combat spell. During the special attacks phase of the combat round, the necromancer invokes various demonic powers to temporarily raise from the dead fallen soldiers. Up to 10 soldiers on the necromancer’s side of the battle are reanimated. The raised soldiers can be any sort of unit (e.g., archers, soldiers, etc.) but this spell has no effect on dead nobles. At the conclusion of the battle any reanimated soldiers will fall lifeless once again. The raised soldiers return to the inventory of their original owner.

, Raise soldiers [3236], Necromancy [3200],]

E.24.3 Runes of Evil [3230]

Usage: USE 3230 <skill> [# days instruction] [book] Requires: blank scrolls [84], 100 gold Required aura: two points per page Time to learn: 28 days. Time to use: 7 days.

Create a book for the given skill, or add days of instruction to an existing book. The book can be used as a source of instruction. Books may be created only for category skills and teachable subskills.

If the number of number of days argument is missing, a default of 7 days is assumed.

This skill consumes two points of aura for every day of instruction written. The skill must be known in order to be written. The character will still

know the skill after writing it.

This skill requires one day for every day of instruction written.

The optional third argument is the id number of an existing book in the noble’s possession. If this is a book for the given skill, then additional pages will be added to the book. In this case, there is no 100 gold fee to begin a new book.

Only Necromancy [3200] spells may be scribed with this skill.

Examples:

USE 3230 7 Creates a new book with 7 pages of instruction in 3230. USE 3230 3 j3k2 Adds three days of instruction to book [j3k2], provided that book teaches

3230.

, Runes of Evil [3230], Necromancy [3200],]

E.24.4 Summon ghost warriors [3201]

Usage: MAKE 18 [days] Time to learn: 28 days.

Each day of summoning requires one point of aura, and yields two ghost warriors.

Ghost warriors [18] are spirits of dead soldiers summoned to fight again on the battlefield. They have no substance, and thus are ineffectual attackers. Opponents will still select them in combat, however, so they are useful for absorbing enemy hits.

Ghost warriors have a attack and defense values of zero. One hit by any attacker will cause them to vanish.

One ghost warrior returns to the Land of the Dead at the end of each month.

, Summon ghost warriors [3201], Necromancy [3200],]

E.24.5 Raise undead [3202]

Usage: COLLECT 31 [days] Time to learn: 28 days. Time to use: Variable.

Undead [31] are dead bodies animated with foul magic, summoned from their graves. Being already dead, they fight with no regard for their personal safety, making them dangerous to fight. In battle they fight with combat values of (25, 1).

One undead may be summoned per day from a graveyard. Graveyards furnish up to 30 undead each month.

One or two undead may decompose at the end of each month. , Raise undead [3202], Necromancy [3200],]

E.24.6 Summon demon lord [3231]

Usage: USE 3231 [aura] Time to learn: 28 days. Time to use: 7 days.

Summon a demon lord unit from a graveyard. The demon lord will appear with 15-25 undead and stack with the summoner.

Each graveyard will support the summoning of one demon lord. No further demon lord units may be summoned from a graveyard until the graveyard’s demon lord has been destroyed or banished.

The amount of aura invested in the summoning determines the strength of the demon lord. The default aura is three points.

aura attack/defense —- ————– 3 (100,100) 4 (150,150) 5 (190,190) 6 (220,220) 7 (240,240) 8 (250,250)

Each undead accompanying the demon lord fights with attack-25, defense-1.

The demon lord’s service is secured with a bond which lasts five months, unless renewed. Demon lords should be banished or have their bonds renewed before the five months expires, or they may exercise their wrath on the summoner.

Note that bonds expire on the 30th day of the month, regardless of when they were cast. The remainder of the month in which the spell was cast counts as the first month. So a summoning on the 15th day of the first month of the year will expire after the 30th day of the fifth month.

Other Necromatic spells provide for banishment of demon lords and renewal of their loyalty bonds.

, Summon demon lord [3231], Necromancy [3200],]

E.24.7 Renew demon bond [3232]

Usage: USE 3232 <target> Time to learn: 28 days. Time to use: 7 days. Required aura: 3

Renew the loyalty bond for a demon lord. This spell will cause the demon lord to remain stacked with its summoner for an additional four months from the time of the casting. Demon lord loyalty bonds accumulate to a maximum of eight months.

Note that bonds expire on the 30th day of the month, regardless of when they were cast.

It is wise to keep a demon lord’s loyalty bond renewed, lest it erode completely, and the demon unit turns on its master.

, Renew demon bond [3232], Necromancy [3200],]

E.24.8 Banish demon lord [3233]

Usage: USE 3233 <target> Time to learn: 28 days. Time to use: 7 days. Required aura: 6

Return an demon lord to the Land of the Dead. The demon lord unit must be present in the caster’s location.

Once banished, a summoner may again call the demon lord from his home graveyard.

, Banish demon lord [3233], Necromancy [3200],]

E.24.9 Eating of the dead [3234]

Usage: USE 3234 <dead body> Time to learn: 28 days. Time to use: 14 days. Aura required: 5 Requires: one ratspider venom [63]

A complex ritual is executed whereby the mage ingests the ratspider venom together with the brains of the dead noble. The mage will learn every skill (other than religion skills) previously known by the dead noble. (Note that the mage does not incur the normal NP costs of learning skills, should those apply.)

There is a 25% chance that the toxic venom of the ratspider will cause the mage to become ill.

, Eating of the dead [3234], Necromancy [3200],]

E.24.10 Aura blast [3203]

Usage: USE 3203 <target> [aura] [have-left] Time to learn: 28 days. Time to use: 1 day. Required aura: as given

The mage directs a blast of raw energy at the target character. The target will suffer two points of damage for each point of aura used in the blast. For example, USE 3203 5201 25 would send a blast of 25 aura at [5201], causing a 50 point wound.

If aura is zero, all available current aura will be used. If have-left is given, at least that much current aura will be retained, and not used for the blast

, Aura blast [3203], Necromancy [3200],]

E.24.11 Absorb aura blast [3235]

Usage: USE 3235 <flag> Time to learn: 28 days.

Knowledge of 3235 will, by default, cause the mage to absorb any blasts of aura instead of being harmed by them. The mage absorbing the blast gains 1/2 of the aura originally used in the blast.

Instead of absorbing aura, the blast may be reflected back to the attacker. To set this option, specify a flag of 1:

use 3235 1 # reflect aura blasts back at attacker use 3235 0 # absorb aura blasts into current aura Aura blasts may only be reflected once. If a blast is reflected back to an

attacker with knowledge of 3235, the attacker will be harmed, and will not absorb or reflect the blast.

, Necromancy [3200],]

E.24.12 Create Flesh Golem [3237]

Usage: USE 3237 <body> Time to learn: 21 days. Time to use: 14 days. Aura to use: 15.

By use of this spell, the magician creates a magical, animated being out of a corpse. The flesh golem weighs 750 units, can carry 2000 units of weight and fights at (250,250). <body> should be the id number of the corpse to be animated.

Golems will only serve their creator and cannot be transferred between nobles. Golems are not counted when determining if a noble is "alone" (e.g., ⟨undefined⟩ [Meditate [2801]], page ⟨undefined⟩). Golems are commanded in battle as if they were men (i.e., they are not "beasts").

Because flesh can only contain the powerful magic for so long before being destroyed, flesh golems last only about twelve months before suddenly dissolving.

, Necromancy [3200], Appendix - Skill Dictionary,]

E.25 Advanced sorcery [3300]

The blackest of the black arts, the tool of the most powerful mages, Advanced sorcery brings deep knowledge of the workings of the mortal and spiritual realms, and the ability to alter them at will.

Advanced sorcery may only be learned by forging an auraculum.

The following unteachable skills may be studied directly once Advanced sorcery is known:

    num    skill                              time to learn
    ---    -----                              -------------
    3330   Trance                             four weeks
    3331   Teleport items                     four weeks
    3332   Fireball                           four weeks
    3333   Create Iron Golem                  four weeks

, Advanced sorcery [3300], Advanced sorcery [3300],]

E.25.1 Fireball [3332]

Usage: Automatic in combat Time to learn: 28 days. Time to use: Automatic in combat Aura to use: 10. This is a combat spell.

Fireball is a combat spell. During the special attacks phase of the combat round, the wizard creates a massive fireball which arcs high above the battle to fall on the hapless opposing troops. The fireball acts as fifty 500 point attacks against the opposing front line.

, Fireball [3332], Advanced sorcery [3300],]

E.25.2 Trance [3330]

Usage: TRANCE Time to learn: 28 days. Time to use: 28 days.

Causes the mage enters a deep trance which lasts nearly an entire month. Upon completition, the mage is cured of any illness or injury, and his current aura will be restored to its maximum.

, Trance [3330], Advanced sorcery [3300],]

E.25.3 Teleport items [3331]

Usage: USE 3331 <to-who> <item> [qty] [have-left] Time to learn: 28 days. Time to use: 3 days. Required aura: three points + one per 250 weight

Teleport quantity of an item in inventory to another character. Specifying zero for quantity will cause the unit to teleport all of its inventory of the item. An amount specified by have-left will be retained.

For example, if Feasel has 100 gold [1]:

           USE 3331 5499 1 50          #  Send 50 gold to [5499]
           USE 3331 5499 1 0           #  Send all 100 gold to [5499]
           USE 3331 5499 1 100         #  Send all 100 gold to [5499]
           USE 3331 5499 1 100 25      #  Send 75 gold to [5499]

            USE 3331 5499 1 0 25        #  Send 75 gold to [5499]

Three points of aura are required to teleport items, plus one extra point of aura for each 250 units of weight to be sent.

Use of 3331 is subject to the same permission checks as ⟨undefined⟩ [give], page ⟨undefined⟩.

, Advanced sorcery [3300],]

E.25.4 Create Iron Golem [3333]

Usage: USE 3333 Time to learn: 28 days. Time to use: 14 days. Aura to use: 20. Requires: 100 iron [79]

By use of this spell, the magician creates a magical, animated being out of a purified. The iron golem weighs 2500 units, can carry 5000 units of weight and fights at (500,500).

Golems will only serve their creator and cannot be transferred between nobles. Golems are not counted when determining if a noble is "alone" (e.g., ⟨undefined⟩ [Meditate [2801]], page ⟨undefined⟩). Golems are commanded in battle as if they were men (i.e., they are not "beasts").

By means of a special purifying rite, the magician using this spell creates ultra-pure iron that can contain magical energies nearly indefinitely, so that iron golems do not normally self-destruct.

, Appendix - Skill Dictionary,]

E.26 Heroism [3400]

Time to learn: 14 days. Cost: 1 noble point

Heroism gives a noble the skills he needs to survive as a deadly one-man combat force.

The following teachable skills may be studied directly once Heroism is known:

num    skill                              time to learn
---    -----                              -------------
3401   Swordplay   two weeks
3402   Defense   two weeks
3403   Personal Fight to the Death   two weeks

The following guild skills may be studied by guildmembers:

num    skill                              time to learn
---    -----                              -------------
3480   Avoid Wounds   two weeks
3481   Avoid Illness   two weeks
3482   Improved Recovery   two weeks
3483   Survive fatal wound   two weeks
3484   Forced March   two weeks
3485   Extra Melee Attacks   two weeks
3486   Extra Missile Attacks   two weeks
3487   Acute Senses   two weeks
3488   Improved Explore   two weeks
3489   Uncanny Accuracy   two weeks
3490   Blinding Speed   two weeks

, Heroism [3400], Heroism [3400],]

E.26.1 Swordplay [3401]

Usage: USE 3401 Time to learn: 14 days. Time to use: 3 days. Requires:

Use of Swordplay will cause the noble to spend time and effort practicing swordplay, for use in personal combat. Every three-day use will increase the noble’s attack rating by one point.

, Swordplay [3401], Heroism [3400],]

E.26.2 Defense [3402]

Usage: USE 3402 Time to learn: 14 days. Time to use: 3 days. Requires:

The noble will spend time and effort practicing skills and techniques of personal and combat defense. Each use of Defense raises the noble’s defense rating by one point.

, Defense [3402], Heroism [3400],]

E.26.3 Survive Fatal Wound [3483]

Usage: Automatic Time to learn: 14 days.

This skill permits a noble to survive a wound which would be normally be fatal. If the noble’s health reaches zero, either through combat or by other harm, instead of dying the noble will be restored to full health.

Once the skill has been used, it is forgotten. To survive a second fatal wound, the noble must learn the skill again.

, Survive Fatal Wound [3483], Heroism [3400],]

E.26.4 Avoid Wounds [3480]

Usage: Automatic Time to learn: 14 days. Time to practice: 7 days. Practice cost: $10

Avoid Wounds enables a hero to reduce the damage he takes when hit in a combat. For each week of practice, the hero reduces the damage he will take when hit in combat by 2%, to a maximum reduction of 80%.

The hero may only use this skill when he is alone. If he has any other men or fighting beasts in his inventory (even if they are not involved in the battle or have been killed in earlier combat rounds) he cannot concentrate on avoiding wounds. However, he is permitted to be stacked with other nobles.

, Avoid Wounds [3480], Heroism [3400],]

E.26.5 Avoid Illness [3481]

Usage: Automatic Time to learn: 14 days. Time to practice: 7 days. Practice cost: $10

Avoid Illness enables a hero to reduce the change of becoming ill when injured in a combat. For each week of practice, the hero reduces the chance of illness by 2%, to a maximum reduction of 80%.

, Avoid Illness [3481], Heroism [3400],]

E.26.6 Improved Recovery [3482]

Usage: Automatic Time to learn: 14 days. Time to practice: 7 days. Practice cost: $10

Improved Recovery increases the chance that a hero will recover from an illness. For each week of practice, the hero improves his chance of recovery by 1%, to a maximum improvement of 50%.

, Improved Recovery [3482], Heroism [3400],]

E.26.7 Personal Fight to the Death [3403]

Time to learn: 14 days. Time to use: 0 days.

Personal Fight to the Death permits a hero to continue battle even after he has been hit. Instead, the hero may specify the level of health at which he will quit the battle.

For example,

    USE 3403 40

indicates that the hero will fight until his health falls below 40.

The hero may only use this skill when he is alone. If he has any other men or fighting beasts in his inventory (even if they are not involved in the battle or have been killed in earlier combat rounds) he will quit the battle when he receives his first wound. However, he is permitted to be stacked with other nobles.

, Personal Fight to the Death [3403], Heroism [3400],]

E.26.8 Forced March [3484]

Time to learn: 14 days. Time to use: 1 day. Cost to use: $10.

The Forced March skill prepares a hero to move faster than normal. After using this skill, the hero’s next non-zero time move will be at a faster pace.

If the hero does not have a mount, then his next move will be at riding speed. For each day he marches at riding speed, the hero will lose 1 point of health. For example, if the hero was moving into a plains province (normally an 8 day move), with Forced March he would make the move in 4 days, at a cost of 4 health points.

If the hero does have mount, then his next move will be at twice riding speed. For each day he forces his mount, there is a 20% chance that the mount will die (although in any case the mount will not die until the move completes).

The forced march succeeds only if the noble is completely alone. If he has any men or fighting beasts in his inventory, or if he is stacked with someone when he moves, then the forced march will fail and the preparation will be used up.

, Forced March [3484], Heroism [3400],]

E.26.9 Extra Melee Attacks [3485]

Usage: Automatic Time to learn: 14 days. Time to practice: 7 days. Cost to practice: $50

For every eight weeks a noble spends practicing Extra Melee Attacks, he receives one extra attack in each melee phase of combat. For example, a hero who had practiced Extra Melee Attacks for 27 weeks would receive four attacks in each melee phase.

The hero may only use extra attacks when he is alone. If he has any other men or fighting beasts in his inventory (even if they are not involved in the battle or have been killed in earlier combat rounds) he cannot concentrate on using extra attacks. However, he is permitted to be stacked with other nobles.

Note that the hero receives his attacks only if he is selected to perform an attack. If, for example, the attacking force greatly outnumbers the defending force, then only some attackers will receive chances to attack.

, Extra Melee Attacks [3485], Heroism [3400],]

E.26.10 Extra Missile Attacks [3486]

Usage: Automatic Time to learn: 14 days. Time to practice: 7 days. Cost to practice: $50

For every eight weeks a noble spends practicing Extra Missile Attacks, he receives one extra attack in each missile phase of combat. For example, a hero who had practiced Extra Missile Attacks for 8 weeks would receive two attacks in each missile phase.

The hero may only use extra attacks when he is alone. If he has any other men or fighting beasts in his inventory (even if they are not involved in the battle or have been killed in earlier combat rounds) he cannot concentrate on using extra attacks. However, he is permitted to be stacked with other nobles.

, Extra Missile Attacks [3486], Heroism [3400],]

E.26.11 Acute Senses [3487]

Usage: Automatic Time to learn: 14 days. Time to practice: 7 days.

Acute Senses enables a hero to avoid traps.

The hero has a 15% chance plus 2% for every week Acute Senses is practiced (to a maximum of 80%) of avoiding a trap when entering a province.

, Acute Senses [3487], Heroism [3400],]

E.26.12 Improved Explore [3488]

Usage: Automatic Time to learn: 14 days. Time to practice: 7 days.

This skill improves the hero’s chances of discovering a hidden object or feature when using ⟨undefined⟩ [explore], page ⟨undefined⟩.

Each week of practice in this skill improves a hero’s chance of success by 1%, to a maximum of +50%.

, Improved Explore [3488], Heroism [3400],]

E.26.13 Uncanny Accuracy [3489]

Usage: Automatic Time to learn: 14 days.

A hero with this skill will target the non-noble missile forces in the back row of the opposing force. If there are no missile forces, the hero will target his missile attack against the melee troops.

, Heroism [3400],]

E.26.14 Blinding Speed [3490]

Usage: Automatic Time to learn: 14 days. Prerequisites: 3485 3486 Requires: 1 warmount [53]

This skill grants a hero an extra melee attack during the special attacks phase. The hero rides out across the battlefield on his warmount and attacks the front row of the opposing forces with his normal melee attack.

The hero may only use this attack when he is alone. If he has any other men or fighting beasts in his inventory (even if they are not involved in the battle or have been killed in earlier combat rounds) he cannot use this attack. However, he is permitted to be stacked with other nobles.

Note that the hero does not receive Extra Melee Attacks during this attack.

Appendix F Appendix - Magical Artifacts

In ages past, mighty wizards created a variety of magical artifacts that are beyond the capability of today’s mages. Nearly indestructible, most of these artifacts retain their powers even unto the current times. They may be found in ruins, in the possession of monsters, and other such places.

Most of these artifacts are marked with tiny runes indicating their purpose. These runes may be read with the ⟨undefined⟩ [identify], page ⟨undefined⟩ command. Some artifacts (about 20%) are unmarked. Some artifacts (about 20%) have lost their powers. (Note that the name of an artifact will not reveal what it does, although you may change the name of an artifact with the ⟨undefined⟩ [name], page ⟨undefined⟩ command.)

Most artifacts grant a power (e.g., immunity to attacks from chimeras) or provide a specific bonus to a noble’s abilities (e.g., +5 to attack). The amount of the bonus will be shown by the identify command. If a noble carries more than two artifacts of the same type (e.g., a +5 to attack artifact and a +8 to attack artifact), then only the better of the two artifacts applies (e.g., in this case, the noble would receive +8 to his attack, not +13).

Most artifacts have unlimited uses and are applied automatically whenever appropriate. Some artifacts have a limited number of charges and must be explicitly used with the use command (e.g., an orb of scrying).

Unlike previous versions of Olympia, where a noble displayed to other nobles the combat artifacts he carried (e.g., "...wielding the Sword of Arnasuss, +25 attack"), in TAG, artifacts are not visible to other nobles.

The following sections describe the types of artifacts that can be found.

F.1 Combat Artifacts

Combat artifacts grant bonuses to attack or defense during combat. Some artifacts grant specific bonuses (e.g., a +5 bonus to defense against missile attacks), while others grant general bonuses (e.g., a +5 bonus to defense against all attacks). Some artifacts grant the bonus to the noble bearing the artifact, while others grant the bonus to the men he directly commands.

When a combat artifact is identified, it will reveal the bonuses it grants. For example:

8: > identify atq

8: [atq] Combat Artifact, +5 affects: personal melee attack The bonus is added to the noble’s attack or defense value, as appropriate.

For artifacts that apply to the men he commands, the bonus is added to their attack or defense values, as appropriate.

It is possible to have artifacts with multiple bonuses:

8: > identify atq 8

8: [atq] Combat Artifact, +5 bonus affects: personal melee attack, 8: commanded men missile defense

This artifact would give the men commanded by the noble +5 on both his personal melee attacks and on defense for the men in his command against missile attacks.

Only the best artifact for a particular bonus is applied. For example, if a noble had two artifacts, one of which provided a +5 bonus for missile attacks, and the other a +8 bonus for missile attacks, then only the latter would apply.

Here is the list of possible bonuses in a combat artifact:

personal melee attack This bonus improves the user’s melee attack.

personal missile attack This bonus improves the user’s missile attack. If he has no missile attack, it has no effect.

personal special attack This bonus improves the user’s special attacks. If he has no special attack, it has no effect. The bonus only applies to special attacks that use the "attack versus defense" resolution.

personal melee defense This bonus improves the user’s defense against melee attacks.

personal missile defense This bonus improves the user’s defense against missile attacks.

personal special defense This bonus improves the user’s defense against special attacks. The bonus only applies to special attacks that use the "attack versus defense" resolution.

commanded men melee attack This bonus improves the melee attacks of the men the noble directly commands.

commanded men missile attack This bonus improves the missile attacks of the men the noble directly commands.

commanded men special attack This bonus improves the special attacks of the men the noble directly commands.

commanded men melee defense This bonus improves the defense against melee attacks of the men the noble directly commands.

commanded men missile defense This bonus improves the defense against missile attacks of the men the noble directly commands. commanded men special defense

This bonus improves the defense against special attacks of the men the noble directly commands.

Rumors tell of an exceedingly rare artifact called the "Ring of Heroes" that gives a +50 bonus to every aspect of combat.

F.2 Leadership of Men

This artifact increases the number of men a noble can control in combat.

F.3 Beast Enchantment

This artifact increases the number of beasts a noble can control in combat.

F.4 Safety from Attack

This artifact protects the owner from attack by a specific type of monster, e.g., orcs. It only protects the owner from random wilderness attacks, not from attacks by other players, or from elves in Faery.

F.5 Improved Attack

This artifact improves the attack of a specific type of soldier in the owner’s control, e.g., +5 attack for pikemen.

F.6 Improved Defense

This artifact improves the defense for a specific type of soldier in the owner’s control, e.g., +5 attack for pikemen.

F.7 Safety at Sea

If the captain of a boat possesses this artifact, the boat will be protected from all the dangers of sailing, i.e., running aground, capsizing, becoming lost, etc.

F.8 Defensive Terrain Enchantment

This artifact improves the defense of a noble and any men he commands in a particular terrain. The artifact improves their defense by a percentage, e.g., a +15 bonus Defensive Terrain Enchantment (woods) will improve a noble and his men’s defensive values by 15% while they are in woods.

F.9 Fast Terrain

The owner of this artifact and the men he commands move more quickly into the specified terrain. Each bonus point reduces travel time by 1 day, a +2 bonus Fast Terrain (woods) artifact speeds travel into woods by 2 days. This artifact affects only movement with the ⟨undefined⟩ [move], page ⟨undefined⟩ and ⟨undefined⟩ [attack], page ⟨undefined⟩ commands, not flying or sailing.

F.10 Speed Use

This artifact reduces the amount of time required to use a particular skill, e.g., -1 day to use Extract venom from ratspider [1530].

F.11 Hellring

This artifact grants the noble personal safety in Hades. The noble will not be attacked by daemons or undead, and he will not suffer the usual lifedraining effects of Hades. This protection does not extend to nobles stacked with the protected noble. The noble (but not others in his stack) is also given free entrance to Hades.

F.12 Elfstone

This artifact grants the noble personal safety in Faery. This protection does not extend to men in his control or to nobles stacked with the protected noble. This artifact may also be used (⟨undefined⟩ [use], page ⟨undefined⟩) to open a faery hill. This artifact has a limited number of uses, and when exhausted, no longer provides protection in Faery.

F.13 Hard Workers

This artifact improves the output of any workers controlled by the owner. The artifact will improve their production by a percentage, e.g., a +15 bonus Hard Workers ring will improve the output of workers by 15%.

F.14 Improved Income

This artifact will improve the income generated by a structure (e.g., a temple or an inn) by a percentage.

F.15 Fast Learning

This artifact reduces the time a noble requires to learn a new skill by one or two days.

F.16 Fast Teaching

This artifact reduces the time a noble requires to teach a new skill to students by one or two days (that is, the students reduce the time they require to learn the skill if the teacher possesses this artifact).

F.17 Fast Training

This artifact allows a noble to train two of a particular unit each day rather than one, e.g., a noble with Fast Training (elite archers) would train two elite archers each day instead of one. Note that the noble must still have the necessary ingredients (e.g., two archers and two elvish arrows).

F.18 Destroy Monster

When used (⟨undefined⟩ [use], page ⟨undefined⟩), this artifact destroys all the monsters of the designated type in the user’s location, e.g., an artifact of Destroy Monster orc would destroy all orcs. This artifact has a limited number of uses.

F.19 Riding

This artifact permits the owner to move at riding speed. It is rated for capacity (between 100 and 500 units of weight).

F.20 Flying

This artifact permits the owner to fly. It is rated for capacity (between 100 and 500 units of weight).

This artifact can only operate continuously for fifteen days, so only fifteen days in a row can be spent flying over the ocean. At the end of the fifteenth day, the artifact will cease to function, and the entire stack will plunge into the sea, killing all of the stack’s nobles.

Any arrival at a land or ship location will allow the artifact to be stopped and restarted, renewing the fifteen day limit.

F.21 Carry Great Loads

This artifact permits the owner to carry additional weight at walking speeds. It is rated for capacity between 100 and 500 units of weight.

F.22 Protection from Skill

This artifact protects the owner from uses of a particular skill. For example, a noble in possession of "Protection from Skill (Receive vision [2434])" could not be visioned.

F.23 Shield Location

This artifact prevents any magical scrying or visions in the owner’s location.

F.24 Power Jewel

When used (⟨undefined⟩ [use], page ⟨undefined⟩), this artifact grants the user additional aura or mana. This artifact has a limited number of uses.

F.25 Summon Aid

When used (⟨undefined⟩ [use], page ⟨undefined⟩), this artifact summons aid of the indicated type and number, adding those units to the owner’s inventory. For example, an artifact "Summon Aid, 12 soldiers" would add 12 soldiers to the user’s inventory. This artifact has a limited number of uses.

F.26 Reduced Maintenance

This artifact reduces the maintenance costs of a noble by the specified percentage, e.g., a +20 bonus Reduced Maintenance artifact reduces maintenance costs by 20%.

F.27 Improve Bargaining

This artifact reduces the price a noble pays for goods in a city or guild market by the specified percentage, e.g., a +20 bonus Improve Bargaining artifact reduces prices by 20%. This applies only to goods sold by the city or guild (not other players) and the discount is applied only at the end of the sale (i.e., not when the market is determining the highest bidder).

F.28 Weightlessness

This artifact reduces the weight of a noble and his inventory by the indicated amount.

F.29 Increased Healing

This artifact increases the amount a noble heals by the indicated amount.

F.30 Protection from Sickness

A noble possessing this artifact will never fall ill.

F.31 Restore Life

One of the most powerful of the ancient artifacts, this artifact restores a noble to life after his death. The noble will be resurrected in the starting city of his nation, and will have no possessions. This artifact has a limited number of charges.

F.32 Teleport

When used (⟨undefined⟩ [use], page ⟨undefined⟩), this artifact teleports the user and his possessions to a location designated by the user, e.g., if [atq] was an artifact of Teleport, then the command "use atq h24" would attempt to teleport the user to location [h24]. This artifact has an indicated weight limited and a limited number of charges.

F.33 Orb of Scrying

This artifact may be used to gain visions of characters, items, or places. Issue the order USE <artifact> <target> to scry the given target with the artifact. The artifact has a limited number of charges.

Unlike magical and religious scrying skills, the Orb of Scrying work across oceans.

F.34 Crown of Control

This artifact may be used to gain control of a stack of monsters. Each artifact is tuned to one particular type of monster and has a limited number of charges. For example, a Crown of Control (orcs) may be used to seize control of a stack of orcs. To seize control of a stack of monsters, the user must be in the same location as the stack and issue the order USE <artifact> <target>.

Once a player has seized control of a stack of monsters, the new unit will appear in his turn report and can be treated in many ways like any other noble. The new unit will be of loyalty "NPC" and will not require pay or any other incentives. However, the commands that controlled monster units may execute are limited to the following: accept admit attack banner behind contact default defend drop execute flag get give hostile move maxpay name neutral pay pillage promote raze seek stack stop take terrorize unstack wait

In combat, a controlled monster unit can command the same number of men as a garrison unit (i.e., 20 men on attack and 40 on defense). A controlled monster unit can command any number of beasts of the same type, but none of any other type (e.g., an orc unit could command any number of orcs, but no centaurs).

F.35 Staff of the Sun

The Staff of the Sun is a mythical artifact of great power. Broken into eight pieces in a long-forgotten time, it has the power in the hands of priest to utterly destroy all magicians, and in the hands of a magician the power to utterly destroy all priests. Consequently both priests and magicians are actively seeking to rebuild the Staff of the Sun.

If a priest holds all eight pieces of the Staff of the Sun at the end of a month, then the power of the staff will destroy all magicians. Likewise, if a magician holds all eight pieces of the Staff of the Sun at the end of a month, then the staff will destroy all priests. Once it destroys all priests or magicians, the Staff of the Sun destroys itself, although legend holds that it can never be completely destroyed.

The pieces of the Staff of the Sun also grant piety bonuses to priests, help mages regenerate aura, and can be used to discover other pieces of the Staff.

Each month the Times prints a summary of the pieces of the Staff of the Sun:

      Staff of the Sun Summary
      ------------------------
      * Five pieces of the Staff of the Sun are held by monsters.
        * One piece is held somewhere on Mandoria.
        * One piece is held somewhere on Catarack Island.
        * One piece is held somewhere on Imperia.
        * One piece is held somewhere on Imperia.
        * One piece is held somewhere on Mandoria.
      * One piece of the Staff of the Sun is held by priests.
        * One piece is held somewhere on Imperia.
      * One piece of the Staff of the Sun is held by magicians.
        * One piece is held somewhere on Mandoria.

Appendix G Appendix - Command Preprocessing

When you submit orders to TAG, the game can run those orders through a preprocessor for you. The preprocessor can convert orders like

        give Maldor peasants

into (for example)

        give 5499 10

This permits you to write your orders in a more English-like (and perhaps, more error-free) format.

To use this facility, you must include the keyword "preprocess" on the subject line of the mail you send to TAG.

The preprocessor currently includes the following aliases:

#define gold 1
#define golds 1
#define peasant 10
#define peasants 10
#define worker 11
#define workers 11
#define soldier 12
#define soldiers 12
#define archer 13
#define archers 13
#define knight 14
#define knights     14
#define elite_guard 15
#define elite_guards 15
#define paladin 15
#define paladins 15
#define pikeman 16
#define pikemans 16
#define blessed_soldier 17
#define blessed_soldiers 17
#define ghost_warrior 18
#define ghost_warriors 18
#define sailor 19
#define sailors     19
#define swordsman 20
#define swordsmen 20
#define crossbowman 21
#define crossbowmen 21
#define elite_archer 22
#define elite_archers 22

#define angry_peasant 23
#define angry_peasants 23
#define pirate 24
#define pirates 24
#define elf 25
#define elves 25
#define spirit 26
#define spirits 26
#define postulant 27
#define postulants 27
#define fanatic 28
#define fanatics 28
#define ninja 29
#define ninjas 29
#define angel 30
#define angels 30
#define corpse 31
#define corpses 31
#define savage 32
#define savages 32
#define skeleton 33
#define skeletons 33
#define barbarian 34
#define barbarians 34
#define wagon 35
#define wagons 35
#define skirmisher 36
#define skirmishers 36
#define hvy_foot 37
#define hvy_foots 37
#define hvy_feet 37
#define hvy_xbowman 38
#define hvy_xbowmans 38
#define hvy_xbowmen 38
#define elvish_arrow 39
#define elvish_arrows 39
#define hvy_xbow 40
#define hvy_xbows 40
#define horse_archer 41
#define horse_archers 41
#define cavalier 43
#define cavaliers 43
#define new_wagon 44
#define new_wagons 44

#define wild_horse 51
#define wild_horses 51
#define riding_horse 52
#define riding_horses 52
#define warmount 53
#define warmounts 53
#define pegasus 54
#define pegasi 54
#define pegasuses 54
#define nazgul     55
#define nazguls     55
#define flotsam     59
#define flotsams 59
#define battering_ram 60
#define battering_rams 60
#define catapult 61
#define catapults 61
#define siege_tower 62
#define siege_towers 62
#define ratspider_venom 63
#define ratspider_venoms 63
#define lana_bark 64
#define lana_barks 64
#define avinia_leaf 65
#define avinia_leafs 65
#define avinia_leaves 65
#define spiny_root 66
#define spiny_roots 66
#define farrenstone 67
#define farrenstones 67
#define yew 68
#define yews 68
#define elfstone 69
#define elfstones 69
#define mallorn_wood 70
#define mallorn_woods 70
#define pretus_bones 71
#define longbow 72
#define longbows 72
#define plate 73
#define plates 73
#define longsword 74
#define longswords 74
#define pike 75

#define pikes 75
#define ox 76
#define oxs 76
#define oxes 76
#define oxen 76
#define lumber 77
#define lumbers 77
#define stone 78
#define stones 78
#define iron 79
#define irons 79
#define leather 80
#define leathers 80
#define ratspider 81
#define ratspiders 81
#define mithril 82
#define mithrils 82
#define gate_crystal 83
#define gate_crystals 83
#define blank_scroll 84
#define blank_scrolls 84
#define crossbow 85
#define crossbows 85
#define rug 86
#define rugs 86
#define fish 87
#define fishs 87
#define fishes 87
#define pepper 88
#define peppers 88
#define pipeweed 89
#define pipeweeds 89
#define ale 90
#define ales 90
#define wool 91
#define wools 91
#define jewel 92
#define jewels 92
#define opium 93
#define opiums 93
#define basket 94
#define baskets 94
#define pot 95
#define pots 95

#define fish_oil 97
#define fish_oils 97
#define drum 98
#define drums 98
#define hide 99
#define hides 99
#define lead 102
#define leads 102
#define fine_cloak 103
#define fine_cloaks 103
#define chocolate 104
#define chocolates 104
#define ivory 105
#define ivorys 105
#define ivories 105
#define cardamom 106
#define cardamoms 106
#define honey 107
#define honeys 107
#define ink 108
#define inks 108
#define licorice 109
#define licorices 109
#define soap 110
#define soaps 110
#define old_book 111
#define old_books 111
#define jade_idol 112
#define jade_idols 112
#define purple_cloth 113
#define purple_cloths 113
#define rose_perfume 114
#define rose_perfumes 114
#define silk 115
#define silks 115
#define incense 116
#define incenses 116
#define ochre 117
#define ochres 117
#define jeweled_egg 118
#define jeweled_eggs 118
#define obsidian 119
#define obsidians 119
#define orange 251

#define oranges 251
#define cinnabar 252
#define cinnabars 252
#define myrhh 253
#define myrhhs 253
#define saffron 254
#define saffrons 254
#define dried_fish 255
#define dried_fishs 255
#define tallow 256
#define tallows 256
#define candles 257
#define candle 257
#define wax 258
#define waxs 258
#define waxes 258
#define sugar 259
#define sugars 259
#define salt 260
#define salts 260
#define glue 261
#define glues 261
#define linen 262
#define linens 262
#define beans 263
#define bean 263
#define walnuts 264
#define walnut 264
#define flax 265
#define flaxs 265
#define flaxes 265
#define flutes 266
#define flute 266
#define cassava 267
#define cassavas 267
#define plum_wine 268
#define plum_wines 268
#define vinegar 269
#define vinegars 269
#define tea 270
#define teas 270
#define centaur 271
#define centaurs 271
#define minotaur 272

#define minotaurs 272
#define spider 278
#define spiders 278
#define rat 279
#define rats 279
#define lion 280
#define lions 280
#define bird 281
#define birds 281
#define lizard 282
#define lizards 282
#define bandit 283
#define bandits 283
#define chimera 284
#define chimeras 284
#define harpie 285
#define harpies 285
#define dragon 286
#define dragons 286
#define orc 287
#define orcs 287
#define gorgon 288
#define gorgons 288
#define wolf 289
#define wolfs 289
#define wolves 289
#define orb 290
#define orbs 290
#define cyclops 291
#define cyclopses 291
#define giant 292
#define giants 292
#define faery 293
#define fairy 293
#define faerys 293
#define fairys 293
#define faeries 293
#define fairies 293
#define seagrass 295
#define seagrasss 295
#define seagrasses 295
#define firewort 296
#define fireworts 296
#define beastnip 297

#define beastnips 297
#define elf_poppy 298
#define elf_poppys 298
#define ironwood 299
#define ironwoods 299
#define kings_fern 300
#define kings_ferns 300
#define moon_palms 301
#define moon_palm 301
#define otter 302
#define otters 302
#define mole 303
#define moles 303
#define bull 304
#define bulls 304
#define eagle 305
#define eagles 305
#define monkey 306
#define monkeys 306
#define monkies 306
#define hare 307
#define hares 307
#define wardog 308
#define wardogs 308
#define sand_rat 309
#define sand_rats 309
#define balrog 310
#define balrogs 310
#define dirt_golem 311
#define dirt_golems 311
#define flesh_golem 312
#define flesh_golems 312
#define iron_golem 313
#define iron_golems 313
#define lesser_demon 314
#define lesser_demons 314
#define greater_demon 315
#define greater_demons 315
#define green_rose 316
#define green_roses 316
#define elf_ear 317
#define elf_ears 317
#define savage_ear 318
#define savage_ears 318

#define nazgul_tail 319
#define nazgul_tails 319
#define centaur_hide 320
#define centaur_hides 320
#define minotaur_hide 321
#define minotaur_hides 321
#define spider_eye 322
#define spider_eyes 322
#define rat_tail 323
#define rat_tails 323
#define lion_mane 324
#define lion_manes 324
#define bird_feather 325
#define bird_feathers 325
#define lizard_tail 326
#define lizard_tails 326
#define bandit_ear 327
#define bandit_ears 327
#define chimera_eye 328
#define chimera_eyes 328
#define harpie_feather 329
#define harpie_feathers 329
#define dragon_scale 330
#define dragon_scales 330
#define orc_scalp 331
#define orc_scalps 331
#define gorgon_liver 332
#define gorgon_livers 332
#define wolf_hide 333
#define wolf_hides 333
#define cyclops_eye 334
#define cyclops_eyes 334
#define giant_tongue 335
#define giant_tongues 335
#define balrog_horn 336
#define balrog_horns 336
#define shipcraft 1000
#define combat 1100
#define stealth 1200
#define beast 123
#define persuasion 1300
#define construction 1400
#define alchemy 1500
#define forestry 1600

#define mining 1700
#define trading 1800
#define ranger 1900
#define religion 3500
#define eres 2000
#define anteus 2100
#define dol 2200
#define timeid 2300
#define ham 2400
#define kireus 2500
#define halon 2600
#define domingo 2700
#define basic 2800
#define weather 161
#define scry 2900
#define gate 3000
#define artifact 3100
#define necromancy 3200
#define adv_sorcery 3300
#define heroism  3400
#define basic_religion  3500
#define forge_aura 3130
#define mage_menial 2802
#define appear_common 2803
#define view_aura 2804
#define quick_cast 2830
#define fortify_castle 1491
#define detect_artifacts 3101
#define reveal_artifacts 3102
#define mutate_artifact 3131
#define conceal_artifacts 3132
#define teleport 3030
#define obscure_artifact 3133
#define strengthen_castle 1492
#define detect_gates 3001
#define jump_gate 3002
#define seal_gate 3031
#define unseal_gate 3032
#define notify_unseal 3033
#define rem_seal 3034
#define reveal_key 3035
#define notify_jump 3036
#define reveal_mage 2831
#define fierce_wind 2030

#define transcend_death 3238
#define tap_health 2832
#define moat_castle 1493
#define deepen_mine 1703
#define wooden_shoring 1790
#define iron_shoring 1791
#define forge_weapon 3134
#define forge_armor 3135
#define forge_bow 3136
#define bind_storm 9134
#define lightning_bolt 2840
#define foresee_defense 2340
#define drain_mana 3039
#define raise_soldiers 3236
#define fireball 3332
#define conceal_nation 1291
#define scry_region 2901
#define shroud_region 2930
#define dispel_region 2931
#define remove_obscurity 3137
#define spot_hidden 1901
#define protect_noble 1990
#define write_basic 2833
#define assassinate 1292
#define find_food 1903
#define write_scry 2932
#define write_gate 3037
#define write_art 3138
#define write_necro 3230
#define bar_loc 2933
#define unbar_loc 2934
#define destroy_art 3103
#define rev_jump 3038
#define prot_blast_8 2735
#define locate_char 2935
#define deep_identify 3104
#define shroud_abil 2834
#define detect_abil 2835
#define detect_scry 2936
#define proj_cast 2937
#define dispel_abil 2836
#define adv_med 2837
#define hinder_med 2838
#define forge_palantir 3139

#define save_proj 2938
#define save_quick 2839
#define summon_ghost 3201
#define raise_corpses 3202
#define undead_lord 3231
#define renew_undead 3232
#define banish_undead 3233
#define eat_dead 3234
#define aura_blast 3203
#define absorb_blast 3235
#define summon_rain 2036
#define summon_wind 2037
#define summon_fog 2038
#define direct_storm 2039
#define dissipate 2031
#define renew_storm 2032
#define lightning 2033
#define seize_storm 2034
#define death_fog 2035
#define banish_corpses 2939
#define trance 3330
#define teleport_item 3331
#define summon_water_elemental 2040
#define write_weather 2041
#define find_mtn_trail 2130
#define obscure_mtn_trail 2131
#define improve_mining 2132
#define conceal_mine 2133
#define protect_mine 2134
#define bless_fort 2135
#define weaken_fort 2136
#define boulder_trap 2137
#define write_anteus 2138
#define detect_beasts  2234
#define snake_trap 2235
#define write_dol 2236
#define find_forest_trail 2330
#define obscure_forest_trail 2331
#define improve_forestry 2332
#define reveal_forest 2333
#define improve_fort 2334
#define create_deadfall 2335
#define recruit_elves 2336
#define write_timeid 2337

#define reveal_vision 2434
#define enchant_guard 2430
#define urchin_spy 2431
#define draw_crowds 2432
#define arrange_mugging 2433
#define write_ham 2435
#define pr_shroud_loc 2436
#define improve_quarry 2530
#define improve_smithing 2531
#define edge_of_kireus 2532
#define create_mithril 2533
#define quicksand_trap 2534
#define write_kireus 2535
#define calm_ap 2630
#define improve_charisma 2631
#define mesmerize_crowd 2632
#define improve_taxes 2633
#define guard_loyalty 2634
#define instill_fanaticism 2635
#define write_halon 2636
#define find_hidden 2730
#define conceal_loc 2731
#define mists_of_conceal 2732
#define create_ninja 2733
#define write_domingo 2734
#define survive_fatal 1193
#define pilot_ship 1001
#define shipbuilding 1002
#define bird_spy 2231
#define fight_to_death 1102
#define capture_beasts 2232
#define use_beasts 1195
#define breed_beasts 2233
#define petty_thief 1201
#define deep_sea 1094
#define bribe_noble 1301
#define catch_horse 1930
#define spy_inv 1202
#define spy_skills 1203
#define spy_lord 1204
#define find_rich 1230
#define torture 1231
#define train_wild 1931
#define train_warmount 1932

#define persuade_oath 1330
#define raise_mob 1302
#define rally_mob 1303
#define incite_mob 1331
#define make_ram 1601
#define make_catapult 1194
#define make_siege 1401
#define extract_venom 1530
#define brew_slave 1531
#define brew_heal 1501
#define brew_death 1502
#define brew_weightlessness 1590
#define add_ram 1095
#define cloak_trade 1232
#define mine_iron 1701
#define mine_gold 1702
#define mine_mithril 1730
#define quarry_stone 1402
#define mine_crystal 1731
#define harvest_lumber 1602
#define harvest_yew 1603
#define record_skill 1532
#define sneak_build 1233
#define archery 1902
#define train_soldiers 1105
#define weaponsmith 1106
#define fishing 1004
#define collect_foliage 1604
#define collect_elem 1503
#define summon_savage 1332
#define keep_savage 1333
#define harvest_mallorn 1630
#define harvest_opium 1605
#define improve_opium 1631
#define lead_to_gold 1533
#define hide_lord 1205
#define train_angry 1304
#define hide_self 1290
#define control_battle  1107
#define attack_tactics  1131
#define defense_tactics  1132
#define combat_discipline 1108
#define train_armor 1192
#define smuggle_goods 1830

#define smuggle_men 1831
#define avoid_taxes 1832
#define build_wagons 1801
#define increase_demand 1802
#define decrease_demand 1803
#define increase_supply 1804
#define decrease_supply 1805
#define hide_money 1890
#define hide_item 1891
#define grow_pop 1390
#define build_city 1490
#define train_knight 1190
#define train_paladin 1191
#define add_sails 1005
#define add_forts 1091
#define add_ports 1006
#define add_keels 1090
#define remove_sails 1096
#define remove_forts 1093
#define remove_ports 1097
#define remove_keels 1092
#define remove_ram 1098
#define brew_fiery 1591
#define create_dirt_golem 2841
#define create_flesh_golem 3237
#define create_iron_golem 3333
#define swordplay 3401
#define defense 3402
#define survive_fatal2 3483
#define avoid_wounds 3480
#define avoid_illness 3481
#define improved_recovery 3482
#define personal_fttd 3403
#define forced_march 3484
#define extra_attacks 3485
#define extra_missile_attacks 3486
#define acute_senses 3487
#define improved_explore 3488
#define uncanny_accuracy 3489
#define blinding_speed 3490
#define heal 3501
#define last_rites 3502
#define resurrect 3530
#define create_holy 3503

#define dedicate_temple 3504
#define pray 3505
#define bless 3506
#define gather_holy_plant 3507
#define write_religion 3508
#define proselytise 3509
#define banish_undead_r 3510
#define prot_blast 3531
#define hinder_med_r 3532
#define scry_r 3533
#define armor  plate
#define clay_pot pot
#define clay_pots  pot
#define lt_foot  swordsman
#define light_foot swordsman
#define heavy_foot hvy_foot
#define sword  longsword
#define swords longsword
#define woven_basket basket
#define woven_baskets  basket
#define control_men  control_battle
#define control_men_in_battle control_battle
#define armour plate
#define forge_armour forge_armor
#define train_armour train_armor
#define defence  defense
#define defence2 defense2
#define foresee_defence foresee_defense
#define defence_tactics defense_tactics
#define light_wagon             44
#define light_wagons            44
#define heavy_wagon             45
#define heavy_wagons            45
#define war_wagon               46
#define war_wagons              46
#define warhorse                warmount
#define warhorses               warmount
#define war_horse               warmount
#define war_horses              warmount
#define mine_gate_crystal       mine_crystal

These aliases replace the first item with the second item. For instance, the alias:

#define prot_blast 3531

causes every occurence of "prot blast" in your commands to be replaced with "3531".

You may add your own aliases at the beginning of your commands. For instance, if you have a noble "Maldor" with the id "5499" you could add the alias

        #define Maldor 5499

at the beginning of your orders to cause "Maldor" to be replaced with "5499".

There are a couple of problems you should be aware of if you use the preprocessor.

First, the aliases for collecting fish, wood, etc., will not work. If you try to give the command "fish" it will be replaced with "87" and you’ll get the error:

    >line 105: unrecognized command: 87

Second, the preprocessor works on your entire turn, which means if you submit the command:

    rumor 1
    All that glitters is not gold.

what you’ll actually see as a rumor is something like:

    rumor 1
    All that glitters is not  1 .

because the preprocessor will replace "gold" with "1".

For advanced users, the preprocessor used is the Gnu C preprocessor.

Appendix H Appendix - Output Formats

The Age of Gods can produce its output in several formats. The output format is controlled by the ⟨undefined⟩ [format], page ⟨undefined⟩ command.

TEXT format is the standard, Olympia-style output. It contains plain text and can be read by any mail reader.

HTML format uses HTML to enhance the turn report with active links and better formatting. It can be read by many mail readers, and is also the format used for the Web-based turn reports.

TAGS format is similar to TEXT format, but contains embedded XMLlike markers. This format is intended read and processed by a computer program.

Finally, the RAW format is completely unprocessed and is intended only for people who want to build their own post-processing tools.

This appendix describes the RAW format and the markers embedded in the TAGS format.

H.1 RAW Format

RAW format is similar to TEXT format with several exceptions.

First, RAW format includes the embedded tags of TAGS format. Second, RAW format is not "line-wrapped". Lines in RAW format may be longer than 80 characters. Third, the RAW format uses the tilde character (~) to indicate a non-breaking space. Finally, each line in RAW format starts with a header. The format of the header is:

    <unit identifier>:<indent>/<secondary indent>:<day>:<line>

<unit identifier> The unit which produced this line of output. Currently the unit identifier is only used in the garrison log.

<indent> The number of spaces to indent this line when displaying it in TEXT format.

<secondary indent> The number of spaces to indent any continuation lines if this line must be "wrapped".

<day> The day of the month for this line of output.

Here is a (contrived) example of a line of RAW format output:

    :0/3:17:Eristan Guardenriver~[n6l], Imperial Empire, duke, with five~workers, eight~sailors, arrived from [x1x] Imperial Plains.

and here is how it might be rendered in TEXT format:

    17: Eristan Guardenriver [n6l], Imperial Empire, duke, with, five workers,
    17:   eight sailors, arrived from [x1x] Imperial Plains.

H.2 TAGS Format

In TAGS format, XML-like "tags" are inserted into the turn report. Each tag is enclosed in angle brackets and begins with "tag" or "/tag". For example:

<tag type=game_totals players=32 controlled=171 other=10,999>
Game totals:
Players:                 32
Controlled units:       171
Other units:          10,999
</tag type=game_totals>

This example contains two tags, and illustrates some important features:

1. Note the matching "<tag...>" and "</tag...>". The matching tags enclose the text description of the same information. Not every "<tag...>" has a balancing "</tag...>".

2. The type of tag is indicated by the "type=game totals" within the tag. Every tag has a type keyword.

3. Additional keywords ("players", "controlled" and "other") provide additional information, corresponding to the information in the enclosed text block.

4. Game identifiers in tags are printed in integer format. For example, the identifier "[aa0b]" will print as "36121". To convert back and forth between integer and "license plate" format, see the conversion tools available on the web site.

The following table summarizes the current tags:

Tag  Encloses? Description

tab No Tab to indicated column in text output.

turn Yes Entire turn report.

box Yes A reference to a game entity.

unit section Yes Unit summary section of report.

player summary Yes Player summary section of report.

account summary Yes Account summary section of report.

game totals Yes Game total section of report.

faction summary Yes Faction summary section of report. faction entry No

Entry  in  faction

summary.

header Yes A header line for text output.

loc summary Yes Location summary section of report.

capacity summary Yes Capacity summary section of report.

capacity unit No Entry in capacity summary.

ship report  Yes  Ship report section.
ship  Yes  Entry in ship report.
garr report  Yes  Garrison report section.
garrison  Yes  Entry in garrison report.
garr inv  Yes  Garrison inventory.

hrule No Insert a horizontal rule across page.

unit report  Yes  Unit report.
loc report  Yes  Location report.
order template  Yes  Order template.

command Yes Use of command in unit report.

Table of Contents

Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

1.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.2 Important terms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 Game Mechanics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

1.3.1 Scenarios. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.3.1.1 The Age of Gods Scenario . . . . . . . . . . . . . . 5 1.3.1.2 Novice Scenario . . . . . . . . . . . . . . . . . . . . . . . . 6 1.3.1.3 The Never-Ending Kingdom Scenario …. 6 1.3.1.4 The Twelve Kings Scenario . . . . . . . . . . . . . 6

1.3.2 Alternate Rules. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.3.2.1 Mage/Priest Antipathy . . . . . . . . . . . . . . . . . 6 1.3.2.2 Turn Limit . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 1.3.2.3 Autodrop. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

1.3.3 Winning the Game . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 1.3.4 Submitting orders . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

1.3.4.1 Interrupting orders . . . . . . . . . . . . . . . . . . . . . 9 1.3.4.2 Units not controlled by you (yet) . . . . . . . 10 1.3.4.3 Use the order template . . . . . . . . . . . . . . . . 10 1.3.4.4 Be careful . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 1.3.4.5 Failed orders. . . . . . . . . . . . . . . . . . . . . . . . . . 12 1.3.4.6 Player Entity . . . . . . . . . . . . . . . . . . . . . . . . . 12 1.3.4.7 Quitting the Game . . . . . . . . . . . . . . . . . . . . 12 1.3.4.8 More examples. . . . . . . . . . . . . . . . . . . . . . . . 13 1.3.4.9 Order of nobles in a location . . . . . . . . . . . 13 1.3.4.10 Command priority . . . . . . . . . . . . . . . . . . . 14 1.3.4.11 Command precedence within a location

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 1.3.4.12 At the same time . . . . . . . . . . . . . . . . . . . 15

1.3.5 Turn reports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 1.3.6 Turn schedule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 1.3.7 Game Newsletter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 1.3.8 Compensation for errors. . . . . . . . . . . . . . . . . . . . . . . . 17 1.3.9 Cheating . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 1.3.10 Game additions and rule changes . . . . . . . . . . . . . . 18 1.3.11 How do I join? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 1.3.12 How do I quit? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 1.3.13 Email address, etc. . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

2 The Olympia World. . . . . . . . . . . . . . . . . . . . . . . 20

2.1 Olympian geography . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.1.1 Inner Locations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 2.1.2 Inside a City . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 2.1.3 Who else is here? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 2.1.4 More about geography . . . . . . . . . . . . . . . . . . . . . . . . . 22 2.1.5 Holes in the Map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 2.1.6 Hidden routes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 2.1.7 Ocean ports . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 2.1.8 Port Cities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 2.1.9 An example city description . . . . . . . . . . . . . . . . . . . . 25 2.1.10 Wilderness and civilization . . . . . . . . . . . . . . . . . . . . 25 2.1.11 Common Dangers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 2.1.12 Monsters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 2.1.13 Hades . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 2.1.14 Faery . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 2.1.15 The Cloudlands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 2.1.16 Guard monsters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

2.2 Political Geography . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 2.2.1 The Imperial Empire. . . . . . . . . . . . . . . . . . . . . . . . . . . 29 2.2.2 The Mandor Confederacy. . . . . . . . . . . . . . . . . . . . . . . 30 2.2.3 The Pen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

3 Nobles & What They Can Do . . . . . . . . . . . . . 32

3.1 Health . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

3.1.1 Death . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 3.2 Loyalty bonds. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

3.2.1 Converting Nobles to a New Faction . . . . . . . . . . . . 34 3.2.2 When a Noble Deserts . . . . . . . . . . . . . . . . . . . . . . . . . 35 3.2.3 When a Noble Dies . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 3.2.4 When a Player Quits. . . . . . . . . . . . . . . . . . . . . . . . . . . 36

3.3 Skills . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 3.3.1 Skill Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 3.3.2 Studying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 3.3.3 Instruction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 3.3.4 Diminishing returns from study . . . . . . . . . . . . . . . . . 41 3.3.5 Learning Bonuses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 3.3.6 Costs to begin study . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 3.3.7 Experience and Practice. . . . . . . . . . . . . . . . . . . . . . . . 43 3.3.8 Skills not rated for experience . . . . . . . . . . . . . . . . . . 43 3.3.9 Guild Skills . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 3.3.10 Study Points . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 3.3.11 Summary of study . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 3.3.12 Religion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

3.3.12.1 The Eight Religions . . . . . . . . . . . . . . . . . . 45 3.3.12.2 Becoming a Priest . . . . . . . . . . . . . . . . . . . 47 3.3.12.3 Miracles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 3.3.12.4 Piety. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 3.3.12.5 Followers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 3.3.12.6 Special Religious Troops . . . . . . . . . . . . . . 49 3.3.12.7 Holy Symbols. . . . . . . . . . . . . . . . . . . . . . . . 50 3.3.12.8 Holy Plants. . . . . . . . . . . . . . . . . . . . . . . . . . 50 3.3.12.9 Temples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 3.3.12.10 Miscellaneous Notes . . . . . . . . . . . . . . . . . 51

3.3.13 Magic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 3.3.13.1 Aura . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 3.3.13.2 Magician status . . . . . . . . . . . . . . . . . . . . . . 52 3.3.13.3 Required Items . . . . . . . . . . . . . . . . . . . . . . 53 3.3.13.4 Study of Spells . . . . . . . . . . . . . . . . . . . . . . 53 3.3.13.5 Schools of Magic . . . . . . . . . . . . . . . . . . . . . 53 3.3.13.6 Magical Combat . . . . . . . . . . . . . . . . . . . . . 54

3.4 Training men . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 3.4.1 Maintenance cost . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 3.4.2 Making weapons and armor . . . . . . . . . . . . . . . . . . . . 57

4 Movement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

4.1 Stacking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

4.1.1 Allowing stacking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 4.2 Carrying capacity and Travel. . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

4.2.0.1 Swamps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

4.3 Ships . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 4.3.1 Making ships . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 4.3.2 Ship Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

4.3.2.1 Rowing Ports . . . . . . . . . . . . . . . . . . . . . . . . . 62 4.3.2.2 Sails . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 4.3.2.3 Keels. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 4.3.2.4 Ship Fortifications. . . . . . . . . . . . . . . . . . . . . 65

4.3.3 Repairing a Ship. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 4.3.4 How to sail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 4.3.5 Operating a ferry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 4.3.6 Ferry synchronization . . . . . . . . . . . . . . . . . . . . . . . . . . 67

5 Battles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

5.1 Resolution of battle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 5.2 Front and rear . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 5.3 Limits to Missile Troops . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 5.4 Limits to Melee Troops . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 5.5 Allies. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 5.6 Who Participates in Combat? . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 5.7 Declared attitudes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 5.8 Combat Bonuses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

5.8.1 Fortifications . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 5.8.2 Terrain Effects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 5.8.3 Item bonuses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 5.8.4 Skill Bonuses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 5.8.5 Other Bonuses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

5.9 Combat Spells . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 5.9.1 Selection of Spells . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 5.9.2 Round Two Rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 5.9.3 Effects of Combat Spells . . . . . . . . . . . . . . . . . . . . . . . 80

5.10 Panic in Battle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 5.11 Prisoners and Looting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 5.12 Prisoner escapes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 5.13 Combat and Movement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 5.14 Combat Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

6 Locations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

6.1 Buildings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 6.1.1 The first tower you build is free . . . . . . . . . . . . . . . . . 85 6.1.2 Guild Halls. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 6.1.3 Inns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 6.1.4 Castles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 6.1.5 Mines and Mining . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

6.1.5.1 Mining . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 6.1.5.2 The Dangers of Mining . . . . . . . . . . . . . . . . 88

6.2 Control of Locations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88

6.2.1 Control of Sublocations . . . . . . . . . . . . . . . . . . . . . . . . 89 6.2.1.1 Permission to Enter . . . . . . . . . . . . . . . . . . . 89 6.2.1.2 Entrance Fees. . . . . . . . . . . . . . . . . . . . . . . . . 90

6.2.2 Control of Provinces . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 6.2.2.1 Garrisons. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 6.2.2.2 Garrison reports . . . . . . . . . . . . . . . . . . . . . . 92 6.2.2.3 Referring to garrisons . . . . . . . . . . . . . . . . . 93 6.2.2.4 Status. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 6.2.2.5 Province Control . . . . . . . . . . . . . . . . . . . . . . 93 6.2.2.6 Province Entrance Fees . . . . . . . . . . . . . . . . 94 6.2.2.7 Taxes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 6.2.2.8 Province Tax . . . . . . . . . . . . . . . . . . . . . . . . . 94 6.2.2.9 Market Taxes . . . . . . . . . . . . . . . . . . . . . . . . . 94

7 Economics & Growth . . . . . . . . . . . . . . . . . . . . . 96

7.1 Peasants. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 7.1.1 Population Growth . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 7.1.2 Creating Wealth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 7.1.3 Taxing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 7.1.4 Pillaging . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 7.1.5 Opium . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98

7.2 Markets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 7.2.1 Market report . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 7.2.2 City purchasing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 7.2.3 Resolution of trades . . . . . . . . . . . . . . . . . . . . . . . . . . 101 7.2.4 Partial trades . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 7.2.5 Sales Tax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 7.2.6 More market examples . . . . . . . . . . . . . . . . . . . . . . . . 101 7.2.7 Trade goods. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102

8 Miscellaneous . . . . . . . . . . . . . . . . . . . . . . . . . . . 104

8.1 Hints. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 8.2 Offenses against realism . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 8.3 Credits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

Appendix H Appendix - Reference Sheets… 106

Appendix H Appendix - Mythology of Olympia . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110

Appendix H Appendix - History . . . . . . . . . . . 113 Appendix H Appendix - Listing of all commands

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115

accept . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 admit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 attack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 banner . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 behind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 board . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 border . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 breed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 bribe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 build . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 buy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 catch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 claim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 collect. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 contact. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 decree . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 dedicate. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 default. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 defend . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 drop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 email . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 execute. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 explore. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 fee . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 ferry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 fish . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 134 flag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 fly . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 forget . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 form . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 garrison. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 give . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 guard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 guild . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 honor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 hostile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 identify. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 make . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 message. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 move . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 maxpay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 neutral. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 notab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 oath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 password. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 149 pay . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 pillage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 players. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 post . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 press . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 practice. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 promote. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 quarry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 raze . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 155 recruit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 repair . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 resend . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 rumor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 sail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 seek . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 sell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 split . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 stack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 stop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 study . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 take . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 tax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 teach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 terrorize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 train . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 ungarrison . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 unload . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 unstack. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 use . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 email . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 wait . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 wood . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 yew . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

Appendix H Appendix - Skill Dictionary …. 174

H.1 Shipcraft [1000] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 180 H.1.1 Sailing [1001] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 181 H.1.2 Build Hull [1002] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 H.1.3 Navigate Deep Seas [1094] . . . . . . . . . . . . . . . . . . . . 182 H.1.4 Add ram to ship [1095] . . . . . . . . . . . . . . . . . . . . . . . 183 H.1.5 Fishing [1004]. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 H.1.6 Add Sails to Ship [1005] . . . . . . . . . . . . . . . . . . . . . . 184 H.1.7 Add Fortification to Ship [1091] . . . . . . . . . . . . . . . 184 H.1.8 Add Rowing Ports to Ship [1006] . . . . . . . . . . . . . . 185 H.1.9 Add Keel to Ship [1090] . . . . . . . . . . . . . . . . . . . . . . 185 H.1.10 Remove Sails from Ship [1096] . . . . . . . . . . . . . . . 186 H.1.11 Remove Fortification from Ship [1093] . . . . . . . . 186 H.1.12 Remove Rowing Ports from Ship [1097] . . . . . . . 187 H.1.13 Remove Keel from Ship [1092] . . . . . . . . . . . . . . . 187 H.1.14 Remove Ram from Ship [1098] . . . . . . . . . . . . . . . 187

H.2 Combat [1100]. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 H.2.1 Fight to the death [1102] . . . . . . . . . . . . . . . . . . . . . 188 H.2.2 Use beasts in battle [1195] . . . . . . . . . . . . . . . . . . . . 189 H.2.3 Construct catapult [1194] . . . . . . . . . . . . . . . . . . . . . 190 H.2.4 Train Swordsmen [1105] . . . . . . . . . . . . . . . . . . . . . . 191 H.2.5 Weaponsmithing [1106] . . . . . . . . . . . . . . . . . . . . . . . 191 H.2.6 Control Men in Battle [1107] . . . . . . . . . . . . . . . . . . 192 H.2.7 Combat Tactics [1131] . . . . . . . . . . . . . . . . . . . . . . . . 192 H.2.8 Defense Tactics [1132] . . . . . . . . . . . . . . . . . . . . . . . . 193 H.2.9 Instill Combat Discipline [1108] . . . . . . . . . . . . . . . 193 H.2.10 Train Armored Troops [1192] . . . . . . . . . . . . . . . . 193 H.2.11 Train Mounted [1190] . . . . . . . . . . . . . . . . . . . . . . . 194 H.2.12 Train Paladin [1191]. . . . . . . . . . . . . . . . . . . . . . . . . 194

H.3 Stealth [1200] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 H.3.1 Conceal Nation [1291] . . . . . . . . . . . . . . . . . . . . . . . . 195 H.3.2 Assassinate Noble [1292] . . . . . . . . . . . . . . . . . . . . . . 196 H.3.3 Petty thievery [1201] . . . . . . . . . . . . . . . . . . . . . . . . . 197 H.3.4 Determine inventory of character [1202] . . . . . . . . 198 H.3.5 Determine skills of character [1203] . . . . . . . . . . . . 198 H.3.6 Determine character’s lord [1204] . . . . . . . . . . . . . . 198 H.3.7 Learn of richest nearby noble [1230] . . . . . . . . . . . 199 H.3.8 Torture prisoner [1231] . . . . . . . . . . . . . . . . . . . . . . . 199 H.3.9 Conceal identity of trader [1232]. . . . . . . . . . . . . . . 200 H.3.10 Sneak into structure [1233]. . . . . . . . . . . . . . . . . . . 200 H.3.11 Conceal faction [1205] . . . . . . . . . . . . . . . . . . . . . . . 201 H.3.12 Conceal self [1290] . . . . . . . . . . . . . . . . . . . . . . . . . . 201

H.4 Persuasion [1300] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 H.4.1 Bribe noble [1301]. . . . . . . . . . . . . . . . . . . . . . . . . . . . 203 H.4.2 Persuade oathbound noble [1330] . . . . . . . . . . . . . . 204 H.4.3 Raise peasant mob [1302] . . . . . . . . . . . . . . . . . . . . . 204 H.4.4 Rally peasant mob [1303] . . . . . . . . . . . . . . . . . . . . . 204 H.4.5 Incite mob violence [1331] . . . . . . . . . . . . . . . . . . . . 205 H.4.6 Summon wild men [1332] . . . . . . . . . . . . . . . . . . . . . 205 H.4.7 Persuade wild men to remain [1333] . . . . . . . . . . . 206 H.4.8 Train angry peasants [1304] . . . . . . . . . . . . . . . . . . . 207 H.4.9 Encourage Population Growth [1390] . . . . . . . . . . 207

H.5 Construction [1400] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 H.5.1 Fortify Castle [1491]. . . . . . . . . . . . . . . . . . . . . . . . . . 208 H.5.2 Strengthen Castle Walls [1492] . . . . . . . . . . . . . . . . 208 H.5.3 Add Moat to Castle [1493] . . . . . . . . . . . . . . . . . . . . 209 H.5.4 Construct siege tower [1401] . . . . . . . . . . . . . . . . . . 210 H.5.5 Stone quarrying [1402]. . . . . . . . . . . . . . . . . . . . . . . . 210 H.5.6 Build City [1490]. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211 H.5.7 Widen Entrance [1494] . . . . . . . . . . . . . . . . . . . . . . . 211

H.6 Alchemy [1500] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 H.6.1 Extract venom from ratspider [1530] . . . . . . . . . . . 212 H.6.2 Make potion of slavery [1531] . . . . . . . . . . . . . . . . . 213 H.6.3 Brew healing potion [1501]. . . . . . . . . . . . . . . . . . . . 213 H.6.4 Brew potion of death [1502]. . . . . . . . . . . . . . . . . . . 214 H.6.5 Make potion of weightlessness [1590]. . . . . . . . . . . 214 H.6.6 Make potion of fiery death [1591] . . . . . . . . . . . . . . 215 H.6.7 Record skill on scroll [1532] . . . . . . . . . . . . . . . . . . . 215 H.6.8 Collect rare elements [1503] . . . . . . . . . . . . . . . . . . . 216 H.6.9 Turn lead into gold [1533]. . . . . . . . . . . . . . . . . . . . . 216

H.7 Forestry [1600] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 H.7.1 Construct battering ram [1601]. . . . . . . . . . . . . . . . 217 H.7.2 Harvest lumber [1602] . . . . . . . . . . . . . . . . . . . . . . . . 218 H.7.3 Harvest yew [1603] . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 H.7.4 Collect rare foliage [1604] . . . . . . . . . . . . . . . . . . . . . 219 H.7.5 Harvest mallorn wood [1630] . . . . . . . . . . . . . . . . . . 220 H.7.6 Harvest opium [1605] . . . . . . . . . . . . . . . . . . . . . . . . . 220 H.7.7 Improve opium production [1631]. . . . . . . . . . . . . . 221

H.8 Mining [1700] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 H.8.1 Build Mine Shaft [1703] . . . . . . . . . . . . . . . . . . . . . . 222 H.8.2 Add Wooden Shoring [1790] . . . . . . . . . . . . . . . . . . 223 H.8.3 Add Iron Shoring [1791] . . . . . . . . . . . . . . . . . . . . . . 223 H.8.4 Mine iron [1701] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 H.8.5 Mine gold [1702] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 224 H.8.6 Mine mithril [1730]. . . . . . . . . . . . . . . . . . . . . . . . . . . 224 H.8.7 Mine gate crystal [1731] . . . . . . . . . . . . . . . . . . . . . . 225

H.9 Trading [1800] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 225 H.9.1 Smuggle Goods [1830] . . . . . . . . . . . . . . . . . . . . . . . . 226 H.9.2 Smuggle Men [1831] . . . . . . . . . . . . . . . . . . . . . . . . . . 227 H.9.3 Avoid Taxes [1832] . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 H.9.4 Build Wagons [1801] . . . . . . . . . . . . . . . . . . . . . . . . . 228 H.9.5 Increase Demand [1802] . . . . . . . . . . . . . . . . . . . . . . 229 H.9.6 Decrease Demand [1803] . . . . . . . . . . . . . . . . . . . . . . 229 H.9.7 Increase Supply [1804] . . . . . . . . . . . . . . . . . . . . . . . . 230 H.9.8 Decrease Supply [1805] . . . . . . . . . . . . . . . . . . . . . . . 231 H.9.9 Hide Money [1890] . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 H.9.10 Hide Item [1891] . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232

H.10 Ranger [1900] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232 H.10.1 Spot Hidden Noble [1901]. . . . . . . . . . . . . . . . . . . . 233 H.10.2 Protect Noble [1990] . . . . . . . . . . . . . . . . . . . . . . . . 234 H.10.3 Find Food in Wilderness [1903] . . . . . . . . . . . . . . 234 H.10.4 Catch wild horses [1930] . . . . . . . . . . . . . . . . . . . . . 235 H.10.5 Train wild horse to riding horse [1931] . . . . . . . . 235 H.10.6 Train wild horse to warmount [1932]. . . . . . . . . . 236 H.10.7 Archery [1902] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236

H.11 Basic Religion [3500] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 H.11.1 Heal [3501] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 H.11.2 Last rites [3502]. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 238 H.11.3 Create Holy Symbol [3503]. . . . . . . . . . . . . . . . . . . 238 H.11.4 Dedicate Temple [3504] . . . . . . . . . . . . . . . . . . . . . . 239 H.11.5 Devotional Prayer [3505]. . . . . . . . . . . . . . . . . . . . . 239 H.11.6 Bless follower [3506] . . . . . . . . . . . . . . . . . . . . . . . . . 240 H.11.7 Gather Holy Plant [3507] . . . . . . . . . . . . . . . . . . . . 240 H.11.8 Create Holy Instruction [3508] . . . . . . . . . . . . . . . 240 H.11.9 Proselytise [3509] . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 H.11.10 Banish undead [3510]. . . . . . . . . . . . . . . . . . . . . . . 242 H.11.11 Resurrect dead noble [3530] . . . . . . . . . . . . . . . . . 242 H.11.12 Protection from Aura Blast [3531] . . . . . . . . . . . 242 H.11.13 Hinder meditation [3532] . . . . . . . . . . . . . . . . . . . 243 H.11.14 Vision of Holy Terrain [3533]. . . . . . . . . . . . . . . . 243

H.12 Church of Eres [2000] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 244 H.12.1 Fierce wind [2030]. . . . . . . . . . . . . . . . . . . . . . . . . . . 245 H.12.2 Summon rain [2036] . . . . . . . . . . . . . . . . . . . . . . . . . 245 H.12.3 Summon wind [2037] . . . . . . . . . . . . . . . . . . . . . . . . 246 H.12.4 Summon fog [2038] . . . . . . . . . . . . . . . . . . . . . . . . . . 246 H.12.5 Direct storm [2039]. . . . . . . . . . . . . . . . . . . . . . . . . . 247 H.12.6 Dissipate storm [2031] . . . . . . . . . . . . . . . . . . . . . . . 247 H.12.7 Renew storm strength [2032] . . . . . . . . . . . . . . . . . 248 H.12.8 Storm fury [2033] . . . . . . . . . . . . . . . . . . . . . . . . . . . 248 H.12.9 Seize control of storm [2034] . . . . . . . . . . . . . . . . . 248 H.12.10 Fog of death [2035]. . . . . . . . . . . . . . . . . . . . . . . . . 249 H.12.11 Summon Water Elemental [2040] . . . . . . . . . . . . 249

H.13 Church of Anteus [2100]. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 H.13.1 Find Mountain Trail [2130] . . . . . . . . . . . . . . . . . . 250 H.13.2 Obscure Mountain Trails [2131] . . . . . . . . . . . . . . 251 H.13.3 Improve Mining [2132]. . . . . . . . . . . . . . . . . . . . . . . 251 H.13.4 Conceal Mine [2133]. . . . . . . . . . . . . . . . . . . . . . . . . 252 H.13.5 Protect Mine [2134] . . . . . . . . . . . . . . . . . . . . . . . . . 252 H.13.6 Bless Building [2135] . . . . . . . . . . . . . . . . . . . . . . . . 253 H.13.7 Weaken Building [2136]. . . . . . . . . . . . . . . . . . . . . . 253 H.13.8 Boulder Trap [2137] . . . . . . . . . . . . . . . . . . . . . . . . . 254

H.14 Church of Dol [2200] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254 H.14.1 Detect Beasts [2234]. . . . . . . . . . . . . . . . . . . . . . . . . 255 H.14.2 Snake Trap [2235] . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 H.14.3 Bird spy [2231] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 256 H.14.4 Capture beasts [2232] . . . . . . . . . . . . . . . . . . . . . . . 256 H.14.5 Breed beasts [2233]. . . . . . . . . . . . . . . . . . . . . . . . . . 257

H.15 Church of Timeid [2300]. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 H.15.1 Find Forest Trail [2330]. . . . . . . . . . . . . . . . . . . . . . 258 H.15.2 Obscure Forest Trails [2331] . . . . . . . . . . . . . . . . . 259 H.15.3 Improve Forestry [2332] . . . . . . . . . . . . . . . . . . . . . 259 H.15.4 Reveal Forest [2333] . . . . . . . . . . . . . . . . . . . . . . . . . 260 H.15.5 Create Dense Thicket [2334] . . . . . . . . . . . . . . . . . 260 H.15.6 Create Deadfall [2335] . . . . . . . . . . . . . . . . . . . . . . . 261 H.15.7 Recruit Elves [2336] . . . . . . . . . . . . . . . . . . . . . . . . . 261

H.16 Church of Ham [2400]. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 H.16.1 Receive vision [2434] . . . . . . . . . . . . . . . . . . . . . . . . 262 H.16.2 Enchant City Guards [2430]. . . . . . . . . . . . . . . . . . 263 H.16.3 Urchin Spy [2431] . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 H.16.4 Draw Crowds [2432] . . . . . . . . . . . . . . . . . . . . . . . . . 264 H.16.5 Arrange Mugging [2433] . . . . . . . . . . . . . . . . . . . . . 264 H.16.6 Shroud location from magical scry [2436]. . . . . . 265

H.17 Church of Kireus [2500] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 265 H.17.1 Improve Quarrying [2530]. . . . . . . . . . . . . . . . . . . . 266 H.17.2 Improve Smithing [2531] . . . . . . . . . . . . . . . . . . . . . 266 H.17.3 Edge of Kireus [2532]. . . . . . . . . . . . . . . . . . . . . . . . 267 H.17.4 Create Mithril [2533] . . . . . . . . . . . . . . . . . . . . . . . . 267 H.17.5 Set Quicksand Trap [2534] . . . . . . . . . . . . . . . . . . . 268

H.18 Church of Halon [2600]. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 268 H.18.1 Calm Angry Peasants [2630] . . . . . . . . . . . . . . . . . 269 H.18.2 Improve Charisma [2631] . . . . . . . . . . . . . . . . . . . . 269 H.18.3 Mesmerize Crowd [2632] . . . . . . . . . . . . . . . . . . . . . 270 H.18.4 Improve Taxes [2633] . . . . . . . . . . . . . . . . . . . . . . . . 270 H.18.5 Guard Loyalty [2634] . . . . . . . . . . . . . . . . . . . . . . . . 271 H.18.6 Instill Fanaticism [2635] . . . . . . . . . . . . . . . . . . . . . 271

H.19 Church of Domingo [2700] . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 H.19.1 Find Hidden Location [2730] . . . . . . . . . . . . . . . . . 272 H.19.2 Conceal Wilderness Location [2731] . . . . . . . . . . 273 H.19.3 Mists of Concealment [2732] . . . . . . . . . . . . . . . . . 273 H.19.4 Train Ninjas [2733] . . . . . . . . . . . . . . . . . . . . . . . . . . 273

H.20 Magic [2800] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 H.20.1 Meditate [2801] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 H.20.2 Perform common tasks for gold [2802] . . . . . . . . 275 H.20.3 Appear common [2803] . . . . . . . . . . . . . . . . . . . . . . 276 H.20.4 View current aura level of others [2804] . . . . . . . 277 H.20.5 Speed casting of next spell [2830]. . . . . . . . . . . . . 277 H.20.6 Reveal abilities of another mage [2831]. . . . . . . . 277 H.20.7 Tap health for aura [2832] . . . . . . . . . . . . . . . . . . . 278 H.20.8 Lightning bolt [2840] . . . . . . . . . . . . . . . . . . . . . . . . 278 H.20.9 Modern magic script [2833] . . . . . . . . . . . . . . . . . . 279 H.20.10 Shroud abilities from scry [2834] . . . . . . . . . . . . 279 H.20.11 Detect ability scry [2835] . . . . . . . . . . . . . . . . . . . 280 H.20.12 Dispel ability shroud [2836] . . . . . . . . . . . . . . . . . 280 H.20.13 Advanced meditation [2837]. . . . . . . . . . . . . . . . . 281 H.20.14 Hinder meditation [2838] . . . . . . . . . . . . . . . . . . . 281 H.20.15 Save speeded cast [2839] . . . . . . . . . . . . . . . . . . . . 282 H.20.16 Create Dirt Golem [2841] . . . . . . . . . . . . . . . . . . . 282

H.21 Scrying [2900] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 283 H.21.1 Foresee Defense [2940] . . . . . . . . . . . . . . . . . . . . . . . 283 H.21.2 Scry location [2901] . . . . . . . . . . . . . . . . . . . . . . . . . 284 H.21.3 Shroud location from magical scry [2930]. . . . . . 284 H.21.4 Dispel location shroud [2931]. . . . . . . . . . . . . . . . . 285 H.21.5 Ciphered writing of Areth-Pirn [2932]. . . . . . . . . 285 H.21.6 Create magical barrier [2933]. . . . . . . . . . . . . . . . . 286 H.21.7 Remove magical barrier [2934] . . . . . . . . . . . . . . . 286 H.21.8 Locate character [2935] . . . . . . . . . . . . . . . . . . . . . . 287 H.21.9 Detect location scry [2936] . . . . . . . . . . . . . . . . . . . 287 H.21.10 Farcasting [2937] . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 H.21.11 Save farcast state [2938] . . . . . . . . . . . . . . . . . . . . 289 H.21.12 Banish undead [2939]. . . . . . . . . . . . . . . . . . . . . . . 289

H.22 Gatecraft [3000] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 289 H.22.1 Teleport [3030]. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 H.22.2 Detect gates [3001] . . . . . . . . . . . . . . . . . . . . . . . . . . 291 H.22.3 Jump through gate [3002]. . . . . . . . . . . . . . . . . . . . 291 H.22.4 Seal gate [3031] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 291 H.22.5 Unseal gate with key [3032] . . . . . . . . . . . . . . . . . . 292 H.22.6 Notify if gate unsealed [3033] . . . . . . . . . . . . . . . . 292 H.22.7 Forcefully unseal gate [3034] . . . . . . . . . . . . . . . . . 293 H.22.8 Reveal gate key [3035] . . . . . . . . . . . . . . . . . . . . . . . 293 H.22.9 Notify of gate jumps [3036] . . . . . . . . . . . . . . . . . . 293 H.22.10 Drain Mana [3039] . . . . . . . . . . . . . . . . . . . . . . . . . 294 H.22.11 Language of the Ancients [3037]. . . . . . . . . . . . . 294 H.22.12 Reverse jump through gate [3038] . . . . . . . . . . . 295

H.23 Artifact construction [3100]. . . . . . . . . . . . . . . . . . . . . . . . . . . 296 H.23.1 Forge auraculum [3130] . . . . . . . . . . . . . . . . . . . . . . 296 H.23.2 Detect Artifact [3101] . . . . . . . . . . . . . . . . . . . . . . . 298 H.23.3 Reveal Artifacts [3102] . . . . . . . . . . . . . . . . . . . . . . 298 H.23.4 Mutate Artifact [3131]. . . . . . . . . . . . . . . . . . . . . . . 298 H.23.5 Conceal Artifacts [3132] . . . . . . . . . . . . . . . . . . . . . 299 H.23.6 Obscure Artifact [3133] . . . . . . . . . . . . . . . . . . . . . . 299 H.23.7 Forge magical weapon [3134] . . . . . . . . . . . . . . . . . 299 H.23.8 Forge magical armor [3135] . . . . . . . . . . . . . . . . . . 300 H.23.9 Forge magical bow [3136] . . . . . . . . . . . . . . . . . . . . 300 H.23.10 Remove Obscurity [3137] . . . . . . . . . . . . . . . . . . . 301 H.23.11 Arcane symbols [3138]. . . . . . . . . . . . . . . . . . . . . . 301 H.23.12 Destroy artifact [3103]. . . . . . . . . . . . . . . . . . . . . . 302 H.23.13 Deep Identify [3104] . . . . . . . . . . . . . . . . . . . . . . . . 302 H.23.14 Forge palantir [3139] . . . . . . . . . . . . . . . . . . . . . . . 303

H.24 Necromancy [3200]. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 303 H.24.1 Transcend death [3238] . . . . . . . . . . . . . . . . . . . . . . 304 H.24.2 Raise soldiers [3236] . . . . . . . . . . . . . . . . . . . . . . . . . 305 H.24.3 Runes of Evil [3230] . . . . . . . . . . . . . . . . . . . . . . . . . 305 H.24.4 Summon ghost warriors [3201] . . . . . . . . . . . . . . . 306 H.24.5 Raise undead [3202] . . . . . . . . . . . . . . . . . . . . . . . . . 306 H.24.6 Summon demon lord [3231] . . . . . . . . . . . . . . . . . . 307 H.24.7 Renew demon bond [3232] . . . . . . . . . . . . . . . . . . . 308 H.24.8 Banish demon lord [3233] . . . . . . . . . . . . . . . . . . . . 308 H.24.9 Eating of the dead [3234] . . . . . . . . . . . . . . . . . . . . 308 H.24.10 Aura blast [3203] . . . . . . . . . . . . . . . . . . . . . . . . . . 309 H.24.11 Absorb aura blast [3235]. . . . . . . . . . . . . . . . . . . . 309 H.24.12 Create Flesh Golem [3237] . . . . . . . . . . . . . . . . . . 310

H.25 Advanced sorcery [3300]. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 H.25.1 Fireball [3332] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 H.25.2 Trance [3330] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 H.25.3 Teleport items [3331] . . . . . . . . . . . . . . . . . . . . . . . . 311 H.25.4 Create Iron Golem [3333] . . . . . . . . . . . . . . . . . . . . 312

H.26 Heroism [3400] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 H.26.1 Swordplay [3401] . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 H.26.2 Defense [3402] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 H.26.3 Survive Fatal Wound [3483] . . . . . . . . . . . . . . . . . . 314 H.26.4 Avoid Wounds [3480] . . . . . . . . . . . . . . . . . . . . . . . . 314 H.26.5 Avoid Illness [3481] . . . . . . . . . . . . . . . . . . . . . . . . . 314 H.26.6 Improved Recovery [3482]. . . . . . . . . . . . . . . . . . . . 315 H.26.7 Personal Fight to the Death [3403] . . . . . . . . . . . 315 H.26.8 Forced March [3484]. . . . . . . . . . . . . . . . . . . . . . . . . 316 H.26.9 Extra Melee Attacks [3485] . . . . . . . . . . . . . . . . . . 316 H.26.10 Extra Missile Attacks [3486] . . . . . . . . . . . . . . . . 317 H.26.11 Acute Senses [3487] . . . . . . . . . . . . . . . . . . . . . . . . 317 H.26.12 Improved Explore [3488] . . . . . . . . . . . . . . . . . . . . 317 H.26.13 Uncanny Accuracy [3489] . . . . . . . . . . . . . . . . . . . 318 H.26.14 Blinding Speed [3490] . . . . . . . . . . . . . . . . . . . . . . 318

Appendix H Appendix - Magical Artifacts . . 319

H.1 Combat Artifacts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 319 H.2 Leadership of Men . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 H.3 Beast Enchantment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 H.4 Safety from Attack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 H.5 Improved Attack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 H.6 Improved Defense. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 H.7 Safety at Sea . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 321 H.8 Defensive Terrain Enchantment . . . . . . . . . . . . . . . . . . . . . . . . 321 H.9 Fast Terrain. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 H.10 Speed Use . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 H.11 Hellring. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 H.12 Elfstone. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 H.13 Hard Workers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 H.14 Improved Income . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 H.15 Fast Learning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 322 H.16 Fast Teaching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 H.17 Fast Training. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 H.18 Destroy Monster. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 H.19 Riding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 H.20 Flying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 H.21 Carry Great Loads. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 H.22 Protection from Skill. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 H.23 Shield Location. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 H.24 Power Jewel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 H.25 Summon Aid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 H.26 Reduced Maintenance. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 H.27 Improve Bargaining . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 H.28 Weightlessness. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 H.29 Increased Healing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 H.30 Protection from Sickness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 H.31 Restore Life . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 H.32 Teleport . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 H.33 Orb of Scrying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 H.34 Crown of Control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 325 H.35 Staff of the Sun . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 327

Appendix H Appendix - Command Preprocessing . . . . . . . . . . . . . . . . . . . . . . . . . . 328

Appendix H Appendix - Output Formats … 345

H.1 RAW Format. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 345 H.2 TAGS Format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 346